A Complete Bibliography of Publications in the Journal of Statistical Planning and Inference:

2010–2019

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

08 November 2023 Version 1.01

Title word cross-reference

(3t, 3, 1) [BG10]. (83, 55, 12) [FK19]. (83, 56, 12) [FK19]. (k_{1}, k_{2}) [SM13]. (p, q) [VT18]. (q, k, λ, t) [WWC10]. 128 [XM10]. 2 [BT19a, BLP11, RH13]. 2^{m} [CKG19, ETMJ11]. 2^{n−m} [CZ10]. 2^{n_{1}} × 3^{n_{2}} [TL11]. 2m [ETMJ11]. 3 [ITA^{+}09]. 32 [WMB10]. 3^{n} [SC10]. 7 ≤ m ≤ 10 [CKG19]. ^{2} [TN14, YU13]. _{1} [BL10b, GF11]. _{2} [CKM11, GJ12, HS11b, Moj12]. _{4} [BGR11]. A [CKM11, Gra12a, SD16]. α [Miy18, WYY^{+}10, WZ10, ZL12]. AR(p) [SZ19]. β [CWcIC11, mFS12]. C [BGR11]. C_{p} [CH11a, LJ12, Oga17]. D [BSH13, BSKS18, BS16, CL13a, CT14, DPA10, DL11, DH11a, DDS13, FK19, Gaf19, GVV10, HKN18, Har14, HHJ13, RS19, Sek12, SX11, TL18b, WZ11b]. δ [AF12]. E [FR13, FHK11]. E(f_{NOD}) [CKMS12]. E(s^{2}) [CDST18, GM12]. E_{max} [WY14]. ℓ_{1} [BC13, LL19, vdG13b]. F [KF12]. G [Imh14, WM18b, XGK16]. h [Nag15]. j [MSTT11]. K [Bar17, Qiu10, Aya12, Beu10, CPH19, DL12, Ery11, GHM11, GM12, GZH11, jLP11, LSWY16, LYLZ12, RS19, SAE11, TSO10, TV12]. K2 × J [HY10a]. k = 4, 5, 6 [WWC10]. L [Ani10, BL10b, GF11, GJ12, HS11b, JWY16, MA08, Moj12, TN14, WMH12, WN11a]. L^{2} [SvdVvZ15]. L_{1} [Qin10, Kol13, Lec13]. L_{2} [EQ15]. L_{p} [Wu19, AJM19]. l_{q} [FN17]. λ = 1 [IT10, ITA^{+}09]. λ = 2 [HYH10]. λ = 3, 4, 6 [CY10]. M [Bra10, CLL11, Cla19, FYX14, RLLT17, SZ19, GK12]. m = 4λ± [ETMJ11]. L_{2} [CCAGV14]. MφE [JMW13]. N [HM13, Beu10, DL12, DLZ14, Ery11, Miy18, SAE11, TV12]. N/4 + 1 ≤ n ≤ 9N/32 [CZ10]. ν = 2n [HYH10]. P [Par10b, TS10, AE16, BS16, Dic13, DLZ14, EO12, FC10, Hab15, MSBM16, RP11a, ZCP13, ZZ14a]. P (Y < X) [Won12]. P [Y < X] [RTM10]. πps [LO11a]. q [Cha10b, Cha10a, Cha12, KV10, QWC12]. R [HY19, LYC14, LYC15, LYXC16, YU13]. s [GCDK12]. S − −S [STS10]. T [Ruk14, ÁG12a, AM11, AAV13, COP10, CPM14, CSC14, HLCW12, JD16, LN15, LVK10, LW11b, MS12b, WSL17]. t_{0} [FPLR11]. × [AM12a, KBPW12, Nik12]. × ⋯ × [KBPW12]. U [BK16, CR16a, Dör10, EQ15, JA17, TJ12, YQC11]. U(3, 3) [CCS14]. UE(s^{2}) [CDST18]. V [DH11a]. φ [JMW13]. W [LZL^{+}10]. Z [Nis11]. Z_{p^{n}} [Fan10].

-ADFs [WWC10]. -Bernstein [Cha10b]. -binomial [Cha10b]. -Boosting [CCAGV14]. -Brownian [ZL12]. -Charlier [KV10]. -circulant [GHM11, GM12]. -consistent [Miy18]. -D [BLP11]. -designs [JA17]. -dimensional [RS19]. -discrepancy [EQ15]. -Discrete [KV10]. -distributions [Cha10a, MS12b]. -divergences [JMW13]. -efficient [SD16]. -error [BL10b]. -estimates [SZ19, WMH12]. -estimation [Nis11, Qin10]. -estimator [CLL11, RLLT17]. -estimators [Bra10, Cla19, FYX14]. -factor [AM11, BT19a]. -formula [LZL^{+}10]. -frame [EO12]. -generalized [GK12]. -kurtosis [WN11a]. -level [GCDK12, QWC12]. -max [RP11a]. -means [RH13, Qiu10]. -Meixner [KV10]. -mixing [mFS12, WYY^{+}10, WZ10]. -mode [Nag15]. -nearest [Aya12]. -neighborhoods [AF12]. -norm [SvdVvZ15]. -optimal [ySD10, BSH13, Bar17, BSKS18, BS16, CL13a, CT14, CKMS12, CDST18, DPA10, DL11, DH11a, DDS13, FK19, FHK11, GVV10, Gra12a, GM12, HKN18, Har14, HHJ13, Imh14, LYC14, LYC15, LYXC16, RS19, Sek12, SX11, SD16, TL18b, WZ11b]. -optimality [CKM11, FR13, Gaf19, HY19]. -out-of- [Beu10, DL12, Ery11, SAE11, TV12]. -penalty [Kol13, Lec13, vdG13b]. -Pólya [Cha12]. -prior [WM18b]. -priors [XGK16]. -process [WSL17]. -projection [GJ12]. -protection [CWcIC11]. -regularized [LL19]. -run [WMB10, XM10]. -sample [LSWY16]. -skewness [WN11a]. -statistic [Ani10, MA08]. -Statistics [TJ12, BK16, CR16a, Dör10, JWY16]. -Step [TSO10]. -systems [BGR11, Ruk14]. -type [EQ15, YQC11]. -value [MSBM16, ZCP13, ZZ14a, Par10b]. -values [AE16, Dic13, Hab15].

0 [Nik12].

1 [Nik12]. 1-100 [Ano16-37, Ano16-38, Ano17-37]. 1-102 [Ano15-34]. 1-104 [Ano15-35]. 1-106 [Ano15-36]. 1-110 [Ano15-37, Ano17-38]. 1-114 [Ano14-34]. 1-116 [Ano17-39]. 1-120 [Ano14-35, Ano16-39]. 1-122 [Ano19h]. 1-124 [Ano15-38, Ano17-40]. 1-126 [Ano15-39]. 1-128 [Ano14-36]. 1-136 [Ano15-40, Ano18m]. 1-140 [Ano14-37, Ano16-40]. 1-142 [Ano17-41]. 1-144 [Ano14-38]. 1-146 [Ano16-41, Ano19i]. 1-158 [Ano14-39]. 1-168 [Ano16-42]. 1-170 [Ano19j]. 1-180 [Ano18n]. 1-184 [Ano18o]. 1-188 [Ano18p]. 1-190 [Ano16-43]. 1-200 [Ano16-44]. 1-218 [Ano14-40]. 1-220 [Ano13-37]. 1-222 [Ano14-41]. 1-224 [Ano19k]. 1-228 [Ano15-41]. 1-238 [Ano19l]. 1-248 [Ano14-42]. 1-334 [Ano10-35]. 1-342 [Ano19m]. 1-378 [Ano12-37]. 1-402 [Ano19n]. 1-48 [Ano15-42]. 1-52 [Ano16-45]. 1-54 [Ano15-43]. 1-58 [Ano17-42]. 1-60 [Ano16-46, Ano17-43, Ano17-44]. 1-610 [Ano11-37]. 1-62 [Ano17-45]. 1-64 [Ano18q]. 1-74 [Ano16-47]. 1-80 [Ano17-46]. 1-94 [Ano16-48, Ano17-47, Ano17-48]. 1001-1108 [Ano13-38]. 1031-1278 [Ano12-38]. 1077-1354 [Ano11-38]. 1101-1352 [Ano10-36]. 1109-1256 [Ano13-39]. 1249-1416 [Ano13-40]. 1273-1626 [Ano12-39]. 1353-1618 [Ano10-37]. 1355-1632 [Ano11-39]. 138 [Ani10]. 139 [IT10]. 140 [BP13a, BP19, BK11c, dLJ12]. 141 [AA11c, GS12a, YS11b]. 1417-1620 [Ano13-41]. 142 [BRB18]. 150 [KZ15]. 16- [WMB10]. 1619-2154 [Ano10-38]. 1621-1834 [Ano13-42]. 1627-2240 [Ano12-40]. 1633-2006 [Ano11-40]. 1835-2038 [Ano13-43]. 1986-1987 [Ano19n]. 1986/87 [Ano19a].

2007-2174 [Ano11-41]. 2010 [Gho11]. 2039-2170 [Ano13-44]. 2175-2502 [Ano11-42]. 221-456 [Ano13-45]. 2241-2484 [Ano12-41]. 2355-2798 [Ano10-39]. 2485-2734 [Ano12-42]. 2503-2986 [Ano11-43]. 2735-2870 [Ano12-43]. 2871-3060 [Ano12-44]. 2937-3576 [Ano10-40]. 2987-3292 [Ano11-44]. 2f [BT19b].

3061-3254 [Ano12-45]. 3293-3366 [Ano11-45]. 335-584 [Ano10-41]. 3367-3680 [Ano11-46]. 36 [KS11b]. 3681-3876 [Ano11-47]. 379-612 [Ano12-46].

457-636 [Ano13-46].

585-850 [Ano10-42].

611-1076 [Ano11-48]. 613-780 [Ano12-47]. 637-866 [Ano13-47]. 6th [Ano10a].

781-1030 [Ano12-48].

851-1100 [Ano10-43]. 867-1000 [Ano13-48].

91 [TLBJ11].

=2 [HM13]. =3 [TS10].

abandonments [DEF11]. ABC [PBRT16]. Aberration [FRR16, HM13, LMT14, LST12, TLM12, XM10]. ability [GO11]. abrupt [WK13]. absolute [SK12]. absolutely [Bou11]. Accelerated [Pas10, LP12, LT10, Wan10a, WLCK11, XT15, ZY11, Zho10]. acceleration [LT10]. acceptance [JM10b, MR11a, SS19]. according [PMW^{+}16]. Accounting [AS11b, LSP12a, SHYH16]. accuracy [CL15b, CMCSC12, GP10, NdCJ10, TXLV11, Woo19]. Accurate [Oga17]. acreage [CP19]. active [BT19a, DL12]. activity [Kon12]. adaptation [MA17, Scr15]. Adaptive [BP13b, CCR17, Efr13, GLT16, HK10, HC11a, KU14, Lia14a, Mab16, MDT^{+}12, NU19, PR12, Pla13, RBRTM11, WZW13, WY14, XI12, YH19, ZGK13, BBB11, BK17, CWH10, CYZL12, CZHC14, Dio14, DK17, GZH11, HJ16, HLH13, LYF14, LW12b, LBH17, LSW13, MS10, SGF12, SCHG19, YY14, YW16, YBB10]. Adaptivity [PY11]. adding [HP10b]. Addition [GSKP10, GCDK12]. additional [FK11b, HK19]. Additive [XSJ14, BLP11, DN13, FZS16, FKMR19, KWL14, LHH16, LLL18, MY11, MSW12, QRGLDMA^{+}12, WX15, YYX19, ZYZL14, ZN13, vdGM15]. additive-multiplicative [QRGLDMA^{+}12]. additivity [SS13c]. ADF [MYW11]. ADFs [WWC10]. Adjusted [WLCK11, Beh12, CZHC14, CT11, ES12, Hab15, Hir19, HLH13]. Adjusting [FRF12]. adjustment [Sch10a, SA10, YH12]. adjustments [TW11]. Admissibility [HP10c]. affine [AS10b, BDLP14]. after [KF12, TCC12, Wu16]. against [Ani10, Ani13, GW11, GG15, MA08]. ageing [CZ12, HKR11]. Aggregation [FM19, Lav11]. aging [FPLR11, MPRG10]. ahead [MF10]. AIDS [AS10a]. al [IT10]. algorithm [BLP11, CLP13, DB19, GS19, Kim16, MK18, MCC12, NPK10, PL14, Qiu10, Sek12, SL19a, SdL10, VB10a, ZCX11]. Algorithmic [LMT14, LYXC16, PZ14]. algorithms [CCAGV14, FRC11, GB19, ZG12]. aliasing [Cha19a]. all-bias [DM10]. allocation [BMPR11, DL12, HDC14, JM13, LBH17, MS10, RQ15, WCZ17]. allows [CSC14]. Almost [Vas13, DDL14, Ros10, WW11a]. alpha [MFABG13, YW10]. alpha-integrability [YW10]. alpha-power [MFABG13]. alphabets [CXCK13]. alternate [CND10]. alternating [Gas11a]. Alternative [NS10b, BHS12, CK13, Jac11, LZL^{+}10]. alternatives [Ani10, Ani13, BS10a, BH15, Ing10, MA08, NN10, PPB16, QM11, Wu19]. American [KK12]. amid [SL14b]. among [BDKM10, Jon13b, jLP11, LZL16]. among-group [LZL16]. analogy [HLY12]. analyses [GS11b, Kat19, TTB11]. Analysis [DDD18, FK11b, GB12, Lin12, Sah11, Alb10a, AM11, BR14, BHS12, BCMRMP11, Beh10, BS11a, Beh12, BLT16, BP14b, BM12b, CLM12, COP10, CCGL10, CHT17, CPSY10, CL15b, CSN19, DH11b, FZS16, Geo14, GC19, GZ14b, Hab13, Han14, HH15, HLS10, HH14c, Huc14, HY10b, HYS12, JBB10, JGJR16, KPR14, Kri10, KSKE10, LCC^{+}17, LD14, LJK12, LC16, Lia12, Lia14b, LL15, LQ16, MA10, RSH12, Ros10, RP11b, SAE11, SS16, Shu12, SSZ11, TH14, TB11, TG16, VN10, WF11, WFG^{+}15, ZLL17]. analytic [DKY17]. Analytical [BE14]. analyzers [AM11]. Analyzing [KMMP11]. ancillary [SF10]. Anisotropic [BPR19]. ANOPOW [SHQG10]. ANOVA [WTZ12, Zha12]. Antieigenvalues [Kha10]. Application [DGCS16, PR12, Ver13, Alo11, AS10a, BT16, BM15, BS10c, CF12a, CC11, CHT17, DDD18, Fon11, FJK11, GLZ14, Hil11, KBPW12, KBK15, Lew18, Lon12, Maj14, OG10, PL10, RF13, SR14, TB11, Wan11, XZ19, ZS10, dLJ10, dLJ12]. Applications [CSR15, Que10, TP10, BP10b, BP13a, BP19, BAHS11, BU10, BP14b, BTV12, BK16, CQ11, CKM11, CPL18, DM10, DG14, DF12b, EGS10a, GRD16, GRM14, HP19, HM11a, Huc14, KCZ13, LJK12, Lon17, Moj12, Oga17, Wag11, WMH12, dOCG12, Ano10a, Ano10-44]. applied [AGPS12, SB19, YW16, Yan17]. Applying [ABK^{+}11]. approach [AVML11, Ani10, BMW16, BS10a, BK11b, BOvSS10, BLW16, Cha19a, Cha19b, Che10a, DDW16, DLZ14, DELS16, FS14, GD10, GRS^{+}12, HH15, HG11, HBP12, KS12, KM10, KL10a, LP12, LBR12, LZL16, LSZ16, LD12, MA08, MO10b, MA10, PMW^{+}16, RBS10, sS10a, TN14, WH12, WT10, WFG^{+}15, XWY18, ZG12]. approaches [DKY17, Kli11, ZGST12]. appropriate [MS10]. Approximate [FHK11, WQ11, GS19, HOK19, Llo10, Zha12, ZK15]. Approximated [CS12a]. approximately [KW14]. Approximating [FJC12]. approximation [AS11a, Alo11, EZS11, KZ14, KZ15, MYW11, RS16, Shu12, Vag10]. Approximations [AP14a, Wu17, BD18, IKU11, MO10a, Miy10, Miy18, MZ13]. approximator [SL19a]. April [Ano10-43, Ano11-39, Ano12-48, Ano13-47, Ano14-42, Ano15-35, Ano16-43, Ano17-45]. arbitrary [BSG16, FKMR19, SLC18]. ARCH [CSW11, SZ12, ZLW10]. Archimedean [GP15a]. Area [FT10, CWcIC11, Chr14, DRT10, DMR10b, Hir19, HC11b, PG12, SM11, TN10]. ARFIMA [PC13]. ARFIMA-GARCH [PC13]. ARMA [ACC18, Bao18, CT10b, WW12b]. armed [Hun12]. Array [HHMO15, BS16, CGLW18, LMW12]. Array-based [HHMO15]. arrays [BP10b, BP13a, BP19, CSWY11, GKM10, LT14a, Lin14, RLZ12, SGS12, Tsa17, WTZ12, YLL14, fZyZD16]. arrivals [PL16]. ARTFIMA [SMM19]. Article [YDZA13, De 13]. ascertainment [GH10]. aspects [AK10, SE19]. assays [MP12b]. Assessing [HG11, LC16, Grö18, Unk17]. assessment [IR19, Ruk17]. assignments [ZW11]. assisted [GT10]. associate [ABW13]. associated [BGK^{+}11, CM16, GD18, LQ11, LQL12, LGG12, PG12, QLL11, SM13, SY18, YW10]. Association [CL10b, Lom11, NS10a, Seb19, sS13b, Unk17, XL12]. associations [Mar18]. assumption [DY19, LM11b, YO11, YKH13, YW10]. assumptions [WYY^{+}10]. asymmetric [BS16, BM15, GZ14a, HY19, IK15, KKP10, LVK10, SZ12, fZyZD16]. Asymptotic [BP10c, BBKP19, BWW19, BP15, BL10b, Bro10, CFT10, CLL11, Cla19, Dem12, DKMdUÁ16, FYX14, Fre14, GMP15, HZ11a, Hu13, HY10b, IYA16, Jar10, KF10, KCZ13, KV10, LWZ13, LNdDLdCH11, MBB11, MO10a, NCC11, Oga13, QM11, SZ12, Shu11, SZ19, SS13d, SdL10, TV12, WC10, Zha13, ZWM12, AF12, Bir15, CS12b, DJ19, DKY15, DKL13, FO13, GB19, HOK19, KM12b, Lem12, Lia14b, LW10, MS11, Miy10, Ogi19, Pav12, QRGLDMA^{+}12, Que10, Shu12]. Asymptotically [Bou11, CR16a, CE11, Men12, RS16, YW10]. Asymptotics [ALT10, Bin14, CSW11, EGS10b, MSBM16, Zho19, GF18, LLJK13, MV15, SR14]. atomic [Sum17]. attainable [Rüs19]. attribute [JM10b]. attributes [JM11b, KGV10, TCE^{+}12]. AUC [CH10]. audit [Göb12]. augmentation [EZ12, GSZ11]. Augmented [SM16, YZZ17, OQ10]. augmenting [ZDLW19]. August [Ano11-43, Ano12-41, Ano13-40, Ano14-35, Ano15-42, Ano16-37, Ano17-41, Ano18o]. August-September [Ano14-35]. Author [Ano19a]. auto [DF10]. auto-regression [DF10]. autocorrelation [Lwi11, NR11, SR10]. autocorrelations [CL13b]. autocovariance [BC19]. autocovariances [LZ13a]. autoregression [Akn15, Lee18]. autoregressions [GP16, Kre16, NL16, PP16a, PP16b]. autoregressive [Cav18, Che11b, GLZ14, Hud13, KMW16, KG10c, LÁPMF19, Lav11, Lwi11, Mac10, MSP10, MR10, PR12, PVB^{+}19, RS16, SN12, UP14, XXWL12, ZK17, ZLL16b]. auxiliary [Bra10, LdUÁ11, LT10, LLZ11b, LLS19, VS12]. Availability [SS00, TLBJ11, TLBJ13]. available [LH14]. average [AG19, BC19, CL13b, DF10, LLX12, PVB^{+}19, Wae10, YSW10]. averaged [BCZ19, GB19]. averages [KS93, vdMC10]. averaging [AS11b, FJKW13, SSL^{+}17, YTHL13, ZX19]. aversion [HC11a]. avoiding [NS10c, Nik12].

B [Fre11a, JKL17]. B-spline [Fre11a, JKL17]. backfitted [MY11]. backward [KZ14, KZ15]. Bahadur [WHY11, WYY^{+}10]. Baker [DKL13]. Balakrishnan [FKV11, KG10b]. balance [FHK11, Gra12a, HYH10]. Balanced [ALLL13, SR10, AJGG10, AA11b, AA11c, ABW13, BC11b, CKM11, CBD11, CKG19, DD13, HP19, HP10c, Nil11, PMK19, RSS19, TY10, VB11, vdMC10]. ball [FGN10, RS19]. ballistic [AMNS11]. ballot [NS10c]. balls [AJM19]. band [YS11b, YS11c]. bandit [Hun12]. bands [CWWT12, CLO17, Fre12, LAk10, SY12, SZ13, WQ10]. bandwidth [Dut11, FWS18, Lin11, Oua13, RS13a]. bandwidth-free [RS13a]. bandwidths [KKP10]. bargraphs [DMR10b]. BART [WK10]. Bartlett [BCN13]. based [AM11, AMRP18, BTKN10, BGK^{+}11, BL10a, Bed10, BET17, BL10b, BP14b, BF13, BC12, BM12b, BK10, BK11c, CCGL10, CND10, CR10, CF12a, CM10, CW10, CRW13, Che10a, CGLW18, CR16a, Coa14, CE11, CMO13, DMTZ15, DDW16, DS10, DM12, DKY15, DLZ14, FRK14, Gas11b, GD10, GMSK10, GBP15, GT10, HHMO15, HGH12, HLY12, Huc14, HOK19, IB11, JB14, JR12, JM10b, Kak18, KDJ13, Kom11, Kon12, KL10a, KX11b, Kri10, KS11b, KV10, LGLO12, LCC^{+}17, LW12a, LZZ15, LLZ11a, LST12, LYLZ12, LYL13, Llo10, LMW12, LW15, LA14, LGG12, MR11a, MSBM16, MBB11, McN10, MRY11, Müh10, NB12b, NB10, Ozt12, PL10, PR11, RM11, SHW16, SDNH13, SD10, Shu11, Shu12, SN10, SM11, SF10, SR14, SZ13, SSDM14, SR13b, TSO10, TNWW12, TJ12, VCA10]. based [VV10, Ver17, VGV11, VSK^{+}11, VGH13, VDW13, VB10b, WTZ12, WY14, WSSG19, WW11b, YSW10, YL10b, YKH13, YLL17, ZZOY13, ZX12, Zha13, ZGST12, ZJ10]. baseline [GLT16, KT17, LMT14, LM17b]. baselines [WK13]. basic [Ste11a]. basis [Cha10b, MLa11]. Basu [Nog13]. Batch [LMW10, KM10]. batches [vdMC10]. Bayes [BP13b, Bir15, CY11, CR15, DRT10, JGJR16, KLW10, KK19, KKP10, LM11b, MS13a, Ogi19, PG10, Ruk14, Ruk17, SG12b, Tor12, Tsu10, VDW13, WS14, XGK16]. Bayes-type [Ogi19]. Bayesian [De 13, HW13, AVML11, Abr12, AAKR12, AS11b, BFT15, BS10a, BHS12, BJQ15, Bas11, BGS10, BSvZ15, BW10b, BW12, BRBM19, BDG14, CLM12, COP10, Cha19a, Che10a, CM15, CO10, CXCK13, DPC11, DH11b, FBHZ14, Fus10, GMSK10, GKvdVvZ15, Gri11, GR18, HvdM12, HW13, Hun12, JGJR16, KL12, KW12, KLKO15, Kli11, Kom11, KKL15, KR12, Lau12, LD14, Lew18, Lia11a, Lia14a, LCY12, LW11b, Mac10, MA17, MDT^{+}12, MRW11, MO10b, NBSB13, O’H13, PR16, PHW15, PBRT16, RSH12, SBCK11, Scr15, SC11, SL14a, SN12, SL19b, SvdVvZ15, Thu14, TN10, TCE^{+}12, VM18, Wal13a, Wal13b, WF11, WW13, WFG^{+}15, XW13a, YQC11, ZPS15, ZSL14, dOCG12]. be [BF12, LS12, WLS13]. behavior [BP10c, BBKP19, BWW19, GG10, HW17, JSV12, KX11a, Per10, YW10]. behaviour [Bir15, Jar10, KV10, PR11, Vas13]. Behrens [CPH19]. Benefits [PR16]. bent [ZL17]. Beran [SL19b]. Berkson [DM11, KS16]. Bernoulli [BTT15, BW10b, Cha10a, Hun12, KK19]. Bernshtein [HH15]. Bernstein [BJQ15, Cha10b, JSV12, Leb12, MPN^{+}17, WSSG19]. Berry [Bou11, HWZ11, YHWL12, ZL11a]. Berry-Esséen [YHWL12]. Besov [AP14b]. Best [MR11b, AS10b, Che10a, Son11a, Tak10]. Beta [PVB^{+}19, BCN13, BK11a, DFM10, GF18, RM12, ZCX11, OJMRCSRA11]. beta-mixing [ZCX11]. beta-Stacy [RM12]. beta-transformed [GF18]. Better [PS14, Kon13, KK19, Tak10]. between [Bas11, BSG16, BDG14, CK13, DKY17, EGS10a, HP10a, MF10, PR16, PBND11, Sar14, Say12, Sch10b]. BH [SCHG19]. Bhattacharya [Ber14]. Bi [DKW17]. Bi-log-concave [DKW17]. Bias [CM16, DMP10, IK15, Kim16, LPY10, SRBS11, SA10, WN11a, WG12, YH12, ÁG12a, BL10a, Bao18, BMNY13, DS18, DM10, FHK11, FRF12, GD15, HG11, LGG12, NYA17, Ruk17, Sai10, SM16]. Bias-adjustment [SA10, YH12]. Bias-corrected [SRBS11, BL10a]. Bias-correction [CM16, Kim16]. Bias-reduced [WN11a, BMNY13]. biased [Azr15, CLL12, CR16b, FSS18, FCC10, GLW11, GZH11, LLZ11a, RV10, WW15b, WZ18]. BIC [LC13, SG11]. bilateral [TQ12]. bilateral-sample [TQ12]. bilinear [FO13, MK18]. binary [AZ13, BM15, CYCZ12, FGH14, GN16, GW11, HKN18, HYH10, HM11b, HH14b, Hud13, KK11, MP12a, MZ13, NdCJ10, Pan10, PD11, Sah11, SM13, SSC11, WW12a, YCB16]. Bingham [LD14]. Binomial [KL10a, Alb10b, CT10a, Cha10b, Che11a, Gam13, Gua12, KP11, Men12, PS14, RAB14c, Sei14, Wan10b, Wan11, XZ19, ZSL14, ZJ10, Zhu12, OJMRCSRA11]. binomials [TRWP10]. binormal [DN12]. biodiversity [BR14]. bioequivalence [HOK19]. biology [BS10c]. biomarker [HCZ19]. bipartite [GR18]. Birnbaum [CLB12, BGK^{+}11, COP10, LF11, Pat12, PK13]. Birnbaum-Saunders [BGK^{+}11, LF11, Pat12, PK13]. Bivariate [LSP12b, Qiu10, AE16, DF12a, FKV11, GMSK10, HvdM12, HN12, HZ18, KG10b, LM16, LH11, NS10a, NS10b, OJMRCSRA11, RR12, sS11, sS13b, SK12, Sum17, Unk17]. Block [GS11a, GS12a, Par10a, AP14a, AM12a, ABW13, BSH13, CY10, CYW10, DIAT11, DD13, FF12, FR13, FMZ10, GW11, GG15, HC16, ITA^{+}09, IT10, ITA^{+}12, Jac19, KME10, LC15, Pav12, SDNH13, SR10, TL18b, Wag11, ZC11, ZW11]. Block-circulant [GS11a, GS12a]. block-decreasing [Pav12]. block-factor [AP14a]. Blocked [YL14, Jac11, Lin14, ZLZK13, ZLL16a, vdVW14]. blocking [OQL11, XM10]. blocks [AA11b, AA11c, GW11, GG15, SYA11a, TL13, ZSA10]. BLUEs [HP10b, Tia13]. Board [Ano10n, Ano10o, Ano10p, Ano10q, Ano11o, Ano10r, Ano10i, Ano10j, Ano10k, Ano10l, Ano10m, Ano11p, Ano11m, Ano11n, Ano17l, Ano17m, Ano17n, Ano17o, Ano18d, Ano18e, Ano18f, Ano18g, Ano18h, Ano19b, Ano19c, Ano19d, Ano19e, Ano19f, Ano19g]. Bon [Sri10]. Boosting [CCAGV14, Zha13]. Bootstrap [ADPW10, GP16, GMMC11, Kre16, NL16, PP16a, PP16b, BPP10, CR13, DKY17, Fus10, KHN16, Kim16, MCC12, RL17, SL14b, WMM^{+}10]. Bootstrapping [GP19, Loh10, OVG13]. bootstraps [BK16]. both [DST10, HY10b]. Botswana [AS10a]. Bound [LL11c, CKM11, EQ15, FRC11, Kha10, LZL^{+}10, YHWL12]. boundaries [CS16a]. boundary [Leb12, Wan10b]. bounded [BR10, CC15, KM12a, WQ11]. Bounding [Kim10]. Bounds [JR12, Bou11, BK10, BK11c, Fro14, Göb12, GR11, HWZ11, LLP14, RRS11, TCC12, ZL11a]. Bowden [DJ19]. brain [CKP17]. Branch [FRC11]. branching [CH11b, GP15b, Kas17]. breakdown [LT14b, TL13]. breaking [CS12b, MRW11]. breaks [CGL11]. Bridge [CK13, PY11, Tan13, ZL12]. brief [SZZ16]. broadband [HM10]. Brownian [ES16, FLL17b, HS11a, SvdV15, SY18, ZL11b, ZL12]. Bühlmann [RS13b, SS13a]. building [Wah10]. Burman [SM16, QL15]. business [SKFMR14]. BVN [NCC11].

calculating [Sek12]. calculations [BCG10]. calculus [NS10c]. Calibrated [RS10, Tan13, FGV12]. calibration [GT10, HLB^{+}16, LL19, SA10, YH12]. call [GRM14]. Can [WLS13]. canonical [Lia14b, Oga13]. capability [vdMC10]. capture [RAB14c, SHYH16]. capture-recapture [RAB14c, SHYH16]. care [RS15]. Carlo [AT10a, AHO13, BCG10, JGK11, Kim10]. CARMA [BH10]. carry [PBND11, Sai10]. carry-over [PBND11, Sai10]. carryover [DT14]. case [BE14, FK11b, GD15, HC16, Lia15, PBRT16, Qiu10, SGL13, ZZOY13, vdMC10]. case-control [GD15]. case-family [ZZOY13]. cases [CWcIC11, KL10a]. Catalan [Kra10]. catalogues [LST12]. categorical [DMP10, GD10, PR16, WFG^{+}15, Wei11]. categorizing [PMW^{+}16]. Cauchy [MV11a]. causal [BVG10, CFW14, KW12, Sch10a, SG12a]. causality [GLZ14, PR12]. cause [HLS12, sS11]. cause-specific [sS11]. caused [GDHS12]. cautionary [HCMP12]. CDT [EMS11]. CDT-optimum [EMS11]. cell [AM12b]. censored [AHD12, BTKN10, BAHS11, Che10a, CC11, CL15b, Che12, CL10b, DF12a, DP11, HN12, IB11, KLS11, KWL14, KX11b, KR12, KKP10, LZ12a, Liu11a, dUJC12, LSP12b, Mad10, Miy11, MRY11, Oua13, PK13, RM11, RH11, Say12, SBCK11, sS10a, sS11, She11, sS12b, sS12c, sS13b, SDNH13, SLZ10, SM11, SZ10, SSL^{+}17, SdL10, VdUÁMM11, VB10b, Wan10a, WTL11, WDRB18, WG12]. censoring [BM11, BCG13, Bur11b, CCR17, CE11, DMP18, GMSK12, GT12b, HZ18, LP12, LW12a, LCY12, MCSC10, NDD16, NCC11, OPP19, PK13, SSZ11, YYZ14, ZZW12]. censorship [Aly19, BWW19, DKMdUÁ16]. center [GRM14]. Centered [KCF12, EQ15]. Central [CHF11, SY18, BO11, CL13b, Dem12, Di 16, EG16, RL17, MZ13]. certain [MP12a]. Cesàro [YW10]. chain [AT10a, AHO13, CS11b, YL10b]. chaining [BC13, Kol13, Lec13, vdG13c, vdG13b]. Change [FJK11, LM17a, BH15, CHH13, CHT17, Ciu13, CHH18, Dac10, Dör10, Dör11, FGH14, Hab13, Jar10, Jar13, KSN18, KK15, MLa11, MGB10, PC14, SG11, SL19a, SR14, TN14, WG13, WK13, Wu16]. Change-point [FJK11, LM17a, BH15, CHH13, Ciu13, Dac10, Dör10, SR14]. change-points [CHH18, SG11]. changes [BD12, Buc14, Hud13, LZ13a, ST11]. characteristic [MSA14, SR14, WMH12]. Characteristics [DPM13, GS10]. Characterization [HP19, Bar12, FRR14]. Characterizations [Aok19]. Characterizing [Jon13b]. Charlier [GJ10, KV10]. chart [CM11]. charts [Alb10b, Alb11, Alb12, BS11b, JM11b, LSZ13, Ver13, YC11a]. Chebyshev [BW10a, BEJS10]. check [LYL11, LLS19]. Checking [NGXZ16, Beu10, DY19, GK17, KS10a, KS16]. checks [MLP17]. Cheng [CK12]. chi [CS12a, MYW11, YY10]. chi-square [MYW11]. chi-squared [CS12a, YY10]. chirp [RGW10]. Choice [CC12b, Alb10b, BBS10, DYV11, DDS13, GVV10, GGHS13, SD16, VGV11, Wei11]. choice-based [VGV11]. Cholesky [CT10c]. Chopra [CKG19]. circle [CL13a, ES16, SL19b]. circulant [GS11a, GS12a, GHM11, GM12]. circulants [HM11a]. Circular [AZ13, Ery11, AA11b, AA11c, DPT16, DT12, HYH10, LM11a, LM16, SYA11a, ZSA10]. claims [CM15]. class [AS15, AMR11a, Ani12, BCG10, BRB12, BRB18, Bos10, DFM10, DLDC12, FHP11, Fra14, GD10, GF13, GN16, GC17, HZ11a, HGH12, Hir19, JWY16, QOK10, SB11, SZ10, SL19b, SR13b, Wie19b]. classes [AJGG10, AHL12, ABW13, CZ12, DPC11, DPM13, Seb19]. classical [AT10b, DM11]. Classification [BOvSS10, LÖ11b, MC11, RH13, SGS12, AM11, BP10d, BM12b, CXCK13, DPC11, Hoc19, McN10, Rüs19, TH14, TP10, Zha13]. classifiers [FBHZ14]. Clatworthy [RR10]. clear [ZC12]. climate [CHT17, Lew18]. clinical [BBB11, CS13, LJK12, TNWW12, WCZ17, YBB10, ZL11b]. clinically [SM17]. cloning [HG12]. closed [KP11]. closed-form [KP11]. closeness [BDKM10, RA12]. closest [ZN12]. CLT [Fro14, Yan15]. cluster [Alb10a, Alo11, WAL17, ZLZ^{+}16]. clustered [CYCZ12, CPSY10, dUJC12, PA12, Sah11, YS11a, YSH11]. clustered/longitudinal [CPSY10]. Clustering [BW13, BRvdGZ13a, MS12b, RS13b, SMPF10, SS13a, AM11, Bha14, BM12b, DGCS16, JB14, PR16, Qiu10, ZPS15, ZGW^{+}17]. code [PMX12]. Coding [Grö18]. coefficient [ALT10, AMRP18, DZZ18, DELS16, EEL10, FLW13, GMSK10, Gri10, HP10c, LCL11, LZ12a, LM13, LMZ18, LL15, LLS19, LZH11, NCC11, RVF10, SBAC19, WZW13, WL16, Wei12, ZS15, ZLZL16, ZL18]. coefficients [CCR11, GDHS12, HH14a, KV11, Ni12, NS13, OP12]. Coherence [MZ10]. coherent [BP14c, NR12, Sad11, Zha10]. coin [Azr15, GZH11]. Collapsibility [VV10]. collector [JM11a]. colour [BSH13]. colourable [Seb19]. Coloured [RSS19]. Column [SDGK15, BL15, BM18, CKM11, QOK10, Qu11, SE19, YLL17]. Column-orthogonal [SDGK15, YLL17]. columns [Kri10, NIHY11]. combinations [Abr12, KS11a]. Combinatorial [BEJS10, CCS14, Fro14, Grö18, KM10]. Combinatorially [BW10a]. Combinatorics [Ano10a, Ano10-44, Eri10]. combined [Cha13, CL13a, MBB11, WYH10, ZCP13]. Combining [Lew18, ZWY11]. Comment [Koe13, Lec13, Xia13, YDZA13, Huc14]. Commentary [RRN14]. Comments [BM12a, Chr10, Lia11c, Yua13, YH12]. Common [GF13, VN10, COS12, FVB11, LZL12, Ruk11, Tsu10]. compact [PZ14]. compactly [DM13, SSDM14]. comparability [AMNS11]. comparative [Qiu10, RP11b]. Compare [TDW10, CSR15, JBB10, LNdDLdCH11]. Comparing [DST10, LD12, MH10, Jac19, KME10, YL10a]. Comparison [BVG10, NJ16, NdCJ10, WMM^{+}10, Cha13, Chr10, GSG12, MCSC10, OL19, Ste11b, WY10, WTL11, ZN13, vB19]. Comparisons [NR12, Bas11, HFI^{+}18, KFK11, MPRG10, WH12]. compatibility [Ruk17]. competing [HLS12, Pas10, sScY10, sS11, ZN13]. Competition [EGS10a]. complete [ABFG12, FR13, Wie19b]. completely [Lin11]. complex [BSV10, BSKS18, CO10, LD14, LL11b, PL10, SLC15]. complexity [DYV11, KK12]. compliance [MR11a]. component [IYA16, LZL16, NGXZ16, SS16]. components [BMPR11, Bur11a, CCGL10, DDL10, DKN12, Ery11, Gup12, JM10a, LC16, LZL12, MZP12, NR12, Sad11, SAE11, Zha10, ZB11, vdGM15]. composing [BW10a]. Composite [BCZ19, ZLZL16, Cab10, CGLW18, CHF11, Dic13, SGL13, XR11, Xu17]. composition [CHP11]. compositional [BJQ15]. Compound [Ste11a, YBB10, MDT^{+}12]. Compromise [Fou11, KT17]. computation [Bar12, Che11a, Fre11c, Gam13]. Computational [AK10, Kat19, HG11]. compute [JBB10]. computer [AZ10, CO10, Geo11, LMW10, LMW12, Mor14, PS15]. Computing [FRM11, Sag11, Har14]. concave [DKW17, PZ14]. concentration [HHMO15]. concentrations [Ver17]. concept [NQ10]. conceptual [CND10]. concerning [Nku10]. conclusions [Jon13b]. Concomitants [JB12, WN10]. condition [BK12, CDd11, Ton13, YO11]. Conditional [CPS15, KD12, KS10a, LMZ18, LdUÁ11, MZ10, TL18a, Cav18, CHH19, DE16, DF12b, GGS12, HK19, HCT16, KLS11, Lee18, LVK10, NDD16, OVG13, PD11, Pla13, SKFMR14, Sør19, Tan10, Unk17, VV10, XW13b, ZK17, Zha10, ZCLZ17, ZLL16b]. conditionally [KG10c, MM12]. conditioned [AP14b, JR10]. conditioning [Llo10]. conditions [BK16, CK12, CHF11, GW11, Li11, RP11a]. cone [Sag11]. cones [DB19]. Conference [Ano10-27, Ano10a, Ano10-44, Ano14v]. Confidence [BFCGR12, Bur11a, CHH18, DS18, HX16, HS18, Kon12, KL10b, LQL12, Ruk12, Sch18, Woo19, ADPW10, BI12, Beh10, CWWT12, CLO17, CWH10, Coa14, CL11b, Gua12, HK10, HvdM12, KF12, KHN16, Lew18, Lia12, LAk10, LYL13, LLP14, LZL16, MH10, SY12, SLgX18, SS13d, SZ13, SvdVvZ15, TXLV11, TCC12, VM18, VB10b, WQ10, Wan10b, WW11b, YS11b, YS11c]. confidentiality [HG12]. confirmatory [LW19]. conflict [BDG14]. confounding [CZ10, HZ11b, Jac11, SB19, ZC10, ZLZK13, ZLL16a, ZZ14b]. conjecture [FGN10]. conjoint [DYV11, GVV10, KGV10, VGV11]. conjugate [HG13]. conjunction [Alo11]. connected [FGN10]. connection [MV11a]. connections [Sar14, SP12a]. Consecutive [GKM10, Ery11, GJM18, SAE11]. Considerations [CHP11, QL15]. consistencies [LL11a]. Consistency [DM12, LM11b, Nag15, RWB15, SC11, XGK16, ZCLZ17, ACP16, CL15a, CR15, FLTV10, FN17, KLS11, KK11, Le13, LLLP19, Mei11, WT10, WS14, WM18b, XW13a]. Consistent [GLW11, KK16, LLS19, FJKW13, LLZ11a, Miy18, NB12a, sS12c, SL19a, SlG12]. constant [AX19, FHK11, HH15, KGV10, LL11c, LT10, TLBJ13]. constant-stress [LT10]. Constrained [WX15, BHM16, MKW18, MHK10, SP12b]. constraint [LLL18]. constraints [Abr12, CL15a, KS19b, LMKS17, OK11, SC13, ZGST12]. constructed [CCS14]. Constructing [DDS13, GM12, GS11a, GS12a]. Construction [AJGG10, JA17, LC15, LQ17, LL11d, SLL10, SYL13, Tsa17, ZLZK13, CKMS12, CLFZ13, CZ10, Geo14, Jac11, Jac19, LYXC16, Mar18, PZ14, Seb19, ySDM13, TQ12, TX13, YHS19, ZC10, fZyZD16, ZLL16a]. constructions [PMK19, TY10]. Cont [Ano17-35, Ano10t, Ano10u, Ano10v, Ano10w, Ano10x, Ano10y, Ano10z, Ano11q, Ano11r, Ano11s, Ano11t, Ano11u, Ano11v, Ano11w, Ano11x, Ano11y, Ano11z, Ano12m, Ano12n, Ano12o, Ano12p, Ano12q, Ano12r, Ano12s, Ano12t, Ano12u, Ano12v, Ano12w, Ano12x, Ano13m, Ano13n, Ano13o, Ano13p, Ano13q, Ano13r, Ano13s, Ano13t, Ano13u, Ano13v, Ano13w, Ano13x, Ano14k, Ano14l, Ano14m, Ano14n, Ano14o, Ano14p, Ano14q, Ano14r, Ano14s, Ano14t, Ano14u, Ano15v, Ano15w, Ano15x, Ano15y, Ano15z, Ano15-27, Ano15-28, Ano15-29, Ano15-30, Ano15-31, Ano15-32, Ano16y, Ano16z, Ano16-27, Ano16-28, Ano16-29, Ano16-30, Ano16-31, Ano16-32, Ano16-33, Ano16-34, Ano16-35, Ano16-36, Ano17-27, Ano17-28, Ano17-29, Ano17-30, Ano17-31, Ano17-32, Ano17-33, Ano17-34, Ano17-36]. cont [Ano18k, Ano18l]. contamination [SHYH16]. Contents [Ano10b, Ano10c, Ano10d, Ano10e, Ano10f, Ano10g, Ano10h, Ano11a, Ano11b, Ano11c, Ano11d, Ano11e, Ano11f, Ano11g, Ano11h, Ano11i, Ano11j, Ano11k, Ano11l, Ano12a, Ano12b, Ano12c, Ano12d, Ano12e, Ano12f, Ano12g, Ano12h, Ano12i, Ano12j, Ano12k, Ano12l, Ano13a, Ano13b, Ano13c, Ano13d, Ano13e, Ano13f, Ano13g, Ano13h, Ano13i, Ano13j, Ano13k, Ano13l, Ano14a, Ano14b, Ano14c, Ano14d, Ano14e, Ano14f, Ano14g, Ano14h, Ano14i, Ano14j, Ano15a, Ano15b, Ano15c, Ano15d, Ano15e, Ano15f, Ano15g, Ano15h, Ano15i, Ano15j, Ano15k, Ano16a, Ano16b, Ano16c, Ano16d, Ano16e, Ano16f, Ano16g, Ano16h, Ano16i, Ano16j, Ano16k, Ano16l, Ano17a, Ano17b, Ano17c, Ano17d, Ano17e, Ano17f, Ano17g, Ano17h, Ano17i, Ano17j, Ano17k, Ano18a, Ano18b, Ano18c]. context [IYA16]. contiguous [Wu19]. contingency [BSV10, KD12, KBPW12, NBSB13, VV10, Zho19]. contingent [RF13]. continual [CL15a]. continuation [GRM14]. continued [Ano10n, Ano10o, Ano10p, Ano10q, Ano10r, Ano11o, Ano11p]. Continuous [dUJC12, ACC18, AG19, BBB11, Bha11, Bou11, CH11b, CL13b, DDL14, HFW10, HCZ19, Jon15, KLW10, LA14, MRW11, Mon17, Nik13, PMW^{+}16, Reg11, SK12, SS13d, Wie19b]. continuous-time [CH11b, DDL14, MRW11]. contoured [AT10b, JD16, WM18a]. contraction [Lia14a]. contrast [Aok19, HHJ13]. Contrasting [Mac13]. contrasts [PL10]. Control [Alb12, JM11b, Alb11, AM12a, BL15, CM11, CC15, DZ18, GD15, GR15, HJ16, KME10, LSZ13, Ver13, WM12, YL10a, YC11a, YH19, dVGS11]. controlling [BS14, SGF12, SCHG19, ZZ14a]. controls [CWcIC11, DST10]. conventional [MS10]. converge [Cla19]. Convergence [BP10b, BP13a, BP19, Dör11, EEL10, Lia15, MKW18, VB10a, Xin11, Aya12, JWY16, LL11c, RW18, RP11a, SP12a, Wu19, Zha13]. convex [DB19, Kat19, WM18a, Zho10]. Convolution [CGC12, CMLBSM13]. convolutions [Ban12, KA11, KX11a]. coordinate [DF12a]. copula [BC12, JSV12, Nik13, OVG13, Seg15, YL10b]. copulas [GS12b, GP15a, JDB13, Sum17, TS12, TS13, Ver13]. coregionalization [KKL15]. Cornish [DJ19]. correct [LM11b, Moj12, SC10]. Corrected [CWH10, Coa14, BL10a, SRBS11, WG12]. correcting [WLS13]. correction [CM16, DMP10, DS18, Kim16, NYA17]. corrections [BCN13, IK15, SG12b]. Correlated [BW13, BRvdGZ13a, AVML11, ABK^{+}11, BRB12, BRB18, CS12a, DDL14, EE17, Gra12b, KME10, Sai10, TRWP10, Tsa19, ZLZ^{+}16, RS13b, SS13a]. correlation [CCR11, Dhi18, FKMR19, GMSK10, LC15, LM16, LMZ18, Lia14b, LH11, MR10, NCC11, OP12, SSDM14, WG13, Xu17]. correlations [SDZ19]. correspondence [Beh10, BS11a, Beh12]. corresponding [KMW15]. Corrigendum [Ani10, BRB18, GS12a, KZ15]. cosine [MLa11]. count [AKK12, Lin10, PA12, SP12b, TRWP10]. counterparts [JBB10]. Counting [MP12a, MSTT11, Böh10]. counts [DHH10, DDD18, Sør19, Ste11a, XZ19]. coupled [HLSL17]. Coupon [JM11a]. course [MS12b]. covariables [DW17]. Covariance [CG14, BMW16, BMPT19, BF13, CT10c, CNL17, DMP11, DPM13, DP13, DM13, Fis12, FO13, HCMP12, HBP12, IYA16, IYA19, Jar13, JJY12, LR11, LCQ^{+}13, LB15, LL15, LLZS17, MJS12, MYW11, NHSP13, QZZ11, RV16, Sch10b, She14, VB10a, ZW12b]. covariances [BMW15, FCC10, Ros10]. Covariate [CZHC14, SG12a, Fre11b, HLH13, LBH17, dUJC12, MC11, Moj12, OPP19, SXX11, WG12, XSJ14]. covariate-adaptive [LBH17]. Covariate-adjusted [CZHC14, HLH13]. covariate-dependent [OPP19]. covariates [Bha11, CYCZ12, CYZL12, DMP10, DKMdUÁ16, FRF12, FC12, GLT16, GX12, HDC14, HLS10, Hua11, MQ10, NDD16, SRBS11, SHYH16, SSZ11, TN10, XI12]. Coverage [GP10, KF12, Wan10b]. covering [CSWY11, GKM10, WYH10]. Cox [CMT17, CC11, FN17, GLT16, HH14a, LHH16, LM17b, LLL18, RH11, WZ18, WWZ19, YLC^{+}16]. Cramér [Gas11b]. Cramér-von [Gas11b]. Creating [LST12]. Credible [SvdV15, Bos10, Thu14]. criteria [CSR15, GZ14a, LST12, Luk14, PZ14, QWC12]. criterion [Bha14, CH11a, DHM18, HY19, HLL12, HYS12, KHS12, WZ11b, XT15, ZZ14b, CS16a]. crop [CP19]. Cross [Jan15, AM19, AJGG10, BMW15, BDvdA15, CR10, DHM18, DT14, IR19, LLJK13]. cross-over [AJGG10, DT14]. cross-sectionally [BDvdA15]. cross-validation [AM19, CR10, DHM18, IR19]. crossings [AG19]. crossmatch [ACP16]. Crossover [YL10a, BS10b, PBND11]. cryptology [Sar14]. cube [BC11b]. CUMAX [Alb12]. Cumulants [Leh11, BM12a, Oga13, WN11b]. cumulative [GMP15, NdAA10, NB10, Pla13, sScY10, ZWM12]. Cun [SS13a]. Cun-Hui [SS13a]. curable [CF12b]. cure [CLB12]. current [OZC16, Pla13, Unk17]. curse [RBRTM11]. Curve [CD15, CWcIC11, Chr14, DN12, HX16, HC11b, LZ12b, Son11a, SC16, WSSG19]. curves [FRF12, TDW10, TvS15, ZPS15]. CUSUM [BS11b, CM11, Wu16]. cut [AHO13]. cutoff [MFW11]. CV [GC17]. cyclic [BG10, CLFZ13, LÖ11b]. cylindrical [Maj14].

D [BLP11, NL16, vdG13a]. Dantzig [AH12, FN17]. Darbellay [Le13]. Darling [CHH13]. Data [CM10, Fre13, GPRY11, HG12, MA10, APQ10, AS15, ABK^{+}11, Alo11, AMRP18, BFZ10, BM11, BJQ15, BS10b, BLT16, BFCGR12, BR10, BM12b, BCG13, BH15, CW19, CPH19, CHH13, CLL12, Cha19b, Che10a, CC11, CYC11, CYCZ12, CW17, CGLW18, CPSY10, Che12, CR16b, Cop11, CXCK13, CN12, CS11b, Cue14, CDNS14, DF12a, DMTZ15, DP11, DGCS16, DW17, Día17, DELS16, DK17, FZS16, FJK11, GMR10, GD10, GK17, GSZ11, GRM14, Gri10, Gui19, Han14, HK19, HW17, Hon10, HLS10, HFQ11, HCMP12, HYS12, HLS12, JB14, JP12, Kak18, KBPW12, KK11, KSL11, Kon12, Kon13, Kre18, KSKE10, KKP10, LÁPMF19, LL11a, LRB11, LR11, Li11, LZ12a, LZ13b, LM16, Lia12, LdUÁ11, Lin10, LW11b, Lin11, LLV15, LW10, LLZ11b, Liu11a, LZ12b, LL15, LLZS17, dUJC12, LQ16, Lwi11, Mad10, Maj14]. data [MCSC10, MS12b, Miy11, MRY11, NS10a, Nis11, NHSP13, OG10, OZC16, Oua13, PD11, PA12, Pla13, Qin10, RM11, RH11, RAB14c, Sah11, Say12, SXX11, sS10a, sS10b, sS11, She11, sS12b, sS12c, sS13b, SDNH13, Shu12, SN10, SM11, STS10, SP12b, SLC15, SZ10, SSZ11, SdL10, TH14, TRWP10, TB11, TWLQ15, Tor15, TW11, Unk17, VBLN17, VGH13, VB10b, Wah10, WK10, WZ11a, WTL11, WH11, WS15, WDRB18, WZ18, WQ11, WW12a, YSH11, ZW12a, ZS15, ZLZ^{+}16, ZZW12, ZZW19, ZGST12, ZGK13, ZWM12]. Data-dependent [CM10, SN10]. Data-driven [Fre13, Lin11]. Data-transformation [MA10]. datastreams [DZ18]. De-aliasing [Cha19a]. Deal [MYJ11]. death [PP10]. December [Ano11-47, Ano12-45, Ano13-44, Ano14-37, Ano15-41, Ano16-46, Ano17-38, Ano18n, Ano19k]. Deciding [HP10a]. Decision [Thu14, BS10a, BP10d, HP10a, Jon13b, YC11b]. Decision-theoretic [Thu14]. Decomposable [BTV12]. Decomposing [ABFG12]. Decomposition [WV12, HS14, KDJ13]. decompositions [DKY15, Lom11]. deconvolution [BKPS14, BPR19, Mab16, Mei11, Var15, Wis13]. Decoupling [SR14]. decreasing [Pav12]. Default [VCR10]. defense [AMNS11]. defining [WH11]. definiteness [DMP11]. definitions [NS10b]. deformation [Fra14]. deformed [Sei14]. degeneracy [GRD16]. degenerate [BK16]. degradation [TTB11]. degree [RP11b]. degrees [CSC14, GF11, Zha12]. delay [BP15, Fre14]. delayed [JRM10]. Deletion [CT11]. Deletions [Pan10]. Delta [Lom11]. demand [KZW10]. densities [Kom11, LVK10, Pav12, Wen15]. Density [KKP10, AC11, Ban12, BJQ15, BK17, BFKK17, BR10, CM10, CSW11, CGC12, CDNS14, DW17, DPT11, DHM18, Dut11, Efr13, GZ10, HM10, Kak18, LMKS17, Leb12, LQL12, Lu13, Mül12, QLL11, RV10, RBRTM11, SP12a, SL14b, Tan10, THG12, Ton13, VGH13, WMH12, XL12, ZCLZ17]. density-based [VGH13]. density-quantile [HM10]. Dependence [DKL13, AS11b, BGT16, BKPS14, FCC10, FRM11, GR15, KCF12, Kra12, MPN^{+}17, PR16, PC14, SLC15, TSZ13, Unk17]. dependency [Le13]. Dependent [BK16, AP14a, BP10b, BP13a, BP19, BDvdA15, BMPR11, BLT16, BK17, Bha14, Bin14, BFKK17, Bro10, CM10, CLL12, CL15b, CH10, Ery11, FYX14, Gup12, HW17, Hon10, HC11b, JR12, Kro19, LdUÁ11, MZ13, OPP19, SN10, SS19, Sum17, SSZ11, Wis13, YC11a, Yan15, YLC^{+}16]. depth [AR11, DS10, DM12, KMW16, Nag15, PS15]. depth-based [DS10]. depth-design [PS15]. derivation [BP14c, BE14]. derivatives [CWWT12, LLX12]. DerSimonian [JBB10]. Desarguesian [BBPW14]. description [Cla19]. Design [Ano14v, BR14, GT10, HDC14, LW19, LH14, LM14, OL19, ZLM12, AAKR12, AFHZ14, Azr15, BBB11, BCMAW14, BTT15, BM15, CSR15, CCGL10, CS13, DYV11, DD13, EZ12, GS19, GZ14a, GN16, Göb12, Gra12a, GDHS12, GGHS13, GSKP10, GCDK12, GZH11, HJ14, JGJR16, Kou11, KL11, LYF14, LM19a, LL11d, LYL13, MZP12, MS10, MGD14, NS13, NZF11, PP13, PS15, PL11, RRN14, SvdV15, SDV14, TNWW12, Tsa11, VGV11, Wag11, WZ11b, WT14b, YLC^{+}16, YCB16, YBB10, YQC11, ZZOY13]. Design-based [GT10, CCGL10]. designed [AT10a, Coa14]. Designs [AA11b, AA11c, GSG12, ALN10, AJGG10, AG10, AHL12, ALLL13, AM12a, AZ10, AHO13, Aok19, ABW13, AG12b, AB14, BSH13, BL15, Bar17, BSKS18, BT19a, BET17, BS16, BM18, BRBM19, BD14, BBS10, CY10, CYW10, CL13a, CT14, Cha19a, CQ11, CKM11, CKMS12, CLFZ13, CGLW18, CZ10, CKP17, CDST18, CC19, CY19, CZHC14, CKG19, CHF11, DPA10, DL11, DH11a, DM10, DMS11, DT12, DT14, DD13, EMS11, EQ15, FK19, FR13, FHK11, FRR14, FRR16, GIS19, Gaf19, GS11a, Geo11, GS12a, Geo14, GF13, GC17, GW11, GG15, GVV10, Grö18, GHM11, GM12, HKN18, HYH10, HS14, HS11a, Har14, HLSL17, HZY14, HZ11b, HLL12, HHJ13, HLH13, HH14b, HC16, HM13, HYF11, Imh14, ITA^{+}09, IT10, ITA^{+}12, Jac11, Jac19, JA17, KW14, KT17, Kat12, KDJ13, KGV10, KBK15]. designs [KMMP11, KME10, KM11, Lac14, LO11a, LB11, LMT14, LC15, LZZ15, LJA15, LST12, LYL11, LYC14, LYC15, LYXC16, LMW10, LMW12, LM19b, LSW13, LNdDLdCH11, MS12a, MZWF17, MP12b, MDT^{+}12, MFW11, Mor14, Müh10, NIHY11, Nil11, OQ10, OQL11, OQC13, OZQ19, PP10, PL10, PBND11, PMX12, PBRT16, Pro10, PZ14, PMK19, QCG15, QOK10, Qu11, QL15, RF13, RS19, RR10, RSS19, Sag11, Seb19, Sek12, SE19, SYA11a, SYA11b, SM16, SHW16, ST19, SD10, SR10, SX11, SS19, SDGK15, ySD10, ySDM13, SLL10, SYL13, SD16, TB12, TQ12, TX13, TY10, TLM12, TL11, TL13, TL18b, TS10, TKB11, WMB10, WY14, WA16, WLL^{+}17, WZZ19, WL14, Wie19a, Wie19b, WZ11b, XM10, YHS19, YL10a, YL14, YZZ17, YL13, YLL14, YAW18, YLL17, ZSA10, ZC10, ZC11, ZDLW19, ZC12, ZLZK13, ZLL16a, ZZ14b, vB19, vdVW14]. desirable [Hir19]. desired [HK10]. detecting [GP12, MX15, MB11, Ros14, SG11]. Detection [CGL11, BT19a, CHH13, Ciu13, FGH14, Jar10, KSN18, PC14, SR14, YCY15]. detectors [AHKS12]. determinant [Qui10]. Determinantal [LM19b]. Determinants [Kra10, Wag11]. determination [Jac19, PBRT16, RVF10, TQ12]. determined [Dem12]. Determining [Tso11, MP12b]. deterministic [ZW11]. Developing [SG11]. deviations [GR14, JZ11, Mac10, ZL12]. diagnosing [YC11a]. Diagnostic [CFW14, CL15b, CMCSC12, LNdDLdCH11, TDW10, TXLV11]. diagnostics [COP10, CT11, LGLO12, ZSH12]. diagonal [JDB13]. Difference [DMTZ15, BLW16, BG10, Fan10, LÖ11b, PS14, Sch10b, SYL13, SM17, TQ12, TXLV11, WW11a, YYZ14, YS11b, YS11c, ZSM10]. Difference-based [DMTZ15]. differences [HX16]. different [CSC14, LJA15, vdGM15]. differentiable [PZ14]. Differential [CB19, BP15, DGCS16, KZ14, KZ15, SP12a]. differentiate [CK13]. differentiating [TP10]. diffusion [ALS12, AG19, BP10c, Bro10, DI10, Gas11b, KU14, Kon13, LS10a, SLW13]. diffusions [FS14, JZ11, NU19]. dilemma [AMR11b]. dilution [MP12b].

Dimension [DDW16, DW17, HFQ11, Mei13, Dhi18, HY10b, IYA16, KWS19, Lia15, Moj12, NYA17, NHSP13, PA13, XWY18, YY10, Yua13, Zha13, vdG13a]. dimensional [ABK^{+}11, AP14a, AC11, Bar12, BP10d, CPH19, CYZL12, CS16b, GLW11, GP12, GR18, GLT16, HYS12, IYA19, JJY12, LLL^{+}12, LCQ^{+}13, LB15, LJA15, LMZ18, LLZ11a, LLZS17, LQ16, LA14, MY13, MPT13a, MPT13b, NPK10, Nan12, Ni12, Nkw10, Obo13, PD11, PS15, RS19, SL14a, TS13, WTZ12, WX19, YW16, YYX19, ZGK13, ZK15, ZGZ17]. dimensionality [Fis12]. Dimensionally [MV11b]. dimensions [BSG16, BMPT19, FK11a, MC11]. Dimitris [GP16]. Direct [WMH12, BMW15, LLM11, PBND11]. directional [GR15, Jon13b, KG10c, LM11a]. Dirichlet [GS11b, Gri11, MS12c]. discontinuity [BRBM19]. discovery [DZ18, LS10b]. Discrepancy [NZF11, CND10, CQ11, EQ15, SlG12, Zho10]. Discrete [Cha10a, KV10, AP14a, ACC18, CQ11, DP11, DDS13, EGS10a, GGHS13, Hab15, HFW10, JM13, KU14, KK12, KL10a, Nik13, PL11, SD16, WH11, Wie19b]. discrete-time [JM13]. discretely [CH11b, DI10]. Discriminant [CHT17, AM11, BM12b, HY10b, HYS12, Shu12, TH14]. discrimination [ALN10, EMS11, LR11, Luk14, Shu11]. Discussion [Ber14, BP14a, BC13, BW13, De 13, GP16, HL13, HJ14, Jon13a, Kol13, Kre16, KL13, MY13, Mam13, Mei13, Obo13, SS13a, SDV14, vdG13a, vdGM13, Goo19, HW13, NL16, Wal13b]. disease [NdCJ10]. disjoint [BG10]. disparities [MBB11]. disparity [MBB11]. dispersion [ALT10, CT10a, EA10, JR12, PA12, SCR10, SRBS11]. dispersions [Sah11]. dissimilarities [TvS15]. dissimilarity [CMO13, Wah10]. distance [Dal11, Hab10, HS11b, KL12, Kou11, KS16, MKT15, Mor14, NJ16, TvS15, XW13a]. distances [SR13b]. distortion [ZYZL14]. distributed [BP15, JR12, NR12]. Distribution [DE16, DG13a, LLZ11b, Son11b, WN10, Alo11, ÁG12a, BP10a, BAHS11, BGK^{+}11, dZBK10a, dZBK10b, BSG16, BMNY13, Cha10b, Che10a, CS12b, DR13, DH11b, DG13b, DY19, Dör11, DKW17, FP12a, FKV11, Fre14, FJC12, GMSK12, GT12b, GRS^{+}12, GP15b, HN12, HGH12, HKR11, Ho10, HLCW12, HFQ11, IB11, IKU11, KG10a, KG10b, KR12, Leb12, LD14, Lew18, LCQ^{+}13, Lia11a, LW11b, LZL16, LSP12b, MRY11, MSW12, NB12a, NB12b, NB10, NCC11, OJMRCSRA11, Pla13, PK13, RM11, Reg11, RTM10, Rüs19, She14, sS10a, sS11, SK12, SC13, Vag10, VBLN17, WQ10, Wan11, Won12, YO11, ZX12, Zou14]. Distribution-free [DE16, Son11b, BSG16, HGH12]. distributional [KNM10]. Distributions [DAP10, IAH11, SM13, AVML11, AJM19, Ani12, ALP10, AAV13, BT16, BCMRMP11, Bas11, BK11b, Bou11, Bur11a, COS12, Cha10a, Cha12, CPM14, CSC14, CS12a, CL11b, CH11c, CHH18, Dal11, DS18, DDNN10, DKL13, DLPW14, DF12b, EGGG12, FHP11, FGV12, FVB11, Fre11a, GGG11, Hab13, Hab15, HvdM12, HS18, HG13, IF10, JB12, JJY12, JD16, JM11a, Jon15, KLW10, Kem10, KL10a, KP11, KV10, LGLO12, LM11a, LVY13, LP11a, LN15, LVK10, LYH12, Lia14a, LCY12, LH11, LT18, LS12, LLLP19, MS12b, MO10a, MO10b, MV11a, NS10b, Oga17, OS10, Pat12, RP11a, SBCK11, SGL13, VCR10, VM18, VSK^{+}11, WW11b, Wu17, Xin11, ZZ10]. divergence [GBP15, SN10]. Divergences [BK12, DI10, Che12, JMW13]. diverging [FLL17a, LC13, WW15a]. diversity [ZZ10]. divisible [CY10, CYW10, RR10, RSS19, Vag10, ZC11]. dodecahedra [ABFG12]. does [JBB10]. domain [FRM11, KLKO15]. dominance [Kub14]. dose [AMR11b, HC13, LLP14, MDT^{+}12, Pro10, WY14, YCB16]. dose-finding [HC13, Pro10, WY14]. dose-response [LLP14, YCB16]. doses [XI12]. double [GLZ14, HTZ14, Kub14, VS12]. doubly [sS11, sS13b]. down [DB12]. downtime [TLBJ13]. Downton [HN12]. downturn [HYF11]. drawn [BC11b]. drift [HS11a, ST11, XZZ18]. driven [BC19, CL13b, Fre13, GPRY11, Lin11]. drug [TP10]. Dual [Che12, Kak18, RBS10]. duality [BK12]. due [TLBJ13]. duration [JM13, ZLL16b]. durations [DHH10]. Dutter [Hu13]. Dyck [DMR10b, DMR10a, MSTT11]. Dynamic [CSN19, HJ16, ND11, CO10, DN13, DLDC12, GMR10, LL11c]. dynamical [Kut11].

early [CMCSC11]. EBLUP [Hir19]. ecology [BR14]. economics [DDD18]. ECV [GC17]. EDF [Fre12]. Editorial [Ano10n, Ano10o, Ano10p, Ano10q, Ano11o, Ano10r, Ano10i, Ano10j, Ano10k, Ano10l, Ano10m, Ano11p, Ano11m, Ano11n, Ano17l, Ano17m, Ano17n, Ano17o, Ano18d, Ano18e, Ano18f, Ano18g, Ano18h, Ano19b, Ano19c, Ano19d, Ano19e, Ano19f, Ano19g]. Editors [Ano10t, Ano10u, Ano10v, Ano10w, Ano10x, Ano10y, Ano10z, Ano10-28, Ano10-29, Ano10-30, Ano10-31, Ano10-32, Ano10-33, Ano10-34, Ano11q, Ano11r, Ano11s, Ano11t, Ano11u, Ano11v, Ano11w, Ano11x, Ano11y, Ano11z, Ano11-27, Ano11-28, Ano11-29, Ano11-30, Ano11-31, Ano11-32, Ano11-33, Ano11-34, Ano11-35, Ano11-36, Ano12m, Ano12n, Ano12o, Ano12p, Ano12q, Ano12r, Ano12s, Ano12t, Ano12u, Ano12v, Ano12w, Ano12x, Ano12y, Ano12z, Ano12-27, Ano12-28, Ano12-29, Ano12-30, Ano12-31, Ano12-32, Ano12-33, Ano12-34, Ano12-35, Ano12-36, Ano13m, Ano13n, Ano13o, Ano13p, Ano13q, Ano13r, Ano13s, Ano13t, Ano13u, Ano13v, Ano13w, Ano13x, Ano13y, Ano13z, Ano13-27, Ano13-28, Ano13-29, Ano13-30, Ano13-31, Ano13-32, Ano13-33, Ano13-34, Ano13-35, Ano13-36, Ano14k, Ano14l, Ano14m, Ano14n]. Editors [Ano14o, Ano14p, Ano14q, Ano14r, Ano14s, Ano14t, Ano14u, Ano14w, Ano14x, Ano14y, Ano14z, Ano14-27, Ano14-28, Ano14-29, Ano14-30, Ano14-31, Ano14-32, Ano14-33, Ano15v, Ano15w, Ano15x, Ano15y, Ano15z, Ano15-27, Ano15-28, Ano15-29, Ano15-30, Ano15-31, Ano15l, Ano15m, Ano15n, Ano15o, Ano15p, Ano15q, Ano15r, Ano15s, Ano15t, Ano15u, Ano15-32, Ano15-33, Ano16y, Ano16z, Ano16-27, Ano16-28, Ano16-29, Ano16-30, Ano16-31, Ano16-32, Ano16-33, Ano16-34, Ano16-35, Ano16-36, Ano16m, Ano16n, Ano16o, Ano16p, Ano16q, Ano16r, Ano16s, Ano16t, Ano16u, Ano16v, Ano16w, Ano16x, Ano17-27, Ano17-28, Ano17-29, Ano17-30, Ano17-31, Ano17-32, Ano17-33, Ano17-34, Ano17-35, Ano17-36, Ano17p, Ano17q, Ano17r, Ano17s, Ano17t, Ano17u, Ano17v, Ano17w, Ano17x, Ano17y, Ano17z, Ano18k, Ano18l, Ano18i, Ano18j]. Effect [DYV11, HP10b, LM16, BVG10, FZS16, HHMO15, Hay16, LL13, LT14b, Ros14, Ruk12, dLJ10, dLJ12]. Effective [LZZ15, TX13, XI12]. effectiveness [AM19]. Effects [TH14, AA11b, AA11c, AMRP18, BR14, BT19a, BGS10, Bur11a, CLM12, CC12b, CSC14, DGCS16, DKY15, DT14, EA10, ETMJ11, HC16, JBB10, KK11, KPZ13, LT14a, LZ13b, LW10, LZL16, MP11, MR10, OG10, PBND11, RSH12, Ruk11, Ruk19a, SB19, SG12a, Tsa19, Wae10, WT10, YL10a, YS11a, YSW10, ZS15, ZC12, vdMC10]. efficacy [RF13]. efficiencies [BM18]. Efficiency [Che10b, CS11a, MS18, VGV11, APQ10, AM12a, Azr15, CC19, DYV11, HC13, Ing10, Lac14, PMK19, TWLQ15, TY10, Ver17, WLS13, XL11, ZLM12, ZG12]. efficiency-balanced [PMK19]. Efficient [Bra10, Cho17, Día17, DT12, GH10, JP12, LN15, LCY12, LLL18, MY10, PBND11, Ban12, BT19b, BLP11, CS13, CKP17, DB11, JBB10, KS19b, KBNB11, NPK10, PS15, PBRT16, RS16, Soi13, SD16, XW13b]. Efron [Azr15, LZL^{+}10]. eigenmaps [CC12a]. eigenspaces [BLT16]. eigenvalue [IYA19]. eigenvalues [LB15, She14]. eighth [PMX12]. Eisenthal [DJ19]. Elastic [CYZL12]. Elastic-Net [CYZL12]. elasticities [KZW10]. electromagnetic [Zha15]. element [Kat11]. Elemental [AFHZ14]. Elfving [Bar12, Qi11]. Elicitation [MO10b]. Elliptical [Beh10, DG13a, DIAT11, JB12, Kra12, LVY13]. elliptically [AT10b, JD16, WM18a]. Embedding [ACC18]. Emmanuel [Ano10s]. Empirical [Alb11, AMRP18, CY11, CS16b, HZ18, LQ11, LP11b, Lia12, LW15, QLL11, SS12a, VCA10, WQ10, WZ11a, WW11b, XL12, XXC11, ZY11, BFZ10, BP13b, CM10, Chr14, CG14, DRT10, FSS18, FLL17a, FP12a, Gho13, GMP15, GF18, KLW10, Kon12, KS93, LYLZ12, LZ12b, LX18, NL14, PG10, PHW15, QL10, Rab14a, RW18, SDZ19, Sch18, VBLN17, VSK^{+}11, VGH13, VDW13, VZH17, WLCK11, WL11, WWZ19, YZ12, YYZ14, Yan15, ZX12]. empirical-type [CM10]. empty [BF12]. emulation [CO10, JGK11]. Encoding [Wah10]. encryption [HG12]. endogenous [Kim16]. endpoints [JM13, WCZ17]. Energy [SR13b]. entropic [KMMP11]. entropies [GMP15]. entropy [AA11a, DF17, DGS19, GLR11, Gui19, KNM10, NdAA10, NMS12, Reg11, SP12a]. enumeration [Hum10]. environment [MX15]. environmental [TG16]. enzyme [AB14]. epidemic [Buc14]. epidemics [PP13, PBRT16]. epistemic [Sch18]. EPMC [HY10b, Shu12]. equal [CM10, GO11]. equal-tailed [CM10]. Equalities [SC16]. Equality [IYA19, MHK10, AB14, HP10b, IYA16, KS19b, OP12, SDZ19]. equals [Fis12]. equation [BP15, KZ14, KZ15, Lin10, Lin12, WF11]. Equations [CB19, DGCS16, Gro13a, Jon13a, LP11b, LJ12, Mam13, WLCK11, WL11, vdGM13]. equicorrelated [LR11]. equidistant [HS11a]. equilibrium [NS10b]. Equivalence [Kat12, CY11, Grö18, KDJ13]. equivariant [Mad10, ZN12]. Erdos [CHH13]. ergodic [FJKW13, KU14, LL11a, LLV15, NU19]. Erratum [AA11c, BK11c, IT10, KS11b, TLBJ11, YS11b, dLJ12]. erroneously [LG12]. Error [PKL12, BS14, BL10b, CPM14, CC15, CSC14, DRT10, DR13, He11, Hir19, Hua11, HZ13, HCZ19, LM16, LZL16, LLW16, Lu13, LZH11, LGG12, Mai11, MS13a, MSW12, SGF12, SXX11, SGL13, TH14, Tar10, TRL13, Var15, WZW13, WG12, YH19, ZYZL14]. error-in-variable [LZH11]. error-prone [Hua11]. errors [AVML11, AD11, BLT16, BM15, CLL11, CNdGAL11, DRT10, DM11, FLW13, FYX14, FK11b, GZ14a, GK17, GR15, HHMO15, HK19, HY19, Hon10, HG11, Kau14, KS16, KL11, KME10, LCT10, LLS19, Nan12, Pan10, PG12, Sai10, SS16, SR10, Wei12, WZ10, Wis13, WQ11, WW12b, XSJ14, YAW18]. errors-in-variable [HG11]. errors-in-variables [DM11, FLW13, Hon10, LLS19, Wei12]. Esseen [Bou11, Fro14, HWZ11, ZL11a, YHWL12]. Esseen-type [Bou11]. Estimable [Tsa19]. estimate [CWcIC11, Dut11, RS16, XR11, ZCLZ17]. estimated [CM11, CM16]. estimates [ÁG12a, BM12a, Dal11, FJKW13, FLTV10, FK11b, HWZ11, KK12, Kre18, SDNH13, SZ19, WMH12, WN11a, WN11b, XL11]. Estimating [CT10a, CL11a, GGS12, Mül12, MSW12, PD11, Ruk11, Ruk19a, SSDM14, AG12b, BTT15, BLP11, CC12b, CC15, CHF11, DF12a, GT12a, HFW10, HBP12, KK15, LT14a, Lew18, LP11b, Lin10, Lin12, LS10b, LSW13, NPK10, TRL13, VS12, WMM^{+}10, WLCK11, WL11, WTZ12]. Estimation [APQ10, AG19, BP10a, BSV10, BWW19, BMPT19, BFKK17, BF13, CT10c, CGCL10, CP19, DR13, Dut11, EGGG12, EZS11, FT10, FZ13, FLL17b, FKLL13, GS10, GRCGL12, HK19, HCT16, HLY12, JRM10, Kas17, KMW15, LGLO12, Li11, LM13, LCQ^{+}13, Liu11a, LX18, Nan12, NMS12, PG12, Reg11, RH11, RTM10, SXX11, sS11, sS13b, SSS12, WSHW10, WW15b, YSW10, ZS15, AHD12, AC11, AS10b, AAV13, BMW16, Bao18, BLBL10, BDLP14, BJQ15, BET17, BP13b, dZBK10a, dZBK10b, BK17, BW13, BWYZ11, BK11b, BR10, BC11b, BK12, BTV12, BCG13, BH15, BRvdGZ13a, Bur11b, CUW10, COP10, CR10, CCR17, CPS15, Che10b, Che11a, Che11b, Che12, CH10, Cho17, CB19, CHH19, CGC12, CH11c, CA10, CHR13, CPL18, CDNS14, DMTZ15, DM14a, DRT10, DM11, DW17, DKN12, DGS19, DPT11, Día17]. estimation [DZZ18, Dio14, DELS16, DHM18, Dör10, DMP18, DD13, DK17, Efr13, EMS11, FCC10, Fra14, GLR11, Gam13, GZ10, GD18, Gas11a, GT12a, GLR10, Gho14, GC11, GB19, GH10, GS12b, GT10, GLT16, GSS10, Hab10, Hir19, Ho10, HM10, HFQ11, HC13, HH14b, IF10, JZ13, JP12, Kak18, KMT10, KV11, KWS19, KS19b, KM12a, KWL14, KBNB11, KKP10, KG10c, LMKS17, LÁPMF19, LBL12, Lia11b, LdUÁ11, Lia11c, LM17a, LHH16, LSWY16, LLV15, LLZ11b, LLX12, LLL18, LW15, LZH11, Lwi11, MY10, MK18, Mad10, MR11b, MSBM16, MBB11, MPN^{+}17, MN16, Mei11, Men12, Men13, vMMA14, Mon17, NB12a, NB12b, NDD16, Ni12, NS13, Nik13, Nis11, Oga17, Oua13, Ozt11, PT13, PKL12, PH13, PR12, Pla13, Qin10, RV16, RV10, RRS11, RBRTM11, RS13b, Ruk17]. estimation [SMM19, SBCK11, Sch10b, SS13a, sS10a, sScY10, SLZ10, SN10, SP12a, SK10, SM11, SL14b, SLW13, SSC11, SC13, SGL13, SSL^{+}17, SY18, SBAC19, SlG12, Tan10, THG12, TWLQ15, TXLV11, Tor12, VAM10, Wae10, Wan10a, WMH12, WX15, WW15a, WZ18, WSSG19, WZ10, Won12, WQ11, WG12, WW12b, XI12, XZZ18, XZ19, XW13b, YSGN10, YKH13, ZH12, ZS10, Zho10, ZSL14, vdGM15]. estimator [BP10c, Bin14, BMNY13, CLL12, CLL11, CSW11, DJ19, DN12, DF17, Dhi18, FO13, GGS12, Han14, HS11b, HCT16, JSV12, JKL17, KLS11, Kra12, LL11a, LLJK13, Liu11b, LSP12b, Lu13, MYJ11, MJS12, MS10, NB10, NCC11, Oga13, PC13, RV16, RLLT17, RAB14c, RWB15, SSZ16, She11, SHW16, SA10, Son11a, Tar10, WYY^{+}10, WW12a, YH12, ZX19, ZGW^{+}17]. Estimators [WN13, ÁY12, AT10b, Aya12, BVG10, Ban12, BBKP19, BK11a, BP14c, Bir15, Bra10, BDG14, CK13, COS12, CH11b, CW17, CPSY10, Cla19, CR16b, Cop11, CM16, DM12, Dör11, EEL10, FYX14, Fou11, FRM11, GGG11, GP19, GBdW10, HZ11a, Han11, HY10a, HP10c, Hu13, IK15, KF10, KS12, KS11a, Kub14, KS93, KS11b, LRB11, Le13, Leb12, LM11b, LW12a, LM17b, LGG12, Mac10, MS13a, MKT15, Miy11, Miy18, Now12, Ogi19, PG10, PS14, QZZ11, RS10, Ruk14, Ruk19b, Seg15, sS12c, SCR10, SRBS11, SC16, TSO10, Ton13, Tsu10, VY14, VBLN17, WV12, YL10b, ZL12, ZN12, ZWM12]. Euclidean [KS19b]. Euro [BT16]. Euro-Mediterranean [BT16]. European [BS11a]. EV [Liu11a]. Evaluating [PPB16, KPZ13, VZH17]. evaluation [BTKN10, Ing10]. event [APQ10, AS15, PG12, ZZW12]. events [DAP10, FC10, GC19, SM13]. every [ES16]. Exact [AZ10, BCG10, Che11a, GMSK12, HHJ13, IB11, KM12b, TXLV11, VB10b, BSG16, BK10, BK11c, CR16a, Gaf19, Gam13, HLB^{+}16, Imh14, Sai10, VGH13]. exceedance [SB11]. exceedance-type [SB11]. exceedances [FP12b]. exceeds [Fis12]. excellent [Kat11]. excess [ZZW12]. exchange [AK10]. exchangeable [AZ13, CM15, Ery11, IAH11, Sad11, TRWP10, Zha10]. exemplified [LS12]. existence [FK11a, WWC10, Zhi10]. exists [RBS10]. expansion [CLL11, HOK19]. expansion-based [HOK19]. Expansions [Wen15, MS11, Shu11]. Expectation [BU10, LSWY16, ZJ10]. expectations [GR11]. expected [FRC11, JZ13, MKT15, VB10a]. experiment [CMCSC11, HHJ13, SC10]. experimental [AAKR12, AFHZ14, CSR15, CQ11, HDC14, JGJR16, Lac14, MGD14, PP13, PBRT16]. experimentation [AMR11b]. Experiments [Ano14v, AD11, AZ10, AT10a, BSH13, BR14, BL15, BT19b, BM18, BBS10, CC12b, CKP17, Coa14, DYV11, DL11, DB12, DDS13, EA10, ETMJ11, FRR14, Geo11, GC17, GGHS13, GSG12, Har14, HH14b, HJ14, LW19, LH14, LMW10, LMW12, Mor14, NZF11, OL19, PS15, RRN14, RLZ12, Sag11, SDV14, Ste11b, SD16, TL11, VGV11, Wie19b, WT14b, YCB16]. expert [BSV10]. explanatory [AKK12]. Exploring [PR16]. explosion [GP15b]. Explosive [Akn15, Lee18]. exponent [HM10]. Exponential [WTL11, BZ12, BLP11, BK11b, CGCL10, DLDC12, FJK11, GMSK12, GT12b, HdGT12, HN12, IF10, IKU11, JWY16, Lem11, LCY12, Mad10, Oga13, PC13, SBCK11, Sei14, SK12, Wan10a]. exponentiality [Ani10, Ani13, MA08]. exponentials [KPQ11]. exposure [BVG10, HG11]. expression [LCC^{+}17, MS12b]. Extended [LC13, PT13, QCG15, WSL17, DGG19]. Extension [CK12, CS16a, FKV11, Nik13, VZH17]. extensions [OQ10]. extinction [GP15b]. Extremal [Hil11, AS11b, Per10]. Extreme [LB15, BWW19, CDd11, EGGG12, FJC12, GBdW10, GS12b, Ho10, LS12, Luk14, MV11a, NDD16, YSGN10]. extreme-value [GS12b, MV11a, NDD16]. extremes [BT16, BGT16, LT13].

Factor [GLZ14, AP14a, AM11, BT19a, BDLP14, CT10c, CR15, Cho17, LT10, LYC14, ST19, SLC18, WS14, ZK15]. factorial [AHO13, Aok19, BL15, BT19b, BM18, Cha19a, CKG19, EA10, ETMJ11, FRR14, GF13, GC17, Grö18, Jac11, Kat12, KDJ13, LST12, LYL11, OQC13, OZQ19, PL10, SC10, SD10, ySDM13, TL11, WMB10, WZZ19, WZ11b, ZC12, vB19]. factorials [LQ17, QWC12]. factors [Aok19, BL15, BDvdA15, CSWY11, GSG12, HLSL17, HLL12, Kat12, VDW13, WZZ19, Wie19b, XGK16]. failure [AZ13, DGG19, HLS12, TLBJ13, TNWW12, ZY11, Zho10]. false [DZ18, LS10b]. Families [BT19a, Kem10, BI12, BG10, HG13, KX11b, Sei14, WW11a]. family [Alb10a, Ani13, BZ12, DDNN10, Kem10, Lem11, Oga13, ZZOY13]. familywise [BS14, SGF12]. fast [BLP11, NPK10]. faster [SS13d]. favored [FWS18]. Fay [Hir19, SM11]. FDR [HJ16, SCHG19, ZZ14a]. feasible [WX19]. feature [DPC11, LC13, Xu17, Zha15]. features [CMO13, LC13]. February [Ano10-41, Ano11-48, Ano12-46, Ano13-45, Ano14-41, Ano15-40, Ano16-38, Ano17-46, Ano18p, YS11b]. February-March [Ano15-40]. few [CSWY11]. fewer [HYS12]. FFT [FCC10]. Fiducial [VM18, WHI12, ZXD12, KL10a, TL18a, WH12]. field [AX19, KCF12, Per10]. fields [Alo11, BMW15, CFT10, DM13, FP12b, GRD16, KF10, Lav11, Ton13, WC10, WM18a]. filling [BHM16, LJA15, Müh10, SYL13, TB12]. filter [CM16, LÁPMF19, PL11]. filters [Zha15]. financial [LT13, TSZ13, YW16]. Finding [Kat11, AMR11b, HC13, MDT^{+}12, Pro10, WY14]. finitary [Bas11]. Finite [Hab13, MYJ11, SG12b, BWW19, Cha19b, CXCK13, HJ16, JJ12, LLLP19, MR11b, Miy11, NS10c, RS10, Tor15, XLJ11, YK15]. Finite-sample [MYJ11]. first [GSG12, IYA16, Lwi11, Mac10, WW11b]. first-order [GSG12, Lwi11, Mac10]. Fisher [ALS12, CPH19, CE11, HN12, HG13, LS12]. Fisher-Snedecor [ALS12]. fit [BT16, CD15, DLZ14, DF12b, GG10, GX12, KS10b, Kou11, Kut11, Li12, LC11, MSA14, Mir10, RH11, SS12a, VT18, VSK^{+}11, Wie19a]. Fitting [JL11, DB11, GN16, Kon13, ZG12]. fixed [AMRP18, DKL13, Kou11, MN16, SB19, Sek12, SvdV15, TCC12, VB10a, Woo19, Zho19]. fixed-sequence [TCC12]. flats [TL11, Tsa17, TL18b]. Flexible [LVK10, BGT16, LQ17, SLL10]. flow [LA14]. FMRI [Alo11, ABK^{+}11, CSN19]. folding [YSGN10]. foldover [LQ17, OQL11, WMB10]. food [BS11a]. forces [AX19]. forecasting [FJKW13, JL11, KG10a, MF10]. foreign [AK10]. form [CG14, CNdGAL11, KP11]. formal [DKY17]. forms [BU10, DG13a]. formula [LZL^{+}10, ZW11]. forum [Goo19]. four [CY10, CDZ11, HHJ13, ZC12]. four-level [ZC12]. Fourier [YL13]. Fourier-polynomial [YL13]. fourth [YHS19]. fractal [TS12]. fraction [MN16, PMX12]. Fractional [ZL11b, AHO13, Aok19, Bro10, CUW10, CKG19, GF13, Jac11, KDJ13, LST12, LYL11, LG12, Men13, OQC13, OZQ19, SD10, ySDM13, SY18, WMB10, WH11, WZZ19, WZ11b, XZZ18, ZC12]. fractionally [PVB^{+}19, SZ12]. Fractions [Fon11]. Fragility [FF12]. frailty [GC19, Lin12, LZL12, dUJC12, MV11a, MPRG10, Unk17]. frame [EO12]. framework [BWW19, BP14c, KSKE10, RV16, SLgX18]. Fréchet [KPR14, ZLL16b]. Fredholm [YAW18]. Fredholm-type [YAW18]. free [BP14c, BSG16, DDW16, DE16, HGH12, Que10, RS13a, SD10, Son11b, WV12, XW13b, Xu17]. freedom [CSC14, GF11, Zha12]. frequencies [BLP11]. frequency [CFT10, KLKO15, Kon12, Kon13, KBNB11]. Frequentist [SSL^{+}17, BDG14, Kli11]. fronts [CSR15]. Full [COP10, WW13, KBK15, TRWP10, Wan10b]. function [BC12, Bur11b, CT11, CHH19, DHM18, EMS11, FSS18, FBHZ14, GF11, Gho13, GC11, GRM14, GLT16, HH15, HP10c, HFQ11, HZ18, HBP12, IF10, JZ13, KKKP19, KMW15, LL11a, LSP12b, MPN^{+}17, MS13b, MSA14, Moj12, MSW12, NMS12, NB10, PL10, PH13, Pla13, Sad11, SS12a, sS10a, sScY10, sS11, sS12c, WH11, ZCLZ17, ZX12, Zho10, ZWM12]. Functional [Lia15, BPR19, BLT16, CWWT12, CCGL10, CLO17, CKP17, CH19, Cue14, DRT10, DZZ18, FLTV10, FKLL13, FRM11, GMMC11, KHN16, KSL11, KSKE10, LL11a, Lia12, Lia14b, LLV15, Lon12, OG10, SLC15, ZW12a]. functionals [CG14, Mab16]. functions [Aok19, Aya12, BW12, DF12a, DR13, DPM13, DM13, DKW17, FP12a, FKMR19, FRC11, GD18, GJ10, GT12a, Hab15, HRR10, HZ13, HYF11, JKL17, JM10a, KK15, KL10a, LRB11, Leb12, LQL12, Men12, MZ10, Ogi19, QLL11, Ruk19b, Ste11a, SR14, SZ13, SSDM14, TRL13, VB10a, WQ10, WMH12, XL12]. further [Han14, Wes16]. fused [SL19a]. Fusion [HFI^{+}18]. future [Fre12]. fuzzy [AS10b].

G [De 13]. gained [WLS13]. Galton [GP15b, SSZ16]. Gamma [KX11a, LMKS17, MV11a, DH11b, GIS19, HMW12, LS10a, NB12b, TTB11]. GARCH [HP10a, HZ11a, PC13, XZ19, Zhu12]. Gauss [HP10c]. Gauss-Markoff [HP10c]. Gaussian [Alo11, BS11b, CSK17, CH11c, Dhi18, DG19, Gho14, GRD16, HS14, JGK11, Kak18, KKKP19, KSN18, KKL15, LWZ13, Maj14, McN10, RS16, SSDM14, VSK^{+}11]. GDDs [Seb19]. GEE [ZLL17]. Geer [RS13b, SS13a, BC13, Kol13, Lec13]. gene [LCC^{+}17, MX15, MS12b]. General [Kub14, LZ12a, LBH17, SBAC19, ZC10, AS15, BHS12, Bao18, CKMS12, CZ10, CSC14, FZ13, GX12, Hir19, HP10c, HZ11b, MS12a, MCC12, Mei11, Men12, NJ16, Ogi19, PL10, PMW^{+}16, SLgX18, Son11a, SC16, SGL13, Tia13, VS12, Zha10, ZW12b, ZLZK13, ZLL16a, ZZ14b, ZGZ17]. Generalised [GSZ11, BK13, EZ12, Kra10, PT13]. Generalization [KHS12, Gra10, Nog13, OJMRCSRA11, dC12]. Generalized [BB14, CQ11, CLL12, Cíz16, EA10, GC19, ITA^{+}09, IT10, ITA^{+}12, Seb19, TLM12, Wei11, AJM19, AFHZ14, AF12, BFZ10, dZBK10a, dZBK10b, BH10, CZ12, CW19, Cha13, CCR11, CC19, CZHC14, CT11, DM14a, EZS11, FMZ10, FVB11, FC12, GF11, GK12, Gua12, He11, Ho10, HLY12, HCMP12, JM11a, LH11, LST12, Lon12, Mar18, vMMA14, MV11a, Par10b, RM11, RP11a, RTM10, RGW10, SYA11b, SK12, SX11, Tso11, VY14, WMM^{+}10, Won12, WG12, XL11, XXC11, YYX19, ZSA10, ZWZ10, ZZ14b, Jan15]. Generalizing [RR10, Zou14]. generating [GJ10, Gho13, Ste11a]. Generation [TB12, AJM19, PS15, SD10]. Generic [Kol13, Lec13, vdG13c, vdG13b, BC13]. genetic [AX19, ZZOY13]. genotyping [Rab14a, Rab14b]. Geometric [GRS^{+}12, MRW11, Che10a, DGG19, DEF11, KPQ11, TNWW12, CLB12]. Geometrical [She14]. geometrically [Cha10a, ZCX11]. Geometry [sCK11]. Gibbs [SM17]. Gini [SDZ19]. given [JDB13]. GLM [LP12]. GLMM [LJK12]. GLMs [EE17]. global [LHH16, WTZ12]. GMANOVA [HLY12]. GMM [Hay16, LW15]. good [RS16, YLL17]. Goodness [DLZ14, GG10, GX12, KS10b, Kut11, BT16, CD15, DF12b, LC11, MSA14, Mir10, RH11, VT18, VSK^{+}11]. Goodness-of-fit [DLZ14, GG10, GX12, KS10b, Kut11, BT16, CD15, DF12b, LC11, MSA14, Mir10, RH11, VT18, VSK^{+}11]. GPU [YBB10]. gradient [GB19, LH14]. gradients [TG16]. gradual [ST11]. Graph [SD10, ABFG12, AMR11a, CC19, RP11b]. graphical [CSK17, KKKP19]. graphs [HGH12]. Graybill [MYJ11]. Greedy [Giu16, CLP13]. Greenwood [Mir10]. grid [DLDC12]. Groeneboom [Jon13a, Mam13, vdGM13]. Group [jLP11, LLW16, SCHG19, BE14, CY10, CYW10, CSN19, CCS14, Fan10, HHMO15, LZL16, PY11, RR10, RSS19, THG12, TW11, Tsa19, WTL11, YHS19, ZC11, Zhi10]. grouped [DK17, LSZ16, ZZ14a]. Grouping [Gui19, MY13, Mei13, MPT13a, MPT13b, Obo13, Yua13, vdG13a]. groups [HDC14, LÖ11b, TXLV11]. growing [SC11, WS14, WM18b, XGK16]. growth [DM14b, HX16, Kas17, KMW16, Son11a, SC16]. Gumbel [HCZ19]. GWP [BK11c, BK10].

Hadamard [Seb19]. Haenszel [HY10a]. Hamburger [SLD14]. hand [FJC12]. Handling [PL14]. harder [JBB10]. Härdle [HL13, KL13, Xia13, YDZA13]. having [SXX11]. hazard [BK11b, CCR17, DDL10, DL12, DMP18, GJ12, GLT16, JM10a, KWL14, LM17b, ND11, QRGLDMA^{+}12, WK13]. hazards [Bha11, DY19, FN17, HLS12, LHH16, YLC^{+}16]. health [DDD18, RS15, dLJ10, dLJ12]. heavy [BT16, BP10a, BMNY13, EGGG12, Nan12, WW11b]. heavy-tailed [BP10a, BMNY13, Nan12, WW11b]. hedging [PL16]. height [BK13, MSTT11]. Hellinger [KL12]. Herriot [Hir19, SM11]. Heston [BBKP19]. heterogeneity [CLM12, Ruk14, Ruk19a]. heterogeneous [CH11c, Ruk11, ZB11]. heteroscedastic [CZ11b, Giu16, GZH11, LGLO12, MS18, WL14, ZK17, ZL11a, Zha12, ZSM10]. heteroscedasticity [GDHS12, LSP12a, ZGZ17]. heteroskedastic [CGL11]. heteroskedasticity [Cav18, CNdGAL11, Lee18]. hidden [ABMP15, TLBJ13]. Hierarchical [Tor12, BF13, PHW15, SLC18, ZGW^{+}17]. High [Mei13, ZGZ17, ABK^{+}11, Alb10b, Alb11, Alb12, Bar12, BP10d, CPH19, CYZL12, CS16b, EZS11, GLW11, GP12, GR18, GLT16, HCT16, HYS12, IYA16, IYA19, JJY12, Kon12, Kon13, LLL^{+}12, LPY10, LLZ11a, LLZS17, LQ16, LC13, MY13, MPT13a, MPT13b, NYA17, Ni12, NHSP13, Obo13, PD11, PMX12, PS15, SMPF10, SL14a, WMM^{+}10, WTZ12, WX19, YW16, Yua13, ZGK13, ZK15, vdG13a]. high-dimension [IYA16, NYA17, NHSP13]. High-dimensional [ZGZ17, ABK^{+}11, BP10d, CYZL12, CS16b, GLW11, GLT16, HYS12, IYA19, JJY12, LLL^{+}12, LLZ11a, LLZS17, LQ16, Ni12, SL14a, WTZ12]. high-quality [Alb10b, Alb11, Alb12]. Higher [IKU11, MV15, Nkw10]. higher-dimensional [Nkw10]. Higher-order [MV15]. highest [CM10]. highly [CT10a]. Hilbert [CC18]. Hill [HCT16]. history [Hum10, KSL11]. holes [DvB15, SKS10]. Homogeneity [TQ12, Nku10, PA12, Ver17, YSH11]. homogeneous [ABMP15]. homoscedastic [Qiu10]. Honest [SvdVvZ15]. horizon [Tor15]. Hsieh [DN12]. Huber [Hu13]. Hui [SS13a]. human [OG10]. Hunting [FBHZ14]. hurdle [CZ11a]. Hybrid [Seg15, DZZ18, GMSK12, MRY11, TNWW12]. hyper [XGK16]. hypercube [AHL12, DD12, GS11a, GS12a, SLL10, YL13, YLL14]. hypercubes [TB12]. hyperoctahedral [YHS19]. hypervolume [CSR15]. hypotheses [BS14, Cab10, CR10, DB12, Dic13, GBP15, KL12, LSZ16, NHSP13, ZZ14a]. Hypothesis [ALS12, BPR16, PH17, BS10a, BHS12, GIP17, GRCGL12, Kim10, Kli11, LZL^{+}10, LNdDLdCH11, MCC12, NN10, RA10, Thu14, ZCP13, ZGZ17].

i.i.d [RL17]. IBC [Ano10t, Ano10u, Ano10v, Ano10w, Ano10x, Ano10y, Ano10z, Ano11q, Ano11r, Ano11s, Ano11t, Ano11u, Ano11v, Ano11w, Ano11x, Ano11y, Ano11z, Ano12m, Ano12n, Ano12o, Ano12p, Ano12q, Ano12r, Ano12s, Ano12t, Ano12u, Ano12v, Ano12w, Ano12x, Ano13m, Ano13n, Ano13o, Ano13p, Ano13q, Ano13r, Ano13s, Ano13t, Ano13u, Ano13v, Ano13w, Ano13x, Ano14k, Ano14l, Ano14m, Ano14n, Ano14o, Ano14p, Ano14q, Ano14r, Ano14s, Ano14t, Ano14u, Ano15-32]. ICODOE [Ano14v]. ICODOE-2011 [Ano14v]. ideals [AHO13]. Idempotent [TS12]. identical [SAE11]. identically [NR12]. identifiability [AMR11a, LSWY16]. Identification [SR13a, KKK17, SC10, WL16]. identified [LM19a]. identifiers [DS10, WS15]. identify [GF13]. identifying [RS15, SG12a]. identities [Eri10]. identity [Fis12]. IDMRL [Ani12]. IFR [Ani10, Ani13, MA08]. IFRA [IK12]. ignorance [BZ12]. II [BS14, dZBK10b, CL10b, GMSK12, IB11, KR12, MRY11, Say12, SBCK11, TNWW12, VB10b, Wan10a]. II/III [TNWW12]. III [BBB11, TNWW12]. ill [AP14b, JR10]. ill-conditioned [AP14b, JR10]. illustration [BM11]. image [BP14b, Huc14, TKB11]. images [WY10]. imaging [CKP17, Zha15]. imbalanced [Sei14]. impact [AS10a]. implementation [Kim10, ZCLZ17]. Implementing [WCZ17]. implicitly [Sum17]. important [SM17, WZZ19]. imprecise [Hab10]. improper [LT18]. Improved [BK11b, BC11b, BK10, BK11c, CW10, COS12, CRW13, GW11, JM10b, KP11, Liu11b, MS13a, SK10, YS11a, CH11a, CNdGAL11, MY10, SM16]. Improvement [MYW11, DL11, FRC11, VB10a]. Improving [DN12, KS12, RS13a, TWLQ15, YK15, YSGN10]. imputation [CH19, GD10, MH16]. inactivity [JZ13]. inadmissibility [KS93, KS11b]. INAR [ZWZ10]. incidence [sScY10]. inclusion [CBD11, Fre11c]. incomplete [ALLL13, ABW13, BS10b, CYC11, CYCZ12, DD13, HFQ11, JP12, LW10, SR10, ZGST12]. incorporate [BBS10, HHMO15]. incorporating [TN10]. increasing [HCZ19]. Independence [Sør19, BSG16, FWS18, GR15, HGH12, LS12, NBSB13, Nog13, Que10, YYX19]. independent [CLM12, DDL10, KA11, Kro19, Lew18, MH10, Mai11, MP19, MM12, Pas10]. index [Ano19a, BWW19, CW19, CY10, CYW10, FF12, FRK14, GGS12, HCT16, Hua11, Kon12, LW11a, Lin11, LX18, Lom11, MN16, NDD16, PH11, SBAC19, WSHW10, WW15a, YY14, YTHL13, ZC11, ZLZL16, ZCLZ17, ZLL17, vdMC10]. Indicator [PL10, Aok19, CSR15, SSS12]. indicators [BCG13, LW12a]. indices [Alb10a, FKLL13, SSDM14, WTZ12, ZZ10]. indirect [LLM11]. individual [LZZ15]. Indivisible [EW11]. Indonesian [FJK11]. induced [DFM10, RQ15]. inequalities [AP11, DE16, LYH12]. inequality [GK12, MHK10, ZGST12]. Inference [AA11c, AX19, Ani10, BP13a, BP19, BRB18, BK11c, GS12a, GP15a, HMW12, IT10, KL10a, KZ15, KS11b, MRY11, PK13, RS15, RM11, She14, TLBJ11, TRL13, Wu16, YS11b, ZWZ10, AMRP18, BFT15, BM11, BAHS11, BGK^{+}11, BGS10, BP15, BLT16, BP14c, CG11, CZ11b, CW10, CRW13, CGCL10, CP12, Chr14, CS11b, DGG19, DP11, De 13, DKY17, FSS18, FLW13, GLW11, GMSK12, GIP17, GMSK10, Gri11, HTZ14, Hil11, HW13, IB11, KL12, KR12, LT11, Lau12, LS10a, LN15, LCL11, LM13, LW11b, LLZ11a, LYLZ12, LZ12b, MZ10, NR11, O’H13, Ozt12, PZ12, Rab14b, RS13a, RP11b, Sch10a, SLgX18, SLZ10, SM17, TN10, Wal13a, Wal13b, WW13, Wei12, XL12, YS11a, YZ12, YYZ14, YLC^{+}16, ZZOY13, ZXD12, ZSM10, dLJ10, dLJ12, dOCG12, CMCSC12]. Inferences [CW19, BZ12, GSZ11, Hua11, LCC^{+}17]. inferential [Mar18]. inferiority [WLL^{+}17]. inferring [LZL16]. infinite [BMNY13, CFW14, Hun12, LS12, MV11a, SZ19, Tar10, XXWL12]. infinite-armed [Hun12]. Infinitely [Sei14, Vag10]. inflated [Zhu12]. inflation [DDD18]. Influence [BCMRMP11, HH14c, BDvdA15, COP10, HRR10, Wie19b, WQ11, ZSH12]. influencing [KV11]. Information [CF12a, CS16a, LYF14, Le13, RRS11, SN10, AS10b, AFHZ14, BE14, Bra10, CE11, EE17, HN12, Kom11, KBK15, KHS12, LH14, LdUÁ11, LS12, LLZ11b, MP19, VS12, WYH10]. Information-based [CF12a]. informative [APQ10, BD14, FZS16, MP19, SSZ11, ZLZ^{+}16, ZZW12]. informatively [CYC11]. Informed [PL16]. ingredients [HHJ13]. inhibition [AB14]. inhomogeneous [BP10c, BSvZ15, CG14, JZ11, Loh10, Wu17]. innovation [CSW11]. innovations [SZ19]. input [FM19]. input-output [FM19]. inputs [Mor14]. inspected [SS00, TLBJ11, TLBJ13]. inspection [TLBJ13]. instrumental [JVV13]. Integer [MSP10, BK10, BK11c, HdGT12, XZ19, ZLW10, Zhu12]. Integer-valued [MSP10, HdGT12, XZ19, ZLW10, Zhu12]. integrability [YW10]. integral [Gro13a, Jon13a, LWZ13, Mam13, vdGM13]. integrals [DKMdUÁ16, Leh11, MLa11]. integrated [BCG10, BK11b, LLJK13, Lu13, PVB^{+}19, SZ12]. integrating [GGS12]. intensity [BSvZ15, CG14, GC11]. inter [HFI^{+}18]. inter-laboratory [HFI^{+}18]. interaction [BT19a, FRK14, HKN18, LYL11, PR16]. interactions [CC12b, GSG12, LT14a, LG12, MX15, PBND11, ST19]. intercept [DH11a]. Interdisciplinary [Ano10-27]. interest [AK10]. interference [DT12, FR13, KM11, Rab14a]. interlaced [Ton13]. Intermediate [Ing10, BO11]. International [Ano10-27, Ano10a, Ano10-44, Ano14v]. intersection [DB19, Nil11]. intertrade [DHH10]. Interval [PS14, Wan10a, Won12, BFCGR12, Gua12, KCZ13, Kon12, LYC12, LP12, MH10, Say12, SS13d, TQ12, TXLV11]. intervals [BI12, Bur11a, CWH10, CR13, Coa14, CL11b, Fre13, GP16, HLB^{+}16, HK10, HvdM12, KF12, KHN16, Kim16, Kre16, KP11, KX11b, LQL12, Lia12, MH10, NL16, PP16a, PP16b, Ruk12, TLBJ13, VB10b, Wan10b, WHI12, Woo19]. interventions [SG12a]. Intrinsic [Lau12, DLPW14]. introduction [DG14]. invariance [Grö18]. invariances [GRD16]. invariant [MR11b, NB12b, TDW10, YHS19]. invariants [RP11b]. Inverse [VSK^{+}11, AGPS12, BF13, Cha12, Kak18, LLM11, Men12, sS10a, WN13, YY10]. inverse-probability-weighted [sS10a]. inverse-Wishart [BF13]. invertibility [CT10b]. invertible [AC11]. Investigating [BS11a]. investigation [Say12, Wes16]. involving [KPQ11, KL10b]. irregular [HY10a]. isomorph [SD10]. isomorph-free [SD10]. Isomorphism [LYL11]. Isotonic [GJ12, ÁY12, WMO15, XI12]. Isotonized [LM17b]. isotropic [Lav11]. isotropy [MS12a]. Issue [Det14, GKvdVvZ15]. issues [HDC14]. item [WFG^{+}15]. iterated [Leh11]. iterations [Cla19]. Iterative [JVV13, vMMA14].

J [AA11c, BP13a, BP19, BRB18, BK11c, Det14, GS12a, IT10, KZ15, TLBJ11]. J. [MGD14]. Jackknife [FP12a, SDZ19, sS10b, LBR12, SA10, YH12].

Jagdish [Gho11]. January [Ano10-35, Ano11-37, Ano12-37, Ano13-37, Ano15-34, Ano16-39, Ano17-40, Ano18q, Ano19j]. Jeffreys [HG13]. Joint [FZS16, BS10b, DM14a, DDD18, LW11b, LSP12b, MZ10, ZLZ^{+}16]. jointly [LR11]. Jonckheere [VB11]. Journal [Ani10, KS11b, YS11b, dLJ12]. July [Ano10-38, Ano11-42, Ano12-40, Ano13-39, Ano14-39, Ano15-37, Ano16-42, Ano17-42, Ano19i, dLJ12]. jump [BBKP19, Kon12, Kon13, SLW13]. jump-diffusion [Kon13, SLW13]. jump-type [BBKP19]. jumps [LWZ13]. June [Ano10-37, Ano11-41, Ano12-39, Ano13-38, Ano14-40, Ano15-38, Ano16-41, Ano17-47, Gho11]. jury [CHP11]. justification [LC16]. justifications [Thu14].

k1 [DAP10]. k2 [DAP10]. Kalman [CM16, LÁPMF19]. Kaplan [DKMdUÁ16, NB10]. Kernel [DPT11, ES12, GBdW10, THG12, Ton13, BK11a, CLL12, CL10a, CLP13, CC18, CPSY10, Dhi18, Dut11, GZ10, GGS12, HFQ11, HWZ11, IK15, KF10, KLS11, MY11, Oua13, WYY^{+}10]. kernels [CGC12, Kak18, KKP10, Nog13]. kinds [IAH11]. kinetic [AB14]. known [BP10a, HDC14]. Koenker [Zou14]. Kozachenko [DF17]. Krawtchouk [DG14]. Kriging [AZ10]. Kullback [BHS12, SlG12, WV12]. kurtosis [WN11a].

L [NV10, NL16]. L-moments [NV10]. l1 [vdG13c, YY14]. L1-penalized [YY14]. l1-penalty [vdG13c]. L2 [CSW11]. L2-norm [CSW11]. labeled [LQ16]. laboratory [HFI^{+}18]. Lack [Li12, Kou11, SS12a, Wie19a]. Lack-of-fit [Li12, Kou11, SS12a]. LAD [YY14]. Lag [ZK17, CK13, Kim16]. Laird [JBB10]. landmark [Huc14]. landmarks [BP14b]. Laplace [BPR19, CN12, EZS11, IB11, Miy10, Miy18, Var15]. Laplacian [CC12a]. Large [BI12, IR19, JSV12, JZ11, Mac10, ZL12, BMPT19, CNL17, DPC11, DLZ14, DvB15, GR14, HG11, HY10b, LCQ^{+}13, LB15, LL11d, PA13]. large- [DLZ14]. Large-sample [BI12]. Lasso [Kau14, SL19a, CT10c, TMK10, YW16, AHD12, AH12, GP19, WZW13]. lasso-type [TMK10, GP19]. Latent [GD10, BCG10, KMW15, Kom11, McN10]. Latin [GS12a, DvB15, EW11, AHL12, ALLL13, DD12, GS11a, Soi13, SLL10, YL13, YLL14]. Lattice [Ano10a, Ano10-44, Böh10, Eri10, CW17, FGN10, Hum10, Nkw10, Qui10, YLL17]. Läuter [Tsa11]. Läuter-type [Tsa11]. law [ACC18, Vag10, dOCG12]. laws [HdGT12, He11, KZ16, RP11a]. leading [WS15]. Learning [CL10a, CLP13, mFS12, ZCX11, BW10b, BW12, DG19, HFI^{+}18, HXZ12]. Least [FMZ10, XZZ18, Dhi18, DB19, mFS12, FJKW13, FK11b, Han11, HLY12, Kat11, KK12, LW12a, Ni12, PKL12, PH11, Ros14, ZG12, ZWY11, vdGM15]. least-squares [mFS12, Ni12, ZWY11]. leave [YLL17]. leave-one-out [YLL17]. left [BM11, KWL14, LdUÁ11, sS10b, sScY10, sS12b, SM11]. left-censored [KWL14, SM11]. left-truncated [LdUÁ11, sS10b, sScY10].

Leibler [BHS12, SlG12, WV12]. lemma [LVY13]. length [FSS18, HK10, LZZ15, WW15b, WZ18]. length-biased [FSS18, WW15b, WZ18]. Leonenko [DF17]. Lerch [Kem10]. letter [LYL11]. level [AHO13, Aok19, Cha19a, CKM11, CKMS12, Dhi18, GSG12, GSKP10, GCDK12, HX16, HLSL17, HZ11b, HM13, Jac11, KGV10, LR11, LT14a, LQ17, LT14b, LST12, OQL11, OQC13, QWC12, QCG15, RLZ12, SGS12, SLC18, SD10, TX13, TLM12, TL18b, WZ11b, YL14, ZC12, ZLZK13, ZLL16a]. levels [EQ15, LJA15, LL11d, OL19, SLC18]. Lévy [BC19, CL13b, GS10]. Li [GP16]. lies [Zhi10]. life [Ani12, CZ12, DN13, FPLR11, Gup12, KA11, LP12, LT10, NV10, Pas10, Sad11, SZ10, Wan10a, WLCK11, XSJ14, XT15]. life-testing [Wan10a]. lifetime [DL11, MCSC10]. lifetimes [MA10, MPRG10, SAE11]. light [EGGG12]. like [BP10d]. Likelihood [BM11, JJY12, LRB11, ZGST12, ÁG12a, AMRP18, AAV13, BFZ10, BMW16, BP10c, BBKP19, BWYZ11, CW10, CRW13, CH11b, CW17, CT11, sCK11, Chr14, CS16b, Cop11, CH11c, Dac10, DKY15, FSS18, FLL17a, FP12a, GMSK10, Gro13a, GH10, He11, Hir19, HZ18, IB11, JL11, Jon13a, Kon12, KBK15, KWL14, KG10c, LBL12, LQ11, LF11, LP11b, Lia12, LHH16, LYLZ12, LZ12b, LX18, LW15, LNdDLdCH11, Mam13, Miy11, NS13, NCC11, Nik13, NL14, PC13, PHW15, QL10, QLL11, RS16, RWB15, Ruk12, SDZ19, She11, SHW16, SCR10, Sør19, SC13, TRWP10, Tar10, VBLN17, VCA10, VSK^{+}11, VGH13, VDW13, VZH17, WQ10, WLCK11, WZ11a, WL11, WZ18, WWZ19, WW11b, XL12, XZ19, XR11, XXC11, YZ12, YYZ14, Yao10, YCY15, ZY11, vdGM13]. Likelihood-based [ZGST12, CW10, CRW13, DKY15, GMSK10]. likelihoods [BCG10, CM10, DS18, VCR10]. Limit [BO11, Lee18, MZ13, CHH13, CT14, CL13b, KZ16, Lon17, MO10a, RL17, Sai10, SY18]. limitations [KCF12, PR16]. limited [PG10]. Limiting [YW10, Dac10, DLPW14, GD18, LZL^{+}10, MP12b]. limits [KL10b]. line [CC19, DZ18, KK16, Mab16, RR12, TCE^{+}12, ZL17]. Linear [LCT10, LR11, Mei13, WDRB18, AS10b, BFZ10, BM12a, BFCGR12, BLW16, CLM12, CW19, CWH10, CYZL12, CZHC14, CT11, CR15, CP12, CB19, CM16, CH19, DDW16, DRT10, DS12, Dio14, DB11, EA10, EZS11, EZ12, FT10, FYX14, FKLL13, FC12, Gaf19, GS19, GLW11, Gho14, GC17, Giu16, GP16, GMMC11, GDHS12, Grö18, GX12, Han11, HP10b, HTZ14, HP10c, Hu13, Hua11, HCMP12, HHJ13, HZ13, HY10b, IF10, JGJR16, KHN16, KSL11, KKL15, KM12a, KS16, Kre16, KS11a, KHS12, LW11a, Lem12, Li11, LZ12a, LZ13b, LH14, LM19a, LL11c, LLZ11a, LSWY16, LAk10, Liu11a, Liu11b, LYL13, LYC15, Lon12, Lu13, LLL18, LC13, MY11, Mab16, MY13, MS13a, MYW11, vMMA14, MPT13a, MPT13b, MB11, MHK10, Ni12, NHSP13, NGXZ16, NL16, Obo13, PP16a, PP16b, PR16, PR12, QL10, RS19]. linear [RS10, RGW10, SB19, SC11, SLC18, Shu12, SX11, Son11a, Ste11b, SGL13, SSL^{+}17, TRL13, Tia13, Tsa19, Tso11, VY14, WZW13, WS14, WL16, WYY^{+}10, Wei12, WN11b, WW12b, Wu19, XL11, XXC11, YS11a, YZ12, YY14, YS11b, YS11c, YTHL13, YAW18, Yua13, ZW12a, ZX19, ZW12b, ZGST12, ZGZ17, vdG13a, vdVW14]. linearity [BPP10, MO10a]. lines [jLP11]. link [FKMR19]. linkage [LC16, ZGW^{+}17]. Liu [Liu11b]. load [DDNN10, LL11b]. load-sharing [LL11b]. loaded [LT10]. Local [AR11, AF12, CW17, HL13, Koe13, KL13, Lem11, Lem12, Pav12, SWH13a, SWH13b, Xia13, FJKW13, Gas11b, Gho14, GJ12, HTZ14, QM11, SL14b, SGL13, SBAC19, TvS15]. Locally [GIS19, NN10, PP10, RS19, HH15, Pal10, PL11]. location [BI12, Cla19, IF10, KX11b, KS11a, LW11b, LLLP19, NB12b, Ozt11, VdUÁMM11, ZN12]. location-scale [BI12, KX11b, LLLP19, VdUÁMM11, ZN12]. locus [Rab14a]. log [COP10, sCK11, DH11b, DKW17, FO13, HM10, HHJ13, LD12, MV11a, PR16, PH13, RP11a, ZSL14]. log-bilinear [FO13]. log-binomial [ZSL14]. log-Birnbaum [COP10]. log-gamma [DH11b]. log-likelihood [sCK11]. log-linear [PR16]. log-logistic [MV11a]. log-rank [LD12]. log-variance [PH13]. Logistic [BMW15, HLS10, BAHS11, DMP10, DELS16, HKN18, LL19, LC11, MV11a]. logistic-normal [LC11]. logit [GVV10, WT10, ZH12]. loglinear [JMW13]. lognormal [BM11, HvdM12, NB12a, Pas10]. Long [BH15, DHH10, BKPS14, CLL11, FRM11, Gho13, God10, Kau14, KS10b, KL11, MGB10, Pal10, PV10, RLLT17, WC10, Wis13, ZL11b]. long-memory [CLL11, Gho13, Pal10, PV10, RLLT17]. long-range [BKPS14, FRM11, Wis13]. Long-run [BH15]. Longitudinal [HLH13, BFZ10, CW19, CYC11, CYCZ12, CPSY10, DELS16, FZS16, HCMP12, KSL11, Li11, LZ13b, LW11b, LL15, PA12, Qin10, SSZ11, TWLQ15, TP10, WZ11a, WQ11, ZLL17]. look [GGG11]. Looking [CDd11]. Loss [KKKP19, BW12, DGS19, FBHZ14, GF11, GW11, GG15, HP10c, Men12, NMS12, Ogi19, XLJ11, YC11b]. Lot [MR11a]. Lotka [SSZ16]. Low [PL11, IYA16, LQ16, NYA17]. Low-pass [PL11]. low-sample [LQ16]. low-sample-size [IYA16, NYA17]. Lower [EQ15, GR11, JDB13, BK10, BK11c, CKM11, CZ10, CHP11, FK11a, HZ11b, ZC10, ZLZK13, ZLL16a, ZZ14b]. lower-order [HZ11b]. LROC [TB11]. LRT [CNL17]. LS [HZ11a].

M [ÁY12, PC14, Zho10]. M-estimation [Zho10]. M-estimators [ÁY12]. M-procedures [PC14]. M. [vdG13a]. MA [CRW13]. Machine [WYH10, NYA17, PKMK12]. macro [CS11b]. macro-data [CS11b]. main [BT19a, CC12b, ETMJ11, LT14a, LT14b]. main-effect [LT14b]. maintenance [MZP12]. Malliavin [Yan17]. Mallows [Oga17]. Manifolds [Ber14, BP14b, Huc14]. Mantel [HY10a]. manufacturing [jLP11]. mapping [CP12, KCZ13, Rab14a]. March [Ano10-42, Ano11-38, Ano12-47, Ano13-46, Ano14-38, Ano15-40, Ano16-44, Ano17-37, Ano18m, Ano19m]. margin [WLL^{+}17]. Marginal [CYCZ12, DY19, BK17, LS12, Que10, YSH11]. marginal-free [Que10]. Marginally [HLSL17]. marginals [DKL13, Zho19]. marine [BS10c]. Markoff [HP10c]. Markov [ABMP15, AT10a, AHO13, CH11b, CS11b, KCF12, KLK11, Lon17, MK18, MZ13, Nog13, Reg11, SM13, YL10b]. Markov-dependent [MZ13].

Markov-switching [ABMP15]. Markovian [BPR16]. Masking [OK11, DS10, WS15]. mass [MZ10]. massive [ZZW19]. Matched [GD15, YSH11]. matched-pair [YSH11]. Matching [MP19, CM10, Has19, Ho10]. Mathematical [Ano10-27]. matrices [AFHZ14, BF13, CNL17, GS11a, GS12a, HP19, IYA16, IYA19, JJY12, LÖ11b, LB15, Lia11b, NHSP13, Nkw10, Sch10b, Seb19, SYL13, Wag11]. matrix [AP11, BMW16, BMPT19, CT10c, Di 16, DG13a, DM13, FKMR19, Fis12, FO13, GS19, Hay16, LCQ^{+}13, LYL11, LLZS17, MJS12, MYW11, MLa11, PA14, She14, Tsu10, XLJ11]. max [CDd11, HdGT12, RP11a, ZS10, Alb12]. max-semistability [CDd11]. max-semistable [HdGT12]. max-stable [ZS10]. maxima [JM10b, JM11b, KPQ11, TSZ13, Vas13]. Maximin [Mor14, Wie19a]. maximization [BE14, Wag11]. Maximizing [SKFMR14, FRC11]. Maximum [AAV13, BWYZ11, CH11c, GLR11, GJM18, KWL14, KG10c, LH11, Miy11, SC13, ÁG12a, BP10c, BBKP19, CH11b, GH10, HdGT12, Hir19, JL11, LRB11, LBL12, MR11a, NCC11, PC13, RWB15, RP11b, Ruk12, She11, SHW16, SCR10, Tar10, WZ18, XR11, YW10]. maximum-likelihood [Tar10]. May [Ano10-36, Ano11-40, Ano12-38, Ano13-48, Ano14-36, Ano15-36, Ano16-45, Ano17-39, Ano19l, BF12, LS12]. MCMC [RRS11, SL14a]. Mean [DLPW14, GSS10, AVML11, ADPW10, BLBL10, BLP11, BS11b, BMNY13, Buc14, COS12, DN13, GP10, Gup12, HFW10, Hir19, Jar10, KA11, Kli11, KZ16, KM12a, LM11a, LL15, LLZS17, LSW13, LGG12, MR11a, MKT15, NHSP13, Pal10, PA13, PR12, Sad11, SSZ16, SK10, Son11b, SZ10, Tso11, Tsu10, VB10a, VS12, WX19, WN13, XSJ14, YK15, ZK15, ZWM12]. mean-covariance [LL15]. mean-variance [Tso11]. Means [Alb10a, COS12, CC15, DRT10, HK10, HvdM12, LCT10, LLL^{+}12, LR11, Lia11c, Miy10, PD11, Qiu10, RH13, TN10, VAM10, ZSM10]. measurable [BET17]. measure [CND10, CKM11, Lom11, NS10a, NJ16]. measurement [DMTZ15, DRT10, DR13, GK17, HK19, HZ13, HCZ19, KS16, SXX11, SS16, WZW13, WQ11, WG12, XSJ14, ZYZL14]. measurements [MSBM16]. measures [BI12, BM15, BDG14, GR14, KMMP11]. measuring [TvS15]. Median [sS12b, YC11b, DJ19, DS18, SLZ10]. medians [LYLZ12]. medicine [LW19]. Mediterranean [BT16]. Meier [DKMdUÁ16, NB10]. Meixner [KV10]. membership [THG12]. Memory [Det14, Che10b, CLL11, DHH10, Gho13, God10, Hil11, Kau14, KS10b, KL11, MGB10, Pal10, PV10, RLLT17, WC10]. MEP.1 [BT19b]. MEP.2 [SC10]. meta [JBB10, Kra12]. meta-analysis [JBB10]. meta-elliptical [Kra12]. metamodelling [AZ10]. metamodels [SSDM14]. method [Bar12, BC11b, CL15a, Chr10, Cíz16, Dal11, Hab12, HTZ14, Kra12, LÁPMF19, LC11, SZZ16, SDNH13, STS10, TX13, YW16, Yao10, ZDLW19, ZY11, dC12]. Methodology [DB11, SR14, BRBM19, HH14c, HC11b, LJK12]. methods [AHO13, CYCZ12, CC15, DB12, Har14, Hir19, Kri10, NL14, OK11, OP12, RS13a, SDZ19, sS10b, WMM^{+}10]. microarray [BSH13]. microbiome [KPR14]. microdata [OK11]. middle [DP11, sS10a, She11]. middle-censored [DP11, sS10a, She11]. Milankovitch [PPB16]. Minimal [OQ10, SYA11a, Qu11]. Minimax [AP14b, BK11a, BLBL10, BP13b, BKPS14, GM12, MR11b, Pav12, Tsu10, WZ11b]. minimax-optimal [GM12]. minimaxity [DIAT11, KS11a]. Minimization [EO12]. minimizing [MKT15]. Minimum [Hab10, Kou11, KS16, LT14b, MBB11, TL13, XM10, Che11a, CZ10, Gaf19, Gam13, HS11b, HZ11b, HM13, KF12, KL12, LMT14, MR11a, SM17, TLM12, XI12, ZC10, ZLZK13, ZLL16a, ZZ14b, vdVW14]. minimum-support [vdVW14]. minimum/maximum [MR11a]. mirror [TKB11]. Mis [TTB11]. Mis-specification [TTB11]. misclassification [BVG10, WLS13]. misclassifications [BF12]. MISE [Dut11]. Mises [HG13, Gas11b, HH15, RP11a]. mismatch [CP12]. mismeasurement [LW12b]. missed [GX12]. missile [AMNS11]. missing [AG10, BCG13, CGLW18, CH19, DMP10, DW17, Día17, Efr11, Efr13, FC12, GD10, Han14, HLS10, HLS12, KV11, LW11a, Li12, LW12a, LLV15, LLZ11b, LZ12b, LLX12, LW15, Mül12, MS18, NGXZ16, QL10, Shu11, Shu12, STS10, THG12, WQ10, WSHW10, WL11, XXC11]. misspecification [HG11, QZZ11, ZLZ^{+}16]. misspecified [sCK11, CB19, De 13, HW13, HG11, O’H13, Wal13a, Wal13b]. Mixed [CSWY11, Mon17, ZC12, BM12b, CLM12, CKM11, CSC14, Dio14, DB11, DT14, EQ15, EZS11, EZ12, FT10, FZS16, FR13, GT10, Han11, Li11, LZ13b, LW10, MP11, QCG15, RAB14c, SB19, SX11, TL11, Vag10, Wan11, Wes16, YS11a]. mixed-effects [LW10, YS11a]. mixed-level [CKM11, QCG15]. mixing [AD11, AP14b, BK16, mFS12, Kre18, Ton13, Vag10, WHY11, WYY^{+}10, WZ10, YHWL12, ZCX11]. mixing-weights [AP14b]. Mixture [AD11, AMR11a, AP14b, BGK^{+}11, Cav18, Dal11, FBHZ14, GS11b, HS11b, HHJ13, KCZ13, LS10b, McN10, MM12, Miy11, MS12c, Pat12, Qiu10, Yao10, YCY15, ZLW10]. mixture-models [AP14b]. Mixtures [AM11, DGCS16, AVML11, CH11c, FVB11, GP15a, KZ16, LGLO12, LVK10, LLLP19, MS12b, MV11b, NZF11, RD12, SdL10]. mle [FJK11, Bro10, LLLP19]. mod [HM13]. modal [FS14]. mode [CHH19, KLS11, Mei11, Nag15, WZ10]. Model [BET17, BM12b, CC11, FC12, GK17, GN16, JB14, LYC15, McN10, SZ13, Xu17, YTHL13, YLC^{+}16, ALN10, AAKR12, AX19, AM11, AS11b, AM12b, BGT16, BDLP14, BBKP19, BS10b, BZ12, Beu10, Bha11, BCG10, BFCGR12, BCZ19, BLW16, Bur11a, CLB12, CZ11a, CW10, CRW13, Cha12, CLL11, CGL11, CSK17, CSC14, Ciu13, CS16b, CHF11, CM16, CH19, DDW16, DN13, DRT10, DGG19, DH11a, DDNN10, DKMdUÁ16, DEF11, DZZ18, DY19, DT12, DT14, EMS11, FZS16, FR13, FKMR19, FHK11, FKLL13, FN17, GLW11, Gas11a, GVV10, GLT16, HKN18, Han11, Hir19, HG11, HP10c, HLY12, Hu13, HHJ13, HYS12, HLS12, IYA19, KKKP19, KW12, KFK11, KS12, KZW10, KCZ13, KK16, KKL15, KS10a, KS16, KPQ11, KWL14, Kro19, KM11, LCT10, LW11a, LM11b, LBR12, LF11, LZ12a, Li12, LL11c, LLZ11a]. model [LM14, LHH16, LS10b, Liu11a, Liu11b, LT13, LZL16, LLS19, Lon12, LM17b, LJ12, LD12, LLL18, Lwi11, MY11, MK18, Mar18, MFABG13, MV11a, Mon17, MHK10, NGXZ16, PA12, Qin10, Qiu10, QRGLDMA^{+}12, RRS11, RH11, RSH12, RS16, Ruk19a, Sai10, Sch10a, SC11, sS12b, SM11, SvdV15, SL14b, Son11a, Son11b, SLW13, SC16, SSC11, SSL^{+}17, TCE^{+}12, Wae10, Wah10, WSHW10, WT10, WTZ12, WZ18, WWZ19, WZ10, Wu16, XGK16, XZ19, XL11, XW13b, XSJ14, YS11a, YSW10, ZK17, ZL11a, ZLZL16, ZCLZ17, ZX19, ZZW12, ZN13, Zho10, ZSL14, ZK15, ZLW10, vdGM15, vdMC10]. model-averaged [BCZ19]. Model-based [BET17, BM12b, JB14, McN10, SZ13, AM11]. Model-free [Xu17, XW13b]. Model-robust [LYC15]. Modeling [DM14b, FC10, Fre11a, LSWY16, AS11b, BS10c, DELS16, JD16, LVK10, MRW11, OG10, PL14, RBS10, TMK10, YW16, YCY15, ZLZ^{+}16, ZLM17]. modelling [FS14, GRD16, HH15, LW11b, PR16, TRWP10, UP14]. Models [Mei13, AVML11, AS15, AHD12, ABMP15, AMR11a, AP14a, AT10b, AMRP18, AFHZ14, AB14, AP14b, BFZ10, BGK^{+}11, Bao18, BPR16, BCMAW14, BCN13, BGS10, BK12, BD14, CLM12, CW19, COP10, CK13, CF12a, CT10b, CZ11b, CLO17, CWH10, CGCL10, CYC11, CYZL12, CPM14, CS16a, Che11b, CZHC14, CT11, CR13, sCK11, CR15, Cho17, CP12, CO10, CA10, Dac10, DM14a, DR13, De 13, DMP11, DP13, DLDC12, DM10, DMS11, Día17, Dio14, DB11, DF12b, EEL10, EA10, EZS11, EZ12, FT10, FLW13, FYX14, FLL17a, FGH14, FMZ10, FO13, Fra14, FC12, GIS19, GZ14a, GF13, Gho14, GN16, GC17, GS11b, Giu16, GP19, GC19, GDHS12, Grö18, GT10, GX12, GLZ14, GZ14b, GSS10, Hab10, HS14, HX16, HS11b, Has19, HP10b, HTZ14, HY19, HZY14, HW13, HH14a, Hua11, HCMP12]. models [HZ13, Hud13, HCZ19, JR10, JMW13, JL11, JAI11, KCF12, KW14, KD12, KK11, KSL11, KBK15, Kra12, KL10b, KHS12, KG10c, LMKS17, LGLO12, LL13, Lem11, Lem12, Li11, LB11, LCL11, LM13, LZ13b, LZ13a, LH14, LMZ18, Lin11, LC11, Lin12, LW10, LYL13, LYC14, LYC15, LYXC16, LX18, dUJC12, LC13, LZH11, LA14, MY10, Mai11, MY13, MZWF17, MS13a, MF10, McN10, MP11, vMMA14, MZ10, MM12, Miy11, MV11b, MR10, MPT13a, MPT13b, MHK10, MPRG10, MQ10, MSW12, MS12c, ND11, NJ16, NR11, Ni12, Nik13, O’H13, Obo13, OZC16, PC13, PT13, PR12, PH11, PHW15, PVB^{+}19, QL10, RF13, RS19, RQ15, RL17, Ruk11, Ruk17, SR13a, Say12, SB19, SLC18, SZ12, SCR10, SRBS11, SX11, SP12b, SHYH16, SDGK15, SZ10, SN12, SBAC19, Tan10, TSZ13, TL18a]. models [Tia13, Tsa11, Tsa19, Tso11, Unk17, VY14, VdUÁMM11, VT18, VV10, Wal13a, Wal13b, WLCK11, WF11, WZ11a, WTL11, WTZ12, WZW13, WY14, WS14, WX15, WW15b, WW15a, WL16, WSL17, WDRB18, Wei12, Wei11, WK13, XW13a, XXWL12, XXC11, YZ12, YY14, YW16, YYX19, YL10b, YL13, YS11b, YS11c, YTHL13, YAW18, Yua13, ZH12, ZK17, ZSH12, ZW12a, ZYZL14, ZS15, ZW12b, ZY11, ZGST12, ZG12, ZLL16b, ZN12, ZL18, ZJ10, Zhu12, vB19, vdVW14, CSN19, vdG13a]. moderate [JZ11, PR11]. modes [Miy10]. Modification [SS13c]. modifications [Oga17]. Modified [KG10b, sS12c, SL19a, AM11, FKV11, He11, HY10b, LJ12, Xu17]. modifying [Hab12]. Moment [Has19, BMNY13, BK12, DE16, Gho13, GP10, Now12, SLD14, YY10]. moments [BAHS11, BS11a, BM12a, CG11, Cíz16, Dhi18, DG13b, Kra12, MV11a, NV10, VZH17, WN11b, Yan17]. Monitoring [WG13, YC11a, APQ10, Alb10b, BS11b, CM11, GO11, HW17, KK15, LSZ13, Lwi11, TCE^{+}12]. monotone [HCZ19, LLP14, LM17b, Shu11, Shu12, Wan10b, WX15]. Monotonicity [AMNS11, CCR11, GJ12, LLL18]. Monte [AT10a, AHO13, BCG10, JGK11, Kim10]. Moran [Gri10]. most [NN10, Ros14]. motion [ES16, FLL17b, HS11a, SvdV15, SY18, ZL11b]. Motzkin [BK13]. Mougeot [vdG13a]. movement [BS10c]. moving [AHKS12, BC19, CL13b, DF10, PVB^{+}19, TSZ13, Vas13]. moving-average [DF10]. MRL [FSS18, ZB11]. Multi [LT14a, Aok19, BE14, CSR15, CKM11, CKMS12, CL10a, CYC11, CLP13, Coa14, CO10, FS14, FJC12, GM12, LR11, LYC14, LD12, MF10, MZP12, TX13, VS12, VB10b]. multi-auxiliary [VS12]. multi-criteria [CSR15]. multi-factor [LYC14]. multi-group [BE14]. multi-kernel [CL10a, CLP13]. Multi-level [LT14a, Aok19, CKM11, CKMS12, LR11, TX13]. multi-modal [FS14]. multi-objective [GM12]. multi-output [CO10]. multi-parameter [Coa14]. multi-sample [VB10b]. multi-stage [LD12]. multi-state [CYC11, FJC12, MZP12]. multi-step-ahead [MF10]. multicenter [LJK12]. Multichannel [BKPS14]. multiclass [XW13b]. multidimensional [BD18, DD12, FKLL13, VM18]. Multilevel [CSK17, FRR16]. Multilinear [GJ10]. multinomial [DM14a, DDD18, JMW13, STS10, WT10, ZH12, ZZ10]. multiparameter [Buc14]. Multiple [Dör10, Hab15, Hoc19, MH16, BFT15, BS14, Cab10, CL15a, CY19, CMS13, DST10, DB12, Dic13, Dör11, DZ18, Eri10, GD10, Hab12, HJ16, LM19a, Lia15, LSZ16, LNdDLdCH11, NdCJ10, TW11, TCC12, WH12, WA16, WCZ17, WFG^{+}15, ZCP13]. multiple-objective [CY19]. Multiplicative [Har14, BLP11, CCR17, LX18, QRGLDMA^{+}12]. multiplicity [Akn15]. multiplier [BK16]. multiply [Han14, SBCK11]. multipower [LWZ13]. Multiregression [CSN19]. multiresponse [LYL13, Sag11, YQC11]. Multistage [SS19, Bur11b]. Multivariate [DG19, FKV11, Mai11, MPN^{+}17, PG10, ABK^{+}11, BS11b, BOvSS10, BR10, BLW16, CPM14, DS10, DH11b, DG14, DG13a, GT12b, GS12b, Han11, HGH12, HLCW12, HOK19, JB14, JB12, JD16, JP12, Kli11, KKL15, KM12a, KPQ11, KHS12, LVY13, LR11, LN15, LM19a, MFABG13, MS12b, MO10b, MHK10, MPRG10, OS10, Par10b, PD11, SG12b, She14, TSZ13, TW11, TS12, Ver13, Wen15, ZY11, ZGST12, Zho19]. mutations [ZZOY13]. mutual [BE14]. Mutually [DvB15]. MV [Jac19]. MV-optimal [Jac19].

N [Gho11, MGD14, NL16]. Nadaraya [SLW13]. Nagaev [SSZ16]. Near [BM18, BL15, TS10, Vas13]. Near-factorial [BM18, BL15]. nearest [Aya12]. Nearly [BTT15, YLL17, SDGK15]. necessity [CK12]. need [YO11].

Negative [MPRG10, ZJ10, Alb10b, CT10a, DST10, DP13, IF10, IKU11, TRWP10, Wan11, XZ19, Zhu12]. negatively [BP10b, BP13a, BP19, LQ11, QLL11, YW10]. negativity [CP12]. negligible [TLBJ13]. neighbor [AA11b, AA11c, Aya12, HYH10, ITA^{+}12, SYA11a, SYA11b, ZSA10]. neighborhoods [AF12]. neighbour [DT12]. neighbourhood [BGS10]. Nested [BL15, BM18, DRT10, GVV10, NJ16, Say12, SYL13, Tsa17, WS14]. Net [CYZL12]. network [DG19]. networked [CSN19]. networks [CSK17, KKK17, SG12a]. neutral [DFM10]. news [PL16]. Newton [Cla19, GS19]. Neyman [Sch10a]. no [GVV10, RBS10]. no-choice [GVV10]. noise [BLP11, Dhi18, Kha10, NU19]. noises [Bro10]. Noisy [Var15, FLL17b]. nominated [JM11b]. nomination [JM10b, JJ12]. Non [ABMP15, CH10, HYH10, KZW10, Kro19, KS11a, Sum17, dLJ10, dLJ12, AM12a, AX19, AG19, BS11a, BET17, Bir15, BOvSS10, BD12, BK16, Bur11a, BD14, CFW14, DP13, Di 16, Gri10, Has19, JD16, Kat19, LL11c, LLZ11a, Mai11, NR12, NJ16, RS19, RL17, SAE11, Say12, SA10, SK10, TLBJ13, VBLN17, WTZ12, WLL^{+}17, WZ10, YH12, vdVW14]. non-asymptotic [Bir15]. Non-atomic [Sum17]. Non-binary [HYH10]. non-causal [CFW14]. non-central [Di 16]. non-constant [AX19, LL11c]. non-convex [Kat19]. non-degenerate [BK16]. non-diffusion [AG19]. non-elliptically [JD16]. Non-homogeneous [ABMP15]. non-i.i.d [RL17]. non-identical [SAE11]. non-identically [NR12]. non-independent [Mai11]. non-inferiority [WLL^{+}17]. non-informative [BD14]. non-linear [RS19, vdVW14]. non-measurable [BET17]. Non-minimaxity [KS11a]. non-negligible [TLBJ13]. non-nested [NJ16, Say12]. non-normal [Bur11a, Gri10]. non-orthogonal [AM12a]. Non-parametric [CH10, Kro19, dLJ10, dLJ12, WTZ12]. non-regular [Has19]. non-respondents [SK10]. non-response [SA10, YH12]. non-separability [DP13]. non-smooth [WZ10]. non-sparse [LLZ11a]. non-standard [VBLN17]. Non-stationary [KZW10, AG19, BOvSS10, BD12]. non-symmetrical [BS11a]. Nonadaptive [Zhi10]. Nonadditivity [JMW13]. nonconvex [YCY15]. nonhomogeneous [LZL12]. nonidentical [Sad11]. nonignorable [PL14]. nonintersecting [Qui10]. Nonlinear [GT12a, TMK10, ZYZL14, BO11, BD14, CT10b, CGCL10, CLL11, Che11b, Ciu13, GP16, HZ11a, HZY14, KW14, Kre16, LGLO12, Lem11, LM13, LSP12a, LW10, MX15, Mül12, NR11, NL16, PP16a, PP16b, SZZ16, TRL13, ZG12, ZL18]. Nonnegative [YW16, Kak18, Mab16]. nonnormal [BU10]. Nonparametric [Beu10, BR10, BCG13, Bur11b, CMT17, CMS13, CR16b, CMCSC11, DS10, DKN12, DPT16, DHM18, DK17, Efr11, FWS18, GO11, Gro13a, GS12b, Hon10, HC11b, JZ13, Kak18, KMT10, LLV15, LZH11, Mam13, MCSC10, NDD16, PH13, RM12, SDNH13, VdUÁMM11, VDW13, WSSG19, vdGM13, Alb11, BFT15, BTKN10, BP13b, BP14b, BRBM19, CLL11, CD15, DMTZ15, DM11, Dio14, Efr13, ES12, FLTV10, Fre13, GLR10, GC11, GP16, Gri11, HH15, HC13, Huc14, JVV13, Jon13a, KW12, Kre16, LZ13a, Lia12, LC11, Luk14, MLP17, MPN^{+}17, MRW11, MO10b, MB11, NGXZ16, NL16, PP16a, PP16b, PKL12, RR12, She11, SS19, TCE^{+}12, VB10b, WC10, WY10, WW13, WS15, WFG^{+}15, WZ10, Xin11, YU13, YQC11, ZK17, Zha15, ZXD12, ZZW19, CMCSC12]. nonparametrics [GKvdVvZ15]. nonregular [LST12, OQL11, ZZ14b]. nonresponse [PL14]. nonstandard [RD12]. nonstationarity [Wes16]. nonstationary [CK13, Cho17, Per10, ZL18]. nonuniform [JWY16]. norm [CSW11, SvdVvZ15]. normal [AVML11, AAV13, BU10, BD18, Bur11a, CLM12, COS12, GMSK10, GF18, GP12, Gri10, GRS^{+}12, HK10, JJY12, KZ16, KM12a, KHS12, LGLO12, LC11, LSW13, MH10, Nik13, Par10b, Per10, Qiu10, RA10, Ruk19b, She14, Tak10, TH14, Tsu10, WN13, Yao10, YKH13, ZSM10]. normal-based [YKH13]. Normality [Gho13, AF12, BL10b, CFT10, Hu13, KF10, QM11, YKH13]. normalization [BO11]. normalized [JWY16]. note [ÁG12a, AT10b, BSH13, Ban12, Cab10, CF12b, CT10b, CR15, CHH19, DM10, DG13b, FGV12, Fre11b, GF11, HCMP12, LVY13, LS10b, LYC14, Lon12, Mad10, MS11, RV16, Sad11, SlG12, Tsa11, TCC12, WMB10, ZSH12, ZLZ^{+}16, ZLL16a]. Notes [Gra12a]. notice [BP13a, BP19]. novel [TvS15]. November [Ano10-40, Ano11-46, Ano12-44, Ano13-43, Ano16-40, Ano17-43, Gho11]. nugget [RSH12]. nuisance [DKY15, LZL^{+}10]. null [Cab10, CR10, Dic13, KL12, KMT10]. number [AG19, DPC11, DKN12, FLL17a, FJC12, LL11d, LC13, MZ13, SC11, SLC18, ySD10, WS14, WW15a, WM18b]. numbers [Dem12, Kra10]. NYSE [DHH10].

OBC [Ano10-28, Ano10-29, Ano10-30, Ano10-31, Ano10-32, Ano10-33, Ano10-34, Ano11-27, Ano11-28, Ano11-29, Ano11-30, Ano11-31, Ano11-32, Ano11-33, Ano11-34, Ano11-35, Ano11-36, Ano12y, Ano12z, Ano12-27, Ano12-28, Ano12-29, Ano12-30, Ano12-31, Ano12-32, Ano12-33, Ano12-34, Ano12-35, Ano12-36, Ano13y, Ano13z, Ano13-27, Ano13-28, Ano13-29, Ano13-30, Ano13-31, Ano13-32, Ano13-33, Ano13-34, Ano13-35, Ano13-36, Ano14w, Ano14x, Ano14y, Ano14z, Ano14-27, Ano14-28, Ano14-29, Ano14-30, Ano14-31, Ano14-32, Ano14-33, Ano15-33]. Objective [MHK10, RSH12, Ruk17, CY19, GM12, VM18, ZSL14]. objects [DM14b]. observation [FZS16, PBRT16, SSZ11, ZWM12]. observations [AG10, BPR16, Bha14, Bin14, BW10b, CA10, Dem12, Efr13, HYS12, JRM10, Kro19, OPP19, RD12, ZCX11, vdMC10]. observed [BDvdA15, Bro10, CH11b, CB19, DI10]. obtained [BS11a]. occupancy [EGS10a]. occurrence [Ste11a]. oceanography [BS10c]. octagon [BGR11]. October [Ano11-45, Ano12-43, Ano13-42, Ano14-34, Ano15-39, Ano16-47, Ano17-44]. odd [ySD10]. odds [FO13, LGG12]. off [FM19]. offspring [GP15b, SSZ16]. omitted [LG12]. omnibus [DF12b]. On-line [DZ18, TCE^{+}12]. One [CNL17, CL11b, Giu16, PMX12, Wan11, Akn15, BL15, BZ12, Bur11a, ETMJ11, HLB^{+}16, Kat11, Kli11, LZL16, NCC11, NN10, Par10b, Ros14, SBCK11, YLL17]. one- [SBCK11]. One-eighth- [PMX12]. one-parameter [BZ12]. One-sided [CL11b, HLB^{+}16, Kli11, NN10, Par10b]. one-sixteenth-fraction [PMX12]. one-way [Bur11a, LZL16]. Online [DMP18, GB19, HXZ12]. only [LZL^{+}10]. onto [DB19]. open [SHYH16]. Operator [GR14, Jar13, NS10c, Var15, ZWZ10]. Operator-valued [GR14]. operators [ZJ10]. Opial [GK12]. Opial-type [GK12]. Optimal [AGPS12, BCMAW14, BM15, BD14, BBS10, CBD11, CKP17, CXCK13, DMS11, DC10, DT14, EZ12, GDHS12, GGHS13, GT12b, GHM11, HRR10, HS14, HZY14, HC16, Hun12, HYF11, JM13, KSN18, KGV10, KBK15, KK19, KME10, KM11, LB11, LM19a, Lin14, LAk10, LYL13, MZWF17, MKT15, MP12b, MZP12, NS13, NIHY11, OQL11, PP13, RF13, SHW16, Soi13, ySDM13, TY10, TL11, WA16, WZZ19, XLJ11, YAW18, YQC11, vB19, vdVW14, Alb10b, AZ10, AFHZ14, BSH13, Bar17, BSKS18, BSvZ15, BMPR11, BS16, Bou11, CK13, CSR15, CL13a, CT14, CPS15, CKMS12, CWcIC11, CC18, CDST18, CY19, CKG19, CE11, DPA10, DL11, DH11a, DM11, DDS13, DL12, Dut11, FK19, FHK11, GIS19, GS19, GZ14a, GS11a, GS12a, GN16, GVV10, Gra12a, GM12, GZH11, HKN18, Har14, HHJ13, HJ14, Imh14, JGJR16]. optimal [LYF14, LC15, LL11d, LYC14, LYC15, LYXC16, Mor14, PP10, PK13, PBRT16, Pro10, PZ14, RS19, RRN14, RLZ12, Sag11, Sek12, SX11, SDV14, ySD10, SD16, TL18b, Wag11, WMB10, WY14, WCZ17, WZ10, WL14, Wie19b, WZ11b, WT14b, Jac19]. Optimality [AM19, HS11a, KS19a, CKM11, FR13, Gaf19, GF13, HY19, HLL12, Jon13b, KKKP19, SL19b, XT15]. Optimally [Bos10]. Optimization [GS19]. Optimum [AB14, DD13, GM19, CHF11, Men12, EMS11]. optional [GSS10]. options [AK10, GVV10, KK12]. Oracle [PKMK12]. ORDANOVA [GB12]. Order [LLL^{+}12, Now12, AG12b, BP10a, BDKM10, BAHS11, BBPW14, BLBL10, BO11, Bed10, BTT15, CMLBSM13, CZ10, CHF11, CL10b, DPA10, Dem12, EZS11, FK11a, FRK14, GBdW10, GK12, GSG12, HZ11b, IKU11, JB12, JR12, JGK11, Kas17, KA11, KFK11, Kim16, KHS12, KR12, LYXC16, Lwi11, Mac10, MV15, NR11, QL15, Sag11, SBCK11, SLW13, SF10, SDGK15, VGV11, WN10, WH11, WW11b, YHS19, ZC10, ZLZK13, ZLL16a, ZZ14b]. order-restricted [KHS12]. ordered [COS12, Ozt12, STS10]. Ordering [FPLR11, Zha10, DDL10, DL12, IK12, NIHY11, WM18a, ZB11]. orders [Bha11, DN13]. ordinal [CHR13, GB12, PMW^{+}16, TXLV11]. Ordinary [CB19]. Ornstein [Bar17, BSKS18, BWYZ11, Gri11, Lac14, LBL12, Men13]. Orthogonal [Geo11, Kre18, WTZ12, YL13, AHL12, AM12a, BS16, CLP13, CGLW18, DvB15, FMZ10, KV10, LT14a, LT14b, Lin14, MFW11, SDGK15, SLL10, Tsa17, YLL14, YLL17, fZyZD16, Giu16]. orthogonal-array [CGLW18]. orthogonality [JAI11]. other [Gri11]. Outcome [YLC^{+}16, CL15b, PMW^{+}16]. Outcome-dependent [YLC^{+}16]. outer [HLY12]. Outlier [YCY15, DS10, WS15, XJ13]. outliers [KV11, LCC^{+}17, SHYH16, WQ11]. outlying [RS15]. outlyingness [MS13b]. output [CO10, FM19]. outputs [Mor14]. over-dispersion [PA12]. over-replacement [AT11]. overall [Ros14]. overdispersed [TRWP10]. overlap [SKFMR14]. overlapping [GC11, MH10]. overview [Cue14, GZ14b, MA10, ZC10].

P [Llo10, BH10, ZWZ10]. p* [Llo10]. P-values [Llo10]. PAC [GR18]. PAC-Bayesian [GR18]. page [Fre14]. Pages [Ano10-43, Ano10-41, Ano10-35, Ano10-37, Ano10-38, Ano10-42, Ano10-36, Ano10-40, Ano10-39, Ano11-39, Ano11-43, Ano11-47, Ano11-48, Ano11-37, Ano11-41, Ano11-42, Ano11-38, Ano11-40, Ano11-46, Ano11-45, Ano11-44, Ano12-48, Ano12-41, Ano12-45, Ano12-46, Ano12-37, Ano12-39, Ano12-40, Ano12-47, Ano12-38, Ano12-44, Ano12-43, Ano12-42, Ano13-47, Ano13-40, Ano13-44, Ano13-45, Ano13-37, Ano13-38, Ano13-39, Ano13-46, Ano13-48, Ano13-43, Ano13-42, Ano13-41, Ano14-42, Ano14-35, Ano14-37, Ano14-41, Ano14-40, Ano14-39, Ano14-38, Ano14-36, Ano14-34, Ano15-35, Ano15-42, Ano15-41, Ano15-40, Ano15-34, Ano15-38, Ano15-37, Ano15-36, Ano15-39, Ano15-43, Ano16-43, Ano16-37, Ano16-46, Ano16-38, Ano16-39, Ano16-41, Ano16-42, Ano16-44, Ano16-45, Ano16-40, Ano16-47, Ano16-48, Ano17-45, Ano17-41, Ano17-38, Ano17-46, Ano17-40, Ano17-47, Ano17-42, Ano17-37, Ano17-39, Ano17-43, Ano17-44, Ano17-48, Ano18o]. Pages [Ano18n, Ano18p, Ano18q, Ano18m, Ano19n, Ano19k, Ano19j, Ano19i, Ano19m, Ano19l, Ano19h]. pair [YSH11]. paired [GPRY11, GSG12, OL19, VGH13, vB19]. pairs [TKB11]. pairwise [RSS19]. Pan [GP16, NL16]. Panel [Tor15, AMRP18, CHH13, GMR10, SXX11, ZS15]. panels [BDvdA15, Wes16]. paper [YH12, vdG13a, vdGM13]. paradigm [Mon17]. parallel [BMPR11, DDL10, Gup12, TL11, Tsa17, TL18b, ZB11]. parallel-flats [TL11, Tsa17, TL18b]. Parameter [AMR11a, BDLP14, dZBK10a, dZBK10b, CUW10, CA10, Gas11a, Lwi11, MK18, NB12b, SMM19, BP10a, BZ12, BK11b, BE14, CT10a, CS16a, CB19, Coa14, DKY15, EMS11, FVB11, FO13, Gam13, GMSK12, GBdW10, HKR11, HMW12, JAI11, LZL^{+}10, LL19, Lia11b, LLLP19, MN16, MS12c, NB12a, NMS12, Oga13, RRS11, RLLT17, RLZ12, SBCK11, SC13, SY18, VM18, WV12, XL11, ZL12]. parameter-free [WV12]. parameterization [KT17, SB19, ZZ10]. parameterizations [KCF12]. parameters [AB14, BT16, BP10a, Bar17, BK11b, CM11, CWH10, Che11a, CM16, DGS19, DD13, FLL17a, FRM11, Hoc19, HW13, HLY12, HBP12, IF10, KS12, KS19b, KL10a, Kub14, LW11b, MY10, MH10, NPK10, Nan12, Ozt11, RS16, Ruk11, Ruk19b, SC11, WS14, WW15a, WM18b, WW11b, Wu16, XZZ18, ZSL14]. Parametric [CP12, Men13, Ozt11, CG11, CH10, CG14, DKY17, EMS11, FZ13, Kro19, LMKS17, PZ12, RBS10, SR13a, Tso11, WTZ12, YSW10, YL10b, ZL11a, ZS15, dLJ10, dLJ12]. parametrized [Di 16]. Pareto [AF12, dZBK10a, dZBK10b, CSR15, Ho10, JM11a, PT13, RP11a, RTM10, Won12]. part [Bou11, dZBK10a, dZBK10b]. Partial [Bha11, CL13a, CWcIC11, CR15, Cue14, GD18, Gho14, HTZ14, HC11b, JB14, Jac11, KBK15, KS16, LHH16, NGXZ16, NdCJ10, SSL^{+}17, TL18b, WL16, WYH10, YW10]. Partially [ABW13, LW11a, OQC13, Bro10, BLW16, CW19, CYZL12, CB19, FLW13, FC12, GLW11, Hua11, LCL11, LZ12a, LM13, LQ17, LQ16, LLL18, MY11, MB11, Ni12, Ozt12, PMK19, SR10, TL11, WZW13, Wei12, YY14, YTHL13, ZW12a, ZL18]. Particle [dVGS11, BS10c]. particular [Kem10]. partition [Le13, MQ10]. Partitioned [WW11a]. partitions [EW11, SN10]. Parzen [Ano10s]. pass [PL11]. past [KNM10]. Path [Ano10a, Ano10-44, Böh10, Eri10, Hum10, SL19a, Tan13]. paths [BK13, DMR10b, DMR10a, FGN10, GRD16, MSTT11, NS10c, Qui10]. pathway [KPZ13]. Patrangenaru [Ber14]. Pattern [NS10c, KM12b, LYL11, QWC12]. Patterned [PMK19]. patterns [BB14, FJC12, IAH11, LZZ15, Ste11a, TvS15]. peak [XI12]. Peakedness [WM18a, ZCN11]. Pearson [XL11]. Penalized [Cop11, FLL17a, GMR10, GT10, JKL17, Ni12, PH11, Pro10, WWZ19, WMO15, vdGM15, AAV13, HZ13, LW12b, YY14, YCY15, ZGK13, PKMK12]. penalties [Fre11a, KK12]. penalty [JKL17, Kol13, Lec13, Par10a, vdG13c, vdG13b, BC13]. penetrances [ZZOY13]. percentile [DM14a, FPLR11, KHN16]. perception [BS11a]. Perfect [BGR11, Fan10, KBPW12, SS00, TLBJ11, VB11]. Performance [Luk14, AG12b, CC18, RH13, YK15, ZCX11]. periodic [Akn15, HdGT12, Maj14, MSP10, Ros10, UP14, Wis13]. periodically [DDL14, SS00, TLBJ11]. permutation [CR16a, OP12]. Permutational [TW11]. permutations [Fon11, GJM18]. permuted [ZW11]. perturbed [Kut11, MS11]. Peter [SS13a]. phase [BBB11, HJ14, LO11a, RRN14, SK10, SDV14, Tan10, TNWW12, WT14b, ZLM12]. phenomenon [Rab14a]. Philipp [SS13a]. photon [Di 16]. phylogenetic [RWB15]. Picard [vdG13a]. Pickands [MPN^{+}17]. piecewise [DLDC12]. Piet [Jon13a, Mam13, vdGM13]. pile [RRS11]. pile-up [RRS11]. Pitfalls [ALP10]. Pitman [Fou11, RA12, ZN12]. placebo [TP10]. Plackett [QL15, SM16]. Plackett-Burman [SM16]. plan [ETMJ11, SC10]. planar [ADPW10]. plane [BBPW14, LA14, Mon17]. Plann [AA11c, BP13a, BP19, BRB18, BK11c, GS12a, IT10, KZ15, TLBJ11]. Planning [Ani10, KS11b, LT10, YS11b, dLJ12, CP19, Pas10, XT15]. plans [CE11, JM10b, LCY12, LT14b, OQL11, WMB10]. plot [AM12a, EGS10b, TLM12, TKB11, WZZ19, ZC12]. plots [Aly19, TKB11]. plug [Reg11]. plug-in [Reg11]. plus [BS16, ETMJ11, NU19]. Point [FP12b, BH15, CF12a, CHH13, Ciu13, CG14, Dac10, Dör10, Dör11, ES16, FJK11, Hab13, KSN18, LM17a, Loh10, Sek12, SR14, TP10, TN14, YKH13, YLL17]. points [BM15, CHH18, KK15, MKT15, MGB10, SG11, SL19a]. Pointwise [BK17]. Poisson [BRB18, Lia11c, AX19, BSvZ15, BRB12, CUW10, CT11, GC11, Gra10, Gra12b, HS11b, KL10a, KP11, Kro19, Lau12, LJK12, NS13, SlG12, Vag10, VAM10, Wu17, Zhu12]. polar [DF12a]. Politis [GP16, NL16]. Pólya [Cha12, GP15a]. polygonal [ITA^{+}09, IT10]. polygons [CFT10]. Polynomial [CPL18, LP11a, SY12, CL13a, CT14, HLB^{+}16, Leb12, PL11, Sek12, Tsa11, WX15, WSSG19, YL13]. polynomials [BJQ15, BW10a, BEJS10, DG14, GJ10, KV10, LG12, MPN^{+}17]. polytopes [PT12]. population [AX19, BDKM10, BHS12, CLM12, CL11b, CMO13, DH11a, HOK19, KS19a, LCQ^{+}13, LSW13, MH10, NMS12, RS10, RA12, Ros14, SG12b, SK10, SSC11, SHYH16, Tak10]. populations [BFT15, Cha19b, GPRY11, GM19, JJ12, KK19, LYLZ12, MR11b, Nku10, OP12, Par10b, STS10, XLJ11, ZSM10]. portfolio [BMPT19]. portmanteau [Mai11]. Positive [DP13, DST10, DMP11, Kub14]. positivity [OK11]. possible [ZX19]. possibly [CB19, DPC11, Mül12]. post [Wu16]. post-change [Wu16]. Posterior [KK11, WM18b, Bas11, CM10, GSZ11, Lia11a, Lia14a, MHK10, SM17, VM18, Xin11]. posteriors [HW13, Lew18]. potential [EO12]. Power [Azr15, WW12a, BTKN10, CGC12, FLL17b, GP15b, HK10, Ing10, Kem10, Lem11, Lem12, MS11, MFABG13, MLa11, Que10, QM11, RP11b, SLC18, SZ12, SS13d, Wie19a, ZWZ10, dOCG12]. Powerful [MLP17, NN10, Ros14]. powers [SLD14, WN13]. pre [QZZ11, WLL^{+}17]. pre-specified [WLL^{+}17]. pre-test [QZZ11]. Precedence [NB10, BTKN10]. Precedence-type [NB10, BTKN10]. precipitation [BT16]. precise [CKG19]. precision [LL13, LW19, Lia11b, MS12c]. predator [Nku10]. predator-prey [Nku10]. Predicting [GRM14, GP15b]. Prediction [CR13, Fre12, God10, SLgX18, BL10a, CHH19, EZS11, FGV12, Fre13, Fus10, GP16, KHN16, Kim16, Kre16, KP11, KR12, LÁPMF19, LSWY16, LYL13, NQ10, NL16, PP16a, PP16b, SBCK11, VB10b, WHI12, WW15a, XLJ11, YQC11]. Predictive [BDG14, CMCSC12, CND10, CMCSC11, CXCK13, EZ12, GO11, Kom11, LMKS17, MCSC10, SLgX18, TN10, dVGS11]. predictor [CL15b, YO11]. predictors [KL11, ZN12]. Preface [GA10, Moh10, PD10]. preference [DDS13]. preliminary [KF12]. presence [BLP11, EA10, FRF12, GDHS12, Lwi11, NdCJ10, PBND11, Rab14a, Sah11, SA10, SSZ11, WL11, XI12, YU13, YH12]. present [BF12, DW17, LZL^{+}10]. Preservation [CZ12]. preserve [OK11]. Preserving [WK10, FK11a, Now12]. presmoothed [DKMdUÁ16]. prewhitening [RS13a]. prey [Nku10]. Pricing [KK12, AK10]. Principal [TP10, CCGL10, IYA16, MKT15, SS16]. principle [HFW10]. prior [AS10b, BZ12, BP14c, BE14, Ho10, MO10b, MS12c, SvdV15, VCR10, WW13, WM18b, YO11]. prior-free [BP14c]. priors [Bos10, BF13, BD14, CM10, DFM10, Has19, HG13, Kom11, MHK10, SP12b, XGK16, ZW12b]. Privacy [BHM16]. Probabilistic [Zhi10, BSV10, EO12, Mac13, PZ12, Yan17]. probabilities [CBD11, CM15, CHP11, Fre11c]. Probability [LYH12, CG11, CM10, Cha10a, DW17, FJC12, Grü18, Hab10, He11, HH14b, JL11, KF12, LQL12, LM16, MSA14, Men12, QLL11, SC10, Sch18, sS10a, SKS10, Unk17, VZH17, XL12, XW13b, ZW11]. probability-weighted [CG11]. probit [KS12, WG12]. problem [BD18, CPH19, CHT17, CNL17, FGN10, Hon10, JM11a, Kat19, LLM11, Ste11b, SB11, SLD14]. problems [BS10a, BHS12, CC11, DM11, DLZ14, EGS10a, KS19a, KS11a, LLM11, TN14]. procedure [CYZL12, Fre14, IKU11, JBB10, LMZ18, SL14b, TRWP10, TCE^{+}12, VS12, YSGN10, YTHL13, Giu16]. procedures [CMS13, GM19, Hab12, JRM10, Jon13b, KK19, PC14, SGF12, ZZ14a]. process [AS11a, BK17, BS11b, BC19, Buc14, CL13b, DM14b, FC10, Gas11b, GS10, GS11b, GRD16, GF18, Gri11, GP15b, HS14, HFW10, HBP12, JGK11, Kas17, Kon13, KKL15, LS10a, Loh10, Lwi11, Men13, MS12c, PP10, RM12, RGW10, SSZ16, SS13d, ST11, SSDM14, TG16, WSL17, YC11a, vdMC10]. processes [Alb10b, Alb11, Alb12, ACC18, AF12, AG19, Bar17, BSKS18, BP10c, BSvZ15, BWYZ11, BRB12, BRB18, BH10, CMT17, CUW10, CF12a, CF12b, Cha13, CH11b, CM15, CG14, Dac10, DFM10, DI10, DDL14, DKN12, FP12b, GIP17, GC11, God10, Gri11, HZ11a, Hu13, KSN18, KU14, KMW16, Lac14, LBL12, LWZ13, Lon17, Mac10, Maj14, MRW11, MSP10, Pal10, PV10, Rab14a, RW18, Reg11, Ros10, SZ19, TG16, TRL13, TTB11, Wu17, XZZ18, Yan15, ZS10, ZWZ10, dOCG12]. product [BTT15, DMP11, Dut11, Eri10, HLY12, JMW13]. product-multinomial [JMW13]. product-sum [DMP11]. production [jLP11]. products [SLD14]. profile [BMW16, He11, KGV10, VCR10, Yao10]. Profiled [CYZL12]. programming [BK10, BK11c, Qi11, Sag11]. progress [SZZ16]. Progressive [CYC11, Bur11b, CE11, MCSC10, Wan10a]. progressively [BTKN10, BAHS11, CL10b, Mad10, MRY11, PK13, RM11, SS13d]. Projection [LMW12, GJ12]. projections [DLPW14, DB19]. projective [BBPW14]. projectivity [TS10]. prolate [PL11]. prone [Hua11]. Proof [FGN10]. Propensity [Wae10, YSW10]. Proper [Tsu10, ZSA10, LT18, SYA11b]. Properties [CCGL10, AA11a, Ani12, BK11a, Bro10, CCR11, DJ19, DS10, Dem12, DKL13, FYX14, FPLR11, FVB11, HZ11a, HY10b, IYA16, IR19, JM10a, KHN16, KPZ13, Kub14, Leb12, Lem12, Lia14b, LWZ13, MYJ11, MM12, Ogi19, PKMK12, QRGLDMA^{+}12, SM16, SdL10, Tia13, VB10a, WC10, YL10b, ZGW^{+}17, ZWM12]. property [AS10b, Now12, Wan10b]. proportion [CR10, Gua12, KL12, Wan10b]. Proportional [HLS12, Bha11, DY19, FN17, LHH16, ND11, Woo19, YLC^{+}16]. proportions [BTT15, PS14, Sah11, TQ12]. protection [CWcIC11]. provide [Kha10]. providers [RS15]. Pseudo [DRT10, MA17, Nil11, OPP19, CH11b, HW13, WL11]. Pseudo-Bayesian [MA17]. Pseudo-empirical [DRT10]. pseudo-likelihood [WL11]. pseudo-maximum [CH11b]. Pseudo-observations [OPP19]. pseudo-true [HW13]. Pseudo-Youden [Nil11]. pseudodistances [BTV12]. pure [Kon13].

q [BH10]. QML [HZ11a]. QTL [KCZ13]. quadrangle [BGR11]. Quadratic [Ruk19b, BU10, DH11a, DG13a, HS11a, HC16, Imh14, vB19]. qualitative [FRR16, HLSL17, Kat12, SSZ16]. quality [Alb10b, Alb11, Alb12, MYW11, TCE^{+}12]. quantification [ZW11]. Quantifying [DKY15, GRCGL12]. Quantile [Aly19, CZ11b, DZZ18, KWS19, OZC16, Ozt12, BCZ19, CPS15, CHT17, DPT16, FZ13, GMR10, GGS12, HL13, Ho10, HM10, HXZ12, Koe13, KL13, LMZ18, LdUÁ11, MLP17, NB12b, SWH13a, SWH13b, SGL13, SSL^{+}17, TWLQ15, WZ11a, WH11, WYY^{+}10, Xia13, Xu17, YSGN10, ZLZL16, ZLL17, ZGK13, ZWY11]. Quantile-slicing [KWS19]. quantiles [BDKM10, Día17, DC10, EGGG12, HK19, HCT16, LQ11, LPY10, MR11b, PH17, RA12, WMM^{+}10, WHY11, YHWL12, YYZ14]. Quantitative [Bas11, Cha19b, HLL12, Kat12, OL19, Rab14a]. quantization [AGPS12, CPS15]. quantum [MA17]. quarter [LQ17]. Quasi [CH11b, GS19, NS13, PC13, SHW16, VCR10, XZ19]. Quasi- [CH11b, VCR10]. quasi-likelihood [NS13, SHW16, XZ19]. quasi-maximum [PC13]. Quasi-Newton [GS19]. quasi-profile [VCR10]. quaternary [PMX12]. quaternion [LB15]. questionnaire [CS11a]. queueing [DEF11, KM10]. queues [Böh10]. quintile [LT11].

R [Ber14]. Random [GZ10, LL13, MQ10, PA14, PV10, AS11a, AJM19, AMR11a, Alo11, Aly19, AX19, AT11, AM12b, BP10b, BP13a, BP19, BMW15, BJQ15, BPR16, BSG16, BH10, BCG13, Bur11a, CLM12, CZ12, CFT10, CC15, CSC14, CS12a, DH11a, DGCS16, DLDC12, Dem12, DW17, DM14b, Di 16, DMP18, DM13, EGS10a, Efr11, FYX14, FP12b, GD18, GRD16, GJM18, GRCGL12, GDHS12, GX12, HFW10, HC16, HBP12, JBB10, JR12, KCF12, KKK17, KF10, KA11, KK11, KX11a, KZ16, KL11, LW11a, Li12, LW12a, LZ13b, LCY12, LLV15, LLX12, LZL16, LW15, Moj12, MR10, MS18, NDD16, NS13, OG10, Pan10, Per10, RP11b, Ruk11, Ruk19a, SB19, SLD14, Sum17, TB11, Ton13, WC10, WSHW10, WT10, WM18a, WG13, YL10a, YW10, ZS15, ZCN11, vdMC10]. Random-effect [LL13]. random-effects [CLM12, ZS15]. random-sum [TB11]. randomisation [Kri10]. randomisation-based [Kri10]. randomised [Kri10]. randomization [LW12b, LBH17, WX19, ZW11]. Randomized [AS10a, Cha19b, Dic13, JJ12, Chr10, GSS10, HLS10, HLH13, LJK12, WAL17]. randomly [DF12a]. randomness [GKM10]. range [BKPS14, FRM11, Wis13]. Rank [LYLZ12, VGV11, Bin14, LP11a, LD12, MA17, vMMA14, Ozt12, Sch10b, SBAC19, Ver17, Wu19, ZLM17]. rank-adaptation [MA17]. Rank-based [LYLZ12, Ver17]. Rank-order [VGV11]. rank-ordered [Ozt12]. ranked [CL11a, DC10, Fre11b, Fre11c, KS93, KS11b, Ozt11, VB11, WAL17]. ranked-set [Fre11b, Fre11c]. ranking [DLZ14, GR18, JB14, VB11]. ranks [TSO10]. Raphson [Cla19]. rapid [MGB10]. rare [FC10, ZZOY13]. RARMAX [FC10]. Rasch [SR13a]. Rasch-type [SR13a]. Rate [BSvZ15, DM11, FLTV10, BK11b, CLB12, CCR17, DDL10, DGG19, DL12, DMP18, FJK11, JWY16, JM10a, Kas17, KLS11, Lau12, LS10b, ND11, SGF12, SP12b]. Rate-optimal [BSvZ15, DM11]. Rates [Aya12, LL11a, AK10, AP14b, BS14, CL10a, CLP13, DZ18, EEL10, Lau12, Lia14a, Lia15, MKW18, PG12, RP11a, Tor12, Wu19, Xin11]. ratio [BK11b, CWcIC11, sCK11, Dac10, FO13, JJY12, Kha10, Kub14, LT11, Lau12, LF11, LBH17, LGG12, VSK^{+}11, VGH13, ZSM10]. ratios [LNdDLdCH11, VDW13]. Rayleigh [RM11]. RCBDs [WM12]. Re [SLW13, ZZ10]. Re-parameterization [ZZ10]. Re-weighted [SLW13]. reaction [AHKS12]. real [Mab16]. realisation [God10]. realized [DHH10, Kha10]. reassessment [CL15a]. recapture [RAB14c, SHYH16]. reciprocal [Kak18, LS10a]. reconfirm [YO11]. record [CN12, GR11, RA12]. records [AA11a, CMLBSM13]. recovery [CCAGV14, SL19a]. recruitment [ZL11b]. rectangular [LYL13]. Recurrence [BAHS11, ZWM12]. recurrences [DS12]. recurrent [APQ10, AS15, GC19, KMT10, ZZW12]. Recursive [Fra14, Fre11c, GT12a]. Reduced [Sch10b, BMNY13, HFQ11, MV11b, WN11a, ZLM17]. Reduced-rank [Sch10b, ZLM17]. reduces [Sag11]. Reducing [LGG12]. reduction [ÁG12a, CPSY10, DDW16, DGG19, DW17, KWS19, LPY10, Lia15, XWY18, YY10, Zha13]. redundancies [DL12]. redundant [BMPR11]. Reference [SP12b, XT15, ZW12b, BE14, HG11, Kem10]. refinements [DKY15]. reflected [BWYZ11, LBL12]. regime [CHH18]. region [BT16, LYL13]. regional [PH17]. regions [ADPW10, Beh10, CM10, Cla19, Dem12, FGV12, HX16, WW11b]. registration [CD15]. Regression [CH19, Gup12, LG12, SSZ11, XJ13, AKK12, Abr12, AHD12, ÁY12, AS10b, AFHZ14, Aya12, AGPS12, BCN13, BW13, BM12a, BFCGR12, BCZ19, BRBM19, BRvdGZ13a, BD14, COP10, CLB12, CCAGV14, CZ11b, CL13a, CT14, CLO17, CLL12, CL10a, CLL11, CGL11, CLP13, CPM14, CHT17, CHR13, DMTZ15, DMP10, DRT10, DH11a, DM10, DMS11, DPT16, DF10, DF12b, DD13, Efr11, ES12, FZ13, mFS12, FKMR19, FK11b, Gaf19, GS19, GMR10, GZ14a, GK17, GLR10, GP19, GMMC11, GDHS12, HRR10, HKN18, HL13, HLB^{+}16, HTZ14, HY19, HM10, Hon10, HH14a, HG11, HM11b, HLS10, HP10c, HXZ12, HC16, HBP12, HCZ19, Imh14, JKL17, JVV13, KF12, KF10, KW14, KMT10, Kat19, Kau14, KHN16, KS12, KPZ13, KK16, KWS19, Koe13, KMW15, KS16, Kro19, KL11, KMW16, KL13, LGLO12, LCT10]. regression [LL11a, LF11, Lem12, Li12, LH14, LM19a, LL19, Lia14a, LM17a, LSP12a, LLZ11a, LSWY16, LLV15, LAk10, Liu11b, LYC15, LL15, LLW16, LSP12b, Luk14, LC13, LGG12, MY10, MX15, MLP17, MZWF17, MKW18, Moj12, MV11b, Mon17, MQ10, Mül12, MSW12, MS18, NR11, Ni12, NS13, Nik13, NGXZ16, OZC16, Par10a, PY11, PKL12, PA12, PL11, RVF10, RM12, RS13b, RAB14c, Sch10a, Sei14, Sek12, SS13a, sS12b, SDNH13, SLZ10, Son11b, SWH13a, SWH13b, SZ10, SSL^{+}17, SlG12, TWLQ15, TMK10, Tia13, Tsa11, WC10, WY10, WZ11a, WSL17, WWZ19, WZ10, WL14, WK13, WN11b, WG12, WW12b, WMO15, Xia13, XW13a, YY14, YW16, YY10, ZW12a, ZLZL16, ZL17, ZX19, ZLL17, ZLM17, ZG12, ZGK13]. regression-based [LGG12]. regressions [DZZ18, ZWY11]. regressors [HP10b, ZL18]. regrowth [LB11]. regular [AHO13, CC19, Has19, HZ11b, Jac11, RLZ12, Seb19, SD10, ZLZK13]. regularisation [JVV13]. Regularization [XXWL12, TMK10, ZCX11].

Regularized [CPM14, CNL17, mFS12, BMW16, CC12a, CL10a, Kat19, LL19]. regulating [BM18]. Rejoinder [BP14a, BRvdGZ13b, Gro13b, PP16b, SWH13b, WT14a, MPT13b, vdG13c]. related [DAP10, HS18, HDC14, KS11a, QWC12]. relating [Shu11]. relation [PR16]. Relations [GK12, BAHS11]. relationship [DKY17, Tso11]. relative [LLW16, ZLM12]. relevant [CMO13, LC13, YU13]. Reliability [FVB11, SAE11, HMW12, KL10b, TV12]. relying [LSP12b]. Remainder [KLK11, Fro14]. Remarks [Lon17, RS13b]. renewal [BC19, DKN12]. Rényi [AA11a]. repair [Gas11a, SS00, TLBJ11, TLBJ13]. repair/replacement [TLBJ13]. repairable [MZP12, dOCG12]. reparameterization [ZG12]. repeated [DMTZ15, MSBM16]. replacement [AT11, TLBJ13]. replicated [LQ17, OQC13, TL11]. replicates [GG15]. replication [RBS10, SM16, TL18b]. replications [GC17]. Reply [Wal13b]. reporting [GD15]. representation [DKMdUÁ16, KLKO15, LL11b, WHY11, WTZ12, WYY^{+}10]. representations [Aok19]. representativeness [SSS12]. reproducing [CC18]. Resampling [BL10a, AM19, Kim10]. Resampling-based [BL10a]. residual [CZ12, DN13, FPLR11, Gup12, KA11, MPRG10, NV10, NdAA10, RQ15, Sad11, SZ10, XL11, XSJ14, YW10]. residuals [Beh12]. resolution [LM14, PMX12]. resolvable [GG15]. respect [Bha11, BD14, DL12]. respondent [PL14]. respondents [SK10]. Response [ZH12, AG10, AS10a, AG12b, BBB11, Cha19b, CZHC14, Chr10, CH19, GN16, GSS10, GZH11, HLS10, HLH13, HH14c, HH14b, HYF11, Li12, LW12b, LM19a, LLP14, LYXC16, LW15, Mül12, NGXZ16, RF13, SSS12, SA10, SSC11, WSHW10, YCB16, YH12]. response-adaptive [BBB11, CZHC14, HLH13, LW12b]. responses [BBB11, Efr11, LW11a, LLX12, MS18, QL10, TP10, WQ10, WL11, WFG^{+}15, WG12, XXC11]. restricted [Gam13, KS11a, KHS12, Li11, LM17a, Ruk12, SC16, Wei12]. restrictions [SN12]. result [Bao18]. resulting [RH13]. results [CHH13, GMP15, HLCW12, HZ11b, KM12b, KL11, KNM10, LW10, MBB11, NdAA10, TS13]. RETRACTED [BP10b]. Retraction [BP13a, BP19]. Retrospective [FGH14]. reversed [ND11]. reversible [Lon17]. review [Geo14, NL14, PA14]. revisited [GGG11]. richness [BFT15]. ridge [KS12]. Riesz [DG13b]. right [BM11, CC11, DFM10, FJC12, LZ12a, Say12, sS12b, VB10b, YYZ14]. right-censored [CC11]. right-hand [FJC12]. Rigorous [CC15]. Riordan [Nkw10]. risk [BI12, Cha13, Kim10, NQ10, RV16, WV12, ZN13]. risks [HLS12, Pas10, sScY10, sS11, ZN13]. RKHS [KSKE10]. Robust [BGS10, CHR13, DB11, FT10, GLR10, GP12, HM11b, HH14b, JD16, KKK17, KW14, KLW10, KV11, UP14, VY14, WL16, XZ19, ZL17, ZLM17, Bos10, CZ11a, DPA10, DL11, HRR10, Han14, Kat19, LYC15, MS13a, MS13b, RLZ12, RVF10, ST19, Tsa11, Tso11, YCY15, GC17]. Robust-efficient [DB11]. Robustness [AG10, CCR11, CGLW18, BSH13, DS10, DM12, GBP15, GW11, GG15, KS19a, SSZ16, WS15, XR11, LM11a]. ROC [CWcIC11, Chr14, DN12, FRF12, HC11b, LZ12b, TDW10, TB11, WSSG19]. root [BDvdA15, CK13, Coa14, CS12b]. roots [BD12]. rotatable [DPA10, YHS19]. rotation [SF10]. rounded [Nis11]. rounds [Kat11]. Row [Qu11, BL15, BM18, QOK10, SE19]. Row-column [Qu11, BL15, BM18, QOK10, SE19]. rows [Kri10]. Rubin [ZX12]. Ruin [CM15]. rules [HP10a, Hoc19, JL11, Mac13, MV15]. Run [QL15, BS16, BH15, SLL10, WMB10, XM10]. runs [AZ13, BC11a, ETMJ11, GMP15, GSKP10, GCDK12, HM13, LQ17, MZ13, ySD10]. Rütimann [RS13b, SS13a].

s [JMW13]. s2 [ySD10]. Saddlepoint [AE16, YO11]. Safe [Grü18]. SAGE [KBPW12]. same [God10]. same-realisation [God10]. sample [AE16, ALT10, BI12, BWW19, BLT16, BD18, BC19, CPH19, Che11a, CC11, CNL17, CR16a, CS16b, CL13b, Fis12, Fre12, Gam13, GIP17, Hab13, HJ16, HG11, HY10b, IYA16, IR19, JSV12, KZ16, LLL^{+}12, LCQ^{+}13, LB15, LSWY16, LQ16, MYJ11, MR11a, MS13b, NYA17, OS10, Pal10, Par10b, PKL12, PR11, RV10, RL17, SB11, TQ12, VB10b, WHY11, YK15, YHWL12]. sampled [BC19, KU14]. samples [BTKN10, BC11b, DC10, HN12, IB11, JM11b, KX11b, LQ11, LQL12, MP19, Ozt12, PA13, PR11, QLL11, SBCK11, Shu11, SKS10]. Sampling [LA14, AT11, BSV10, CBD11, CL11a, DD12, FSS18, Fre11b, Fre11c, GLR11, Göb12, Gra10, Gra12b, GT10, HS11a, HP19, JMW13, JM10b, JJ12, KLK11, KBPW12, KS93, KS11b, LO11a, LCY12, LM19b, LZH11, Men12, MS10, NBSB13, Ozt11, PV10, RS10, RQ15, SG12b, sScY10, SK10, SS19, SS13d, Tan10, Tan13, TN10, VS12, VB11, WW15b, WAL17, YKH13, YLC^{+}16, YC12]. sandwich [HW13]. Sara [Kol13, Lec13, SS13a, BC13]. Sarhan [FKV11, KG10b]. Sarmanov [LH11]. satisfying [Ton13]. saturated [CLFZ13, FRR14, QOK10]. Saunders [BGK^{+}11, COP10, CLB12, LF11, Pat12, PK13]. SB [LM11a]. SB-robustness [LM11a]. SCAD [Par10a, PKMK12]. SCAD-penalized [PKMK12]. scalable [ZZW19]. scalar [ÁG12a]. scale [AVML11, BI12, BK11b, CNL17, KFK11, KX11b, LGLO12, LW11b, LLLP19, Mad10, Ozt11, PH13, Pat12, PMW^{+}16, VdUÁMM11, ZN12]. scale-mixture [Pat12]. scale-space [PH13]. scales [FLL17b]. scalp [CP12]. scan [AP14a, Wu17]. scenarios [HG11]. scheme [LO11a]. schemes [GH10, HHMO15, PK13, XM10]. Schwarz [CS16a]. science [Sch18]. sciences [dLJ10, dLJ12]. score [Gua12, HZ13, Wae10, YSW10]. scoring [CC18, JL11, Mac13, MV15]. Screening [ALN10, FWS18, FRK14, HHMO15, LMZ18, SL14a, Xu17, YYX19]. search [BT19a, GM12, LMT14, Pan10, YC12]. Second [BLBL10, BP10a, BTT15, BMNY13, CHF11, DPA10, FRK14, GBdW10, JGK11, Kas17, LYXC16, Sag11, SLW13, SF10, SDGK15]. second-order [CHF11, FRK14, JGK11, LYXC16, Sag11, SLW13, SDGK15]. section [JDB13]. sectionally [BDvdA15]. seemingly [MY10]. segmentation [Tor15]. segmented [KK16, YS11b, YS11c, ZX19]. seismological [FJK11]. select [CK13, Ruk19a]. selected [HZY14, Lia11c, LSW13, NMS12, SSC11, VAM10, WN10]. Selecting [Che10a, MS12c, Tak10, KK19, Say12]. Selection [He11, MF10, CC11, CMCSC11, DDW16, DPC11, DLZ14, FC12, GLT16, GM19, HH14a, HZ13, HYS12, Jan15, KKKP19, KK16, Kro19, LBR12, Lia11b, Lin11, jLP11, LM14, LT13, LLW16, MHK10, NIHY11, PR16, PY11, PT12, RLZ12, SG12a, SC11, VCA10, WMB10, WW15a, WL16, WM18b, XXWL12, ZK17, ZLZL16]. selective [Rab14a, Rab14b]. selector [FN17, AH12]. Self [Nik12, PC13, DT14, LLJK13, sS12c, WT10]. Self-avoiding [Nik12]. self-consistency [WT10]. self-consistent [sS12c]. Self-weighted [PC13, LLJK13]. Semi [CG11, LZ12b, WL11, FZ13, RBS10, Soi13, SM13, TvS15, ZL11a, ZS15]. semi-distance [TvS15]. Semi-empirical [LZ12b, WL11]. semi-Latin [Soi13]. semi-Markov [SM13]. Semi-parametric [CG11, FZ13, RBS10, ZL11a, ZS15]. semidefinite [Qi11]. semifoldover [OQC13, YL14]. semimartingales [LWZ13]. Semiparametric [DGG19, KG10a, LLX12, MX15, sScY10, WZ18, ZW12a, ZZOY13, ZZW12, AKK12, AS15, BLW16, DKMdUÁ16, FLL17a, FO13, GZ14b, HRR10, KPZ13, LW10, MSW12, PHW15, Qin10, QRGLDMA^{+}12, SL14b, WW15b]. Semiparametrically [KS19b]. semiquadratic [JDB13]. semistability [CDd11]. semistable [HdGT12]. Sensitivity [YKH13, FRK14, GSS10, HJ14, Lew18, QZZ11, RRN14, SDV14, SSDM14, TG16, WTZ12, WT14b]. separability [DP13]. separable [LR11]. September [Ano10-39, Ano11-44, Ano12-42, Ano13-41, Ano14-35, Ano15-43, Ano16-48, Ano17-48, Ano19h]. sequence [AZ13, BC19, CNdGAL11, mFS12, FJC12, Gho13, TCC12, WHY11, WG13, YHWL12]. sequences [FJK11, HdGT12, IAH11, MZ10, MZ13, PL11, RA12, RQ15]. Sequential [BS14, CWcIC11, HW17, IF10, LBL12, LJK12, SX11, ST11, YCB16, YC12, Bed10, BE14, CS13, Ciu13, DB12, Fre14, HC13, Kim10, LM16, LT10, LMW10, MB11, NN10, SCHG19, SBCK11, Wu16, YKH13]. sequentially [Coa14]. serial [MR10]. series [ABMP15, BL10a, BT19b, BMPR11, BP13b, BPP10, BD12, Che11b, CR13, CSN19, CS12b, CFW14, ETMJ11, FJKW13, FGH14, GP15b, Hud13, KMT10, Kem10, KLS11, KK15, Kre18, LSZ13, LZ13a, LT13, NL14, SMM19, SY12, SYA11b, Sør19, TSZ13, UP14, VN10, WF11, Wei11, WW12a, ZSH12, ZWZ10, ZJ10]. server [DEF11]. set [CL11a, DC10, DD13, Fre11b, Fre11c, KS93, KS11b, LM19a, Ozt11, Ozt12, VB11, WAL17]. set-identified [LM19a]. set-up [DD13]. sets [BHM16, BFCGR12, Bos10, BF12, DM14b, Fan10, KGV10, PZ14, SvdV15, SvdVvZ15, Thu14, WW11a, YLL17]. setting [Wis13]. settings [NYA17]. several [BTT15, VN10]. Shannon [BE14, Reg11, SP12a]. shape [ADPW10, FVB11, Hil11, LD14, LM17a, TSO10]. shape-restricted [LM17a]. shapes [DLPW14, ZPS15]. share [LT11]. shared [BFT15, Unk17, YC12]. sharing [CS16a, DDNN10, LL11b, LZL12]. sharp [TvS15]. sheets [Bar17]. shift [HFW10, YK15]. shifted [Bar17]. shifting [Ing10]. shifts [CHH18]. shock [CF12a, CF12b]. shorth [EGS10b]. Shrinkage [Lia11b, PT12, WW12b, AHD12, BP14c, Jan15, Kub14, Oga17, XJ13, ZH12, BOvSS10]. shuffles [TS13]. sided [CM10, CL11b, HLB^{+}16, Kli11, NN10, Par10b]. sieve [CR13]. sigmoid [WY14]. sign [DC10, YH19]. signal [GP12, Kha10, Nan12, SL19a]. signals [BLP11, BOvSS10, NPK10]. Signed [LF11, Bin14, Coa14, ZWZ10]. signed-rank [Bin14]. Significance [LQ16, SLC15, FBHZ14, LC16]. signs [Hoc19, TSO10]. similarity [Alb10a]. Simple [AT11, Beh12, Beh10, BP14c, CY10, CYW10, CS12b, GBP15, IAH11, LAk10, NN10, PP10, RR12, Wes16, Wu19, ZC11]. Simplex [Müh10, FHP11]. simplicial [KMW16]. Simplified [KMW16, ZW11]. simulated [BT16, Nik13]. Simulation [LLZ11a]. Simulation-based [LLZ11a]. simulations [JGK11]. Simultaneous [BDKM10, CLO17, DGS19, Lia11c, LLP14, LLZS17, VAM10, YS11b, YS11c, BT19a, GLZ14, HLB^{+}16, LAk10, RS16]. simultaneously [KV11, Ros14]. single [CW19, DEF11, Hua11, LW11a, Lin11, LX18, MH10, MFW11, OL19, PH11, RLZ12, SBAC19, WSHW10, WW15a, YY14, YTHL13, ZLZL16, ZCLZ17, ZLL17, ZGW^{+}17]. single-index [CW19, Hua11, LW11a, Lin11, LX18, SBAC19, WW15a, YY14, YTHL13, ZLZL16, ZCLZ17, ZLL17]. singular [BMPT19, KDJ13, KG10b, LL13, VN10]. singularity [AS11a, Ste11b]. sinusoidal [NPK10, Nan12]. SIR [PL14]. situations [ZXD12]. six [AA11b, AA11c]. sixteenth [PMX12]. Sixth [Ano10-44]. size [AM19, AA11b, AA11c, Alo11, AX19, CY10, CYW10, CLL12, Che11a, CHP11, Fis12, Gam13, HC16, HY10b, ITA^{+}09, IT10, ITA^{+}12, IYA16, LQ16, NYA17, RV10, RP11b, TQ12, TL13, XGK16, ZC11]. size-biased [CLL12, RV10]. sizes [AM12b, KZ16, LC15, SYA11a, SLL10, ZLZ^{+}16]. Skew [DMR10b, DMR10a, ÁG12a, AAV13, CLM12, CPM14, CSC14, Fre11a, JAI11, LGLO12, RA10]. skew- [AAV13, CPM14, CSC14]. skew-normal [AAV13, CLM12, LGLO12]. skew-symmetric [Fre11a]. skewed [BS10a, CL11b, GRS^{+}12]. Skewness [SCR10, OS10, WN11a]. SLEX [BOvSS10]. SLEX-shrinkage [BOvSS10]. Sliced [LJA15, YY10, YLL14, AGPS12]. slicing [KWS19]. slope [DPA10, Gho14]. slope-rotatable [DPA10]. Small [FT10, SM11, BMPT19, DRT10, DLZ14, Hir19, PG12, PKL12, PA13, TN10]. small- [DLZ14]. Small-area [SM11]. smooth [AC11, BCMAW14, CGCL10, CP12, GPRY11, JZ13, KLS11, LVK10, LM17b, WH11, WZ10]. Smoothed [YYZ14, Gro13a, Jon13a, Mam13, TDW10, vdGM13]. smoother [Aya12]. Smoothing [SHQG10, BSvZ15, CH11a, Dal11, EEL10, FWS18, HG12, HTZ14, LC11, MY11, SS12a, THG12]. smoothness [vdGM15]. Snedecor [ALS12]. Sobol [FKLL13]. social [dLJ10, dLJ12]. Solis [dC12]. solutions [DS12]. Some [AHL12, BTKN10, DN13, HLCW12, HZ11b, JM10a, KL11, KNM10, Lia14b, LW10, MB11, NdAA10, RD12, SYA11b, TS13, AJGG10, AP14a, AM12a, Ani12, BGK^{+}11, BSG16, CCS14, Dac10, DPM13, DS12, GGG11, GR15, HZY14, IK15, IAH11, LRB11, Lem11, Lem12, Nku10, Sar14, SE19, Ver17]. space [BHM16, CA10, CM16, DPM13, DP13, GS19, LÁPMF19, LJA15, Müh10, PH13, SYL13, TB12, TN14]. space-filling [BHM16, LJA15, SYL13, TB12]. space-time [DPM13, DP13, LÁPMF19]. spaces [AP14b, CC18, LC13]. spacings [PR11, TJ12]. Sparse [CCAGV14, KLKO15, SS16, SLZ10, TSZ13, BW13, BRvdGZ13a, CT10c, GP12, HY10a, Kat19, KSL11, LLZ11a, RS13b, SS13a, WWZ19]. Spatial [FCC10, HM11a, TG16, BGT16, BET17, BH15, CW17, CG14, DT12, EZS11, EZ12, GLR10, HBP12, Kre18, Kri10, Lin10, Lin12, MS12a, MS13b, Mon17, MR10, PHW15, RSH12, TSO10, Tor12, VBLN17, ZS15]. Spatially [Gra12b, CP12]. spatio [Maj14]. spatio-temporal [Maj14]. spatiotemporal [FRM11]. Special [Det14, GKvdVvZ15]. species [BFT15, RQ15, YC12]. specific [sS11]. Specification [MP11, Hay16, TTB11, Wae10, ZK17, ZZW19]. specified [KW14, LT14a, WLL^{+}17]. Spectral [Ros10, SB19, FRM11, GR14, HS14, LCQ^{+}13]. spectral-wavelet [FRM11]. spectrum [SS13d, VN10]. spheres [AJM19]. spheroidal [PL11]. spiked [IYA19]. Spline [AS11a, CWWT12, MY11, WW15a, CH11a, DM10, EEL10, Fre11a, JKL17, MZWF17, SY12, SHQG10, WX15]. Spline-backfitted [MY11]. splines [LM17a, MKW18]. Split [TKB11, AM12a, CS11a, TLM12, WZZ19, ZC12]. split-block [AM12a]. Split-plot [TKB11, AM12a, TLM12, WZZ19, ZC12]. Spokoiny [HL13, KL13, Xia13, YDZA13]. spring [FHK11, Gra12a]. square [ALLL13, Dhi18, LW12a, MYW11, OQ10]. squared [CS12a, Hir19, Lu13, LGG12, MKT15, YY10]. squares [DvB15, DB19, EW11, mFS12, FJKW13, FMZ10, FK11b, Han11, HLY12, KK12, Ni12, PKL12, PH11, Soi13, XZZ18, ZG12, ZWY11, vdGM15]. Srivastava [Gho11, CKG19, MGD14]. stable [RP11a, TRL13, ZS10]. Stacy [RM12]. stage [BTT15, DLZ14, GM19, IKU11, LYF14, LD12, LSW13, WLL^{+}17, YW16]. stages [MH16]. standard [AF12, CNdGAL11, Jac19, KS11b, PG12, Tak10, VBLN17]. standby [MZP12]. STAR [BGS10]. state [CYC11, CB19, CA10, CM16, FJC12, MZP12]. stated [BBS10, DDS13]. static [FJKW13, KW14]. stationary [ACC18, AG19, BOvSS10, BD12, CK13, CW10, FJKW13, HBP12, KZW10, LL11a, LLV15, Pal10, PV10, RW18, Reg11]. stationaryprocesses [HWZ11]. Statist [AA11c, BP13a, BP19, BRB18, BK11c, GS12a, IT10, KZ15, TLBJ11]. statistic [Ani10, BC11a, CND10, sCK11, Jar10, LJ12, MA08, TB11, WH11]. Statistical [Ani10, Ano10-27, CC18, DP11, Dhi18, FLW13, GIP17, Goo19, HLB^{+}16, HTZ14, Hua11, KPZ13, KS11b, LCC^{+}17, LT11, LS10a, LCL11, SLC18, Wei12, YL10b, YS11b, ZSM10, ZGW^{+}17, dLJ12, BTV12, Dac10, GBP15, HFW10, Kat19, LBR12, LC16, Lwi11, NQ10, Rab14b, Wah10, ZG12].

Statistics [Ber14, BP14b, Huc14, TJ12, AP14a, BDKM10, BAHS11, BO11, Bed10, BL10b, BK16, CR16a, CL10b, Cue14, DI10, Dem12, DKY15, Di 16, Dör10, FK11a, GBdW10, GK12, Hab13, Hab15, IR19, JB12, JR12, JWY16, KFK11, KU14, KZ16, KR12, LP11a, NB12b, PA14, Sar14, SBCK11, SZZ16, SB11, SR13b, Tor12, WN10, Wu17, Wu19]. status [OZC16, Pla13, Unk17]. Stein [AT10b, Dal11, LVY13, SZZ16]. Stein-type [AT10b]. Step [DB12, Giu16, MFW11, TSO10, DELS16, GT12b, LP12, MF10, Shu11, Wan10a]. step-down [DB12]. step-stress [GT12b, LP12, Wan10a]. Step-up [DB12, MFW11]. Stephen [De 13, HW13]. steps [YC11a]. stepwise [GR15, LT10, Tan13]. stick [MRW11]. stick-breaking [MRW11]. Stochastic [AVML11, BS10c, Cha13, KFK11, MM12, AK10, BP15, Buc14, DN13, DGCS16, DM14b, GIP17, GB19, HP10a, Hil11, KZ14, KZ15, MPRG10, Que10, SG12a, ST11, SdL10]. stochastically [STS10]. stocks [DHH10]. story [Pan10]. strata [CL11a]. Strategies [Mei13, Dio14, Hun12, MY13, MPT13a, MPT13b, Obo13, Yua13, vdG13a]. stratification [VS12]. stratified [CT10a, Har14, MS10, RV10, TQ12]. stratum [CP19]. strength [BB14, CSWY11, HMW12, KL10b, SGS12]. strength-3 [SGS12]. stress [GT12b, HMW12, KL10b, LP12, LT10, Wan10a]. stress-strength [KL10b]. strict [DMP11]. strings [MP12a, MSTT11]. Strong [Le13, Akn15, HP10a, KLS11, Kre18, LL11a, YHWL12]. strongly [IYA19, Tsa19, WHY11]. Structural [BD12, Hud13, CGL11, KZW10, LLLP19, Mai11, WF11]. structure [DG19, FK11a, FMZ10, LÁPMF19, LR11, MYW11, MSP10, SLC15, WL16]. structures [CCS14, DKL13, LP11a, LC15, SC13, ZW12b]. Studden [Det14]. Student [LN15, LVK10, COP10]. Student- [COP10]. studies [AMR11b, CP12, LLP14, MDT^{+}12, RF13, TP10, WY14, WAL17, ZLM12]. Study [QRGLDMA^{+}12, BP13b, BKPS14, GD15, Han14, JM13, PBRT16, Qi11, Qiu10]. Stylometric [GS11b]. sub [GLW11, Men13, TKB11, XZZ18]. sub-fractional [Men13, XZZ18]. sub-model [GLW11]. sub-plots [TKB11]. subcritical [BDLP14]. subject [HCZ19, NCC11, THG12, TLBJ13, YL10a, ZYZL14]. subordinated [FLL17b]. subordination [Gho14]. subpopulation [Ros14, YS11a]. Subsampling [DDL14, SK10, SR14]. subset [AG10, AG12b, Lia11c, VAM10]. subsets [WN10]. success [AZ13, Cha10a]. successive [Sør19, VdUÁMM11]. Sudoku [Fon11]. sufficiency [Son11a]. sufficient [Lia15, XWY18]. sum [AHKS12, CS12a, DMP11, Dhi18, GJM18, TB11, YW10]. sums [BP10b, BP13a, BP19, GD18, KPQ11, PD11, RL17, ZCN11]. Super [CYW10, KBNB11, ZC11, CY10]. Super-simple [CYW10, ZC11, CY10]. supercritical [SSZ16]. superdiagonal [DMR10b]. superimposed [BLP11]. superposition [DFM10, DKN12, GC11]. Supersaturated [Geo14, ST19, TS10, BCMAW14, CKMS12, CLFZ13, CDST18, GSKP10, GHM11, GCDK12, GM12, HLL12, KMMP11, LL11d, NIHY11, QCG15, ySD10, ZDLW19]. Support [NYA17, CCAGV14, Gaf19, HXZ12, Maj14, PKMK12, RBRTM11, TS12, vdVW14]. supported [DM13, Sch10a, SSDM14]. supremum [XW13a]. Sure [YYX19, Vas13]. surface [AG10, AG12b, HH14c, LYXC16]. surfaces [Maj14, WY10]. surprising [AS10b]. Surrogate [JR10, DDW16, NBSB13]. survey [AS10a, BSV10, GLR11, GT10, HP19, Hum10, SZZ16, SSS12, SM11]. surveys [CT10a, Cha19b, CP19, CS11a, SKFMR14]. survival [BCG13, Bur11b, Che12, DF12a, GZ14b, HZ18, JM13, KBK15, KWL14, NS10a, RM12, sS12c, SZ13, SZ10, VdUÁMM11, WZ18, dLJ10, dLJ12]. SVM [BP10d]. SVM-like [BP10d]. swamping [WS15]. switching [ABMP15, Gas11b, MK18]. Symbolic [DS12, LRB11]. symmetric [DDS13, Fre11a, JAI11, KX11b, Lem12, ZCN11]. symmetrical [BS11a, Kak18]. symmetrical-based [Kak18]. symmetrically [JR12]. Symmetries [Jon15]. symmetrized [CND10]. symmetry [BL10b, BC12, DIAT11, EG16, GRCGL12, KD12, RR12]. system [Bro10, Gup12, LZL12, LJ12, MZP12, Sad11, SS00, TLBJ11, TLBJ13]. system-of-equations [LJ12]. systematic [KLK11]. systems [AMNS11, BSV10, BMPR11, BGR11, Beu10, DDL10, DDNN10, DL12, Ery11, Fan10, JR10, JM10a, KM10, Kut11, LL11b, NR12, Ruk14, SAE11, TV12, Zha10, ZB11, dOCG12].

table [NBSB13]. tables [BSV10, CKG19, HY10a, KD12, KBPW12, VV10, Zho19]. tabu [GM12]. Taguchi [TCE^{+}12]. Tail [HM10, MN16, BP10a, CG11, FJC12, GGG11, GGS12, HCT16, Hil11, HW17, KX11a, Kra12, LN15, PT13, TSZ13]. tail-distributions [GGG11]. tailed [ALP10, BT16, BP10a, BMNY13, CM10, EGGG12, FF12, Nan12, WW11b]. tails [HdGT12]. Tangency [BMPT19]. targeting [LBH17]. Techniques [Ano10-27, ABK^{+}11, AS11b, AS10a, Chr10]. temporal [Maj14]. temporally [ABK^{+}11]. tennis [FGN10]. term [DH11a, ZL11b]. termination [CMCSC11]. terms [AG12b, GJM18, Mai11, PR16, SDGK15]. Terpstra [VB11]. Terpstra-type [VB11]. terraces [AJGG10]. Test [BC12, AM19, Ani10, ACP16, BT16, BC11a, BPP10, BL10b, BMPT19, BK12, CDd11, CPH19, CS16b, CS12b, DI10, DC10, DY19, DF12b, Gas11b, GPRY11, GBP15, GKM10, GP12, GJ12, GT12b, Hab15, HHMO15, Hay16, HOK19, Jac19, Jar10, KF12, KCZ13, KU14, LLL^{+}12, LP11b, LZ13b, LC11, LS12, LNdDLdCH11, Mai11, MYW11, MFW11, Mir10, MA08, NU19, NBSB13, NB10, PA13, Pas10, QZZ11, RR12, SS13c, Son11b, TQ12, VT18, Ver17, VSK^{+}11, VGH13, VB11, WX19, Wes16, WK13, Wu16, YL10a, YU13, Zha12, ZCP13]. Testing [BF12, Buc14, EG16, Jar13, LLM11, LZ13a, MS12a, MR10, NHSP13, Nku10, OP12, PA12, YSH11, ZK15, ALS12, BR14, BHS12, BPR16, Cab10, CY11, CMS13, CD15, DB12, Dic13, DLZ14, Fis12, Gho13, GG10, GRCGL12, GLZ14, Hab12, IK12, JZ11, Kli11, KS10b, LP12, LZL^{+}10, Li12, LSZ16, LLZS17, MR11a, MCC12, MSA14, Par10b, PD11, PH17, PR12, RD12, RA10, SDZ19, ST11, SN12, Thu14, TW11, Wan10a, Wie19a, XT15, ZK17, ZCP13, ZZW19, Zhi10, ZGZ17]. testing-based [DLZ14]. Tests [ZN13, AE16, Ani13, AT10a, BTKN10, BS14, BDvdA15, Bed10, Bha11, BSG16, CMT17, CK13, CW17, CR16a, Ciu13, CFW14, DE16, DM12, FP12a, GIP17, GX12, HJ16, HGH12, IYA16, IYA19, JJY12, KS19a, KLW10, Kim10, Kou11, Kut11, Lem11, LF11, Lem12, LM16, LT10, LNdDLdCH11, MS11, MP11, MB11, NdCJ10, NN10, PR11, Que10, QM11, Ros14, SS12a, SLC15, SL19b, TDW10, TCC12, WLCK11, YK15, YY10, ZX12]. their [DG14, Geo14, GF13, GC11, Lon17, OQ10, SDZ19, ySDM13, TG16, TY10]. themes [HS18]. theorem [CT14, CL13b, Qi11, RL17, Yan17, MZ13]. theorems [BO11, HH15, Lon17, Nog13, SY18, Wie19b, Zhi10]. theoretic [BS10a, Thu14]. theoretical [BP10d, LC16]. theory [BP14b, BMPT19, Böh10, Cla19, Cue14, DG13a, DB11, Huc14, Lee18, PA14, Pav12, PPB16, SZ12, SZ19, TV12, Wag11, YC11b, ZC10]. think [PS14]. thinning [CF12a, ZWZ10, ZJ10]. Third [NR11, AG12b, GP10]. Third-order [NR11]. Three [HJ14, Lom11, RRN14, SDV14, WT14b, BTT15, CDZ11, CSWY11, EQ15, HKR11, HHJ13, Jon13b, KD12, MH16, NB12a, NB12b, SGS12, Seb19, TXLV11]. three-decision [Jon13b]. three-level [SGS12]. three-parameter [HKR11, NB12a, NB12b]. Three-phase [HJ14, RRN14, SDV14, WT14b]. three-stage [BTT15]. Three-way [Lom11, KD12]. threshold [LL11b, MSBM16, MP12b, ZX19]. threshold-determining [MP12b]. thresholded [DIAT11]. Thresholding [Mei13, Lia11a, MY13, MPT13a, MPT13b, Obo13, Par10a, SDNH13, Yua13, vdG13a]. Tian [HJ14, SDV14]. ties [BBS10]. Tikhonov [ZCX11]. tilt [WTL11]. tilted [DHM18]. Time [CL15b, ABMP15, ACC18, AHKS12, BL10a, BP10c, BPR16, BP15, BPP10, BFKK17, BD12, CZ12, CGL11, CH11b, Che11b, CH10, CR13, CL13b, CSN19, CFW14, DPM13, DP13, DDL14, DLDC12, ES16, FLW13, FJKW13, FGH14, FPLR11, Fre14, Gas11b, Gri11, HFW10, Hud13, HC11b, JZ13, JZ11, JM13, KMT10, KLS11, KZW10, KLKO15, KK15, KK12, LÁPMF19, LSZ13, LCL11, LZ13a, LT13, MS12b, MRW11, Mor14, NL14, PR12, Reg11, SMM19, SY12, Sør19, SS13d, SSZ11, TSZ13, TNWW12, Tor15, Unk17, UP14, VN10, WF11, Wei11, Yan15, ZSH12, ZY11, Zho10, ZJ10]. Time-dependent [CL15b, BFKK17, CH10, HC11b, SSZ11, Yan15]. time-frequency [KLKO15]. time-varying [CGL11, FLW13, Gri11, KZW10, LCL11, Mor14, PR12]. times [FZS16, KU14, KM12b, KWL14, LD12, Ste11a, SSZ11, VdUÁMM11, dLJ10, dLJ12]. Tobit [GK17, Son11b]. Toeplitz [Wag11]. Tolerance [KX11b, HLB^{+}16, LYC12, VB10b]. tomography [MA17]. tool [Grö18]. tori [BD18]. torus [DPT11]. Total [FRK14, Dal11, DT14, JKL17, Mon17]. toxicity [CL15a, RF13, TNWW12]. toxicity-efficacy [RF13]. toxicodynamic [AAKR12]. toxicokinetic [AAKR12]. toxicokinetic-toxicodynamic [AAKR12]. Tracking [CB19, Say12]. trade [FM19]. trade-off [FM19]. trader [PL16]. training [AM19]. training/test [AM19]. trait [Rab14a]. transaction [XZ19]. Transductive [AH12]. transformation [Coa14, DF12a, FS14, Kak18, LZ12a, MA10, TDW10, WW15b, WDRB18, YZ12]. transformation-invariant [TDW10]. transformations [AC11, EA10, FK11a, Jon15]. transformed [GF18]. transforms [CS11b]. transition [CGCL10]. transitive [CCS14]. translated [DB19]. translation [PG10]. treatment [AMR11b, BL15, BBB11, DD13, HDC14, Jac19, JM13, LT18, PBND11, Ros14, Ruk12, TNWW12, Wae10, YL10a, YSW10, dLJ10, dLJ12]. treatments [AM12a, DST10, GZH11, Jac19, Kri10, KME10, OL19, WA16, YL10a]. tree [RWB15]. treed [KKL15]. trees [GP15a, Pan10]. trend [CR13, MB11, SY12]. trending [CS12b]. trial [HC11a, LW19, TNWW12, ZL11b]. trials [AZ13, BBB11, BS10b, Cha10a, CS13, FJC12, JM13, Kri10, LJK12, SM13, WCZ17, WLL^{+}17, YBB10]. Tribouley [vdG13a]. trigonometric [CL13a, DMS11]. trigonometry [BEJS10]. trimmed [Cíz16, LCT10]. triple [YKH13]. Tripling [OZQ19]. triply [HP19]. trivariate [FKMR19]. true [CR10, HW13, KL12]. Truncated [KM12a, BCMRMP11, Cop11, DF12a, HLCW12, LdUÁ11, dUJC12, sS10b, sScY10, sS12b, SKS10]. truncation [BM11, WDRB18]. Tukey [SS13c]. tumor [LB11]. Tuning [LL19, Lia11b, MN16]. Turnbull [DN12]. twice [sS12c]. Two [AC11, AM12b, Ciu13, FK19, GIP17, HM13, WLL^{+}17, AE16, Alo11, AP14a, AHO13, ABW13, BSH13, BL15, BDLP14, BMPR11, BLT16, BL10b, BSG16, BD18, BK13, CM10, COS12, Cha19a, CC11, CR16a, CS16b, DDL10, DLZ14, DELS16, EQ15, FLL17b, GMSK12, GSG12, GSKP10, GM19, HKN18, HP10b, HLSL17, HDC14, HZ11b, HC16, IF10, IKU11, Jac11, JM10a, KS19a, Kat11, KGV10, KK19, LO11a, LYF14, Lau12, LLL^{+}12, LJA15, LQ17, LT14b, LST12, LSW13, LA14, MY10, MH10, MP19, NPK10, Nan12, NBSB13, Nog13, OP12, OQL11, OQC13, Par10b, RA12, RLZ12, SBCK11, Sch10b, Seb19, ST19, SLC18, SD10, Shu11, SK10, SB11, Tan10, TQ12, TLM12, TS13, TL13, TL18b, WTL11, WZ11b, YL14, YYZ14, YW16, Zha12, ZB11, ZLM12, ZC12, ZLZK13, ZLL16a]. two- [ZC12]. two-colour [BSH13]. Two-dimensional [AC11, AP14a, LJA15, LA14, NPK10, Nan12, TS13]. two-factor [ST19]. two-group [WTL11]. Two-level [HM13, AHO13, Cha19a, GSG12, GSKP10, HLSL17, HZ11b, Jac11, KGV10, LQ17, LT14b, LST12, OQL11, OQC13, RLZ12, SLC18, SD10, TLM12, TL18b, WZ11b, YL14, ZLZK13, ZLL16a]. two-parameter [GMSK12, SBCK11]. two-phase [LO11a, SK10, Tan10, ZLM12]. Two-sample [GIP17, AE16, BD18, CC11, CR16a, CS16b, LLL^{+}12, SB11]. two-scales [FLL17b]. two-sided [CM10]. Two-stage [WLL^{+}17, DLZ14, GM19, IKU11, LYF14, LSW13, YW16]. two-step [DELS16, Shu11]. two-variable [HKN18]. Two-way [AM12b, NBSB13, Zha12]. type [AP14a, AT10b, BTKN10, BBKP19, BS14, Bou11, BM12b, CM10, Che10a, Dac10, EQ15, Fro14, Gas11b, GG10, GK12, HdGT12, KF10, Lac14, LBR12, NB10, Ogi19, RP11a, SR13a, SB11, TMK10, Tsa17, VB11, YHWL12, YAW18, YQC11, ZL11a, Bur11b, CL10b, GMSK12, GT12b, IB11, KR12, MRY11, Say12, SBCK11, VB10b, Wan10a, GP19, Tsa11]. type-I [Che10a, Bur11b, GT12b]. Type-II [CL10b, GMSK12, IB11, KR12, MRY11, SBCK11, VB10b, Wan10a]. types [BK13, RH11].

Uhlenbeck [Bar17, BSKS18, BWYZ11, Gri11, Lac14, LBL12, Men13]. Ultimate [Lac14]. ultra [LC13, YW16]. ultra-high [LC13, YW16]. ultrahigh [LMZ18, YYX19]. ultrahigh-dimensional [LMZ18]. UMPU [Bed10]. Unbalanced [WAL17, Bur11a, HCMP12, SXX11]. Unbiased [BM12a, WN11b, Oga17, RV16, RS10, Son11a]. unbiasedness [KKKP19, NQ10, WV12]. uncensored [KX11b]. uncertainties [SL14b]. uncertainty [AS11b, HFW10, LSP12a]. uncorrelated [Mai11]. undersmoothing [Luk14]. unequal [LZL16, LBH17, MC11, NHSP13, Sah11, SYA11a, Tak10, WA16, Yao10]. unified [Cha19b, HH15, XWY18]. Uniform [AJM19, KLS11, Oua13, CH11c, FLTV10, JA17, KA11, NMS12, NZF11, TB12, TX13, Yan15, YZZ17]. Uniform-in-bandwidth [Oua13]. Uniformity [QWC12, DD12, LJA15, SL19b]. Uniformly [Ros14, AS10b, BP15]. uniqueness [Yan17]. Unit [BDvdA15, BD12, CK13, CS12b, Fan10, TB12]. Unitals [BBPW14]. unitary [CCS14]. units [HDC14, Qu11]. Univariate [Jon15, CA10, HZ18, WS15]. universal [LC15]. universally [FJKW13]. unknown [BT16, CNdGAL11, NB12b, Tsu10]. unlabeled [TH14]. unpaired [LNdDLdCH11]. unrelated [MY10]. unreplicated [EA10, MFW11]. unstable [SZ19]. upcrossings [SMPF10]. upper [BGR11, GR11, LSWY16, Sai10]. urn [Cha12]. use [DH11b, Jac11, KK15, Kri10, TG16]. used [Alb10a]. using [AX19, AC11, AS11b, ALP10, BSV10, BJQ15, BS11a, Beh12, Bur11a, CSR15, CT11, CHP11, CSN19, CS11b, EZS11, FCC10, FM19, Fre11a, Fre11b, GK17, Gho13, GS11b, Gui19, GP15a, HCT16, JKL17, JD16, JMW13, Kak18, KL12, KK12, Kon12, KKP10, LVK10, LM17a, LC11, MK18, Mad10, MPN^{+}17, Mar18, McN10, MS12b, Miy18, MHK10, NBSB13, Nik13, PR16, Par10a, PBRT16, PL11, Reg11, RM12, SG12a, SC13, SYL13, TDW10, Thu14, VS12, YC11a, YY10, YCY15]. utility [HC11a, WK10].

V [Ber14, WL14]. V-optimal [WL14]. vacation [DEF11]. Vajda [Le13]. valid [CR16a]. Validation [LYC12, AM19, CR10, DHM18, GK17, IR19, LZH11, Jan15]. value [BWW19, CDd11, EE17, GBdW10, GS12b, KDJ13, LS12, MSBM16, MV11a, NDD16, Par10b, ZCP13, ZZ14a]. valued [GR14, HdGT12, MSP10, XZ19, ZLW10, Zhu12]. values [AE16, CH19, Dic13, GR11, Hab15, KV11, Llo10, RA12, SMPF10]. Vandermonde [Qui10]. VaR [WYY^{+}10]. Variable [HH14a, HZ13, DDW16, GRCGL12, HKN18, HG11, Jan15, KKK17, LM14, LLW16, LLS19, LZH11, Mül12, NCC11, OL19, PR16, PT12, VCA10, WW15a, WL16, WM18b, XXWL12, ZLZL16]. variables [AKK12, BP10b, BP13a, BP19, BU10, BW13, BRvdGZ13a, CC15, CS12a, DM11, EGS10a, FLW13, FLTV10, FK11b, GD18, Hon10, JR12, KA11, KX11a, KMW15, LLS19, PR16, RS13b, SS13a, SLD14, Sum17, Tsa19, Wei12, WG13, YU13, YW10, ZCN11]. Variance [CPSY10, Che11b, Göb12, Han11, LW12b, AP11, BET17, BC11b, BH15, Bur11a, CM11, CFW14, DMTZ15, FRK14, GD18, GF13, GB19, HC13, KZ16, KS10a, LSZ13, LC16, LZL16, MS13a, PKL12, PH13, PR12, Ruk14, Sai10, SS12a, SA10, SZ19, Tso11, XXWL12, Yao10, YH12]. variance-based [FRK14]. variance-mean [KZ16]. Variance-penalized [LW12b]. variances [Alb10a, COS12, CL11a, CP19, CL11b, HvdM12, LZL16, Miy10, NCC11, Ruk19a, Tak10, Tsu10, WA16]. variate [AJM19]. variation [ALT10, Dal11, GB12, JKL17, LWZ13]. variations [FLL17b]. variogram [LA14]. various [LC15, RH11, SC16]. VARMA [Mai11, VT18]. Varying [LL15, ZL18, AMRP18, AP14b, Cha10a, CGL11, DZZ18, DELS16, EEL10, FLW13, FKMR19, Gri11, HH14a, KZW10, LCL11, LZ12a, LM13, LMZ18, LLS19, LZH11, Mor14, PR12, SBAC19, WZW13, WL16, Wei12, ZS15]. Varying-coefficient [LL15, LZ12a, LLS19, WZW13, ZS15]. Vasicek [XZZ18]. Vector [DM13, BMW15, Dut11, GP12, HXZ12, LLZS17, Moj12, NYA17, PKMK12, PA13, PR12, SN12, UP14, WX19]. vectorial [Que10]. vectors [BSG16, BP10d, FF12, Jar10, MC11, NHSP13, Rüs19, YL10b, ZK15]. verification [FRF12, NdCJ10]. version [CZ11a, CND10]. versions [AH12, SCHG19]. versus [AMR11b, BCZ19, CK13]. via [AZ10, BP13b, BW12, BC11b, CT10c, CR13, CP12, CHH19, DKY15, FJKW13, FWS18, FLL17b, GT12a, GK12, GM12, HP10a, HdGT12, HS14, HH15, HM10, HZ13, LP11b, LQ17, LYL11, OP12, OQC13, PKL12, SL14a, TMK10, TRL13, WK10, WMH12, Xu17, YLL14, YSGN10, ZLM17, ZK15]. vicious [CDZ11]. view [She14]. vision [OG10]. visit [ES16]. visualisation [HG12]. Visualizing [Qui10]. Viterbi [CXCK13]. Volatility [LT13, AVML11, AK10, CGCL10, DHH10, GT12a, HP10a, Hil11, LLJK13]. volume [Ano19a, CPL18]. voyage [Sri10]. vs [NN10]. VSI [YC11a].

waiting [KM12b, LD12, Ste11a]. Walker [De 13, HW13]. walkers [CDZ11]. walks [Nik12, Nkw10]. Wang [HL13, KL13, Xia13, YDZA13]. Watson [GP15b, SSZ16, SLW13]. Wavelet [CDNS14, OG10, RV10, Wis13, BP13b, FRM11, Lia11a, Lia14a, Lu13, MZWF17, Par10a, WW13]. wavelets [DIAT11]. way [AM12b, Bur11a, GRCGL12, KD12, LZL16, Lom11, NBSB13, Tso11, Zha12]. Weak [RW18, CMLBSM13, LC15, PC14]. Weakly [FJKW13, BK17, Kro19]. Weibull [AHD12, ALP10, FVB11, GGG11, GG10, GM19, KL10b, KR12, MRY11, MV11a, WK13]. Weibull-Gamma [MV11a]. Weibull-type [GG10]. weighing [FHK11, Gra12a]. weight [LN15]. Weighted [BFZ10, Chr14, SGL13, ZX19, ZZ14a, BP10b, BP13a, BP19, BC11a, BRB12, BRB18, CG11, CT14, CS12a, CR16b, CDNS14, Dhi18, Gui19, LYH12, LW12a, LLJK13, Lin12, LLX12, LW15, MSA14, PC13, PL11, sS10a, SLW13, SY18, VZH17, Yan15, YY10, ZCN11, ZG12]. Weighted-averaging [ZX19]. weighted-fractional [SY18]. weighting [GZ10, Hay16]. weights [AP14b, BMPT19, TN10]. Wets [dC12]. which [Cla19, HHMO15, MS10, RS16]. Whittle [CW17]. whole [GW11, GG15, WZZ19]. Wiener [TTB11]. Wilcoxon [TB11]. William [Det14]. wind [ABMP15]. window [ZWM12]. window-observation [ZWM12]. winner [WLL^{+}17]. Wishart [BF13, Di 16, SC13]. within [PR16]. without [GP10, Jan15, LM11b]. WLP [BK11c, BK10]. word [LZZ15]. wordlength [BB14]. working [Lin10, LLZ11a]. Wu [HJ14, SDV14, CK12].

yield [jLP11]. yields [RS16]. Youden [Nil11].

Z [Nik12]. Zero [Zhu12, BLP11, DDD18]. Zero-inflated [Zhu12]. zeroinflation [DDD18]. zero-mean [BLP11]. zeros [ZZW12]. Zhang [RS13b, SS13a].

References

Abbasnejad:2011:REP

[AA11a] M. Abbasnejad and N. R. Arghami. Rényi entropy properties of records. Journal of Statistical Planning and Inference, 141(7):2312–2320, July 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

               http://www.sciencedirect.com/science/article/pii/
               S0378375811000346.

Ahmed:2011:DBN

[AA11b] Rashid Ahmed and Munir Akhtar. Designs balanced for neighbor effects in circular blocks of size six. Journal of Statistical Planning and Inference, 141(2):687–691, February 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

               tronic). URL http://www.sciencedirect.com/science/
               article/pii/S0378375810003654. See erratum [AA11c].

Ahmed:2011:EDB

[AA11c] Rashid Ahmed and Munir Akhtar. Erratum to “Designs balanced for neighbor effects in circular blocks of size six” [J. Statist. Plann. Inference 141 (2011) 687–691]. Journal of Statistical Planning and Inference, 141(5):2002–2003, May 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

               1171 (electronic). URL http://www.sciencedirect.com/
               science/article/pii/S0378375810005367. See [AA11b].

Albert:2012:BED

[AAKR12] C. Albert, R. Ashauer, H. R. Künsch, and P. Reichert. Bayesian experimental design for a toxicokinetictoxicodynamic model. Journal of Statistical Planning and Inference, 142(1):263–275, January 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375811002990.

Azzalini:2013:MPL

[AAV13] Adelchi Azzalini and Reinaldo B. Arellano-Valle. Maximum penalized likelihood estimation for skew-normal and skew-t distributions. Journal of Statistical Planning and Inference, 143(2):419–433, February 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375812002352.

Atkinson:2014:ODE

[AB14] Anthony C. Atkinson and Barbara Bogacka. Optimum designs for the equality of parameters in enzyme inhibition kinetic models. Journal of Statistical Planning and Inference, 144(??):47–54, January 2014. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375812002480.

Adams:2012:DCG

[ABFG12] P. Adams, D. E. Bryant, A. D. Forbes, and T. S. Griggs. Decomposing the complete graph into dodecahedra. Journal of Statistical Planning and Inference, 142(5):1040–1046, May 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S037837581100406X.

Adolf:2011:AMT

[ABK^{+}11] Daniela Adolf, Sebastian Baecke, Waltraud Kahle, Johannes Bernarding, and Siegfried Kropf. Applying multivariate techniques to high-dimensional temporally correlated fMRI data. Journal of Statistical Planning and Inference, 141(12):3760–3770, December 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375811002370.

Ailliot:2015:NHH

[ABMP15] Pierre Ailliot, Julie Bessac, Valérie Monbet, and Françoise Pène. Non-homogeneous hidden Markov-switching models for wind time series. Journal of Statistical Planning and Inference, 160(??):75–88, May 2015. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375814002018.

Abraham:2012:BRU

[Abr12] Christophe Abraham. Bayesian regression under combinations of constraints. Journal of Statistical Planning and Inference, 142(9):2672–2687, September 2012. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375812001255.

Arasu:2013:PBI

[ABW13] K. T. Arasu, Pradeep Bansal, and Cody Watson. Partially balanced incomplete block designs with two associate classes. Journal of Statistical Planning and Inference, 143(5):983– 991, May 2013. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375812003436.

Anderes:2011:TDD

[AC11] Ethan Anderes and Marc A. Coram. Two-dimensional density estimation using smooth invertible transformations. Journal of Statistical Planning and Inference, 141(3):1183– 1193, March 2011. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375810004301.

Arratia:2018:ELD

[ACC18] Argimiro Arratia, Alejandra Cabaña, and Enrique M. Cabaña. Embedding in law of discrete time ARMA processes in continuous time stationary processes. Journal of Statistical Planning and Inference, 197(??):156–167, December 2018. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375818300065.

Arias-Castro:2016:CCT

[ACP16] Ery Arias-Castro and Bruno Pelletier. On the consistency of the crossmatch test. Journal of Statistical Planning and Inference, 171(??):184–190, April 2016. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375815300677.

Althubaiti:2011:MEM

[AD11] Alaa Althubaiti and Alexander N. Donev. Mixture experiments with mixing errors. Journal of Statistical Planning and Inference, 141(2):692–700, February 2011. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810003630.

Amaral:2010:BCR

[ADPW10] Getulio J. A. Amaral, Ian L. Dryden, Vic Patrangenaru, and Andrew T. A. Wood. Bootstrap confidence regions for the planar mean shape. Journal of Statistical Planning and Inference, 140(11):3026–3034, November 2010. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810001680.

Abd-Elfattah:2016:SVT

[AE16] Ehab F. Abd-Elfattah. Saddlepoint p-values for two-sample bivariate tests. Journal of Statistical Planning and Inference, 171(??):92–98, April 2016. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375815002013.

Aulbach:2012:LAN

[AF12] Stefan Aulbach and Michael Falk. Local asymptotic normality in δ-neighborhoods of standard generalized Pareto processes. Journal of Statistical Planning and Inference, 142(6):1339–1347, June 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375811004563.

Atkinson:2014:EIM

[AFHZ14] Anthony C. Atkinson, Valerii V. Fedorov, Agnes M. Herzberg, and Rongmei Zhang. Elemental information matrices and optimal experimental design for generalized regression models. Journal of Statistical Planning and Inference, 144(??):81–91, January 2014. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375812003060.

Ahmad:2010:RSR

[AG10] Tanvir Ahmad and Steven G. Gilmour. Robustness of subset response surface designs to missing observations. Journal of Statistical Planning and Inference, 140(1):92–103, January 1, 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375809001955.

Alvarez:2012:NBR

[ÁG12a] B. Lagos Álvarez and M. D. Jiménez Gamero. A note on bias reduction of maximum likelihood estimates for the scalar skew t distribution. Journal of Statistical Planning and Inference, 142(2):608–612, February 2012. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811003429.

Arshad:2012:PSR

[AG12b] Hafiz Muhammad Arshad and Steven G. Gilmour. The performance of subset response surface designs for estimating third order terms. Journal of Statistical Planning and Inference, 142(1):232–247, January 2012. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811002825.

Azais:2019:EAN

[AG19] Romain Azaïs and Alexandre Genadot. Estimation of the average number of continuous crossings for non-stationary non-diffusion processes. Journal of Statistical Planning and Inference, 198(??):119–138, January 2019. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375818300351.

Azais:2012:OQA

[AGPS12] Romain Azaïs, Anne Gégout-Petit, and Jérôme Saracco. Optimal quantization applied to sliced inverse regression. Journal of Statistical Planning and Inference, 142(2):481–492, February 2012. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S037837581100317X.

Alquier:2012:TVL

[AH12] Pierre Alquier and Mohamed Hebiri. Transductive versions of the LASSO and the Dantzig Selector. Journal of Statistical Planning and Inference, 142(9):2485–2500, September 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S037837581200136X.

Ahmed:2012:LSE

[AHD12] S. Ejaz Ahmed, Shakhawat Hossain, and Kjell A. Doksum. LASSO and shrinkage estimation in Weibull censored regression models. Journal of Statistical Planning and Inference, 142(6):1273–1284, June 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375812000055.

Aue:2012:RTM

[AHKS12] Alexander Aue, Lajos Horváth, Mario Kühn, and Josef Steinebach. On the reaction time of moving sum detectors. Journal of Statistical Planning and Inference, 142(8): 2271–2288, August 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375812001000.

Ai:2012:SNC

[AHL12] Mingyao Ai, Yuanzhen He, and Senmao Liu. Some new classes of orthogonal Latin hypercube designs. Journal of Statistical Planning and Inference, 142(10):2809–2818, October 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375812001590.

Aoki:2013:MCM

[AHO13] Satoshi Aoki, Takayuki Hibi, and Hidefumi Ohsugi. Markov chain Monte Carlo methods for the regular two-level fractional factorial designs and cut ideals. Journal of Statistical Planning and Inference, 143(10):1791–1806, October 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375813001468.

Aggarwal:2010:CSC

[AJGG10] M. L. Aggarwal, Mithilesh Kumar Jha, Seema Gupta, and Sarah Garland. Construction of some classes of balanced cross-over designs from terraces. Journal of Statistical Planning and Inference, 140(10):2817–2827, October 2010. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810001163.

Ahmadi-Javid:2019:UDR

[AJM19] Amir Ahmadi-Javid and Asghar Moeini. Uniform distributions and random variate generation over generalized l_{p} balls and spheres. Journal of Statistical Planning and Inference, 201(??):1–19, July 2019. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375818302544.

Ahlip:2010:CAP

[AK10] Rehez Ahlip and Rik King. Computational aspects of pricing foreign exchange options with stochastic volatility and stochastic interest rates. Journal of Statistical Planning and Inference, 140(5):1256–1268, May 2010. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809003619.

Abdous:2012:SRC

[AKK12] Belkacem Abdous, Célestin C. Kokonendji, and Tristan Senga Kiessé. On semiparametric regression for count explanatory variables. Journal of Statistical Planning and Inference, 142(6):1537–1548, June 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375812000158.

Aknouche:2015:ESP

[Akn15] Abdelhakim Aknouche. Explosive strong periodic autoregression with multiplicity one. Journal of Statistical Planning and Inference, 161(??):50–72, June 2015. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375814002079.

Albatineh:2010:MVF

[Alb10a] Ahmed N. Albatineh. Means and variances for a family of similarity indices used in cluster analysis. Journal of Statistical Planning and Inference, 140(10):2828–2838, October 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375810001175.

Albers:2010:OCN

[Alb10b] Willem Albers. The optimal choice of negative binomial charts for monitoring high-quality processes. Journal of Statistical Planning and Inference, 140(1):214–225, January 1, 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375809002122.

Albers:2011:ENC

[Alb11] Willem Albers. Empirical nonparametric control charts for high-quality processes. Journal of Statistical Planning and Inference, 141(9):3151–3159, September 2011. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811001340.

Albers:2012:CCH

[Alb12] Willem Albers. Control charts for high-quality processes: MAX or cumax? Journal of Statistical Planning and Inference, 142(7):1924–1932, July 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375812000237.

Ai:2013:BIL

[ALLL13] Mingyao Ai, Kang Li, Senmao Liu, and Dennis K. J. Lin. Balanced incomplete Latin square designs. Journal of Statistical Planning and Inference, 143(9):1575–1582, September 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375813001043.

Agboto:2010:SDM

[ALN10] Vincent Agboto, William Li, and Christopher Nachtsheim. Screening designs for model discrimination. Journal of Statistical Planning and Inference, 140(3):766–780, March 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809002869.

Alodat:2011:ACS

[Alo11] M. T. Alodat. An approximation to cluster size distribution of two Gaussian random fields conjunction with application to FMRI data. Journal of Statistical Planning and Inference, 141(7):2331–2347, July 2011. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811000358.

Asimit:2010:PUW

[ALP10] Alexandru V. Asimit, Deyuan Li, and Liang Peng. Pitfalls in using Weibull tailed distributions. Journal of Statistical Planning and Inference, 140(7):2018–2024, July 2010. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810000637.

Avram:2012:HTF

[ALS12] F. Avram, N. N. Leonenko, and N. Suvak. Hypothesis testing for Fisher-snedecor diffusion. Journal of Statistical Planning and Inference, 142(8):2308–2321, August 2012. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375812001024.

Albrecher:2010:ASC

[ALT10] H. Albrecher, Sophie A. Ladoucette, and Jef L. Teugels. Asymptotics of the sample coefficient of variation and the sample dispersion. Journal of Statistical Planning and Inference, 140(2):358–368, February 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375809002729.

Aly:2019:QQP

[Aly19] Emad-Eldin A. A. Aly. Quantile–quantile plots under random censorship. Journal of Statistical Planning and Inference, 15(??):123–128, ???? 2019. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
037837588690090X.

Andrews:2011:MMF

[AM11] Jeffrey L. Andrews and Paul D. McNicholas. Mixtures of modified t-factor analyzers for model-based clustering, classification, and discriminant analysis. Journal of Statistical Planning and Inference, 141(4):1479–1486, April 2011. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810004830.

Ambrozy:2012:ESN

[AM12a] Katarzyna Ambrozy and Iwona Mejza. On the efficiency of some non-orthogonal split-plot× split-block designs with control treatments. Journal of Statistical Planning and Inference, 142(3):752–762, March 2012. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811003740.

Arnold:2012:TWM

[AM12b] Steven F. Arnold and Panagis G. Moschopoulos. Two-way model with random cell sizes. Journal of Statistical Planning and Inference, 142(11):2965–2975, November 2012. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375812001796.

Afendras:2019:OTT

[AM19] Georgios Afendras and Marianthi Markatou. Optimality of training/test size and resampling effectiveness in cross-validation. Journal of Statistical Planning and Inference, 199(??):286–301, March 2019. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375818301514.

Al-Mutairi:2011:MCB

[AMNS11] Dhaifalla K. Al-Mutairi, Asok K. Nanda, and Richard M. Soland. Monotonicity and comparability in ballistic missile defense systems. Journal of Statistical Planning and Inference, 141(7):2191–2200, July 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375811000309.

Allman:2011:PIC

[AMR11a] Elizabeth S. Allman, Catherine Matias, and John A. Rhodes. Parameter identifiability in a class of random graph mixture models. Journal of Statistical Planning and Inference, 141(5):1719–1736, May 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375810005227.

Azriel:2011:TVE

[AMR11b] D. Azriel, M. Mandel, and Y. Rinott. The treatment versus experimentation dilemma in dose finding studies. Journal of Statistical Planning and Inference, 141(8):2759–2768, August 2011. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375811000978.

Arteaga-Molina:2018:ELB

[AMRP18] Luis A. Arteaga-Molina and Juan M. Rodriguez-Poo. Empirical likelihood based inference for fixed effects varying coefficient panel data models. Journal of Statistical Planning and Inference, 196(??):144–162, August 2018. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375817302082.

Anis:2010:CSA

[Ani10] M. Z. Anis. Corrigendum to: “An L-statistic approach to a test of exponentiality against IFR alternatives” [Journal of Statistical Planning and Inference 138 (2008) 3144–3148]. Journal of Statistical Planning and Inference, 140(6):1618, June 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375809004005. See [MA08].

Anis:2012:SPI

[Ani12] M. Z. Anis. On some properties of the IDMRL class of life distributions. Journal of Statistical Planning and Inference, 142(11):3047–3055, November 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375812001851.

Anis:2013:FTE

[Ani13] M. Z. Anis. A family of tests for exponentiality against IFR alternatives. Journal of Statistical Planning and Inference, 143(8):1409–1415, August 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375813000578.

Anonymous:2010:ICL

[Ano10a] Anonymous. The 6th International Conference on Lattice Path Combinatorics and Applications. Journal of Statistical Planning and Inference, 140(8):2155, August 2010. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810001618.

Anonymous:2010:Ca

[Ano10b] Anonymous. Contents. Journal of Statistical Planning and Inference, 140(1):iii–iv, January 1, 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375809003000.

Anonymous:2010:Cb

[Ano10c] Anonymous. Contents. Journal of Statistical Planning and Inference, 140(2):iii–iv, February 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375809003292.

Anonymous:2010:Cc

[Ano10d] Anonymous. Contents. Journal of Statistical Planning and Inference, 140(3):iii–iv, March 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375809003401.

Anonymous:2010:Cd

[Ano10e] Anonymous. Contents. Journal of Statistical Planning and Inference, 140(4):iii–iv, April 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S037837580900384X.

Anonymous:2010:Ce

[Ano10f] Anonymous. Contents. Journal of Statistical Planning and Inference, 140(5):iii–iv, May 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S037837581000011X.

Anonymous:2010:Cf

[Ano10g] Anonymous. Contents. Journal of Statistical Planning and Inference, 140(6):iii–iv, June 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S037837581000056X.

Anonymous:2010:Cg

[Ano10h] Anonymous. Contents. Journal of Statistical Planning and Inference, 140(12):iii–iv, December 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375810003381.

Anonymous:2010:EBa

[Ano10i] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 140(7):obc, July 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375810001102.

Anonymous:2010:EBb

[Ano10j] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 140(8):obc, August 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375810001552.

Anonymous:2010:EBd

[Ano10k] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 140(9):obc, September 2010. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810001965.

Anonymous:2010:EBe

[Ano10l] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 140(10):obc, October 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375810002600.

Anonymous:2010:EBf

[Ano10m] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 140(11):obc, November 2010. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810003071.

Anonymous:2010:EBCa

[Ano10n] Anonymous. Editorial board (continued). Journal of Statistical Planning and Inference, 140(7):ibc, July 2010. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810001096.

Anonymous:2010:EBCb

[Ano10o] Anonymous. Editorial board (continued). Journal of Statistical Planning and Inference, 140(8):ibc, August 2010. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810001540.

Anonymous:2010:EBCc

[Ano10p] Anonymous. Editorial board (continued). Journal of Statistical Planning and Inference, 140(9):ibc, September 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810001953.

Anonymous:2010:EBCd

[Ano10q] Anonymous. Editorial board (continued). Journal of Statistical Planning and Inference, 140(10):ibc, October 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810002594.

Anonymous:2010:EBCe

[Ano10r] Anonymous. Editorial Board (continued). Journal of Statistical Planning and Inference, 140(11):ibc, November 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837581000306X.

Anonymous:2010:EP

[Ano10s] Anonymous. Emmanuel Parzen. Journal of Statistical Planning and Inference, 140(12):??, December 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2010:IECa

[Ano10t] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 140(1):ibc, January 1, 2010. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809003012.

Anonymous:2010:IECb

[Ano10u] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 140(2):ibc, February 2010. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809003309.

Anonymous:2010:IECc

[Ano10v] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 140(3):ibc, March 2010. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809003413.

Anonymous:2010:IECd

[Ano10w] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 140(4):ibc, April 2010. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809003851.

Anonymous:2010:IECe

[Ano10x] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 140(5):ibc, May 2010. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810000121.

Anonymous:2010:IECf

[Ano10y] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 140(6):ibc, June 2010. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810000571.

Anonymous:2010:IECg

[Ano10z] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 140(12):ibc, December 2010. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810003538.

Anonymous:2010:ICI

[Ano10-27] Anonymous. International conference on interdisciplinary mathematical and statistical techniques. Journal of Statistical Planning and Inference, 140(10):??, October 2010. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2010:OEa

[Ano10-28] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 140(1):obc, January 1, 2010. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809003024.

Anonymous:2010:OEb

[Ano10-29] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 140(2):obc, February 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375809003310.

Anonymous:2010:OEc

[Ano10-30] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 140(3):obc, March 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375809003425.

Anonymous:2010:OEd

[Ano10-31] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 140(4):obc, April 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375809003863.

Anonymous:2010:OEe

[Ano10-32] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 140(5):obc, May 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375810000133.

Anonymous:2010:OEf

[Ano10-33] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 140(6):obc, June 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375810000583.

Anonymous:2010:OEg

[Ano10-34] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 140(12):obc, December 2010. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837581000354X.

Anonymous:2010:PJa

[Ano10-35] Anonymous. Pages 1-334 (1 January 2010). Journal of Statistical Planning and Inference, 140(1):??, January 1, 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2010:PMb

[Ano10-36] Anonymous. Pages 1101-1352 (May 2010). Journal of Statistical Planning and Inference, 140(5):??, May 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2010:PJb

[Ano10-37] Anonymous. Pages 1353-1618 (June 2010). Journal of Statistical Planning and Inference, 140(6):??, June 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2010:PJc

[Ano10-38] Anonymous. Pages 1619-2154 (July 2010). Journal of Statistical Planning and Inference, 140(7):??, July 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2010:PS

[Ano10-39] Anonymous. Pages 2355-2798 (September 2010). Journal of Statistical Planning and Inference, 140(9):??, September 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2010:PN

[Ano10-40] Anonymous. Pages 2937-3576 (November 2010). Journal of Statistical Planning and Inference, 140(11):??, November 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2010:PF

[Ano10-41] Anonymous. Pages 335-584 (February 2010). Journal of Statistical Planning and Inference, 140(2):??, February 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2010:PMa

[Ano10-42] Anonymous. Pages 585-850 (March 2010). Journal of Statistical Planning and Inference, 140(3):??, March 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2010:PA

[Ano10-43] Anonymous. Pages 851-1100 (April 2010). Journal of Statistical Planning and Inference, 140(4):??, April 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2010:SIC

[Ano10-44] Anonymous. Sixth International Conference on Lattice Path Combinatorics and Applications. Journal of Statistical Planning and Inference, 140(8):??, August 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2011:Ca

[Ano11a] Anonymous. Contents. Journal of Statistical Planning and Inference, 141(1):iii–v, January 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375810003836.

Anonymous:2011:Cb

[Ano11b] Anonymous. Contents. Journal of Statistical Planning and Inference, 141(2):iii–iv, February 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375810004465.

Anonymous:2011:Cc

[Ano11c] Anonymous. Contents. Journal of Statistical Planning and Inference, 141(3):iii–iv, March 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375810004921.

Anonymous:2011:Cd

[Ano11d] Anonymous. Contents. Journal of Statistical Planning and Inference, 141(4):iii–iv, April 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375810005513.

Anonymous:2011:Ce

[Ano11e] Anonymous. Contents. Journal of Statistical Planning and Inference, 141(5):iii–v, May 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375811000073.

Anonymous:2011:Cf

[Ano11f] Anonymous. Contents. Journal of Statistical Planning and Inference, 141(6):iii, June 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375811000486.

Anonymous:2011:Cg

[Ano11g] Anonymous. Contents. Journal of Statistical Planning and Inference, 141(7):iii–iv, July 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375811000875.

Anonymous:2011:Ch

[Ano11h] Anonymous. Contents. Journal of Statistical Planning and Inference, 141(8):iii–iv, August 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S037837581100142X.

Anonymous:2011:Ci

[Ano11i] Anonymous. Contents. Journal of Statistical Planning and Inference, 141(9):iii–iv, September 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375811001819.

Anonymous:2011:Cj

[Ano11j] Anonymous. Contents. Journal of Statistical Planning and Inference, 141(10):iii, October 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375811002138.

Anonymous:2011:Ck

[Ano11k] Anonymous. Contents. Journal of Statistical Planning and Inference, 141(11):iii–iv, November 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375811002515.

Anonymous:2011:Cl

[Ano11l] Anonymous. Contents. Journal of Statistical Planning and Inference, 141(12):iii–iv, December 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375811002874.

Anonymous:2011:EBa

[Ano11m] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 141(1):obc, January 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375810004015.

Anonymous:2011:EBb

[Ano11n] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 141(2):obc, February 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375810004672.

Anonymous:2011:EBCa

[Ano11o] Anonymous. Editorial board (continued). Journal of Statistical Planning and Inference, 141(1):ibc, January 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810004003.

Anonymous:2011:EBCb

[Ano11p] Anonymous. Editorial Board continued. Journal of Statistical Planning and Inference, 141(2):ibc, February 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810004647.

Anonymous:2011:IECa

[Ano11q] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 141(3):ibc, March 2011. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810004970.

Anonymous:2011:IECb

[Ano11r] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 141(4):ibc, April 2011. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810005598.

Anonymous:2011:IECc

[Ano11s] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 141(5):ibc, May 2011. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811000140.

Anonymous:2011:IECd

[Ano11t] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 141(6):ibc, June 2011. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811000553.

Anonymous:2011:IECe

[Ano11u] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 141(7):ibc, July 2011. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811000954.

Anonymous:2011:IECf

[Ano11v] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 141(8):ibc, August 2011. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811001479.

Anonymous:2011:IECg

[Ano11w] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 141(9):ibc, September 2011. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811001868.

Anonymous:2011:IECh

[Ano11x] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 141(10):ibc, October 2011. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811002151.

Anonymous:2011:IECi

[Ano11y] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 141(11):ibc, November 2011. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811002618.

Anonymous:2011:IECj

[Ano11z] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 141(12):ibc, December 2011. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811002965.

Anonymous:2011:OEa

[Ano11-27] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 141(3):obc, March 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375810004982.

Anonymous:2011:OEb

[Ano11-28] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 141(4):obc, April 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375810005604.

Anonymous:2011:OEc

[Ano11-29] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 141(5):obc, May 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375811000152.

Anonymous:2011:OEd

[Ano11-30] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 141(6):obc, June 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375811000565.

Anonymous:2011:OEe

[Ano11-31] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 141(7):obc, July 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375811000966.

Anonymous:2011:OEf

[Ano11-32] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 141(8):obc, August 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375811001480.

Anonymous:2011:OEg

[Ano11-33] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 141(9):obc, September 2011. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837581100187X.

Anonymous:2011:OEh

[Ano11-34] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 141(10):obc, October 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375811002163.

Anonymous:2011:OEi

[Ano11-35] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 141(11):obc, November 2011. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837581100262X.

Anonymous:2011:OEj

[Ano11-36] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 141(12):obc, December 2011. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811002977.

Anonymous:2011:PJa

[Ano11-37] Anonymous. Pages 1-610 (January 2011). Journal of Statistical Planning and Inference, 141(1):??, January 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2011:PMa

[Ano11-38] Anonymous. Pages 1077-1354 (March 2011). Journal of Statistical Planning and Inference, 141(3):??, March 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2011:PAa

[Ano11-39] Anonymous. Pages 1355-1632 (April 2011). Journal of Statistical Planning and Inference, 141(4):??, April 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2011:PMb

[Ano11-40] Anonymous. Pages 1633-2006 (May 2011). Journal of Statistical Planning and Inference, 141(5):??, May 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2011:PJb

[Ano11-41] Anonymous. Pages 2007-2174 (June 2011). Journal of Statistical Planning and Inference, 141(6):??, June 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2011:PJc

[Ano11-42] Anonymous. Pages 2175-2502 (July 2011). Journal of Statistical Planning and Inference, 141(7):??, July 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2011:PAb

[Ano11-43] Anonymous. Pages 2503-2986 (August 2011). Journal of Statistical Planning and Inference, 141(8):??, August 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2011:PS

[Ano11-44] Anonymous. Pages 2987-3292 (September 2011). Journal of Statistical Planning and Inference, 141(9):??, September 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2011:PO

[Ano11-45] Anonymous. Pages 3293-3366 (October 2011). Journal of Statistical Planning and Inference, 141(10):??, October 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2011:PN

[Ano11-46] Anonymous. Pages 3367-3680 (November 2011). Journal of Statistical Planning and Inference, 141(11):??, November 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2011:PD

[Ano11-47] Anonymous. Pages 3681-3876 (December 2011). Journal of Statistical Planning and Inference, 141(12):??, December 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2011:PF

[Ano11-48] Anonymous. Pages 611-1076 (February 2011). Journal of Statistical Planning and Inference, 141(2):??, February 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2012:Ca

[Ano12a] Anonymous. Contents. Journal of Statistical Planning and Inference, 142(1):iii–iv, January 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375811003284.

Anonymous:2012:Cb

[Ano12b] Anonymous. Contents. Journal of Statistical Planning and Inference, 142(2):iii–iv, February 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375811003612.

Anonymous:2012:Cc

[Ano12c] Anonymous. Contents. Journal of Statistical Planning and Inference, 142(3):iii, March 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375811003879.

Anonymous:2012:Cd

[Ano12d] Anonymous. Contents. Journal of Statistical Planning and Inference, 142(4):iii–iv, April 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375811004381.

Anonymous:2012:Ce

[Ano12e] Anonymous. Contents. Journal of Statistical Planning and Inference, 142(5):iii–iv, May 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375811004782.

Anonymous:2012:Cf

[Ano12f] Anonymous. Contents. Journal of Statistical Planning and Inference, 142(6):iii–iv, June 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375812000341.

Anonymous:2012:Cg

[Ano12g] Anonymous. Contents. Journal of Statistical Planning and Inference, 142(7):iii–v, July 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375812001115.

Anonymous:2012:Ch

[Ano12h] Anonymous. Contents. Journal of Statistical Planning and Inference, 142(8):iii–iv, August 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375812001425.

Anonymous:2012:Ci

[Ano12i] Anonymous. Contents. Journal of Statistical Planning and Inference, 142(9):iii–iv, September 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S037837581200167X.

Anonymous:2012:Cj

[Ano12j] Anonymous. Contents. Journal of Statistical Planning and Inference, 142(10):iii, October 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375812001917.

Anonymous:2012:Ck

[Ano12k] Anonymous. Contents. Journal of Statistical Planning and Inference, 142(11):iii–iv, November 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375812002182.

Anonymous:2012:Cl

[Ano12l] Anonymous. Contents. Journal of Statistical Planning and Inference, 142(12):iii–iv, December 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375812002546.

Anonymous:2012:IECa

[Ano12m] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 142(1):ibc, January 2012. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811003351.

Anonymous:2012:IECb

[Ano12n] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 142(2):ibc, February 2012. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811003685.

Anonymous:2012:IECc

[Ano12o] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 142(3):ibc, March 2012. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811003934.

Anonymous:2012:IECd

[Ano12p] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 142(4):ibc, April 2012. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811004435.

Anonymous:2012:IECe

[Ano12q] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 142(5):ibc, May 2012. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811004812.

Anonymous:2012:IECf

[Ano12r] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 142(6):ibc, June 2012. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375812000407.

Anonymous:2012:IECg

[Ano12s] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 142(7):ibc, July 2012. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375812001140.

Anonymous:2012:IECh

[Ano12t] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 142(8):ibc, August 2012. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375812001449.

Anonymous:2012:IECi

[Ano12u] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 142(9):ibc, September 2012. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375812001735.

Anonymous:2012:IECj

[Ano12v] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 142(10):ibc, October 2012. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375812001930.

Anonymous:2012:IECk

[Ano12w] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 142(11):ibc, November 2012. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375812002194.

Anonymous:2012:IECl

[Ano12x] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 142(12):ibc, December 2012. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375812002571.

Anonymous:2012:OEa

[Ano12y] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 142(1):obc, January 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375811003363.

Anonymous:2012:OEb

[Ano12z] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 142(2):obc, February 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375811003697.

Anonymous:2012:OEc

[Ano12-27] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 142(3):obc, March 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375811003946.

Anonymous:2012:OEd

[Ano12-28] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 142(4):obc, April 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375811004447.

Anonymous:2012:OEe

[Ano12-29] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 142(5):obc, May 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375811004824.

Anonymous:2012:OEf

[Ano12-30] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 142(6):obc, June 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375812000419.

Anonymous:2012:OEg

[Ano12-31] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 142(7):obc, July 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375812001152.

Anonymous:2012:OEh

[Ano12-32] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 142(8):obc, August 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375812001450.

Anonymous:2012:OEi

[Ano12-33] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 142(9):obc, September 2012. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375812001747.

Anonymous:2012:OEj

[Ano12-34] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 142(10):obc, October 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375812001942.

Anonymous:2012:OEk

[Ano12-35] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 142(11):obc, November 2012. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375812002200.

Anonymous:2012:OEl

[Ano12-36] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 142(12):obc, December 2012. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375812002583.

Anonymous:2012:PJa

[Ano12-37] Anonymous. Pages 1-378 (January 2012). Journal of Statistical Planning and Inference, 142(1):??, January 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2012:PMb

[Ano12-38] Anonymous. Pages 1031-1278 (May 2012). Journal of Statistical Planning and Inference, 142(5):??, May 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2012:PJb

[Ano12-39] Anonymous. Pages 1273-1626 (June 2012). Journal of Statistical Planning and Inference, 142(6):??, June 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2012:PJc

[Ano12-40] Anonymous. Pages 1627-2240 (July 2012). Journal of Statistical Planning and Inference, 142(7):??, July 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2012:PAb

[Ano12-41] Anonymous. Pages 2241-2484 (August 2012). Journal of Statistical Planning and Inference, 142(8):??, August 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2012:PS

[Ano12-42] Anonymous. Pages 2485-2734 (September 2012). Journal of Statistical Planning and Inference, 142(9):??, September 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2012:PO

[Ano12-43] Anonymous. Pages 2735-2870 (October 2012). Journal of Statistical Planning and Inference, 142(10):??, October 2012.

CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2012:PN

[Ano12-44] Anonymous. Pages 2871-3060 (November 2012). Journal of Statistical Planning and Inference, 142(11):??, November 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2012:PD

[Ano12-45] Anonymous. Pages 3061-3254 (December 2012). Journal of Statistical Planning and Inference, 142(12):??, December 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2012:PF

[Ano12-46] Anonymous. Pages 379-612 (February 2012). Journal of Statistical Planning and Inference, 142(2):??, February 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2012:PMa

[Ano12-47] Anonymous. Pages 613-780 (March 2012). Journal of Statistical Planning and Inference, 142(3):??, March 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2012:PAa

[Ano12-48] Anonymous. Pages 781-1030 (April 2012). Journal of Statistical Planning and Inference, 142(4):??, April 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2013:Ca

[Ano13a] Anonymous. Contents. Journal of Statistical Planning and Inference, 143(1):iii–iv, January 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375812002923.

Anonymous:2013:Cb

[Ano13b] Anonymous. Contents. Journal of Statistical Planning and Inference, 143(2):iii–iv, February 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375812003138.

Anonymous:2013:Cc

[Ano13c] Anonymous. Contents. Journal of Statistical Planning and Inference, 143(3):iii–iv, March 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375812003357.

Anonymous:2013:Cd

[Ano13d] Anonymous. Contents. Journal of Statistical Planning and Inference, 143(4):iii–iv, April 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375812003564.

Anonymous:2013:Ce

[Ano13e] Anonymous. Contents. Journal of Statistical Planning and Inference, 143(5):iii, May 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375813000104.

Anonymous:2013:Cf

[Ano13f] Anonymous. Contents. Journal of Statistical Planning and Inference, 143(6):iii, June 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375813000372.

Anonymous:2013:Cg

[Ano13g] Anonymous. Contents. Journal of Statistical Planning and Inference, 143(7):iii–iv, July 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375813000815.

Anonymous:2013:Ch

[Ano13h] Anonymous. Contents. Journal of Statistical Planning and Inference, 143(8):iii, August 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375813001006.

Anonymous:2013:Ci

[Ano13i] Anonymous. Contents. Journal of Statistical Planning and Inference, 143(9):iii–iv, September 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375813001341.

Anonymous:2013:Cj

[Ano13j] Anonymous. Contents. Journal of Statistical Planning and Inference, 143(10):iii–iv, October 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375813001730.

Anonymous:2013:Ck

[Ano13k] Anonymous. Contents. Journal of Statistical Planning and Inference, 143(11):iii–iv, November 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375813001997.

Anonymous:2013:Cl

[Ano13l] Anonymous. Contents. Journal of Statistical Planning and Inference, 143(12):iii, December 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375813002292.

Anonymous:2013:IECa

[Ano13m] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 143(1):ibc, January 2013. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375812002935.

Anonymous:2013:IECb

[Ano13n] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 143(2):ibc, February 2013. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837581200314X.

Anonymous:2013:IECc

[Ano13o] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 143(3):ibc, March 2013. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375812003370.

Anonymous:2013:IECd

[Ano13p] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 143(4):ibc, April 2013. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375812003588.

Anonymous:2013:IECe

[Ano13q] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 143(5):ibc, May 2013. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375813000177.

Anonymous:2013:IECf

[Ano13r] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 143(6):ibc, June 2013. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375813000414.

Anonymous:2013:IECg

[Ano13s] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 143(7):ibc, July 2013. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375813000827.

Anonymous:2013:IECh

[Ano13t] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 143(8):ibc, August 2013. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375813001018.

Anonymous:2013:IECi

[Ano13u] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 143(9):ibc, September 2013. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375813001316.

Anonymous:2013:IECj

[Ano13v] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 143(10):ibc, October 2013. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375813001717.

Anonymous:2013:IECk

[Ano13w] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 143(11):ibc, November 2013. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375813002000.

Anonymous:2013:IECl

[Ano13x] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 143(12):ibc, December 2013. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837581300222X.

Anonymous:2013:OEa

[Ano13y] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 143(1):obc, January 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375812002947.

Anonymous:2013:OEb

[Ano13z] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 143(2):obc, February 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375812003151.

Anonymous:2013:OEc

[Ano13-27] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 143(3):obc, March 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375812003382.

Anonymous:2013:OEd

[Ano13-28] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 143(4):obc, April 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S037837581200359X.

Anonymous:2013:OEe

[Ano13-29] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 143(5):obc, May 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375813000189.

Anonymous:2013:OEf

[Ano13-30] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 143(6):obc, June 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375813000426.

Anonymous:2013:OEg

[Ano13-31] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 143(7):obc, July 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375813000839.

Anonymous:2013:OEh

[Ano13-32] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 143(8):obc, August 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S037837581300102X.

Anonymous:2013:OEi

[Ano13-33] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 143(9):obc, September 2013. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375813001328.

Anonymous:2013:OEj

[Ano13-34] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 143(10):obc, October 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375813001729.

Anonymous:2013:OEk

[Ano13-35] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 143(11):obc, November 2013. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375813002012.

Anonymous:2013:OEl

[Ano13-36] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 143(12):obc, December 2013. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375813002231.

Anonymous:2013:PJa

[Ano13-37] Anonymous. Pages 1-220 (January 2013). Journal of Statistical Planning and Inference, 143(1):??, January 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2013:PJb

[Ano13-38] Anonymous. Pages 1001-1108 (June 2013). Journal of Statistical Planning and Inference, 143(6):??, June 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2013:PJc

[Ano13-39] Anonymous. Pages 1109-1256 (July 2013). Journal of Statistical Planning and Inference, 143(7):??, July 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2013:PAb

[Ano13-40] Anonymous. Pages 1249-1416 (August 2013). Journal of Statistical Planning and Inference, 143(8):??, August 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2013:PS

[Ano13-41] Anonymous. Pages 1417-1620 (September 2013). Journal of Statistical Planning and Inference, 143(9):??, September 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2013:PO

[Ano13-42] Anonymous. Pages 1621-1834 (October 2013). Journal of Statistical Planning and Inference, 143(10):??, October 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2013:PN

[Ano13-43] Anonymous. Pages 1835-2038 (November 2013). Journal of Statistical Planning and Inference, 143(11):??, November 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2013:PD

[Ano13-44] Anonymous. Pages 2039-2170 (December 2013). Journal of Statistical Planning and Inference, 143(12):??, December 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2013:PF

[Ano13-45] Anonymous. Pages 221-456 (February 2013). Journal of Statistical Planning and Inference, 143(2):??, February 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2013:PMa

[Ano13-46] Anonymous. Pages 457-636 (March 2013). Journal of Statistical Planning and Inference, 143(3):??, March 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2013:PAa

[Ano13-47] Anonymous. Pages 637-866 (April 2013). Journal of Statistical Planning and Inference, 143(4):??, April 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2013:PMb

[Ano13-48] Anonymous. Pages 867-1000 (May 2013). Journal of Statistical Planning and Inference, 143(5):??, May 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2014:Ca

[Ano14a] Anonymous. Contents. Journal of Statistical Planning and Inference, 144(??):iii–iv, January 2014. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375813002413.

Anonymous:2014:Cb

[Ano14b] Anonymous. Contents. Journal of Statistical Planning and Inference, 145(??):iii–iv, February 2014. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S037837581300267X.

Anonymous:2014:Cc

[Ano14c] Anonymous. Contents. Journal of Statistical Planning and Inference, 146(??):iii, March 2014. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375813002863.

Anonymous:2014:Cd

[Ano14d] Anonymous. Contents. Journal of Statistical Planning and Inference, 147(??):iii–iv, April 2014. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375813003194.

Anonymous:2014:Ce

[Ano14e] Anonymous. Contents. Journal of Statistical Planning and Inference, 148(??):iii, May 2014. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375814000172.

Anonymous:2014:Cf

[Ano14f] Anonymous. Contents. Journal of Statistical Planning and Inference, 149(??):iii–iv, June 2014. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375814000494.

Anonymous:2014:Cg

[Ano14g] Anonymous. Contents. Journal of Statistical Planning and Inference, 150(??):iii, July 2014. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375814000652.

Anonymous:2014:Ch

[Ano14h] Anonymous. Contents. Journal of Statistical Planning and Inference, 151–152(??):iii, August/September 2014. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375814000986.

Anonymous:2014:Ci

[Ano14i] Anonymous. Contents. Journal of Statistical Planning and Inference, 153(??):iii, October 2014. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375814001104.

Anonymous:2014:Cj

[Ano14j] Anonymous. Contents. Journal of Statistical Planning and Inference, 155(??):iii, December 2014. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S037837581400161X.

Anonymous:2014:IECa

[Ano14k] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 144(??):ibc, January 2014. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375813002425.

Anonymous:2014:IECb

[Ano14l] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 145(??):ibc, February 2014. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837581300270X.

Anonymous:2014:IECc

[Ano14m] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 146(??):ibc, March 2014. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375813002875.

Anonymous:2014:IECd

[Ano14n] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 147(??):ibc, April 2014. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375813003200.

Anonymous:2014:IECe

[Ano14o] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 148(??):ibc, May 2014. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375814000184.

Anonymous:2014:IECf

[Ano14p] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 149(??):ibc, June 2014. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375814000512.

Anonymous:2014:IECg

[Ano14q] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 150(??):ibc, July 2014. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375814000664.

Anonymous:2014:IECh

[Ano14r] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 151–152(??):ibc, August/September 2014. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375814000913.

Anonymous:2014:IECi

[Ano14s] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 153(??):ibc, October 2014. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375814001116.

Anonymous:2014:IECj

[Ano14t] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 154(??):ibc, November 2014. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375814001311.

Anonymous:2014:IECk

[Ano14u] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 155(??):ibc, December 2014. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375814001499.

Anonymous:2014:ICD

[Ano14v] Anonymous. International Conference on Design of Experiments (ICODOE-2011). Journal of Statistical Planning and Inference, 144(??):1, January 2014. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375813002449.

Anonymous:2014:OEa

[Ano14w] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 144(??):obc, January 2014. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375813002437.

Anonymous:2014:OEb

[Ano14x] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 145(??):obc, February 2014. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375813002711.

Anonymous:2014:OEc

[Ano14y] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 146(??):obc, March 2014. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375813002887.

Anonymous:2014:OEd

[Ano14z] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 147(??):obc, April 2014. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375813003212.

Anonymous:2014:OEe

[Ano14-27] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 148(??):obc, May 2014. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375814000196.

Anonymous:2014:OEf

[Ano14-28] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 149(??):obc, June 2014. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375814000524.

Anonymous:2014:OEg

[Ano14-29] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 150(??):obc, July 2014. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375814000676.

Anonymous:2014:OEh

[Ano14-30] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 151–152(??):obc, August/September 2014. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375814000925.

Anonymous:2014:OEi

[Ano14-31] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 153(??):obc, October 2014. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375814001128.

Anonymous:2014:OEj

[Ano14-32] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 154(??):obc, November 2014. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375814001323.

Anonymous:2014:OEk

[Ano14-33] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 155(??):obc, December 2014. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375814001505.

Anonymous:2014:PO

[Ano14-34] Anonymous. Pages 1-114 (October 2014). Journal of Statistical Planning and Inference, 153(??):1–114, October 2014. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2014:PAS

[Ano14-35] Anonymous. Pages 1-120 (August-September 2014). Journal of Statistical Planning and Inference, 151–152(??):1–120, August/September 2014. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2014:PMb

[Ano14-36] Anonymous. Pages 1-128 (May 2014). Journal of Statistical Planning and Inference, 148(??):1–128, May 2014. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2014:PD

[Ano14-37] Anonymous. Pages 1-140 (December 2014). Journal of Statistical Planning and Inference, 155(??):1–140, December 2014. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2014:PMa

[Ano14-38] Anonymous. Pages 1-144 (March 2014). Journal of Statistical Planning and Inference, 146(??):1–144, March 2014. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2014:PJb

[Ano14-39] Anonymous. Pages 1-158 (July 2014). Journal of Statistical Planning and Inference, 150(??):1–158, July 2014. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2014:PJa

[Ano14-40] Anonymous. Pages 1-218 (June 2014). Journal of Statistical Planning and Inference, 149(??):1–218, June 2014. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2014:PF

[Ano14-41] Anonymous. Pages 1-222 (February 2014). Journal of Statistical Planning and Inference, 145(??):1–222, February 2014. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2014:PA

[Ano14-42] Anonymous. Pages 1-248 (April 2014). Journal of Statistical Planning and Inference, 147(??):1–248, April 2014. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2015:Ca

[Ano15a] Anonymous. Contents. Journal of Statistical Planning and Inference, 156(??):iii, January 2015. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375814001694.

Anonymous:2015:Cb

[Ano15b] Anonymous. Contents. Journal of Statistical Planning and Inference, 157–158(??):iii, February/March 2015. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375814001852.

Anonymous:2015:Cc

[Ano15c] Anonymous. Contents. Journal of Statistical Planning and Inference, 159(??):iii, April 2015. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375814002122.

Anonymous:2015:Cd

[Ano15d] Anonymous. Contents. Journal of Statistical Planning and Inference, 160(??):iii, May 2015. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S037837581500018X.

Anonymous:2015:Ce

[Ano15e] Anonymous. Contents. Journal of Statistical Planning and Inference, 161(??):iii, June 2015. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375815000312.

Anonymous:2015:Cf

[Ano15f] Anonymous. Contents. Journal of Statistical Planning and Inference, 162(??):iii, July 2015. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375815000476.

Anonymous:2015:Cg

[Ano15g] Anonymous. Contents. Journal of Statistical Planning and Inference, 163(??):iii, August 2015. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375815000634.

Anonymous:2015:Ch

[Ano15h] Anonymous. Contents. Journal of Statistical Planning and Inference, 164(??):iii, September 2015. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375815000865.

Anonymous:2015:Ci

[Ano15i] Anonymous. Contents. Journal of Statistical Planning and Inference, 165(??):iii, October 2015. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375815001032.

Anonymous:2015:Cj

[Ano15j] Anonymous. Contents. Journal of Statistical Planning and Inference, 166(??):iii–iv, November 2015. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375815001226.

Anonymous:2015:Ck

[Ano15k] Anonymous. Contents. Journal of Statistical Planning and Inference, 167(??):iii–iv, December 2015. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375815001494.

Anonymous:2015:Ea

[Ano15l] Anonymous. Editors. Journal of Statistical Planning and Inference, 157–158(??):obc, February/March 2015. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375814001864.

Anonymous:2015:Eb

[Ano15m] Anonymous. Editors. Journal of Statistical Planning and Inference, 159(??):obc, April 2015. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375814002158.

Anonymous:2015:Ed

[Ano15n] Anonymous. Editors. Journal of Statistical Planning and Inference, 160(??):obc, May 2015. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375815000208.

Anonymous:2015:Ee

[Ano15o] Anonymous. Editors. Journal of Statistical Planning and Inference, 161(??):obc, June 2015. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375815000336.

Anonymous:2015:Ef

[Ano15p] Anonymous. Editors. Journal of Statistical Planning and Inference, 162(??):obc, July 2015. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375815000518.

Anonymous:2015:Eg

[Ano15q] Anonymous. Editors. Journal of Statistical Planning and Inference, 163(??):obc, August 2015. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375815000658.

Anonymous:2015:Eh

[Ano15r] Anonymous. Editors. Journal of Statistical Planning and Inference, 164(??):obc, September 2015. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375815000907.

Anonymous:2015:Ei

[Ano15s] Anonymous. Editors. Journal of Statistical Planning and Inference, 165(??):obc, October 2015. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375815001068.

Anonymous:2015:Ej

[Ano15t] Anonymous. Editors. Journal of Statistical Planning and Inference, 166(??):obc, November 2015. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S037837581500124X.

Anonymous:2015:Ek

[Ano15u] Anonymous. Editors. Journal of Statistical Planning and Inference, 167(??):obc, December 2015. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375815001536.

Anonymous:2015:ECa

[Ano15v] Anonymous. Editors (cont.). Journal of Statistical Planning and Inference, 157–158(??):ibc, February/March 2015. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375814001888.

Anonymous:2015:ECb

[Ano15w] Anonymous. Editors (cont.). Journal of Statistical Planning and Inference, 159(??):ibc, April 2015. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375814002146.

Anonymous:2015:ECc

[Ano15x] Anonymous. Editors (cont.). Journal of Statistical Planning and Inference, 160(??):ibc, May 2015. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375815000191.

Anonymous:2015:ECd

[Ano15y] Anonymous. Editors (cont.). Journal of Statistical Planning and Inference, 161(??):ibc, June 2015. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375815000324.

Anonymous:2015:ECe

[Ano15z] Anonymous. Editors (cont.). Journal of Statistical Planning and Inference, 162(??):ibc, July 2015. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375815000506.

Anonymous:2015:ECf

[Ano15-27] Anonymous. Editors (cont.). Journal of Statistical Planning and Inference, 163(??):ibc, August 2015. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375815000646.

Anonymous:2015:ECg

[Ano15-28] Anonymous. Editors (cont.). Journal of Statistical Planning and Inference, 164(??):ibc, September 2015. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375815000890.

Anonymous:2015:ECh

[Ano15-29] Anonymous. Editors (cont.). Journal of Statistical Planning and Inference, 165(??):ibc, October 2015. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375815001056.

Anonymous:2015:ECi

[Ano15-30] Anonymous. Editors (cont.). Journal of Statistical Planning and Inference, 166(??):ibc, November 2015. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375815001238.

Anonymous:2015:ECj

[Ano15-31] Anonymous. Editors (cont.). Journal of Statistical Planning and Inference, 167(??):ibc, December 2015. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375815001524.

Anonymous:2015:IEC

[Ano15-32] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 156(??):ibc, January 2015. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375814001669.

Anonymous:2015:OE

[Ano15-33] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 156(??):obc, January 2015. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375814001670.

Anonymous:2015:PJa

[Ano15-34] Anonymous. Pages 1-102 (January 2015). Journal of Statistical Planning and Inference, 156(??):1–102, January 2015. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2015:PAa

[Ano15-35] Anonymous. Pages 1-104 (April 2015). Journal of Statistical Planning and Inference, 159(??):1–104, April 2015. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2015:PM

[Ano15-36] Anonymous. Pages 1-106 (May 2015). Journal of Statistical Planning and Inference, 160(??):1–106, May 2015. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2015:PJc

[Ano15-37] Anonymous. Pages 1-110 (July 2015). Journal of Statistical Planning and Inference, 162(??):1–110, July 2015. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2015:PJb

[Ano15-38] Anonymous. Pages 1-124 (June 2015). Journal of Statistical Planning and Inference, 161(??):1–124, June 2015. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2015:PO

[Ano15-39] Anonymous. Pages 1-126 (October 2015). Journal of Statistical Planning and Inference, 165(??):1–126, October 2015. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2015:PFM

[Ano15-40] Anonymous. Pages 1-136 (February-March 2015). Journal of Statistical Planning and Inference, 157–158(??):1–136, February/March 2015. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2015:PD

[Ano15-41] Anonymous. Pages 1-228 (December 2015). Journal of Statistical Planning and Inference, 167(??):1–228, December 2015. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2015:PAb

[Ano15-42] Anonymous. Pages 1-48 (August 2015). Journal of Statistical Planning and Inference, 163(??):1–48, August 2015. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2015:PS

[Ano15-43] Anonymous. Pages 1-54 (September 2015). Journal of Statistical Planning and Inference, 164(??):1–54, September 2015. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2016:Ca

[Ano16a] Anonymous. Contents. Journal of Statistical Planning and Inference, 168(??):iii, January 2016. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S037837581500169X.

Anonymous:2016:Cb

[Ano16b] Anonymous. Contents. Journal of Statistical Planning and Inference, 169(??):iii, February 2016. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375815001810.

Anonymous:2016:Cc

[Ano16c] Anonymous. Contents. Journal of Statistical Planning and Inference, 170(??):iii–iv, March 2016. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375815002086.

Anonymous:2016:Cd

[Ano16d] Anonymous. Contents. Journal of Statistical Planning and Inference, 171(??):iii, April 2016. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375815002232.

Anonymous:2016:Ce

[Ano16e] Anonymous. Contents. Journal of Statistical Planning and Inference, 172(??):iii, May 2016. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375816000136.

Anonymous:2016:Cf

[Ano16f] Anonymous. Contents. Journal of Statistical Planning and Inference, 173(??):iii, June 2016. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S037837581600032X.

Anonymous:2016:Cg

[Ano16g] Anonymous. Contents. Journal of Statistical Planning and Inference, 174(??):iii, July 2016. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375816300106.

Anonymous:2016:Ch

[Ano16h] Anonymous. Contents. Journal of Statistical Planning and Inference, 175(??):iii, August 2016. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375816300234.

Anonymous:2016:Ci

[Ano16i] Anonymous. Contents. Journal of Statistical Planning and Inference, 176(??):iii, September 2016. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375816300337.

Anonymous:2016:Cj

[Ano16j] Anonymous. Contents. Journal of Statistical Planning and Inference, 177(??):iii, October 2016. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375816300490.

Anonymous:2016:Ck

[Ano16k] Anonymous. Contents. Journal of Statistical Planning and Inference, 178(??):iii, November 2016. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375816300817.

Anonymous:2016:Cl

[Ano16l] Anonymous. Contents. Journal of Statistical Planning and Inference, 179(??):iii, December 2016. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375816300970.

Anonymous:2016:Ea

[Ano16m] Anonymous. Editors. Journal of Statistical Planning and Inference, 168(??):obc, January 2016. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375815001767.

Anonymous:2016:Eb

[Ano16n] Anonymous. Editors. Journal of Statistical Planning and Inference, 169(??):obc, February 2016. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375815001871.

Anonymous:2016:Ed

[Ano16o] Anonymous. Editors. Journal of Statistical Planning and Inference, 170(??):obc, March 2016. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375815002104.

Anonymous:2016:Ee

[Ano16p] Anonymous. Editors. Journal of Statistical Planning and Inference, 171(??):obc, April 2016. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S037837581500227X.

Anonymous:2016:Ef

[Ano16q] Anonymous. Editors. Journal of Statistical Planning and Inference, 172(??):obc, May 2016. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375816000173.

Anonymous:2016:Eg

[Ano16r] Anonymous. Editors. Journal of Statistical Planning and Inference, 173(??):obc, June 2016. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375816000367.

Anonymous:2016:Eh

[Ano16s] Anonymous. Editors. Journal of Statistical Planning and Inference, 174(??):obc, July 2016. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375816300143.

Anonymous:2016:Ei

[Ano16t] Anonymous. Editors. Journal of Statistical Planning and Inference, 175(??):obc, August 2016. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375816300271.

Anonymous:2016:Ej

[Ano16u] Anonymous. Editors. Journal of Statistical Planning and Inference, 176(??):obc, September 2016. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375816300374.

Anonymous:2016:Ek

[Ano16v] Anonymous. Editors. Journal of Statistical Planning and Inference, 177(??):obc, October 2016. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375816300556.

Anonymous:2016:El

[Ano16w] Anonymous. Editors. Journal of Statistical Planning and Inference, 178(??):obc, November 2016. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375816300842.

Anonymous:2016:Em

[Ano16x] Anonymous. Editors. Journal of Statistical Planning and Inference, 179(??):obc, December 2016. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S037837581630101X.

Anonymous:2016:ECa

[Ano16y] Anonymous. Editors (cont.). Journal of Statistical Planning and Inference, 168(??):ibc, January 2016. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375815001755.

Anonymous:2016:ECb

[Ano16z] Anonymous. Editors (cont.). Journal of Statistical Planning and Inference, 169(??):ibc, February 2016. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S037837581500186X.

Anonymous:2016:ECc

[Ano16-27] Anonymous. Editors (cont.). Journal of Statistical Planning and Inference, 170(??):ibc, March 2016. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375815002098.

Anonymous:2016:ECd

[Ano16-28] Anonymous. Editors (cont.). Journal of Statistical Planning and Inference, 171(??):ibc, April 2016. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375815002268.

Anonymous:2016:ECe

[Ano16-29] Anonymous. Editors (cont.). Journal of Statistical Planning and Inference, 172(??):ibc, May 2016. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375816000161.

Anonymous:2016:ECf

[Ano16-30] Anonymous. Editors (cont.). Journal of Statistical Planning and Inference, 173(??):ibc, June 2016. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375816000355.

Anonymous:2016:ECg

[Ano16-31] Anonymous. Editors (cont.). Journal of Statistical Planning and Inference, 174(??):ibc, July 2016. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375816300131.

Anonymous:2016:ECh

[Ano16-32] Anonymous. Editors (cont.). Journal of Statistical Planning and Inference, 175(??):ibc, August 2016. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S037837581630026X.

Anonymous:2016:ECi

[Ano16-33] Anonymous. Editors (cont.). Journal of Statistical Planning and Inference, 176(??):ibc, September 2016. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375816300362.

Anonymous:2016:ECj

[Ano16-34] Anonymous. Editors (cont.). Journal of Statistical Planning and Inference, 177(??):ibc, October 2016. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375816300544.

Anonymous:2016:ECk

[Ano16-35] Anonymous. Editors (cont.). Journal of Statistical Planning and Inference, 178(??):ibc, November 2016. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375816300830.

Anonymous:2016:ECl

[Ano16-36] Anonymous. Editors (cont.). Journal of Statistical Planning and Inference, 179(??):ibc, December 2016. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375816301008.

Anonymous:2016:PAb

[Ano16-37] Anonymous. Pages 1-100 (August 2016). Journal of Statistical Planning and Inference, 175(??):1–100, August 2016. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2016:PF

[Ano16-38] Anonymous. Pages 1-100 (February 2016). Journal of Statistical Planning and Inference, 169(??):1–100, February 2016. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2016:PJa

[Ano16-39] Anonymous. Pages 1-120 (January 2016). Journal of Statistical Planning and Inference, 168(??):1–120, January 2016.

CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2016:PN

[Ano16-40] Anonymous. Pages 1-140 (November 2016). Journal of Statistical Planning and Inference, 178(??):1–140, November 2016. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2016:PJb

[Ano16-41] Anonymous. Pages 1-146 (June 2016). Journal of Statistical Planning and Inference, 173(??):1–146, June 2016. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2016:PJc

[Ano16-42] Anonymous. Pages 1-168 (July 2016). Journal of Statistical Planning and Inference, 174(??):1–168, July 2016. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2016:PAa

[Ano16-43] Anonymous. Pages 1-190 (April 2016). Journal of Statistical Planning and Inference, 171(??):1–190, April 2016. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2016:PMa

[Ano16-44] Anonymous. Pages 1-200 (March 2016). Journal of Statistical Planning and Inference, 170(??):1–200, March 2016. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2016:PMb

[Ano16-45] Anonymous. Pages 1-52 (May 2016). Journal of Statistical Planning and Inference, 172(??):1–52, May 2016. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2016:PD

[Ano16-46] Anonymous. Pages 1-60 (December 2016). Journal of Statistical Planning and Inference, 179(??):1–60, December 2016. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2016:PO

[Ano16-47] Anonymous. Pages 1-74 (October 2016). Journal of Statistical Planning and Inference, 177(??):1–74, October 2016.

CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2016:PS

[Ano16-48] Anonymous. Pages 1-94 (September 2016). Journal of Statistical Planning and Inference, 176(??):1–94, September 2016. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2017:Ca

[Ano17a] Anonymous. Contents. Journal of Statistical Planning and Inference, 180(??):iii, January 2017. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375816301161.

Anonymous:2017:Cb

[Ano17b] Anonymous. Contents. Journal of Statistical Planning and Inference, 181(??):iii, February 2017. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S037837581630132X.

Anonymous:2017:Cc

[Ano17c] Anonymous. Contents. Journal of Statistical Planning and Inference, 182(??):iii, March 2017. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375816301446.

Anonymous:2017:Cd

[Ano17d] Anonymous. Contents. Journal of Statistical Planning and Inference, 183(??):iii, April 2017. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375816301598.

Anonymous:2017:Ce

[Ano17e] Anonymous. Contents. Journal of Statistical Planning and Inference, 184(??):iii, May 2017. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375817300058.

Anonymous:2017:Cf

[Ano17f] Anonymous. Contents. Journal of Statistical Planning and Inference, 185(??):iii, June 2017. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S037837581730023X.

Anonymous:2017:Cg

[Ano17g] Anonymous. Contents. Journal of Statistical Planning and Inference, 186(??):iii, July 2017. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375817300472.

Anonymous:2017:Ch

[Ano17h] Anonymous. Contents. Journal of Statistical Planning and Inference, 188(??):iii, September 2017. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375817300708.

Anonymous:2017:Ci

[Ano17i] Anonymous. Contents. Journal of Statistical Planning and Inference, 189(??):iii, October 2017. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375817300940.

Anonymous:2017:Cj

[Ano17j] Anonymous. Contents. Journal of Statistical Planning and Inference, 190(??):iii, November 2017. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375817301027.

Anonymous:2017:Ck

[Ano17k] Anonymous. Contents. Journal of Statistical Planning and Inference, 191(??):iii, December 2017. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375817301271.

Anonymous:2017:EBa

[Ano17l] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 187(??):iii, August 2017. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375817300605.

Anonymous:2017:EBb

[Ano17m] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 187(??):ibc, August 2017. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375817300630.

Anonymous:2017:EBc

[Ano17n] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 187(??):obc, August 2017. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375817300642.

Anonymous:2017:EBd

[Ano17o] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 191(??):ibc, December 2017. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375817301295.

Anonymous:2017:Ea

[Ano17p] Anonymous. Editors. Journal of Statistical Planning and Inference, 180(??):obc, January 2017. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375816301203.

Anonymous:2017:Eb

[Ano17q] Anonymous. Editors. Journal of Statistical Planning and Inference, 181(??):obc, February 2017. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375816301367.

Anonymous:2017:Ed

[Ano17r] Anonymous. Editors. Journal of Statistical Planning and Inference, 182(??):obc, March 2017. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375816301483.

Anonymous:2017:Ee

[Ano17s] Anonymous. Editors. Journal of Statistical Planning and Inference, 183(??):obc, April 2017. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S037837581630163X.

Anonymous:2017:Ef

[Ano17t] Anonymous. Editors. Journal of Statistical Planning and Inference, 184(??):obc, May 2017. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375817300083.

Anonymous:2017:Eg

[Ano17u] Anonymous. Editors. Journal of Statistical Planning and Inference, 185(??):obc, June 2017. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375817300265.

Anonymous:2017:Eh

[Ano17v] Anonymous. Editors. Journal of Statistical Planning and Inference, 186(??):obc, July 2017. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375817300502.

Anonymous:2017:Ei

[Ano17w] Anonymous. Editors. Journal of Statistical Planning and Inference, 188(??):obc, September 2017. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375817300745.

Anonymous:2017:Ej

[Ano17x] Anonymous. Editors. Journal of Statistical Planning and Inference, 189(??):obc, October 2017. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375817300976.

Anonymous:2017:Ek

[Ano17y] Anonymous. Editors. Journal of Statistical Planning and Inference, 190(??):obc, November 2017. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375817301052.

Anonymous:2017:El

[Ano17z] Anonymous. Editors. Journal of Statistical Planning and Inference, 191(??):obc, December 2017. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375817301301.

Anonymous:2017:ECa

[Ano17-27] Anonymous. Editors (cont.). Journal of Statistical Planning and Inference, 180(??):ibc, January 2017. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375816301197.

Anonymous:2017:ECb

[Ano17-28] Anonymous. Editors (cont.). Journal of Statistical Planning and Inference, 181(??):ibc, February 2017. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375816301355.

Anonymous:2017:ECc

[Ano17-29] Anonymous. Editors (cont.). Journal of Statistical Planning and Inference, 182(??):ibc, March 2017. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375816301471.

Anonymous:2017:ECd

[Ano17-30] Anonymous. Editors (cont.). Journal of Statistical Planning and Inference, 183(??):ibc, April 2017. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375816301628.

Anonymous:2017:ECe

[Ano17-31] Anonymous. Editors (cont.). Journal of Statistical Planning and Inference, 184(??):ibc, May 2017. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375817300071.

Anonymous:2017:ECf

[Ano17-32] Anonymous. Editors (cont.). Journal of Statistical Planning and Inference, 185(??):ibc, June 2017. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375817300253.

Anonymous:2017:ECg

[Ano17-33] Anonymous. Editors (cont.). Journal of Statistical Planning and Inference, 186(??):ibc, July 2017. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375817300496.

Anonymous:2017:ECh

[Ano17-34] Anonymous. Editors (cont.). Journal of Statistical Planning and Inference, 188(??):ibc, September 2017. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375817300733.

Anonymous:2017:ECi

[Ano17-35] Anonymous. Editors (cont.). Journal of Statistical Planning and Inference, 189(??):ibc, October 2017. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375817300964.

Anonymous:2017:ECj

[Ano17-36] Anonymous. Editors (cont.). Journal of Statistical Planning and Inference, 190(??):ibc, November 2017. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375817301040.

Anonymous:2017:PMa

[Ano17-37] Anonymous. Pages 1-100 (March 2017). Journal of Statistical Planning and Inference, 182(??):1–100, March 2017. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2017:PD

[Ano17-38] Anonymous. Pages 1-110 (December 2017). Journal of Statistical Planning and Inference, 191(??):1–110, December 2017. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2017:PMb

[Ano17-39] Anonymous. Pages 1-116 (May 2017). Journal of Statistical Planning and Inference, 184(??):1–116, May 2017. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2017:PJa

[Ano17-40] Anonymous. Pages 1-124 (January 2017). Journal of Statistical Planning and Inference, 180(??):1–124, January 2017. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2017:PAb

[Ano17-41] Anonymous. Pages 1-142 (August 2017). Journal of Statistical Planning and Inference, 187(??):1–142, August 2017. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2017:PJc

[Ano17-42] Anonymous. Pages 1-58 (July 2017). Journal of Statistical Planning and Inference, 186(??):1–58, July 2017. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2017:PN

[Ano17-43] Anonymous. Pages 1-60 (November 2017). Journal of Statistical Planning and Inference, 190(??):1–60, November 2017. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2017:PO

[Ano17-44] Anonymous. Pages 1-60 (October 2017). Journal of Statistical Planning and Inference, 189(??):1–60, October 2017. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2017:PAa

[Ano17-45] Anonymous. Pages 1-62 (April 2017). Journal of Statistical Planning and Inference, 183(??):1–62, April 2017. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2017:PF

[Ano17-46] Anonymous. Pages 1-80 (February 2017). Journal of Statistical Planning and Inference, 181(??):1–80, February 2017. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2017:PJb

[Ano17-47] Anonymous. Pages 1-94 (June 2017). Journal of Statistical Planning and Inference, 185(??):1–94, June 2017. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2017:PS

[Ano17-48] Anonymous. Pages 1-94 (September 2017). Journal of Statistical Planning and Inference, 188(??):1–94, September 2017. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2018:Ca

[Ano18a] Anonymous. Contents. Journal of Statistical Planning and Inference, 192(??):iii, January 2018. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375817301507.

Anonymous:2018:Cb

[Ano18b] Anonymous. Contents. Journal of Statistical Planning and Inference, 193(??):iii, February 2018. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375817301830.

Anonymous:2018:Cc

[Ano18c] Anonymous. Contents. Journal of Statistical Planning and Inference, 194(??):iii, March 2018. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375817302021.

Anonymous:2018:EBa

[Ano18d] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 193(??):ibc, February 2018. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375817301866.

Anonymous:2018:EBb

[Ano18e] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 193(??):obc, February 2018. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375817301878.

Anonymous:2018:EBc

[Ano18f] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 195(??):ii, May 2018. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375818300107.

Anonymous:2018:EBd

[Ano18g] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 196(??):ii, August 2018. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S037837581830020X.

Anonymous:2018:EBe

[Ano18h] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 197(??):ii, December 2018. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375818300405.

Anonymous:2018:Ea

[Ano18i] Anonymous. Editors. Journal of Statistical Planning and Inference, 192(??):obc, January 2018. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375817301568.

Anonymous:2018:Ed

[Ano18j] Anonymous. Editors. Journal of Statistical Planning and Inference, 194(??):obc, March 2018. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375817302069.

Anonymous:2018:ECa

[Ano18k] Anonymous. Editors (cont.). Journal of Statistical Planning and Inference, 192(??):ibc, January 2018. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375817301556.

Anonymous:2018:ECb

[Ano18l] Anonymous. Editors (cont.). Journal of Statistical Planning and Inference, 194(??):ibc, March 2018. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375817302057.

Anonymous:2018:PM

[Ano18m] Anonymous. Pages 1-136 (March 2018). Journal of Statistical Planning and Inference, 194(??):1–136, March 2018. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2018:PD

[Ano18n] Anonymous. Pages 1-180 (December 2018). Journal of Statistical Planning and Inference, 197(??):1–180, December 2018.

CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2018:PA

[Ano18o] Anonymous. Pages 1-184 (August 2018). Journal of Statistical Planning and Inference, 196(??):1–184, August 2018. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2018:PF

[Ano18p] Anonymous. Pages 1-188 (February 2018). Journal of Statistical Planning and Inference, 193(??):1–188, February 2018. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2018:PJ

[Ano18q] Anonymous. Pages 1-64 (January 2018). Journal of Statistical Planning and Inference, 192(??):1–64, January 2018. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2019:AIV

[Ano19a] Anonymous. Author index volume 15 (1986/87). Journal of Statistical Planning and Inference, 15(??):401–402, ???? 2019. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375886901126.

Anonymous:2019:EBa

[Ano19b] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 198(??):ii, January 2019. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375818300582.

Anonymous:2019:EBb

[Ano19c] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 199(??):ii, March 2019. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375818302878.

Anonymous:2019:EBc

[Ano19d] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 200(??):ii, May 2019. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375818303501.

Anonymous:2019:EBd

[Ano19e] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 201(??):ii, July 2019. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375819300102.

Anonymous:2019:EBe

[Ano19f] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 202(??):ii, September 2019. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375819300229.

Anonymous:2019:EBf

[Ano19g] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 203(??):ii, December 2019. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375819300424.

Anonymous:2019:PS

[Ano19h] Anonymous. Pages 1-122 (September 2019). Journal of Statistical Planning and Inference, 202(??):1–122, September 2019. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2019:PJb

[Ano19i] Anonymous. Pages 1-146 (July 2019). Journal of Statistical Planning and Inference, 201(??):1–146, July 2019. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2019:PJa

[Ano19j] Anonymous. Pages 1-170 (January 2019). Journal of Statistical Planning and Inference, 198(??):1–170, January 2019.

CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2019:PD

[Ano19k] Anonymous. Pages 1-224 (December 2019). Journal of Statistical Planning and Inference, 203(??):1–224, December 2019. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2019:PMb

[Ano19l] Anonymous. Pages 1-238 (May 2019). Journal of Statistical Planning and Inference, 200(??):1–238, May 2019. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2019:PMa

[Ano19m] Anonymous. Pages 1-342 (March 2019). Journal of Statistical Planning and Inference, 199(??):1–342, March 2019. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2019:P

[Ano19n] Anonymous. Pages 1-402 (1986-1987). Journal of Statistical Planning and Inference, 15(??):1–402, ???? 2019. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Aoki:2019:CIF

[Aok19] Satoshi Aoki. Characterizations of indicator functions and contrast representations of fractional factorial designs with multi-level factors. Journal of Statistical Planning and Inference, 203(??):91–105, December 2019. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375819300266.

Afendras:2011:MVI

[AP11] G. Afendras and N. Papadatos. On matrix variance inequalities. Journal of Statistical Planning and Inference, 141(11):3628–3631, November 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375811002217.

Amarioarei:2014:ATD

[AP14a] Alexandru Amarioarei and Cristian Preda. Approximations for two-dimensional discrete scan statistics in some blockfactor type dependent models. Journal of Statistical Planning and Inference, 151–152(??):107–120, August/September 2014. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375814000767.

Autin:2014:MRB

[AP14b] Florent Autin and Christophe Pouet. Minimax rates over Besov spaces in ill-conditioned mixture-models with varying mixing-weights. Journal of Statistical Planning and Inference, 146(??):20–30, March 2014. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375813002322.

Adekpedjou:2010:EER

[APQ10] Akim Adekpedjou, Edsel A. Peña, and Jonathan Quiton. Estimation and efficiency with recurrent event data under informative monitoring. Journal of Statistical Planning and Inference, 140(3):597–615, March 2010. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837580900247X.

Agostinelli:2011:LD

[AR11] Claudio Agostinelli and Mario Romanazzi. Local depth. Journal of Statistical Planning and Inference, 141(2):817–830, February 2011. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375810003691.

Arnab:2010:RRT

[AS10a] Raghunath Arnab and Sarjinder Singh. Randomized response techniques: an application to the Botswana AIDS impact survey. Journal of Statistical Planning and Inference, 140(4):941–953, April 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375809003097.

Arnold:2010:SPU

[AS10b] Bernhard F. Arnold and Peter Stahlecker. A surprising property of uniformly best linear affine estimation in linear regression when prior information is fuzzy. Journal of Statistical Planning and Inference, 140(4):954–960, April 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809003085.

Abramowicz:2011:SAR

[AS11a] K. Abramowicz and O. Seleznjev. Spline approximation of a random process with singularity. Journal of Statistical Planning and Inference, 141(3):1333–1342, March 2011. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810004751.

Apputhurai:2011:AUE

[AS11b] P. Apputhurai and A. G. Stephenson. Accounting for uncertainty in extremal dependence modeling using Bayesian model averaging techniques. Journal of Statistical Planning and Inference, 141(5):1800–1807, May 2011. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810005380.

Adekpedjou:2015:GCS

[AS15] Akim Adekpedjou and Russell Stocker. A general class of semiparametric models for recurrent event data. Journal of Statistical Planning and Inference, 156(??):48–63, January 2015. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375814001396.

Aoki:2010:MCM

[AT10a] Satoshi Aoki and Akimichi Takemura. Markov chain Monte Carlo tests for designed experiments. Journal of Statistical Planning and Inference, 140(3):817–830, March 2010. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809002912.

Arashi:2010:NCS

[AT10b] M. Arashi and S. M. M. Tabatabaey. A note on classical Stein-type estimators in elliptically contoured models. Journal of Statistical Planning and Inference, 140(5):1206–1213, May 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375809003553.

Antal:2011:SRS

[AT11] Erika Antal and Yves Tillé. Simple random sampling with over-replacement. Journal of Statistical Planning and Inference, 141(1):597–601, January 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375810003228.

Abanto-Valle:2011:SVM

[AVML11] C. A. Abanto-Valle, H. S. Migon, and V. H. Lachos. Stochastic volatility in mean models with scale mixtures of normal distributions and correlated errors: a Bayesian approach. Journal of Statistical Planning and Inference, 141(5):1875– 1887, May 2011. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375810005392.

Amei:2019:IGF

[AX19] Amei Amei and Jianbo Xu. Inference of genetic forces using a Poisson random field model with non-constant population size. Journal of Statistical Planning and Inference, 203(??):57–69, December 2019. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375818302428.

Alvarez:2012:MEI

[ÁY12] Enrique E. Álvarez and Víctor J. Yohai. M-estimators for isotonic regression. Journal of Statistical Planning and Inference, 142(8):2351–2368, August 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375812000985.

Ayano:2012:RCN

[Aya12] Takanori Ayano. Rates of convergence for the k-nearest neighbor estimators with smoother regression functions. Journal of Statistical Planning and Inference, 142(9):2530– 2536, September 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375812001280.

Antognini:2010:EOD

[AZ10] Alessandro Baldi Antognini and Maroussa Zagoraiou. Exact optimal designs for computer experiments via kriging metamodelling. Journal of Statistical Planning and Inference, 140(9):2607–2617, September 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375810001400.

Atalay:2013:CSF

[AZ13] Sevcan Demir Atalay and Melis Zeybek. Circular success and failure runs in a sequence of exchangeable binary trials. Journal of Statistical Planning and Inference, 143(3):621– 629, March 2013. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375812002649.

Azriel:2015:PEE

[Azr15] David Azriel. Power efficiency of Efron’s biased coin design. Journal of Statistical Planning and Inference, 159(??):15–27, April 2015. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375814001918.

Balakrishnan:2011:RRM

[BAHS11] N. Balakrishnan, E. K. Al-Hussaini, and H. M. Saleh. Recurrence relations for moments of progressively censored order statistics from logistic distribution with applications to inference. Journal of Statistical Planning and Inference, 141(1):17–30, January 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S037837581000282X.

Bandyopadhyay:2012:NED

[Ban12] Soutir Bandyopadhyay. A note on efficient density estimators of convolutions. Journal of Statistical Planning and Inference, 142(11):3056–3060, November 2012. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375812001620.

Bao:2018:GRE

[Bao18] Yong Bao. A general result on the estimation bias of ARMA models. Journal of Statistical Planning and Inference, 197(??):107–125, December 2018. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S037837581830003X.

Bartroff:2012:NCE

[Bar12] Jay Bartroff. A new characterization of Elfving’s method for high dimensional computation. Journal of Statistical Planning and Inference, 142(4):863–871, April 2012. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811003983.

Baran:2017:ODP

[Bar17] Sándor Baran. K-optimal designs for parameters of shifted Ornstein–Uhlenbeck processes and sheets. Journal of Statistical Planning and Inference, 186(??):28–41, July 2017. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375817300162.

Bassetti:2011:QCB

[Bas11] Federico Bassetti. Quantitative comparisons between finitary posterior distributions and Bayesian posterior distributions. Journal of Statistical Planning and Inference, 141(2):787–799, February 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375810003745.

Beder:2014:GWP

[BB14] Jay H. Beder and Jesse S. Beder. Generalized wordlength patterns and strength. Journal of Statistical Planning and Inference, 144(??):41–46, January 2014. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375812002285.

Bandyopadhyay:2011:NRA

[BBB11] Uttam Bandyopadhyay, Atanu Biswas, and Rahul Bhattacharya. A new response-adaptive design for continuous treatment responses for phase III clinical trials. Journal of Statistical Planning and Inference, 141(7):2256–2265, July 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375811000231.

Barczy:2019:ABM

[BBKP19] Mátyás Barczy, Mohamed Ben Alaya, Ahmed Kebaier, and Gyula Pap. Asymptotic behavior of maximum likelihood estimators for a jump-type Heston model. Journal of Statistical Planning and Inference, 198(??):139–164, January 2019. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837581830048X.

Bamberg:2014:UDP

[BBPW14] John Bamberg, Anton Betten, Cheryl E. Praeger, and Alfred Wassermann. Unitals in the Desarguesian projective plane of order 16. Journal of Statistical Planning and Inference, 144(??):110–122, January 2014. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375812003217.

Bush:2010:ODS

[BBS10] Stephen Bush, Leonie Burgess, and Deborah Street. Optimal designs for stated choice experiments that incorporate ties. Journal of Statistical Planning and Inference, 140(7):1712– 1718, July 2010. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375809004066.

Beaujean:2011:TSW

[BC11a] Frederik Beaujean and Allen Caldwell. A test statistic for weighted runs. Journal of Statistical Planning and Inference, 141(11):3437–3446, November 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375811001935.

Breidt:2011:IVE

[BC11b] F. Jay Breidt and G. Chauvet. Improved variance estimation for balanced samples drawn via the cube method. Journal of Statistical Planning and Inference, 141(1):479–487, January 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375810003150.

Bouzebda:2012:TSB

[BC12] Salim Bouzebda and Mohamed Cherfi. Test of symmetry based on copula function. Journal of Statistical Planning and Inference, 142(5):1262–1271, May 2012. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811004538.

Bickel:2013:DSV

[BC13] Peter J. Bickel and Mu Cai. Discussion of Sara van de Geer: Generic chaining and the ℓ_{1} penalty. Journal of Statistical Planning and Inference, 143(6):1013–1018, June 2013. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375812003679. See [vdG13b].

Brandes:2019:SAL

[BC19] Dirk-Philip Brandes and Imma Valentina Curato. On the sample autocovariance of a Lévy driven moving average process when sampled at a renewal sequence. Journal of Statistical Planning and Inference, 203(??):20–38, December 2019. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375819300151.

Biernacki:2010:EMC

[BCG10] C. Biernacki, G. Celeux, and G. Govaert. Exact and Monte Carlo calculations of integrated likelihoods for the latent class model. Journal of Statistical Planning and Inference, 140(11):2991–3002, November 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375810001631.

Brunel:2013:NES

[BCG13] E. Brunel, F. Comte, and A. Guilloux. Nonparametric estimation for survival data with censoring indicators missing at random. Journal of Statistical Planning and Inference, 143(10):1653–1671, October 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S037837581300092X.

Bates:2014:ODS

[BCMAW14] Ron A. Bates, Peter R. Curtis, Hugo Maruri-Aguilar, and Henry P. Wynn. Optimal design for smooth supersaturated models. Journal of Statistical Planning and Inference, 154(??):3–11, November 2014. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375813003005.

Barranco-Chamorro:2011:IAT

[BCMRMP11] I. Barranco-Chamorro, J. L. Moreno-Rebollo, and J. M. Muñoz-Pichardo. Influence analysis in truncated distributions. Journal of Statistical Planning and Inference, 141(5):1862–1874, May 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375810005409.

Bayer:2013:BCB

[BCN13] Fábio M. Bayer and Francisco Cribari-Neto. Bartlett corrections in beta regression models. Journal of Statistical Planning and Inference, 143(3):531–547, March 2013. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375812002868.

Bloznelis:2019:CVM

[BCZ19] Daumantas Bloznelis, Gerda Claeskens, and Jing Zhou. Composite versus model-averaged quantile regression. Journal of Statistical Planning and Inference, 200(??):32–46, May 2019. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375818302568.

Brodsky:2012:SCU

[BD12] Boris Brodsky and Boris Darkhovsky. Structural changes and unit roots in non-stationary time series. Journal of Statistical Planning and Inference, 142(1):327–335, January 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811003065.

Burghaus:2014:ODN

[BD14] Ina Burghaus and Holger Dette. Optimal designs for nonlinear regression models with respect to non-informative priors. Journal of Statistical Planning and Inference, 154(??):12–25, November 2014. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375814000950.

Bourguin:2018:NAT

[BD18] Solesne Bourguin and Claudio Durastanti. On normal approximations for the two-sample problem on multidimensional tori. Journal of Statistical Planning and Inference, 196(??):56–69, August 2018. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375817301933.

Brutti:2014:PMC

[BDG14] Pierpaolo Brutti, Fulvio De Santis, and Stefania Gubbiotti. Predictive measures of the conflict between frequentist and Bayesian estimators. Journal of Statistical Planning and Inference, 148(??):111–122, May 2014. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837581300311X.

Balakrishnan:2010:SCA

[BDKM10] N. Balakrishnan, K. F. Davies, J. P. Keating, and R. L. Mason. Simultaneous closeness among order statistics to population quantiles. Journal of Statistical Planning and Inference, 140(9):2408–2415, September 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375810000868.

Barczy:2014:PES

[BDLP14] Mátyás Barczy, Leif Dor̈ing, Zenghu Li, and Gyula Pap. Parameter estimation for a subcritical affine two factor model. Journal of Statistical Planning and Inference, 151–152(??):37–59, August/September 2014. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375814000561.

Becheri:2015:URT

[BDvdA15] I. Gaia Becheri, Feike C. Drost, and Ramon van den Akker. Unit root tests for cross-sectionally dependent panels: the influence of observed factors. Journal of Statistical Planning and Inference, 160(??):11–22, May 2015. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375814001943.

Bodnar:2014:ADR

[BE14] Olha Bodnar and Clemens Elster. Analytical derivation of the reference prior by sequential maximization of Shannon’s mutual information in the multi-group parameter case. Journal of Statistical Planning and Inference, 147(??):106–116, April 2014. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375813002802.

Bedbur:2010:UTB

[Bed10] S. Bedbur. UMPU tests based on sequential order statistics. Journal of Statistical Planning and Inference, 140(9):2520–2530, September 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375810001345.

Beh:2010:ECR

[Beh10] Eric J. Beh. Elliptical confidence regions for simple correspondence analysis. Journal of Statistical Planning and Inference, 140(9):2582–2588, September 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S037837581000131X.

Beh:2012:SCA

[Beh12] Eric J. Beh. Simple correspondence analysis using adjusted residuals. Journal of Statistical Planning and Inference, 142(4):965–973, April 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375811004095.

Benjamin:2010:CTC

[BEJS10] Arthur T. Benjamin, Larry Ericksen, Pallavi Jayawant, and Mark Shattuck. Combinatorial trigonometry with Chebyshev polynomials. Journal of Statistical Planning and Inference, 140(8):2157–2160, August 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375810000224.

Beran:2014:DSM

[Ber14] Rudolf Beran. Discussion of “Statistics on Manifolds … ” by R. Bhattacharya and V. Patrangenaru. Journal of Statistical Planning and Inference, 145(??):23–24, February 2014. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837581300181X. See [BP14b].

Benedetti:2017:MBV

[BET17] Roberto Benedetti, Giuseppe Espa, and Emanuele Taufer. Model-based variance estimation in non-measurable spatial designs. Journal of Statistical Planning and Inference, 181(??):52–61, February 2017. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375816301070.

Beutner:2010:NMC

[Beu10] Eric Beutner. Nonparametric model checking for k-out-of-n systems. Journal of Statistical Planning and Inference, 140(3):626–639, March 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375809002547.

Bradley:2012:TES

[BF12] Jonathan R. Bradley and David L. Farnsworth. Testing for empty sets when misclassifications may be present. Journal of Statistical Planning and Inference, 142(7):2166–2178, July 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375811004605.

Bouriga:2013:ECM

[BF13] M. Bouriga and O. Féron. Estimation of covariance matrices based on hierarchical inverse-Wishart priors. Journal of Statistical Planning and Inference, 143(4):795–808, April 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837581200300X.

Blanco-Fernandez:2012:CSL

[BFCGR12] Angela Blanco-Fernández, Ana Colubi, and Gil González-Rodríguez. Confidence sets in a linear regression model for interval data. Journal of Statistical Planning and Inference, 142(6):1320–1329, June 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S037837581100454X.

Bott:2017:ETD

[BFKK17] A.-K. Bott, T. Felber, M. Kohler, and L. Kristl. Estimation of a time-dependent density. Journal of Statistical Planning and Inference, 180(??):81–107, January 2017. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375816300878.

Bacallado:2015:BNI

[BFT15] Sergio Bacallado, Stefano Favaro, and Lorenzo Trippa. Bayesian nonparametric inference for shared species richness in multiple populations. Journal of Statistical Planning and Inference, 166(??):14–23, November 2015. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375814001001.

Bai:2010:WEL

[BFZ10] Yang Bai, Wing Kam Fung, and Zhongyi Zhu. Weighted empirical likelihood for generalized linear models with longitudinal data. Journal of Statistical Planning and Inference, 140(11):3446–3456, November 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375810002442.

Buratti:2010:DCD

[BG10] Marco Buratti and Dina Ghinelli. On disjoint (3t, 3, 1) cyclic difference families. Journal of Statistical Planning and Inference, 140(7):1918–1922, July 2010. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810000625.

Balakrishnan:2011:SMM

[BGK^{+}11] N. Balakrishnan, Ramesh C. Gupta, Debasis Kundu, Víctor Leiva, and Antonio Sanhueza. On some mixture models based on the Birnbaum-Saunders distribution and associated inference. Journal of Statistical Planning and Inference, 141(7):2175–2190, July 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S037837581000563X.

Berardi:2011:POQ

[BGR11] Luigia Berardi, Mario Gionfriddo, and Rosaria Rota. Perfect octagon quadrangle systems with upper C _{4}-systems. Journal of Statistical Planning and Inference, 141(7):2249–2255, July 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375811000322.

Beamonte:2010:RBI

[BGS10] Asunción Beamonte, Pilar Gargallo, and Manuel Salvador. Robust Bayesian inference in STAR models with neighbourhood effects. Journal of Statistical Planning and Inference, 140(11):3047–3057, November 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375810001722.

Bacro:2016:FDM

[BGT16] Jean-Noel Bacro, Carlo Gaetan, and Gwladys Toulemonde. A flexible dependence model for spatial extremes. Journal of Statistical Planning and Inference, 172(??):36–52, May 2016. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837581500213X.

Brockwell:2010:CPQ

[BH10] Peter J. Brockwell and Jan Hannig. CARMA(p, q) generalized random processes. Journal of Statistical Planning and Inference, 140(12):3613–3618, December 2010. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810002089.

Bucchia:2015:LRV

[BH15] Béatrice Bucchia and Christoph Heuser. Long-run variance estimation for spatial data under change-point alternatives. Journal of Statistical Planning and Inference, 165(??):104–126, October 2015. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375815000774.

Bhattacharjee:2011:POR

[Bha11] Arnab Bhattacharjee. Partial orders with respect to continuous covariates and tests for the proportional hazards model. Journal of Statistical Planning and Inference, 141(1):243–261, January 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375810002910.

Bharath:2014:CCD

[Bha14] Karthik Bharath. On a clustering criterion for dependent observations. Journal of Statistical Planning and Inference, 147(??):84–94, April 2014. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375813002917.

Benkova:2016:PSC

[BHM16] Eva Benková, Radoslav Harman, and Werner G. Müller. Privacy sets for constrained space-filling. Journal of Statistical Planning and Inference, 171(??):1–9, April 2016. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375815301610.

Bansal:2012:BAH

[BHS12] Naveen K. Bansal, G. G. Hamedani, and Ru Sheng. Bayesian analysis of hypothesis testing problems for general population: a Kullback–Leibler alternative. Journal of Statistical Planning and Inference, 142(7):1991–1998, July 2012. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837581200064X.

Bae:2012:LSC

[BI12] Taehan Bae and Ian Iscoe. Large-sample confidence intervals for risk measures of location-scale families. Journal of Statistical Planning and Inference, 142(7):2032–2046, July 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375812000171.

Bindele:2014:ASR

[Bin14] Huybrechts F. Bindele. Asymptotics of the signed-rank estimator under dependent observations. Journal of Statistical Planning and Inference, 146(??):49–55, March 2014. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375813002334.

Birge:2015:ANA

[Bir15] Lucien Birgé. About the non-asymptotic behaviour of Bayes estimators. Journal of Statistical Planning and Inference, 166(??):67–77, November 2015. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375814001232.

Barrientos:2015:BDE

[BJQ15] Andrés F. Barrientos, Alejandro Jara, and Fernando A. Quintana. Bayesian density estimation for compositional data using random Bernstein polynomials. Journal of Statistical Planning and Inference, 166(??):116–125, November 2015. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375815000105.

Bulutoglu:2010:IWG

[BK10] Dursun A. Bulutoglu and David M. Kaziska. Improved WLP and GWP lower bounds based on exact integer programming. Journal of Statistical Planning and Inference, 140(5):1154– 1161, May 2010. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375809003504. See erra-

tum [BK11c].

Bertin:2011:MPB

[BK11a] Karine Bertin and Nicolas Klutchnikoff. Minimax properties of beta kernel estimators. Journal of Statistical Planning and Inference, 141(7):2287–2297, July 2011. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811000267.

Bobotas:2011:IES

[BK11b] Panayiotis Bobotas and Stavros Kourouklis. Improved estimation of the scale parameter, the hazard rate parameter and the ratio of the scale parameters in exponential distributions: an integrated approach. Journal of Statistical Planning and Inference, 141(7):2399–2416, July 2011. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811000425.

Bulutoglu:2011:EIW

[BK11c] Dursun A. Bulutoglu and David M. Kaziska. Erratum to “Improved WLP and GWP lower bounds based on exact integer programming” [J. Statist. Plann. Inference 140 (2010) 1154–1161]. Journal of Statistical Planning and Inference, 141(7):2500–2501, July 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375811000206. See [BK10].

Broniatowski:2012:DDE

[BK12] Michel Broniatowski and Amor Keziou. Divergences and duality for estimation and test under moment condition models. Journal of Statistical Planning and Inference, 142(9):2554–2573, September 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375812001292.

Brennan:2013:HTT

[BK13] Charlotte Brennan and Arnold Knopfmacher. The height of two types of generalised Motzkin paths. Journal of Statistical Planning and Inference, 143(12):2112–2120, December 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375813001651.

Bucher:2016:DMB

[BK16] Axel Bücher and Ivan Kojadinovic. Dependent multiplier bootstraps for non-degenerate U-statistics under mixing conditions with applications. Journal of Statistical Planning and Inference, 170(??):83–105, March 2016. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375815001603.

Bertin:2017:PAE

[BK17] Karine Bertin and Nicolas Klutchnikoff. Pointwise adaptive estimation of the marginal density of a weakly dependent process. Journal of Statistical Planning and Inference, 187(??):115–129, August 2017. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375817300381.

Benhaddou:2014:MDL

[BKPS14] Rida Benhaddou, Rafal Kulik, Marianna Pensky, and Theofanis Sapatinas. Multichannel deconvolution with longrange dependence: a minimax study. Journal of Statistical Planning and Inference, 148(??):1–19, May 2014. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375813003108.

Bandyopadhyay:2010:RBB

[BL10a] S. Bandyopadhyay and S. N. Lahiri. Resampling-based biascorrected time series prediction. Journal of Statistical Planning and Inference, 140(12):3775–3788, December 2010. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810002223.

Berrahou:2010:ANT

[BL10b] Noureddine Berrahou and Djamal Louani. Asymptotic normality of two symmetry test statistics based on the L _{1}-error. Journal of Statistical Planning and Inference, 140(7):1788– 1804, July 2010. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375810000170.

Bailey:2015:NRC

[BL15] R. A. Bailey and A. Lacka. Nested row-column designs for near-factorial experiments with two treatment factors and one control treatment. Journal of Statistical Planning and Inference, 165(??):63–77, October 2015. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375815000750.

Bar-Lev:2010:SOM

[BLBL10] Shaul K. Bar-Lev, Daoud Bshouty, and Zinoviy Landsman. Second order minimax estimation of the mean. Journal of Statistical Planning and Inference, 140(11):3282–3294, November 2010. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375810002363.

Bian:2011:EFA

[BLP11] Jiawen Bian, Hongwei Li, and Huiming Peng. An efficient and fast algorithm for estimating the frequencies of 2-d superimposed exponential signals in presence of zero-mean multiplicative and additive noise. Journal of Statistical Planning and Inference, 141(3):1277–1289, March 2011. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810004398.

Beran:2016:TSI

[BLT16] Jan Beran, Haiyan Liu, and Klaus Telkmann. On two sample inference for eigenspaces in functional data analysis with dependent errors. Journal of Statistical Planning and Inference, 174(??):20–37, July 2016. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375816000239.

Brown:2016:SMP

[BLW16] Lawrence D. Brown, Michael Levine, and Lie Wang. A semiparametric multivariate partially linear model: a difference approach. Journal of Statistical Planning and Inference, 178(??):99–111, November 2016. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375816300623.

Balakrishnan:2011:LIL

[BM11] N. Balakrishnan and Debanjan Mitra. Likelihood inference for lognormal data with left truncation and right censoring with an illustration. Journal of Statistical Planning and Inference, 141(11):3536–3553, November 2011. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811001996.

Blagouchine:2012:CUE

[BM12a] Iaroslav V. Blagouchine and Eric Moreau. Comments on “Unbiased estimates for moments and cumulants in linear regression”. Journal of Statistical Planning and Inference, 142(4):1027–1030, April 2012. CODEN JSPIDN.

ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375811003971.

Browne:2012:MBC

[BM12b] Ryan P. Browne and Paul D. McNicholas. Model-based clustering, classification, and discriminant analysis of data with mixed type. Journal of Statistical Planning and Inference, 142(11):2976–2984, November 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375812001838.

Bose:2015:ODM

[BM15] Mausumi Bose and Rahul Mukerjee. Optimal design measures under asymmetric errors, with application to binary design points. Journal of Statistical Planning and Inference, 159(??):28–36, April 2015. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375814001785.

Bose:2018:NFE

[BM18] Mausumi Bose and Rahul Mukerjee. Near-factorial experiments in nested row-column designs regulating efficiencies. Journal of Statistical Planning and Inference, 193(??):109–116, February 2018. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375817301428.

Brahimi:2013:BRE

[BMNY13] Brahim Brahimi, Djamel Meraghni, Abdelhakim Necir, and Djabrane Yahia. A bias-reduced estimator for the mean of a heavy-tailed distribution with an infinite second moment. Journal of Statistical Planning and Inference, 143(6):1064– 1081, June 2013. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375812003692.

Belzunce:2011:OAR

[BMPR11] Félix Belzunce, Helena Martínez-Puertas, and José M. Ruiz. On optimal allocation of redundant components for series and parallel systems of two dependent components. Journal of Statistical Planning and Inference, 141(9):3094–3104, September 2011. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375811001285.

Bodnar:2019:TPW

[BMPT19] Taras Bodnar, Stepan Mazur, Krzysztof Podgórski, and Joanna Tyrcha. Tangency portfolio weights for singular covariance matrix in small and large dimensions: Estimation and test theory. Journal of Statistical Planning and Inference, 201(??):40–57, July 2019. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375818303458.

Balakrishnan:2015:LVR

[BMW15] N. Balakrishnan, Chunsheng Ma, and Renxiang Wang. Logistic vector random fields with logistic direct and cross covariances. Journal of Statistical Planning and Inference, 161(??):109–118, June 2015. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375815000063.

Banerjee:2016:RPL

[BMW16] Samprit Banerjee, Stefano Monni, and Martin T. Wells. A regularized profile likelihood approach to covariance matrix estimation. Journal of Statistical Planning and Inference, 179(??):36–59, December 2016. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375816300611.

Barakat:2011:LTI

[BO11] H. M. Barakat and A. R. Omar. Limit theorems for intermediate and central order statistics under nonlinear normalization. Journal of Statistical Planning and Inference, 141(1):524–535, January 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375810003289.

Bohm:2010:LPC

[Böh10] W. Böhm. Lattice path counting and the theory of queues. Journal of Statistical Planning and Inference, 140 (8):2168–2183, August 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375810000248.

Bose:2010:ORC

[Bos10] Sudip Bose. Optimally robust credible sets with a class of priors. Journal of Statistical Planning and Inference, 140(10):2918–2925, October 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375810001278.

Boutsikas:2011:AOB

[Bou11] Michael V. Boutsikas. Asymptotically optimal Berry–Esseentype bounds for distributions with an absolutely continuous part. Journal of Statistical Planning and Inference, 141(3):1250–1268, March 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375810004374.

Bohm:2010:CMN

[BOvSS10] Hilmar Böhm, Hernando Ombao, Rainer von Sachs, and Jerome Sanes. Classification of multivariate non-stationary signals: the SLEX-shrinkage approach. Journal of Statistical Planning and Inference, 140(12):3754–3763, December 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S037837581000220X.

Baek:2010:EPH

[BP10a] Changryong Baek and Vladas Pipiras. Estimation of parameters in heavy-tailed distribution when its second order tail parameter is known. Journal of Statistical Planning and Inference, 140(7):1957–1967, July 2010. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810000704.

Baek:2010:RCW

[BP10b] Jong-Il Baek and Sung-Tae Park. RETRACTED: Convergence of weighted sums for arrays of negatively dependent random variables and its applications. Journal of Statistical Planning and Inference, 140(9):2461–2469, September 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810000947. See [BP13a].

Barczy:2010:ABM

[BP10c] M. Barczy and G. Pap. Asymptotic behavior of maximum likelihood estimator for time inhomogeneous diffusion processes. Journal of Statistical Planning and Inference, 140(6):1576–1593, June 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375809003966.

Bradshaw:2010:SLD

[BP10d] David J. Bradshaw and Marianna Pensky. SVM-like decision theoretical classification of high-dimensional vectors. Journal of Statistical Planning and Inference, 140(3):705– 718, March 2010. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375809002821.

Baek:2013:RNC

[BP13a] Jong-Il Baek and Sung-Tae Park. Retraction notice to “Convergence of weighted sums for arrays of negatively dependent random variables and its applications” [J. Statist. Plann. Inference 140(9) (2010) 2461–2469]. Journal of Statistical Planning and Inference, 143(9):1619, September 2013. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375813001377. See [BP10b].

Benhaddou:2013:ANE

[BP13b] Rida Benhaddou and Marianna Pensky. Adaptive nonparametric empirical Bayes estimation via wavelet series: the minimax study. Journal of Statistical Planning and Inference, 143(10):1672–1688, October 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375813001420.

Bhattacharya:2014:RD

[BP14a] Rabi Bhattacharya and Vic Patrangenaru. Rejoinder to the discussion. Journal of Statistical Planning and Inference, 145(??):42–48, February 2014. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/

S0378375813001857. See [BP14b, Ber14].

Bhattacharya:2014:SML

[BP14b] Rabi Bhattacharya and Vic Patrangenaru. Statistics on manifolds and landmarks based image analysis: a nonparametric theory with applications. Journal of Statistical Planning and Inference, 145(??):1–22, February 2014. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375813001808. See discussion [Ber14]

and rejoinder [BP14a].

Bickel:2014:PFF

[BP14c] David R. Bickel and Marta Padilla. A prior-free framework of coherent inference and its derivation of simple shrinkage estimators. Journal of Statistical Planning and Inference, 145(??):204–221, February 2014. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375813001912.

Benke:2015:AIS

[BP15] János Marcell Benke and Gyula Pap. Asymptotic inference for a stochastic differential equation with uniformly distributed time delay. Journal of Statistical Planning and Inference, 167(??):182–192, December 2015. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375815000944.

Baek:2019:RNC

[BP19] Jong-Il Baek and Sung-Tae Park. Retraction notice to “Convergence of weighted sums for arrays of negatively dependent random variables and its applications” [J. Statist. Plann.

Inference 140(9) (2010) 2461–2469]. Journal of Statistical Planning and Inference, 201(??):146, July 2019. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375818303422.

Berg:2010:BTT

[BPP10] Arthur Berg, Efstathios Paparoditis, and Dimitris N. Politis. A bootstrap test for time series linearity. Journal of Statistical Planning and Inference, 140(12):3841–3857, December 2010. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375810002272.

Barsotti:2016:HTM

[BPR16] Flavia Barsotti, Anne Philippe, and Paul Rochet. Hypothesis testing for Markovian models with random time observations. Journal of Statistical Planning and Inference, 173(??):87–98, June 2016. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375815002189.

Benhaddou:2019:AFL

[BPR19] Rida Benhaddou, Marianna Pensky, and Rasika Rajapakshage. Anisotropic functional Laplace deconvolution. Journal of Statistical Planning and Inference, 199(??):271–285, March 2019. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375818301472.

Bouezmarni:2010:NDE

[BR10] Taoufik Bouezmarni and Jeroen V. K. Rombouts. Nonparametric density estimation for multivariate bounded data. Journal of Statistical Planning and Inference, 140(1):139–152, January 1, 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375809002249.

Bailey:2014:DAE

[BR14] R. A. Bailey and Julia Reiss. Design and analysis of experiments testing for biodiversity effects in ecology. Journal of Statistical Planning and Inference, 144(??):69–80, January 2014. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375812003023.

Bravo:2010:EEA

[Bra10] Francesco Bravo. Efficient M-estimators with auxiliary information. Journal of Statistical Planning and Inference, 140(11):3326–3342, November 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375810002338.

Borges:2012:CCW

[BRB12] P. Borges, J. Rodrigues, and N. Balakrishnan. A class of correlated weighted Poisson processes. Journal of Statistical Planning and Inference, 142(1):366–375, January 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375811003132. See corrigen-

dum [BRB18].

Borges:2018:CCC

[BRB18] P. Borges, J. Rodrigues, and N. Balakrishnan. Corrigendum to “A class of correlated weighted Poisson processes” [J. Statist. Plann. Inference 142 (1) (2012) 366–375]. Journal of Statistical Planning and Inference, 197(??):180, December 2018. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375817302161. See [BRB12].

Branson:2019:NBM

[BRBM19] Zach Branson, Maxime Rischard, Luke Bornn, and Luke W. Miratrix. A nonparametric Bayesian methodology for regression discontinuity designs. Journal of Statistical Planning and Inference, 202(??):14–30, September 2019. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375819300035.

Brouste:2010:APM

[Bro10] Alexandre Brouste. Asymptotic properties of MLE for partially observed fractional diffusion system with dependent noises. Journal of Statistical Planning and Inference, 140(2):551–558, February 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375809002493.

Buhlmann:2013:CVR

[BRvdGZ13a] Peter Bühlmann, Philipp Rütimann, Sara van de Geer, and Cun-Hui Zhang. Correlated variables in regression: Clustering and sparse estimation. Journal of Statistical Planning and Inference, 143(11):1835–1858, November 2013. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375813001225. See discussion [BW13,

RS13b, SS13a].

Buhlmann:2013:R

[BRvdGZ13b] Peter Bühlmann, Philipp Rütimann, Sara van de Geer, and Cun-Hui Zhang. Rejoinder. Journal of Statistical Planning and Inference, 143(11):1869–1871, November 2013. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375813001213.

Bansal:2010:BDT

[BS10a] Naveen K. Bansal and Ru Sheng. Bayesian decision theoretic approach to hypothesis problems with skewed alternatives. Journal of Statistical Planning and Inference, 140(10):2894–2903, October 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375810001254.

Basu:2010:JMI

[BS10b] Sanjib Basu and Sourav Santra. A joint model for incomplete data in crossover trials. Journal of Statistical Planning and Inference, 140(10):2839–2845, October 2010. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810001187.

Brillinger:2010:SMP

[BS10c] David R. Brillinger and Brent S. Stewart. Stochastic modeling of particle movement with application to marine biology and oceanography. Journal of Statistical Planning and Inference, 140(12):3597–3607, December 2010. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810002065.

Beh:2011:IEP

[BS11a] Eric J. Beh and Biagio Simonetti. Investigating the European perception of food using moments obtained from nonsymmetrical correspondence analysis. Journal of Statistical Planning and Inference, 141(8):2953–2960, August 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811001182.

Bodnar:2011:CCM

[BS11b] Olha Bodnar and Wolfgang Schmid. CUSUM charts for monitoring the mean of a multivariate Gaussian process. Journal of Statistical Planning and Inference, 141(6):2055–2070, June 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375810005781.

Bartroff:2014:STM

[BS14] Jay Bartroff and Jinlin Song. Sequential tests of multiple hypotheses controlling type I and II familywise error rates. Journal of Statistical Planning and Inference, 153(??):100–114, October 2014. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375814000962.

Bird:2016:OAO

[BS16] E. M. Bird and Deborah J. Street. D-optimal asymmetric orthogonal array plus p run designs. Journal of Statistical Planning and Inference, 170(??):64–76, March 2016. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375815001561.

Biswas:2016:SED

[BSG16] Munmun Biswas, Soham Sarkar, and Anil K. Ghosh. On some exact distribution-free tests of independence between two random vectors of arbitrary dimensions. Journal of Statistical Planning and Inference, 175(??):78–86, August 2016. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375816000392.

Bailey:2013:NRO

[BSH13] R. A. Bailey, Katharina Schiffl, and Ralf-Dieter Hilgers. A note on robustness of D-optimal block designs for twocolour microarray experiments. Journal of Statistical Planning and Inference, 143(7):1195–1202, July 2013. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375813000220.

Baran:2018:ODC

[BSKS18] Sándor Baran, Csilla Szák-Kocsis, and Milan Stehlík. Doptimal designs for complex Ornstein–Uhlenbeck processes. Journal of Statistical Planning and Inference, 197(??):93–106, December 2018. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375818300028.

Ballin:2010:ECT

[BSV10] Marco Ballin, Mauro Scanu, and Paola Vicard. Estimation of contingency tables in complex survey sampling using probabilistic expert systems. Journal of Statistical Planning and Inference, 140(6):1501–1512, June 2010. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809003899.

Belitser:2015:ROB

[BSvZ15] Eduard Belitser, Paulo Serra, and Harry van Zanten. Rate-optimal Bayesian intensity smoothing for inhomogeneous Poisson processes. Journal of Statistical Planning and Inference, 166(??):24–35, November 2015. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375814000998.

Babu:2016:GFT

[BT16] G. Jogesh Babu and Andrea Toreti. A goodness-of-fit test for heavy tailed distributions with unknown parameters and its application to simulated precipitation extremes in the Euro-Mediterranean region. Journal of Statistical Planning and Inference, 174(??):11–19, July 2016. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375816000240.

Barati:2019:FSD

[BT19a] Fahimeh Barati and Hooshang Talebi. Families of search designs for simultaneous detection of the active main and 2-factor interaction effects. Journal of Statistical Planning and Inference, 203(??):151–159, December 2019. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375819300291.

Barati:2019:NEM

[BT19b] Fahimeh Barati and Hooshang Talebi. A new efficient MEP.1 series for 2f factorial experiments. Journal of Statistical Planning and Inference, 198(??):91–104, January 2019. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375818300338.

Balakrishnan:2010:SNP

[BTKN10] N. Balakrishnan, R. C. Tripathi, N. Kannan, and H. K. T. Ng. Some nonparametric precedence-type tests based on progressively censored samples and evaluation of power. Journal of Statistical Planning and Inference, 140(2):559–573, February 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S037837580900250X.

Benkamra:2015:NSO

[BTT15] Zohra Benkamra, Mekki Terbeche, and Mounir Tlemcani. Nearly second order three-stage design for estimating a product of several Bernoulli proportions. Journal of Statistical Planning and Inference, 167(??):90–101, December 2015.

CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375815001111.

Broniatowski:2012:DPA

[BTV12] Michel Broniatowski, Aida Toma, and Igor Vajda. Decomposable pseudodistances and applications in statistical estimation. Journal of Statistical Planning and Inference, 142(9):2574–2585, September 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375812001358.

Bao:2010:EQF

[BU10] Yong Bao and Aman Ullah. Expectation of quadratic forms in normal and nonnormal variables with applications. Journal of Statistical Planning and Inference, 140(5):1193–1205, May 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375809003565.

Bucchia:2014:TEC

[Buc14] Béatrice Bucchia. Testing for epidemic changes in the mean of a multiparameter stochastic process. Journal of Statistical Planning and Inference, 150(??):124–141, July 2014. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375814000366.

Burch:2011:CIV

[Bur11a] Brent D. Burch. Confidence intervals for variance components in unbalanced one-way random effects model using non-normal distributions. Journal of Statistical Planning and Inference, 141(12):3793–3807, December 2011. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811002400.

Burke:2011:NES

[Bur11b] M. D. Burke. Nonparametric estimation of a survival function under progressive Type-I multistage censoring. Journal of Statistical Planning and Inference, 141(2):910–923, February 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375810004040.

Babanezhad:2010:CCE

[BVG10] Manoochehr Babanezhad, Stijn Vansteelandt, and Els Goetghebeur. Comparison of causal effect estimators under exposure misclassification. Journal of Statistical Planning and Inference, 140(5):1306–1319, May 2010. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809003693.

Benjamin:2010:CCC

[BW10a] Arthur T. Benjamin and Daniel Walton. Combinatorially composing Chebyshev polynomials. Journal of Statistical Planning and Inference, 140(8):2161–2167, August 2010. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810000236.

Bissiri:2010:BLB

[BW10b] Pier Giovanni Bissiri and Stephen G. Walker. On Bayesian learning from Bernoulli observations. Journal of Statistical Planning and Inference, 140(11):3520–3530, November 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810002703.

Bissiri:2012:BLL

[BW12] Pier Giovanni Bissiri and Stephen G. Walker. On Bayesian learning via loss functions. Journal of Statistical Planning and Inference, 142(12):3167–3173, December 2012. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375812002091.

Bien:2013:DCV

[BW13] Jacob Bien and Marten Wegkamp. Discussion of “Correlated variables in regression: Clustering and sparse estimation”. Journal of Statistical Planning and Inference, 143(11):1859–1862, November 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375813001237. See [BRvdGZ13a].

Beirlant:2019:EEV

[BWW19] Jan Beirlant, Julien Worms, and Rym Worms. Estimation of the extreme value index in a censorship framework: Asymptotic and finite sample behavior. Journal of Statistical Planning and Inference, 202(??):31–56, September 2019. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375819300047.

Bo:2011:MLE

[BWYZ11] Lijun Bo, Yongjin Wang, Xuewei Yang, and Guannan Zhang. Maximum likelihood estimation for reflected Ornstein–Uhlenbeck processes. Journal of Statistical Planning and Inference, 141(1):588–596, January 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375810003241.

Benavoli:2012:MPI

[BZ12] Alessio Benavoli and Marco Zaffalon. A model of prior ignorance for inferences in the one-parameter exponential family. Journal of Statistical Planning and Inference, 142(7):1960– 1979, July 2012. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375812000535.

Costa:2010:PES

[CA10] Marco Costa and Teresa Alpuim. Parameter estimation of state space models for univariate observations. Journal of Statistical Planning and Inference, 140(7):1889–1902, July 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375810000479.

Cabras:2010:NMT

[Cab10] Stefano Cabras. A note on multiple testing for composite null hypotheses. Journal of Statistical Planning and Inference, 140(3):659–666, March 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375809002596.

Cavicchioli:2018:MAC

[Cav18] Maddalena Cavicchioli. On mixture autoregressive conditional heteroskedasticity. Journal of Statistical Planning and Inference, 197(??):35–50, December 2018. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375817302148.

Clairon:2019:TPS

[CB19] Quentin Clairon and Nicolas J.-B. Brunel. Tracking for parameter and state estimation in possibly misspecified partially observed linear ordinary differential equations. Journal of Statistical Planning and Inference, 199(??):188–206, March 2019. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375818301095.

Chauvet:2011:OIP

[CBD11] G. Chauvet, D. Bonnéry, and J.-C. Deville. Optimal inclusion probabilities for balanced sampling. Journal of Statistical Planning and Inference, 141(2):984–994, February 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810004167.

Chen:2011:MST

[CC11] Chun-Shu Chen and Yu-Mei Chang. Model selection for two-sample problems with right-censored data: an application of Cox model. Journal of Statistical Planning and Inference, 141(6):2120–2127, June 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375810005811.

Cao:2012:RLE

[CC12a] Ying Cao and Di-Rong Chen. On the regularized Laplacian eigenmaps. Journal of Statistical Planning and Inference, 142(7):1627–1643, July 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375812000699.

Chen:2012:CEE

[CC12b] Jing Chen and Pallavi Chitturi. Choice experiments for estimating main effects and interactions. Journal of Statistical Planning and Inference, 142(2):390–396, February 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811003417.

Chen:2015:REC

[CC15] Zhengjia Chen and Xinjia Chen. Rigorous error control methods for estimating means of bounded random variables. Journal of Statistical Planning and Inference, 157–158(??):54–76, February/March 2015. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375814001578.

Chen:2018:SPO

[CC18] Heng Chen and Di-Rong Chen. Statistical performance of optimal scoring in reproducing kernel Hilbert spaces. Journal of Statistical Planning and Inference, 194(??):122–135, March 2018. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375817301581.

Cheng:2019:ERG

[CC19] Ching-Shui Cheng and Gregory M. Constantine. On the efficiency of regular generalized line graph designs. Journal of Statistical Planning and Inference, 15(??):1–10, ???? 2019. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0378375886900807.

Champion:2014:SRS

[CCAGV14] Magali Champion, Christine Cierco-Ayrolles, Sébastien Gadat, and Matthieu Vignes. Sparse regression and support recovery with L_{2}-boosting algorithms. Journal of Statistical Planning and Inference, 155(??):19–41, December 2014. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375814001207.

Cardot:2010:PDB

[CCGL10] Hervé Cardot, Mohamed Chaouch, Camelia Goga, and Catherine Labruère. Properties of design-based functional principal components analysis. Journal of Statistical Planning and Inference, 140(1):75–91, January 1, 2010. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809001967.

Chen:2011:RMP

[CCR11] Vivian Yi-Ju Chen, Vernon M. Chinchilli, and Donald St. P. Richards. Robustness and monotonicity properties of generalized correlation coefficients. Journal of Statistical Planning and Inference, 141(2):924–936, February 2011. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810004118.

Chagny:2017:AEH

[CCR17] G. Chagny, F. Comte, and A. Roche. Adaptive estimation of the hazard rate with multiplicative censoring. Journal of Statistical Planning and Inference, 184(??):25–47, May 2017. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375816301495.

Crnkovic:2014:STC

[CCS14] Dean Crnković, Vedrana Mikulić Crnković, and Andrea Svob. On some transitive combinatorial structures constructed from the unitary group U(3, 3). Journal of Statistical Planning and Inference, 144(??):19–40, January 2014. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375812001966.

Collier:2015:CRN

[CD15] Olivier Collier and Arnak S. Dalalyan. Curve registration by nonparametric goodness-of-fit testing. Journal of Statistical Planning and Inference, 162(??):20–42, July 2015. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375815000129.

CantoeCastro:2011:LMS

[CDd11] Luísa Canto e Castro, Sandra Dias, and Maria da Graça Temido. Looking for max-semistability: a new test for the extreme value condition. Journal of Statistical Planning and Inference, 141(9):3005–3020, September 2011. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811001170.

Cutillo:2014:WDE

[CDNS14] Luisa Cutillo, Italia De Feis, Christina Nikolaidou, and Theofanis Sapatinas. Wavelet density estimation for weighted data. Journal of Statistical Planning and Inference, 146(??): 1–19, March 2014. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375813002504.

Cheng:2018:OSD

[CDST18] Ching-Shui Cheng, Ashish Das, Rakhi Singh, and Pi-Wen Tsai. E(s^{2})- and UE(s^{2})-optimal supersaturated designs. Journal of Statistical Planning and Inference, 196(??):105– 114, August 2018. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375817301969.

Chen:2011:TFV

[CDZ11] William Y. C. Chen, Donna Q. J. Dou, and Terence Y. J. Zhang. On three and four vicious walkers. Journal of Statistical Planning and Inference, 141(1):94–101, January 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810002831.

Cramer:2011:AOP

[CE11] E. Cramer and M. Ensenbach. Asymptotically optimal progressive censoring plans based on Fisher information. Journal of Statistical Planning and Inference, 141(5):1968–1980, May 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375810005677.

Cha:2012:IBT

[CF12a] Ji Hwan Cha and Maxim Finkelstein. Informationbased thinning of point processes and its application to shock models. Journal of Statistical Planning and Inference, 142(8):2345–2350, August 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375812000961.

Cha:2012:NCS

[CF12b] Ji Hwan Cha and Maxim Finkelstein. A note on ‘curable’ shock processes. Journal of Statistical Planning and Inference, 142(12):3146–3151, December 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375812002339.

Carbon:2010:ANF

[CFT10] Michel Carbon, Christian Francq, and Lanh Tat Tran. Asymptotic normality of frequency polygons for random fields. Journal of Statistical Planning and Inference, 140(2):502–514, February 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375809002444.

Cui:2014:DTN

[CFW14] Yunwei Cui, Thomas J. Fisher, and Rongning Wu. Diagnostic tests for non-causal time series with infinite variance. Journal of Statistical Planning and Inference, 147(??):117– 131, April 2014. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375813002760.

Caeiro:2011:SPT

[CG11] Frederico Caeiro and M. Ivette Gomes. Semi-parametric tail inference through probability-weighted moments. Journal of Statistical Planning and Inference, 141(2):937–950, February 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375810004106.

Coeurjolly:2014:CEF

[CG14] Jean-François Coeurjolly and Yongtao Guan. Covariance of empirical functionals for inhomogeneous spatial point processes when the intensity has a parametric form. Journal of Statistical Planning and Inference, 155(??):79–92, December 2014. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375814001177.

Comte:2012:CPK

[CGC12] F. Comte and V. Genon-Catalot. Convolution power kernels for density estimation. Journal of Statistical Planning and Inference, 142(7):1698–1715, July 2012. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375812000857.

Chen:2010:EIE

[CGCL10] Cathy W. S. Chen, Richard H. Gerlach, S. T. Boris Choy, and Celine Lin. Estimation and inference for exponential smooth transition nonlinear volatility models. Journal of Statistical Planning and Inference, 140(3):719–733, March 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809002833.

Chen:2011:DSB

[CGL11] Cathy W. S. Chen, Richard Gerlach, and Feng-Chi Liu. Detection of structural breaks in a time-varying heteroskedastic regression model. Journal of Statistical Planning and Inference, 141(11):3367–3381, November 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375811002072.

Chen:2018:ROA

[CGLW18] Xue-Ping Chen, Bing Guo, Min-Qian Liu, and Xiao-Lei Wang. Robustness of orthogonal-array based composite designs to missing data. Journal of Statistical Planning and Inference, 194(??):15–24, March 2018. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S037837581730188X.

Chiang:2010:NPE

[CH10] Chin-Tsang Chiang and Hung Hung. Non-parametric estimation for time-dependent AUC. Journal of Statistical Planning and Inference, 140(5):1162–1174, May 2010. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809003498.

Chen:2011:ICS

[CH11a] Chun-Shu Chen and Hsin-Cheng Huang. An improved C_{p} criterion for spline smoothing. Journal of Statistical Planning and Inference, 141(1):445–452, January 2011. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810003101.

Chen:2011:QPM

[CH11b] Rui Chen and Ollivier Hyrien. Quasi- and pseudo-maximum likelihood estimators for discretely observed continuous-time Markov branching processes. Journal of Statistical Planning and Inference, 141(7):2209–2227, July 2011. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811000334.

Coretto:2011:MLE

[CH11c] Pietro Coretto and Christian Hennig. Maximum likelihood estimation of heterogeneous mixtures of Gaussian and uniform distributions. Journal of Statistical Planning and Inference, 141(1):462–473, January 2011. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810003174.

Crambes:2019:RIF

[CH19] Christophe Crambes and Yousri Henchiri. Regression imputation in the functional linear model with missing values in the response. Journal of Statistical Planning and Inference, 201(??):103–119, July 2019. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375818303586.

Charalambides:2010:DDB

[Cha10a] Ch. A. Charalambides. Discrete q-distributions on Bernoulli trials with a geometrically varying success probability. Journal of Statistical Planning and Inference, 140(9):2355–2383, September 2010. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375810001370.

Charalambides:2010:BBB

[Cha10b] Charalambos A. Charalambides. The q-Bernstein basis as a q-binomial distribution. Journal of Statistical Planning and Inference, 140(8):2184–2190, August 2010. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837581000025X.

Charalambides:2012:PUM

[Cha12] Ch. A. Charalambides. A q-Pólya urn model and the q-Pólya and inverse q-Pólya distributions. Journal of Statistical Planning and Inference, 142(1):276–288, January 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811003004.

Cha:2013:SCG

[Cha13] Ji Hwan Cha. Stochastic comparison of generalized combined risk processes. Journal of Statistical Planning and Inference, 143(4):818–826, April 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375812003278.

Chang:2019:ATL

[Cha19a] Ming-Chung Chang. De-aliasing in two-level factorial designs: a Bayesian approach. Journal of Statistical Planning and Inference, 203(??):82–90, December 2019. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375819300254.

Chaudhuri:2019:RRS

[Cha19b] Arijit Chaudhuri. Randomized response surveys of finite populations: a unified approach with quantitative data. Journal of Statistical Planning and Inference, 15(??):157–165, ???? 2019. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375886900947.

Chen:2010:SBG

[Che10a] Lee-Shen Chen. Selecting the best geometric distribution based on type-I censored data: a Bayesian approach. Journal of Statistical Planning and Inference, 140(5):1101–1109, May 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375809003474.

Chen:2010:EEM

[Che10b] Willa W. Chen. Efficiency in estimation of memory. Journal of Statistical Planning and Inference, 140(12):3820–3840, December 2010. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375810002260.

Chen:2011:ECM

[Che11a] Xinjia Chen. Exact computation of minimum sample size for estimation of binomial parameters. Journal of Statistical Planning and Inference, 141(8):2622–2632, August 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811000723.

Cheng:2011:VEN

[Che11b] Fuxia Cheng. Variance estimation in nonlinear autoregressive time series models. Journal of Statistical Planning and Inference, 141(4):1588–1592, April 2011. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810005100.

Cherfi:2012:DDE

[Che12] Mohamed Cherfi. Dual divergences estimation for censored survival data. Journal of Statistical Planning and Inference, 142(7):1746–1756, July 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375812000997.

Coetzer:2011:CCD

[CHF11] R. L. J. Coetzer, L. M. Haines, and L. P. Fatti. Central composite designs for estimating the optimum conditions for a second-order model. Journal of Statistical Planning and Inference, 141(5):1764–1773, May 2011. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810005264.

Chan:2013:DEL

[CHH13] Julian Chan, Lajos Horváth, and Marie Husková. Darling– Erdős limit results for change-point detection in panel data. Journal of Statistical Planning and Inference, 143(5):955– 970, May 2013. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S037837581200345X.

Cunen:2018:CDC

[CHH18] Céline Cunen, Gudmund Hermansen, and Nils Lid Hjort. Confidence distributions for change-points and regime shifts. Journal of Statistical Planning and Inference, 195(??):14–34, May 2018. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375817301660.

Collomb:2019:NPE

[CHH19] G. Collomb, W. Härdle, and S. Hassani. A note on prediction via estimation of the conditional mode function. Journal of Statistical Planning and Inference, 15(??):227–236, ???? 2019. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375886900996.

Choi:2017:EEN

[Cho17] In Choi. Efficient estimation of nonstationary factor models. Journal of Statistical Planning and Inference, 183(??):18–43, April 2017. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375816301264.

Coolen:2011:CJS

[CHP11] Frank P. A. Coolen, Brett Houlding, and Steven G. Parkinson. Considerations on jury size and composition using lower probabilities. Journal of Statistical Planning and Inference, 141(1):382–391, January 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375810002934.

Christofides:2010:CMC

[Chr10] Tasos C. Christofides. Comments on a method of comparison of randomized response techniques. Journal of Statistical Planning and Inference, 140(2):574–575, February 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809002328.

Croux:2013:REO

[CHR13] C. Croux, G. Haesbroeck, and C. Ruwet. Robust estimation for ordinal regression. Journal of Statistical Planning and Inference, 143(9):1486–1499, September 2013. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375813000906.

Chrzanowski:2014:WEL

[Chr14] Michał Chrzanowski. Weighted empirical likelihood inference for the area under the ROC curve. Journal of Statistical Planning and Inference, 147(??):159–172, April 2014. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375813002978.

Chen:2017:DAQ

[CHT17] Cathy W. S. Chen, Yi-Tung Hsu, and Masanobu Taniguchi. Discriminant analysis by quantile regression with application on the climate change problem. Journal of Statistical Planning and Inference, 187(??):17–27, August 2017. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375817300150.

Ciuperca:2013:TTS

[Ciu13] Gabriela Ciuperca. Two tests for sequential detection of a change-point in a nonlinear model. Journal of Statistical Planning and Inference, 143(10):1719–1743, October 2013.

CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375813001134.

Cizek:2016:GMT

[Cíz16] Pavel Cízek. Generalized method of trimmed moments. Journal of Statistical Planning and Inference, 171(??):63–78, April 2016. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375815002037.

Chalikias:2012:ENC

[CK12] Miltiadis Chalikias and Stratis Kounias. Extension and necessity of Cheng and Wu conditions. Journal of Statistical Planning and Inference, 142(7):1794–1800, July 2012. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375812000717.

Caner:2013:AUR

[CK13] Mehmet Caner and Keith Knight. An alternative to unit root tests: Bridge estimators differentiate between nonstationary versus stationary models and select optimal lag. Journal of Statistical Planning and Inference, 143(4):691–715, April 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375812003242.

Chopra:2019:MPT

[CKG19] D. V. Chopra, W. A. K. Kipngeno, and S. Ghosh. More precise tables of optimal balanced 2^{m} fractional factorial designs of Srivastava and Chopra, 7 ≤ m ≤ 10. Journal of Statistical Planning and Inference, 15(??):115–121, ???? 2019. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0378375886900893.

Chatterjee:2011:NLB

[CKM11] K. Chatterjee, C. Koukouvinos, and K. Mylona. A new lower bound to A _{2}-optimality measure for multi-level and mixedlevel column balanced designs and its applications. Journal of Statistical Planning and Inference, 141(2):877–888, February 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375810004076.

Chatterjee:2012:GCO

[CKMS12] K. Chatterjee, C. Koukouvinos, P. Mantas, and A. Skountzou. A general construction of E(f_{NOD})-optimal multilevel supersaturated designs. Journal of Statistical Planning and Inference, 142(5):1092–1107, May 2012. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811004198.

Cheng:2017:OED

[CKP17] Ching-Shui Cheng, Ming-Hung Kao, and Federick Kin Hing Phoa. Optimal and efficient designs for functional brain imaging experiments. Journal of Statistical Planning and Inference, 181(??):71–80, February 2017. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375816301094.

Chen:2010:LRM

[CL10a] Hong Chen and Luoqing Li. Learning rates of multikernel regularized regression. Journal of Statistical Planning and Inference, 140(9):2562–2568, September 2010. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810001333.

Cramer:2010:APT

[CL10b] E. Cramer and U. Lenz. Association of progressively Type-II censored order statistics. Journal of Statistical Planning and Inference, 140(2):576–583, February 2010. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837580900233X.

Chen:2011:EVS

[CL11a] Min Chen and Johan Lim. Estimating variances of strata in ranked set sampling. Journal of Statistical Planning and Inference, 141(8):2513–2518, August 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375811000796.

Cojbasic:2011:OSC

[CL11b] Vesna Cojbasic and Dragan Loncar. One-sided confidence intervals for population variances of skewed distributions. Journal of Statistical Planning and Inference, 141(5):1667–1672, May 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375810005070.

Chang:2013:ODC

[CL13a] Fu-Chuen Chang and Chin-Han Li. D-optimal designs for combined polynomial and trigonometric regression on a partial circle. Journal of Statistical Planning and Inference, 143(7):1186–1194, July 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S037837581300030X.

Cohen:2013:CLT

[CL13b] Serge Cohen and Alexander Lindner. A central limit theorem for the sample autocorrelations of a Lévy driven continuous time moving average process. Journal of Statistical Planning and Inference, 143(8):1295–1306, August 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375813000670.

Cheng:2015:CCR

[CL15a] Bin Cheng and Shing M. Lee. On the consistency of the continual reassessment method with multiple toxicity constraints. Journal of Statistical Planning and Inference, 164(??):1–9, September 2015. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375815000415.

Cheng:2015:TDD

[CL15b] Yu Cheng and Jialiang Li. Time-dependent diagnostic accuracy analysis with censored outcome and censored predictor. Journal of Statistical Planning and Inference, 156(??):90–102, January 2015. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375814001426.

Clarke:2019:ATD

[Cla19] Brenton R. Clarke. Asymptotic theory for description of regions in which Newton–Raphson iterations converge to location M-estimators. Journal of Statistical Planning and Inference, 15(??):71–85, ???? 2019. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0378375886900868.

Cancho:2012:GBS

[CLB12] Vicente G. Cancho, Francisco Louzada, and Gladys D. C. Barriga. The Geometric Birnbaum–Saunders regression model with cure rate. Journal of Statistical Planning and Inference, 142(4):993–1000, April 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375811004125.

Chen:2013:CCS

[CLFZ13] Jie Chen, Min-Qian Liu, Kai-Tai Fang, and Dong Zhang. A cyclic construction of saturated and supersaturated designs. Journal of Statistical Planning and Inference, 143(12):2121–2127, December 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375813001638.

Chen:2011:AEN

[CLL11] Jia Chen, Degui Li, and Zhengyan Lin. Asymptotic expansion for nonparametric M-estimator in a nonlinear regression model with long-memory errors. Journal of Statistical Planning and Inference, 141(9):3035–3046, September 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811001224.

Chaubey:2012:GKR

[CLL12] Yogendra P. Chaubey, Naâmane Laïb, and Jun Li. Generalized kernel regression estimator for dependent sizebiased data. Journal of Statistical Planning and Inference, 142(3):708–727, March 2012. CODEN JSPIDN.

ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375811003545.

Cabral:2012:BAS

[CLM12] Celso Rômulo Barbosa Cabral, Víctor Hugo Lachos, and Maria Regina Madruga. Bayesian analysis of skew-normal independent linear mixed models with heterogeneity in the random-effects population. Journal of Statistical Planning and Inference, 142(1):181–200, January 2012. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811002771.

Chang:2017:SCB

[CLO17] Chung Chang, Xuejing Lin, and R. Todd Ogden. Simultaneous confidence bands for functional regression models. Journal of Statistical Planning and Inference, 188(??):67–81, September 2017. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S037837581730037X.

Chen:2013:LRM

[CLP13] Hong Chen, Luoqing Li, and Zhibin Pan. Learning rates of multi-kernel regression by orthogonal greedy algorithm. Journal of Statistical Planning and Inference, 143(2):276–282, February 2013. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375812002704.

Chang:2010:DDP

[CM10] In Hong Chang and Rahul Mukerjee. Data-dependent probability matching priors for highest posterior density and equal-tailed two-sided regions based on empiricaltype likelihoods. Journal of Statistical Planning and Inference, 140(9):2589–2595, September 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375810001308.

Castagliola:2011:CCC

[CM11] Philippe Castagliola and Petros E. Maravelakis. A CUSUM control chart for monitoring the variance when parameters are estimated. Journal of Statistical Planning and Inference, 141(4):1463–1478, April 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375810004829.

Coen:2015:RPB

[CM15] Arrigo Coen and Ramsés H. Mena. Ruin probabilities for Bayesian exchangeable claims processes. Journal of Statistical Planning and Inference, 166(??):102–115, November 2015. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375815000099.

Costa:2016:BCK

[CM16] Marco Costa and Magda Monteiro. Bias-correction of Kalman filter estimators associated to a linear state space model with estimated parameters. Journal of Statistical Planning and Inference, 176(??):22–32, September 2016. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375816300179.

Coolen-Maturi:2011:NPS

[CMCSC11] Tahani Coolen-Maturi, Pauline Coolen-Schrijner, and Frank P. A. Coolen. Nonparametric predictive selection with early experiment termination. Journal of Statistical Planning and Inference, 141(4):1403–1421, April 2011. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810004878.

Coolen-Maturi:2012:NPI

[CMCSC12] Tahani Coolen-Maturi, Pauline Coolen-Schrijner, and Frank P. A. Coolen. Nonparametric Predictive Inference for diagnostic accuracy. Journal of Statistical Planning and Inference, 142(5):1141–1150, May 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375811004204.

Castano-Martinez:2013:COW

[CMLBSM13] A. Castaño-Martínez, F. López-Blázquez, and B. Salamanca-Miño. On the convolution order of weak records. Journal of Statistical Planning and Inference, 143(1):107–115, January 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375812002376.

Cubedo:2013:DBR

[CMO13] M. Cubedo, A. Miñarro, and J. M. Oller. A dissimilarity based on relevant population features. Journal of Statistical Planning and Inference, 143(2):346–355, February 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375812002674.

Cohen:2013:NMT

[CMS13] Arthur Cohen, Yingqiu Ma, and Harold Sackrowitz. Nonparametric multiple testing procedures. Journal of Statistical Planning and Inference, 143(10):1753–1765, October 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375813001109.

Cadre:2017:NTC

[CMT17] Benoît Cadre, Gaspar Massiot, and Lionel Truquet. Nonparametric tests for Cox processes. Journal of Statistical Planning and Inference, 184(??):48–61, May 2017. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375816301549.

Cramer:2012:LRD

[CN12] Erhard Cramer and Guido Naehrig. Laplace record data. Journal of Statistical Planning and Inference, 142(7):2179– 2189, July 2012. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375812000262.

Cavanaugh:2010:AVC

[CND10] Joseph E. Cavanaugh, Andrew A. Neath, and Simon L. Davies. An alternate version of the conceptual predictive statistic based on a symmetrized discrepancy measure. Journal of Statistical Planning and Inference, 140(11):3389–3398, November 2010. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375810002387.

Cribari-Neto:2011:SIS

[CNdGAL11] Francisco Cribari-Neto and Maria da Glória A. Lima. A sequence of improved standard errors under heteroskedasticity of unknown form. Journal of Statistical Planning and Inference, 141(11):3617–3627, November 2011. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811002084.

Choi:2017:RLL

[CNL17] Young-Geun Choi, Chi Tim Ng, and Johan Lim. Regularized LRT for large scale covariance matrices: One sample problem. Journal of Statistical Planning and Inference, 180(??):108–123, January 2017. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375816300635.

Conti:2010:BEC

[CO10] Stefano Conti and Anthony O’Hagan. Bayesian emulation of complex multi-output and dynamic computer models. Journal of Statistical Planning and Inference, 140(3):640–651, March 2010. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375809002559.

Coad:2014:CCI

[Coa14] D. S. Coad. Corrected confidence intervals based on the signed root transformation for multi-parameter sequentially designed experiments. Journal of Statistical Planning and Inference, 147(??):173–187, April 2014. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837581300253X.

Cancho:2010:EID

[COP10] Vicente G. Cancho, Edwin M. M. Ortega, and Gilberto A. Paula. On estimation and influence diagnostics for log-Birnbaum–Saunders Student-t regression models: Full Bayesian analysis. Journal of Statistical Planning and Inference, 140(9):2486–2496, September 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S037837581000090X.

Cope:2011:PLE

[Cop11] Eric W. Cope. Penalized likelihood estimators for truncated data. Journal of Statistical Planning and Inference, 141(1):345–358, January 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375810002892.

Chang:2012:IEC

[COS12] Yuan-Tsung Chang, Youhei Oono, and Nobuo Shinozaki. Improved estimators for the common mean and ordered means of two normal distributions with ordered variances. Journal of Statistical Planning and Inference, 142(9):2619–2628, September 2012. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S037837581200122X.

Choudhury:2012:PSM

[CP12] Kingshuk Roy Choudhury and Catharine Pettigrew. Parametric scalp mapping and inference via spatially smooth linear models for mismatch negativity studies. Journal of Statistical Planning and Inference, 142(1):12–24, January 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811001637.

Chhikara:2019:ESV

[CP19] Raj S. Chhikara and Charles R. Perry. Estimation of stratum variances in planning of crop acreage surveys. Journal of Statistical Planning and Inference, 15(??):97–114, ???? 2019. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0378375886900881.

Cao:2019:TSB

[CPH19] Ming-Xiang Cao, Junyong Park, and Dao-Jiang He. A test for the k sample Behrens–Fisher problem in high dimensional data. Journal of Statistical Planning and Inference, 201(??):

86–102, July 2019. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375818303562.

Cuevas:2018:PVE

[CPL18] Antonio Cuevas and Beatriz Pateiro-López. Polynomial volume estimation and its applications. Journal of Statistical Planning and Inference, 196(??):174–184, August 2018. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375817302100.

Chen:2014:RMR

[CPM14] Lianfu Chen, Mohsen Pourahmadi, and Mehdi Maadooliat. Regularized multivariate regression models with skew-t error distributions. Journal of Statistical Planning and Inference, 149(??):125–139, June 2014. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375814000081.

Charlier:2015:CQE

[CPS15] Isabelle Charlier, Davy Paindaveine, and Jérôme Saracco. Conditional quantile estimation through optimal quantization. Journal of Statistical Planning and Inference, 156(??):14–30, January 2015. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375814001438.

Cheng:2010:VRK

[CPSY10] Ming-Yen Cheng, Robert L. Paige, Shan Sun, and Ke Yan. Variance reduction for kernel estimators in clustered/ longitudinal data analysis. Journal of Statistical Planning and Inference, 140(6):1389–1397, June 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375809003929.

Chatterjee:2011:GDD

[CQ11] Kashinath Chatterjee and Hong Qin. Generalized discrete discrepancy and its applications in experimental designs. Journal of Statistical Planning and Inference, 141(2):

951–960, February 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S037837581000409X.

Celisse:2010:CVB

[CR10] Alain Celisse and Stéphane Robin. A cross-validation based estimation of the proportion of true null hypotheses. Journal of Statistical Planning and Inference, 140(11):3132–3147, November 2010. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375810001813.

Chlapinski:2013:PIT

[CR13] Grzegorz Chlapiński and Roman Rózański. Prediction intervals for time series models with trend via sieve bootstrap. Journal of Statistical Planning and Inference, 143(2):221–236, February 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375812002819.

Choi:2015:NBF

[CR15] Taeryon Choi and Judith Rousseau. A note on Bayes factor consistency in partial linear models. Journal of Statistical Planning and Inference, 166(??):158–170, November 2015. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375815000737.

Chung:2016:AVE

[CR16a] EunYi Chung and Joseph P. Romano. Asymptotically valid and exact permutation tests based on two-sample U-statistics. Journal of Statistical Planning and Inference, 168(??):97–105, January 2016. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S037837581500138X.

Comte:2016:NWE

[CR16b] Fabienne Comte and Tabea Rebafka. Nonparametric weighted estimators for biased data. Journal of Statistical Planning and Inference, 174(??):104–128, July 2016. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375816000185.

Chang:2013:ILB

[CRW13] Fang Chang, Marie Rekkas, and Augustine Wong. Improved likelihood-based inference for the MA(1) model. Journal of Statistical Planning and Inference, 143(1):209–219, January 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375812002443.

Chipperfield:2011:ESQ

[CS11a] James O. Chipperfield and David G. Steel. Efficiency of split questionnaire surveys. Journal of Statistical Planning and Inference, 141(5):1925–1932, May 2011. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810005458.

Crowder:2011:IMC

[CS11b] Martin Crowder and David Stephens. On inference from Markov chain macro-data using transforms. Journal of Statistical Planning and Inference, 141(9):3201–3216, September 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375811001509.

Chuang:2012:ADW

[CS12a] Li-Ling Chuang and Yu-Shan Shih. Approximated distributions of the weighted sum of correlated chi-squared random variables. Journal of Statistical Planning and Inference, 142(2):457–472, February 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375811003156.

Costantini:2012:ADS

[CS12b] Mauro Costantini and Amit Sen. On the asymptotic distribution of a simple unit root test for trending and breaking series. Journal of Statistical Planning and Inference, 142(7):1690– 1697, July 2012. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375812000833.

Cheng:2013:ESD

[CS13] Yi Cheng and Yu Shen. An efficient sequential design of clinical trials. Journal of Statistical Planning and Inference, 143(2):283–295, February 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375812002637.

Chen:2016:ESI

[CS16a] Le-Yu Chen and Jerzy Szroeter. Extension of the Schwarz Information Criterion for models sharing parameter boundaries. Journal of Statistical Planning and Inference, 174(??): 68–84, July 2016. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375816000203.

Ciuperca:2016:ELT

[CS16b] Gabriela Ciuperca and Zahraa Salloum. Empirical likelihood test for high-dimensional two-sample model. Journal of Statistical Planning and Inference, 178(??):37–60, November 2016. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375816300416.

Choudhary:2014:GSM

[CSC14] Pankaj K. Choudhary, Dishari Sengupta, and Phillip Cassey. A general skew-t mixed model that allows different degrees of freedom for random effects and error distributions. Journal of Statistical Planning and Inference, 147(??):235–247, April 2014. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375813003017.

Cheng:2017:MGG

[CSK17] Lulu Cheng, Liang Shan, and Inyoung Kim. Multilevel Gaussian graphical model for multilevel networks. Journal of Statistical Planning and Inference, 190(??):1–14, November 2017. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375817300836.

Costa:2019:GAU

[CSN19] Lilia Costa, James Q. Smith, and Thomas Nichols. A group analysis using the Multiregression Dynamic Models for fMRI networked time series. Journal of Statistical Planning and Inference, 198(??):43–61, January 2019. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375818300302.

Cao:2015:UHI

[CSR15] Yongtao Cao, Byran J. Smucker, and Timothy J. Robinson. On using the hypervolume indicator to compare Pareto fronts: Applications to multi-criteria optimal experimental design. Journal of Statistical Planning and Inference, 160(??): 60–74, May 2015. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375814002006.

Cheng:2011:ALN

[CSW11] Fuxia Cheng, Shuxia Sun, and Miin-Jye Wen. Asymptotics for L2-norm of ARCH innovation density estimator. Journal of Statistical Planning and Inference, 141(12):3771–3779, December 2011. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375811002382.

Colbourn:2011:MCA

[CSWY11] Charles J. Colbourn, Ce Shi, Chengmin Wang, and Jie Yan. Mixed covering arrays of strength three with few factors. Journal of Statistical Planning and Inference, 141(11):3640–3647, November 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375811002230.

Cadigan:2010:ENB

[CT10a] N. G. Cadigan and Jared Tobin. Estimating the negative binomial dispersion parameter with highly stratified surveys. Journal of Statistical Planning and Inference, 140(7):2138– 2147, July 2010. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S037837581000087X.

Chan:2010:NIN

[CT10b] Kung-Sik Chan and Howell Tong. A note on the invertibility of nonlinear ARMA models. Journal of Statistical Planning and Inference, 140(12):3709–3714, December 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810002168.

Chang:2010:ECM

[CT10c] Changgee Chang and Ruey S. Tsay. Estimation of covariance matrix via the sparse Cholesky factor with lasso. Journal of Statistical Planning and Inference, 140(12):3858–3873, December 2010. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375810002284.

Chien:2011:DDG

[CT11] Li-Chu Chien and Tsung-Shan Tsou. Deletion diagnostics for generalized linear models using the adjusted Poisson likelihood function. Journal of Statistical Planning and Inference, 141(6):2044–2054, June 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375810005744.

Chang:2014:LTO

[CT14] Fu-Chuen Chang and Jhong-Shin Tsai. A limit theorem of Doptimal designs for weighted polynomial regression. Journal of Statistical Planning and Inference, 154(??):26–38, November 2014. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375814000743.

Cuevas:2014:POT

[Cue14] Antonio Cuevas. A partial overview of the theory of statistics with functional data. Journal of Statistical Planning and Inference, 147(??):1–23, April 2014. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375813000748.

Cahoy:2010:PEF

[CUW10] Dexter O. Cahoy, Vladimir V. Uchaikin, and Wojbor A. Woyczynski. Parameter estimation for fractional Poisson processes. Journal of Statistical Planning and Inference, 140(11):3106–3120, November 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375810001837.

Chang:2010:ILB

[CW10] F. Chang and A. C. M. Wong. Improved likelihood-based inference for the stationary AR(2) model. Journal of Statistical Planning and Inference, 140(7):2099–2110, July 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810000789.

Chen:2017:LWL

[CW17] Kun Chen and Man Wang. Local Whittle likelihood estimators and tests for spatial lattice data. Journal of Statistical Planning and Inference, 191(??):25–42, December 2017. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375817301076.

Cai:2019:IGP

[CW19] Quan Cai and Suojin Wang. Inferences with generalized partially linear single-index models for longitudinal data. Journal of Statistical Planning and Inference, 200(??):146–160, May 2019. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375818303008.

Chen:2011:SEP

[CWcIC11] Shih-Chia Chen, Charlotte Wang, and Yuan chin I. Chang. Sequential estimate of partial area under ROC curve with β-protection and optimal ratio of cases to controls. Journal of Statistical Planning and Inference, 141(10):3356–3366, October 2011. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375811001649.

Chen:2010:CCI

[CWH10] Shen-Chien Chen, Ruby C. Weng, and Tzeeming Huang. Corrected confidence intervals for parameters in adaptive linear models. Journal of Statistical Planning and Inference, 140(1):297–309, January 1, 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375809002262.

Cao:2012:SCB

[CWWT12] Guanqun Cao, Jing Wang, Li Wang, and David Todem. Spline confidence bands for functional derivatives. Journal of Statistical Planning and Inference, 142(6):1557–1570, June 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375812000183.

Corander:2013:OVB

[CXCK13] Jukka Corander, Jie Xiong, Yaqiong Cui, and Timo Koski. Optimal Viterbi Bayesian predictive classification for data from finite alphabets. Journal of Statistical Planning and Inference, 143(2):261–275, February 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375812002613.

Cao:2010:SSGa

[CY10] H. Cao and F. Yan. On super-simple group divisible designs with block size four and index λ = 3, 4, 6. Journal of Statistical Planning and Inference, 140(5):1330–1345, May 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809003735.

Chen:2011:EBT

[CY11] Lee-Shen Chen and Ming-Chung Yang. Empirical Bayes testing for equivalence. Journal of Statistical Planning and Inference, 141(8):2670–2681, August 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375811000760.

Cheng:2019:MOO

[CY19] Qianshun Cheng and Min Yang. On multiple-objective optimal designs. Journal of Statistical Planning and Inference, 200(??):87–101, May 2019. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375818302623.

Chen:2011:PMS

[CYC11] Baojiang Chen, Grace Y. Yi, and Richard J. Cook. Progressive multi-state models for informatively incomplete longitudinal data. Journal of Statistical Planning and Inference, 141(1):80–93, January 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S037837581000251X.

Chen:2012:MMC

[CYCZ12] Baojiang Chen, Grace Y. Yi, Richard J. Cook, and Xiao-Hua Zhou. Marginal methods for clustered longitudinal binary data with incomplete covariates. Journal of Statistical Planning and Inference, 142(10):2819–2831, October 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375812001565.

Cao:2010:SSGb

[CYW10] H. Cao, F. Yan, and R. Wei. Super-simple group divisible designs with block size 4 and index 2. Journal of Statistical Planning and Inference, 140(9):2497–2503, September 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375810000935.

Chen:2012:PAE

[CYZL12] Baicheng Chen, Yao Yu, Hui Zou, and Hua Liang. Profiled adaptive Elastic-Net procedure for partially linear models with high-dimensional covariates. Journal of Statistical Planning and Inference, 142(7):1733–1745, July 2012. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375812000821.

Cheng:2010:CGM

[CZ10] Yi Cheng and Runchu Zhang. On construction of general minimum lower order confounding 2^{n−m} designs with N/4 + 1 ≤ n ≤ 9N/32. Journal of Statistical Planning and Inference, 140(9):2384–2394, September 2010. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810000790.

Cantoni:2011:RVH

[CZ11a] Eva Cantoni and Asma Zedini. A robust version of the hurdle model. Journal of Statistical Planning and Inference, 141(3):1214–1223, March 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375810004337.

Chan:2011:QIH

[CZ11b] Ngai Hang Chan and Rong-Mao Zhang. Quantile inference for heteroscedastic regression models. Journal of Statistical Planning and Inference, 141(6):2079–2090, June 2011. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810005768.

Cai:2012:PGA

[CZ12] Nanlian Cai and Yaohui Zheng. Preservation of generalized ageing classes on the residual life at random time. Journal of Statistical Planning and Inference, 142(1):148–154, January 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375811002746.

Cheung:2014:CAR

[CZHC14] Siu Hung Cheung, Li-Xin Zhang, Feifang Hu, and Wai Sum Chan. Covariate-adjusted response-adaptive designs for generalized linear models. Journal of Statistical Planning and Inference, 149(??):152–161, June 2014. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375814000263.

Dachian:2010:LLR

[Dac10] Sergueï Dachian. On limiting likelihood ratio processes of some change-point type statistical models. Journal of Statistical Planning and Inference, 140(9):2682–2692, September 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375810001436.

Daly:2011:SMS

[Dal11] Fraser Daly. On Stein’s method, smoothing estimates in total variation distance and mixture distributions. Journal of Statistical Planning and Inference, 141(7):2228–2237, July 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811000218.

Dafnis:2010:DRK

[DAP10] Spiros D. Dafnis, Demetrios L. Antzoulakos, and Andreas N. Philippou. Distributions related to (k1,k2) events. Journal of Statistical Planning and Inference, 140(7):1691–1700, July 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375809003991.

Dornheim:2011:REF

[DB11] Harald Dornheim and Vytaras Brazauskas. Robust-efficient fitting of mixed linear models: Methodology and theory. Journal of Statistical Planning and Inference, 141(4):1422– 1435, April 2011. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375810004854.

De:2012:SSM

[DB12] Shyamal K. De and Michael Baron. Step-up and stepdown methods for testing multiple hypotheses in sequential experiments. Journal of Statistical Planning and Inference, 142(7):2059–2070, July 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S037837581200050X.

Dykstra:2019:ALS

[DB19] Richard L. Dykstra and James P. Boyle. An algorithm for least squares projections onto the intersection of translated, convex cones. Journal of Statistical Planning and Inference, 15(??):391–399, ???? 2019. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0378375886901114.

Dong:2010:OST

[DC10] Xiaofang Dong and Lirong Cui. Optimal sign test for quantiles in ranked set samples. Journal of Statistical Planning and Inference, 140(11):2943–2951, November 2010. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810002429.

deCarvalho:2012:GSW

[dC12] Miguel de Carvalho. A generalization of the Solis– Wets method. Journal of Statistical Planning and Inference, 142(3):633–644, March 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375811003508.

Deutsch:2012:LHS

[DD12] Jared L. Deutsch and Clayton V. Deutsch. Latin hypercube sampling with multidimensional uniformity. Journal of Statistical Planning and Inference, 142(3):763–772, March 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811003776.

Dutta:2013:ODE

[DD13] Ganesh Dutta and Premadhis Das. Optimum designs for estimation of regression parameters in a balanced treatment incomplete block design set-up. Journal of Statistical Planning and Inference, 143(7):1203–1214, July 2013. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375813000232.

Diallo:2018:AMC

[DDD18] Alpha Oumar Diallo, Aliou Diop, and Jean-François Dupuy. Analysis of multinomial counts with joint zero-inflation, with an application to health economics. Journal of Statistical Planning and Inference, 194(??):85–105, March 2018. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375817301623.

Da:2010:HRO

[DDL10] Gaofeng Da, Weiyong Ding, and Xiaohu Li. On hazard rate ordering of parallel systems with two independent components. Journal of Statistical Planning and Inference, 140(7):2148–2154, July 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375810000832.

Dehay:2014:SCT

[DDL14] Dominique Dehay, Anna Dudek, and Jacek Leśkow. Subsampling for continuous-time almost periodically correlated processes. Journal of Statistical Planning and Inference, 150(??):142–158, July 2014. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375814000342.

Deshpande:2010:FDM

[DDNN10] J. V. Deshpande, Isha Dewan, and U. V. Naik-Nimbalkar. A family of distributions to model load sharing systems. Journal of Statistical Planning and Inference, 140(6):1441–1451, June 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375809003759.

Demirkale:2013:COS

[DDS13] Fatih Demirkale, Diane Donovan, and Deborah J. Street. Constructing D-optimal symmetric stated preference discrete choice experiments. Journal of Statistical Planning and Inference, 143(8):1380–1391, August 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S037837581300044X.

Dai:2016:DRB

[DDW16] Pengjie Dai, Xiaobo Ding, and Qihua Wang. Dimension reduction based linear surrogate variable approach for model free variable selection. Journal of Statistical Planning and Inference, 169(??):13–26, February 2016. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375815001354.

DeBlasi:2013:DAB

[De 13] Pierpaolo De Blasi. Discussion on article “Bayesian inference with misspecified models” by Stephen G. Walker. Journal of Statistical Planning and Inference, 143(10):1634–1637, October 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375813001183. See [Wal13a].

Delgado:2016:DFT

[DE16] Miguel A. Delgado and Juan Carlos Escanciano. Distributionfree tests of conditional moment inequalities. Journal of Statistical Planning and Inference, 173(??):99–108, June 2016. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375815002141.

Dimou:2011:SSV

[DEF11] Spiros Dimou, Antonis Economou, and Demetrios Fakinos. The single server vacation queueing model with geometric abandonments. Journal of Statistical Planning and Inference, 141(8):2863–2877, August 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375811001078.

Dong:2016:TSE

[DELS16] Jun Dong, Jason P. Estes, Gang Li, and Damla Sentürk. A two-step estimation approach for logistic varying coefficient modeling of longitudinal data. Journal of Statistical Planning and Inference, 174(??):38–51, July 2016. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375816000227.

Dembinska:2012:APN

[Dem12] Anna Dembińska. Asymptotic properties of numbers of observations in random regions determined by central order statistics. Journal of Statistical Planning and Inference, 142(2):516–528, February 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S037837581100320X.

Dette:2014:SIM

[Det14] Holger Dette. Special issue in memory of William J. Studden. Journal of Statistical Planning and Inference, 154(??):1–2, November 2014. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375814001256.

Dimitriou-Fakalou:2010:ARM

[DF10] Chrysoula Dimitriou-Fakalou. The auto-regression and the moving-average. Journal of Statistical Planning and Inference, 140(7):1739–1743, July 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375809004029.

Dai:2012:PCT

[DF12a] Hongsheng Dai and Bo Fu. A polar coordinate transformation for estimating bivariate survival functions with randomly censored and truncated data. Journal of Statistical Planning and Inference, 142(1):248–262, January 2012. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811002989.

Ducharme:2012:OTG

[DF12b] Gilles R. Ducharme and Sandie Ferrigno. An omnibus test of goodness-of-fit for conditional distributions with applications to regression models. Journal of Statistical Planning and Inference, 142(10):2748–2761, October 2012. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375812001589.

Delattre:2017:KLE

[DF17] Sylvain Delattre and Nicolas Fournier. On the Kozachenko– Leonenko entropy estimator. Journal of Statistical Planning and Inference, 185(??):69–93, June 2017. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375817300125.

DeBlasi:2010:CNR

[DFM10] Pierpaolo De Blasi, Stefano Favaro, and Pietro Muliere. A class of neutral to the right priors induced by superposition of beta processes. Journal of Statistical Planning and Inference, 140(6):1563–1575, June 2010. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837580900398X.

Diaz-Garcia:2013:DTQ

[DG13a] José A. Díaz-García. Distribution theory of quadratic forms for matrix multivariate elliptical distribution. Journal of Statistical Planning and Inference, 143(8):1330–1342, August 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375813000694.

Diaz-Garcia:2013:NMR

[DG13b] José A. Díaz-García. A note on the moments of the Riesz distribution. Journal of Statistical Planning and Inference, 143(11):1880–1886, November 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375813001663.

Diaconis:2014:IMK

[DG14] Persi Diaconis and Robert Griffiths. An introduction to multivariate Krawtchouk polynomials and their applications. Journal of Statistical Planning and Inference, 154(??):39–53, November 2014. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375814000123.

Du:2019:MGN

[DG19] Xingqi Du and Subhashis Ghosal. Multivariate Gaussian network structure learning. Journal of Statistical Planning and Inference, 199(??):327–342, March 2019. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375818301745.

Delattre:2016:MSD

[DGCS16] Maud Delattre, Valentine Genon-Catalot, and Adeline Samson. Mixtures of stochastic differential equations with random effects: Application to data clustering. Journal of Statistical Planning and Inference, 173(??):109–124, June 2016. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375815302263.

Dauxois:2019:SIE

[DGG19] J.-Y. Dauxois, S. Gasmi, and O. Gaudoin. Semiparametric inference for an extended geometric failure rate reduction model. Journal of Statistical Planning and Inference, 199(??):14–28, March 2019. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375818300508.

Dey:2019:SEP

[DGS19] Dipak K. Dey, Malay Ghosh, and C. Srinivasan. Simultaneous estimation of parameters under entropy loss. Journal of Statistical Planning and Inference, 15(??):347–363, ???? 2019. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375886901084.

Debusho:2011:OPD

[DH11a] Legesse Kassa Debusho and Linda M. Haines. D- and V - optimal population designs for the quadratic regression model with a random intercept term. Journal of Statistical Planning and Inference, 141(2):889–898, February 2011. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810004064.

Demirhan:2011:MLG

[DH11b] Haydar Demirhan and Canan Hamurkaroglu. On a multivariate log-gamma distribution and the use of the distribution in the Bayesian analysis. Journal of Statistical Planning and Inference, 141(3):1141–1152, March 2011. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837581000426X.

Deo:2010:LMI

[DHH10] Rohit Deo, Mengchen Hsieh, and Clifford M. Hurvich. Long memory in intertrade durations, counts and realized volatility of NYSE stocks. Journal of Statistical Planning and Inference, 140(12):3715–3733, December 2010. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837581000217X.

Dhifaoui:2018:SMG

[Dhi18] Zouhaier Dhifaoui. Statistical moments of Gaussian kernel correlation sum and weighted least square estimator of correlation dimension and noise level. Journal of Statistical Planning and Inference, 193(??):55–69, February 2018. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375817301374.

Doosti:2018:NTD

[DHM18] Hassan Doosti, Peter Hall, and Jorge Mateu. Nonparametric tilted density function estimation:a cross-validation criterion. Journal of Statistical Planning and Inference, 197(??):51–68, December 2018. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S037837581730215X.

DiNardo:2016:PSP

[Di 16] E. Di Nardo. On photon statistics parametrized by a non-central Wishart random matrix. Journal of Statistical Planning and Inference, 169(??):1–12, February 2016. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375815001366.

DeGregorio:2010:DTS

[DI10] Alessandro De Gregorio and Stefano M. Iacus. Divergences test statistics for discretely observed diffusion processes. Journal of Statistical Planning and Inference, 140(7):1744– 1753, July 2010. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375810000066.

Diaz:2017:EEQ

[Día17] Iván Díaz. Efficient estimation of quantiles in missing data models. Journal of Statistical Planning and Inference, 190(??):39–51, November 2017. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375817300782.

Doosti:2011:MBT

[DIAT11] H. Doosti, A. Iranmanesh, M. Arashi, and S. M. M. Tabatabaey. On minimaxity of block thresholded wavelets under elliptical symmetry. Journal of Statistical Planning and Inference, 141(4):1526–1534, April 2011. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810005112.

Dickhaus:2013:RVM

[Dic13] Thorsten Dickhaus. Randomized p-values for multiple testing of composite null hypotheses. Journal of Statistical Planning and Inference, 143(11):1968–1979, November 2013. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375813001481.

Dion:2014:NAS

[Dio14] Charlotte Dion. New adaptive strategies for nonparametric estimation in linear mixed models. Journal of Statistical Planning and Inference, 150(??):30–48, July 2014. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375814000421.

Dalgaard:2019:APC

[DJ19] Peter Dalgaard and Søren Johansen. The asymptotic properties of the Cornish–Bowden–Eisenthal median estimator. Journal of Statistical Planning and Inference, 15(??):279–299, ???? 2019. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375886901047.

Duval:2017:NAE

[DK17] C. Duval and J. Kappus. Nonparametric adaptive estimation for grouped data. Journal of Statistical Planning and Inference, 182(??):12–28, March 2017. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375816301240.

Dou:2013:DSA

[DKL13] Xiaoling Dou, Satoshi Kuriki, and Gwo Dong Lin. Dependence structures and asymptotic properties of Baker’s distributions with fixed marginals. Journal of Statistical Planning and Inference, 143(8):1343–1354, August 2013. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375813000645.

Dikta:2016:ARP

[DKMdUÁ16] Gerhard Dikta, René Külheim, Jorge Mendonça, and Jacobo de Uña-Álvarez. Asymptotic representation of presmoothed Kaplan–Meier integrals with covariates in a semiparametric censorship model. Journal of Statistical Planning and Inference, 171(??):10–37, April 2016. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375815002128.

Dewanji:2012:NEN

[DKN12] Anup Dewanji, Subrata Kundu, and Tapan K. Nayak. Nonparametric estimation of the number of components of a superposition of renewal processes. Journal of Statistical Planning and Inference, 142(9):2710–2718, September 2012. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375812001243.

Dumbgen:2017:BLC

[DKW17] Lutz Dümbgen, Petro Kolesnyk, and Ralf A. Wilke. Bilog-concave distribution functions. Journal of Statistical Planning and Inference, 184(??):1–17, May 2017. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375816301379.

DiCiccio:2015:QNP

[DKY15] T. J. DiCiccio, T. A. Kuffner, and G. A. Young. Quantifying nuisance parameter effects via decompositions of asymptotic refinements for likelihood-based statistics. Journal of Statistical Planning and Inference, 165(??):1–12, October 2015. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375815000579.

DiCiccio:2017:FRB

[DKY17] T. J. DiCiccio, T. A. Kuffner, and G. A. Young. The formal relationship between analytic and bootstrap approaches to parametric inference. Journal of Statistical Planning and Inference, 191(??):81–87, December 2017. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375817300873.

Das:2011:ORD

[DL11] Rabindra Nath Das and Dennis K. J. Lin. On D-optimal robust designs for lifetime improvement experiments. Journal of Statistical Planning and Inference, 141(12):3753–3759, December 2011. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375811002369.

Ding:2012:OAA

[DL12] Weiyong Ding and Xiaohu Li. The optimal allocation of active redundancies to k-out-of-n systems with respect to hazard rate ordering. Journal of Statistical Planning and Inference, 142(7):1878–1887, July 2012. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375812000729.

Demarqui:2012:CDP

[DLDC12] Fabio N. Demarqui, Rosangela H. Loschi, Dipak K. Dey, and Enrico A. Colosimo. A class of dynamic piecewise exponential models with random time grid. Journal of Statistical Planning and Inference, 142(3):728–742, March 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811003521.

deLuna:2010:NPI

[dLJ10] Xavier de Luna and Per Johansson. Non-parametric inference for the effect of a treatment on survival times with application in the health and social sciences. Journal of Statistical Planning and Inference, 140(7):2122–2137, July 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810000856. See erratum [dLJ12].

deLuna:2012:ENP

[dLJ12] Xavier de Luna and Per Johansson. Erratum to “Nonparametric inference for the effect of a treatment on survival times with application in the health and social sciences” [Journal of Statistical Planning and Inference 140 (7) (July) (2010) 2122–2137]. Journal of Statistical Planning and Inference, 142(6):1624–1625, June 2012. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375812000134. See [dLJ10].

Dryden:2014:MSP

[DLPW14] Ian L. Dryden, Huiling Le, Simon P. Preston, and Andrew T. A. Wood. Mean shapes, projections and intrinsic limiting distributions. Journal of Statistical Planning and Inference, 145(??):25–32, February 2014. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375813001821.

Ding:2014:GFT

[DLZ14] Xiaobo Ding, Lexin Li, and Lixing Zhu. Goodness-of-fit testing-based selection for large-p–small-n problems: a two-stage ranking approach. Journal of Statistical Planning and Inference, 145(??):148–164, February 2014. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375813002073.

Dette:2010:NAB

[DM10] Holger Dette and Viatcheslav B. Melas. A note on all-bias designs with applications in spline regression models. Journal of Statistical Planning and Inference, 140(7):2037–2045, July 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S037837581000073X.

Delaigle:2011:RON

[DM11] Aurore Delaigle and Alexander Meister. Rate-optimal nonparametric estimation in classical and Berkson errorsin-variables problems. Journal of Statistical Planning and Inference, 141(1):102–114, January 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375810002673.

Denecke:2012:CRT

[DM12] Liesa Denecke and Christine H. Müller. Consistency and robustness of tests and estimators based on depth. Journal of Statistical Planning and Inference, 142(9):2501–2517, September 2012. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375812001486.

Du:2013:VRF

[DM13] Juan Du and Chunsheng Ma. Vector random fields with compactly supported covariance matrix functions. Journal of Statistical Planning and Inference, 143(3):457–467, March 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375812002844.

Das:2014:GMM

[DM14a] I. Das and S. Mukhopadhyay. On generalized multinomial models and joint percentile estimation. Journal of Statistical Planning and Inference, 145(??):190–203, February 2014. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375813002103.

Dey:2014:MGO

[DM14b] Rima Dey and Athanasios C. Micheas. Modeling the growth of objects through a stochastic process of random sets. Journal of Statistical Planning and Inference, 151–152(??):17–36, August/September 2014. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375814000433.

Das:2010:BCL

[DMP10] Ujjwal Das, Tapabrata Maiti, and Vivek Pradhan. Bias correction in logistic regression with missing categorical covariates. Journal of Statistical Planning and Inference, 140(9):2478–2485, September 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375810000911.

DeIaco:2011:SPD

[DMP11] S. De Iaco, D. E. Myers, and D. Posa. On strict positive definiteness of product and product-sum covariance models. Journal of Statistical Planning and Inference, 141(3):1132– 1140, March 2011. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375810004258.

Douma:2018:OEH

[DMP18] Malik Douma, Abdelkader Mokkadem, and Mariane Pelletier. Online estimation of hazard rate under random censoring. Journal of Statistical Planning and Inference, 196(??):87–104, August 2018. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375817301957.

Deutsch:2010:SDPb

[DMR10a] Emeric Deutsch, Emanuele Munarini, and Simone Rinaldi. Skew Dyck paths. Journal of Statistical Planning and Inference, 140(8):2191–2203, August 2010. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810000261.

Deutsch:2010:SDPa

[DMR10b] Emeric Deutsch, Emanuele Munarini, and Simone Rinaldi. Skew Dyck paths, area, and superdiagonal bargraphs. Journal of Statistical Planning and Inference, 140(6):1550–1562, June 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375809003930.

Dette:2011:ODT

[DMS11] Holger Dette, Viatcheslav B. Melas, and Petr Shpilev. Optimal designs for trigonometric regression models. Journal of Statistical Planning and Inference, 141(3):1343–1353, March 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375810004799.

Dai:2015:DBV

[DMTZ15] Wenlin Dai, Yanyuan Ma, Tiejun Tong, and Lixing Zhu. Difference-based variance estimation in nonparametric regression with repeated measurement data. Journal of Statistical Planning and Inference, 163(??):1–20, August 2015. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375815000403.

Davidov:2012:IEH

[DN12] Ori Davidov and Yuval Nov. Improving an estimator of Hsieh and Turnbull for the binormal ROC curve. Journal of Statistical Planning and Inference, 142(4):872–877, April 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837581100396X.

Das:2013:SSO

[DN13] Suchismita Das and Asok K. Nanda. Some stochastic orders of dynamic additive mean residual life model. Journal of Statistical Planning and Inference, 143(2):400–407, February 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375812002686.

deOliveira:2012:BIP

[dOCG12] Maristela Dias de Oliveira, Enrico A. Colosimo, and Gustavo L. Gilardoni. Bayesian inference for power law processes with applications in repairable systems. Journal of Statistical Planning and Inference, 142(5):1151–1160, May 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811004216.

Doring:2010:MCP

[Dör10] Maik Döring. Multiple change-point estimation with U-statistics. Journal of Statistical Planning and Inference, 140(7):2003–2017, July 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375810000649.

Doring:2011:CDM

[Dör11] Maik Döring. Convergence in distribution of multiple change point estimators. Journal of Statistical Planning and Inference, 141(7):2238–2248, July 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375811000292.

Davarzani:2011:SID

[DP11] Nasser Davarzani and Ahmad Parsian. Statistical inference for discrete middle-censored data. Journal of Statistical Planning and Inference, 141(4):1455–1462, April 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810004817.

DeIaco:2013:PNN

[DP13] S. De Iaco and D. Posa. Positive and negative nonseparability for space-time covariance models. Journal of Statistical Planning and Inference, 143(2):378–391, February 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375812002431.

Das:2010:ORS

[DPA10] Rabindra Nath Das, Sung H. Park, and Manohar Aggarwal. On D-optimal robust second order slope-rotatable designs. Journal of Statistical Planning and Inference, 140(5):1269– 1279, May 2010. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375809003668.

Davis:2011:BFS

[DPC11] Justin Davis, Marianna Pensky, and William Crampton. Bayesian feature selection for classification with possibly large number of classes. Journal of Statistical Planning and Inference, 141(9):3256–3266, September 2011. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811001546.

DeIaco:2013:CSC

[DPM13] S. De Iaco, D. Posa, and D. E. Myers. Characteristics of some classes of space-time covariance functions. Journal of Statistical Planning and Inference, 143(11):2002–2015, November 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375813001432.

DiMarzio:2011:KDE

[DPT11] Marco Di Marzio, Agnese Panzera, and Charles C. Taylor. Kernel density estimation on the torus. Journal of Statistical Planning and Inference, 141(6):2156–2173, June 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837581100019X.

DiMarzio:2016:NCQ

[DPT16] Marco Di Marzio, Agnese Panzera, and Charles C. Taylor. Nonparametric circular quantile regression. Journal of Statistical Planning and Inference, 170(??):1–14, March 2016. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375815001548.

Dattner:2013:EDF

[DR13] I. Dattner and B. Reiser. Estimation of distribution functions in measurement error models. Journal of Statistical Planning and Inference, 143(3):479–493, March 2013. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375812002984.

Datta:2010:PEB

[DRT10] Gauri S. Datta, J. N. K. Rao, and Mahmoud Torabi. Pseudo-empirical Bayes estimation of small area means under a nested error linear regression model with functional measurement errors. Journal of Statistical Planning and Inference, 140(11):2952–2962, November 2010. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810001679.

Dang:2010:NDB

[DS10] Xin Dang and Robert Serfling. Nonparametric depthbased multivariate outlier identifiers, and masking robustness properties. Journal of Statistical Planning and Inference, 140(1):198–213, January 1, 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375809002134.

DiNardo:2012:SSS

[DS12] E. Di Nardo and D. Senato. Symbolic solutions of some linear recurrences. Journal of Statistical Planning and Inference, 142(2):423–429, February 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375811003077.

DeBlasi:2018:CDL

[DS18] Pierpaolo De Blasi and Tore Schweder. Confidence distributions from likelihoods by median bias correction. Journal of Statistical Planning and Inference, 195(??):35–46, May 2018. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375817301672.

Dasgupta:2010:CMT

[DST10] Nairanjana Dasgupta, Eleanne Solorzano, and Tiejun Tong. Comparing multiple treatments to both positive and negative controls. Journal of Statistical Planning and Inference, 140(1):180–188, January 1, 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375809002158.

Druilhet:2012:ECN

[DT12] Pierre Druilhet and Walter Tinsson. Efficient circular neighbour designs for spatial interference model. Journal of Statistical Planning and Inference, 142(5):1161–1169, May 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811004228.

Druilhet:2014:OCD

[DT14] Pierre Druilhet and Walter Tinsson. Optimal cross-over designs for total effects under a model with self and mixed carryover effects. Journal of Statistical Planning and Inference, 154(??):54–61, November 2014. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375813001584.

Lopez-de-Ullibarri:2012:CCF

[dUJC12] Ignacio López de Ullibarri, Paul Janssen, and Ricardo Cao. Continuous covariate frailty models for censored and truncated clustered data. Journal of Statistical Planning and Inference, 142(7):1864–1877, July 2012. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375812000912.

Dutta:2011:EMO

[Dut11] Santanu Dutta. Estimation of the MISE and the optimal bandwidth vector of a product kernel density estimate. Journal of Statistical Planning and Inference, 141(5):1817–1831, May 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S037837581000532X.

Dukes:2015:MOL

[DvB15] Peter J. Dukes and Christopher M. van Bommel. Mutually orthogonal latin squares with large holes. Journal of Statistical Planning and Inference, 159(??):81–89, April 2015. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375814001761.

deVilliers:2011:PPC

[dVGS11] J. P. de Villiers, S. J. Godsill, and S. S. Singh. Particle predictive control. Journal of Statistical Planning and Inference, 141(5):1753–1763, May 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375810005252.

Deng:2017:DRE

[DW17] Jianqiu Deng and Qihua Wang. Dimension reduction estimation for probability density with data missing at random when covariables are present. Journal of Statistical Planning and Inference, 181(??):11–29, February 2017. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375816301021.

Dong:2019:MDT

[DY19] Junyi Dong and Qiqing Yu. Marginal distribution test for checking proportional hazards model assumption. Journal of Statistical Planning and Inference, 201(??):58–70, July 2019. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375818301241.

Danthurebandara:2011:ECC

[DYV11] Vishva Manohara Danthurebandara, Jie Yu, and Martina Vandebroek. Effect of choice complexity on design efficiency in conjoint choice experiments. Journal of Statistical Planning and Inference, 141(7):2276–2286, July 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811000255.

Du:2018:LCF

[DZ18] Lilun Du and Changliang Zou. On-line control of false discovery rates for multiple datastreams. Journal of Statistical Planning and Inference, 194(??):1–14, March 2018. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375817301908.

Bermudez:2010:PEGa

[dZBK10a] P. de Zea Bermudez and Samuel Kotz. Parameter estimation of the generalized Pareto distribution — Part I. Journal of Statistical Planning and Inference, 140(6):1353–1373, June 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375809002766.

Bermudez:2010:PEGb

[dZBK10b] P. de Zea Bermudez and Samuel Kotz. Parameter estimation of the generalized Pareto distribution — Part II. Journal of Statistical Planning and Inference, 140(6):1374–1388, June 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375809002778.

Ding:2018:QEH

[DZZ18] Hui Ding, Riquan Zhang, and Jian Zhang. Quantile estimation for a hybrid model of functional and varying coefficient regressions. Journal of Statistical Planning and Inference, 196(??):1–18, August 2018. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375817301891.

Erdbrugge:2010:GLM

[EA10] Martina Erdbrügge and Corinna Auer. Generalized linear models and transformations for unreplicated factorial experiments in the presence of dispersion effects. Journal of Statistical Planning and Inference, 140(1):9–19, January 1, 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837580900278X.

Evangelou:2017:VIC

[EE17] Evangelos Evangelou and Jo Eidsvik. The value of information for correlated GLMs. Journal of Statistical Planning and Inference, 180(??):30–48, January 2017. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837581630091X.

Eggermont:2010:CRS

[EEL10] P. P. B. Eggermont, R. L. Eubank, and V. N. LaRiccia. Convergence rates for smoothing spline estimators in varying coefficient models. Journal of Statistical Planning and Inference, 140(2):369–381, February 2010. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809002754.

Efromovich:2011:NRR

[Efr11] Sam Efromovich. Nonparametric regression with responses missing at random. Journal of Statistical Planning and Inference, 141(12):3744–3752, December 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375811002424.

Efromovich:2013:AND

[Efr13] Sam Efromovich. Adaptive nonparametric density estimation with missing observations. Journal of Statistical Planning and Inference, 143(4):637–650, April 2013. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375812003230.

Einmahl:2016:TCS

[EG16] John H. J. Einmahl and Zhuojiong Gan. Testing for central symmetry. Journal of Statistical Planning and Inference, 169(??):27–33, February 2016. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375815001408.

ElMethni:2012:EEQ

[EGGG12] Jonathan El Methni, Laurent Gardes, Stéphane Girard, and Armelle Guillou. Estimation of extreme quantiles from heavy and light tailed distributions. Journal of Statistical Planning and Inference, 142(10):2735–2747, October 2012. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375812001498.

Eaton:2010:CBD

[EGS10a] Julia Eaton, Anant P. Godbole, and Betsy Sinclair. Competition between discrete random variables, with applications to occupancy problems. Journal of Statistical Planning and Inference, 140(8):2204–2212, August 2010. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810000273.

Einmahl:2010:ASP

[EGS10b] John H. J. Einmahl, Maria Gantner, and Günther Sawitzki. Asymptotics of the shorth plot. Journal of Statistical Planning and Inference, 140(11):3003–3012, November 2010. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837581000162X.

El-Monsef:2011:COD

[EMS11] M. M. E. Abd El-Monsef and M. M. Seyam. CDT-optimum designs for model discrimination, parameter estimation and estimation of a parametric function. Journal of Statistical Planning and Inference, 141(2):639–643, February 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810003575.

Ehler:2012:MPF

[EO12] M. Ehler and K. A. Okoudjou. Minimization of the probabilistic p-frame potential. Journal of Statistical Planning and Inference, 142(3):645–659, March 2012. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811003454.

Elsawah:2015:LBC

[EQ15] A. M. Elsawah and Hong Qin. Lower bound of centered L_{2}- discrepancy for mixed two and three levels U-type designs. Journal of Statistical Planning and Inference, 161(??):1–11, June 2015. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375814002043.

Ericksen:2010:LPC

[Eri10] Larry Ericksen. Lattice path combinatorics for multiple product identities. Journal of Statistical Planning and Inference, 140(8):2213–2226, August 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375810000285.

Eryilmaz:2011:CCS

[Ery11] Serkan Eryilmaz. Circular consecutive k-out-of-n systems with exchangeable dependent components. Journal of Statistical Planning and Inference, 141(2):725–733, February 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810003617.

Eichner:2012:KAN

[ES12] Gerrit Eichner and Winfried Stute. Kernel adjusted nonparametric regression. Journal of Statistical Planning and Inference, 142(9):2537–2544, September 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375812001279.

Ernst:2016:TBM

[ES16] Philip Ernst and Larry Shepp. On the time for Brownian motion to visit every point on a circle. Journal of Statistical Planning and Inference, 171(??):130–134, April 2016. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375815001949.

Esmailzadeh:2011:NSM

[ETMJ11] Nabaz Esmailzadeh, Hooshang Talebi, Koji Momihara, and Masakazu Jimbo. A new series of main effects plus one plan for 2^{m} factorial experiments with m = 4λ±1 and 2m runs. Journal of Statistical Planning and Inference, 141(4):1567– 1574, April 2011. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375810005082.

Egan:2011:IPL

[EW11] Judith Egan and Ian M. Wanless. Indivisible partitions of latin squares. Journal of Statistical Planning and Inference, 141(1):402–417, January 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375810003137.

Evangelou:2012:OPD

[EZ12] Evangelos Evangelou and Zhengyuan Zhu. Optimal predictive design augmentation for spatial generalised linear mixed models. Journal of Statistical Planning and Inference, 142(12):3242–3253, December 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S037837581200198X.

Evangelou:2011:EPS

[EZS11] Evangelos Evangelou, Zhengyuan Zhu, and Richard L. Smith. Estimation and prediction for spatial generalized linear mixed models using high order Laplace approximation. Journal of Statistical Planning and Inference, 141(11):3564–3577, November 2011. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S037837581100200X.

Fan:2010:PDS

[Fan10] Cuiling Fan. Perfect difference systems of sets and the unit group of Z_{p^{n}}. Journal of Statistical Planning and Inference, 140(11):3442–3445, November 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375810002466.

Fuki:2014:HSB

[FBHZ14] Igar Fuki, Lawrence Brown, Xu Han, and Linda Zhao. Hunting for significance: Bayesian classifiers under a mixture loss function. Journal of Statistical Planning and Inference, 154(??):62–71, November 2014. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375814000329.

Ferreira:2010:MRE

[FC10] Marta Ferreira and Luísa Canto e Castro. Modeling rare events through a p RARMAX process. Journal of Statistical Planning and Inference, 140(11):3552–3566, November 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375810002715.

Fu:2012:MSG

[FC12] Ying-Zi Fu and Xue-Dong Chen. Model selection of generalized partially linear models with missing covariates. Journal of Statistical Planning and Inference, 142(1):126–138, January 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375811002722.

Fernandez-Casal:2010:SDE

[FCC10] Rubén Fernández-Casal and Rosa M. Crujeiras. Spatial dependence estimation using FFT of biased covariances. Journal of Statistical Planning and Inference, 140(9):2653–2668, September 2010. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S037837581000145X.

Ferreira:2012:FIB

[FF12] Helena Ferreira and Marta Ferreira. Fragility index of block tailed vectors. Journal of Statistical Planning and Inference, 142(7):1837–1848, July 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375812000444.

Fokianos:2014:RCD

[FGH14] Konstantinos Fokianos, Edit Gombay, and Abdulkadir Hussein. Retrospective change detection for binary time series models. Journal of Statistical Planning and Inference, 145(??):102–112, February 2014. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375813002127.

Fallon:2010:PLP

[FGN10] Joshua Fallon, Shanzhen Gao, and Heinrich Niederhausen. Proof of a lattice paths conjecture connected to the tennis ball problem. Journal of Statistical Planning and Inference, 140(8):2227–2229, August 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375810000297.

Fonseca:2012:NAC

[FGV12] Giovanni Fonseca, Federica Giummolè, and Paolo Vidoni. A note about calibrated prediction regions and distributions. Journal of Statistical Planning and Inference, 142(9):2726–2734, September 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375812001267.

Filova:2011:AOD

[FHK11] Lenka Filová, Radoslav Harman, and Thomas Klein. Approximate E-optimal designs for the model of spring balance weighing with a constant bias. Journal of Statistical Planning and Inference, 141(7):2480–2488, July 2011. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811000681.

Favaro:2011:CDS

[FHP11] Stefano Favaro, Georgia Hadjicharalambous, and Igor Prünster. On a class of distributions on the simplex. Journal of Statistical Planning and Inference, 141(9):2987–3004, September 2011. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375811001121.

Fisher:2012:TIC

[Fis12] Thomas J. Fisher. On testing for an identity covariance matrix when the dimensionality equals or exceeds the sample size. Journal of Statistical Planning and Inference, 142(1):312–326, January 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375811003041.

Fu:2012:AER

[FJC12] James C. Fu, Brad C. Johnson, and Yung-Ming Chang. Approximating the extreme right-hand tail probability for the distribution of the number of patterns in a sequence of multi-state trials. Journal of Statistical Planning and Inference, 142(2):473–480, February 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375811003168.

Fotopoulos:2011:CPM

[FJK11] Stergios B. Fotopoulos, Venkata K. Jandhyala, and Elena A. Khapalova. Change-point mle in the rate of exponential sequences with application to Indonesian seismological data. Journal of Statistical Planning and Inference, 141(1):220–234, January 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S037837581000279X.

Felber:2013:WUC

[FJKW13] Tina Felber, Daniel Jones, Michael Kohler, and Harro Walk. Weakly universally consistent static forecasting of stationary and ergodic time series via local averaging and least squares estimates. Journal of Statistical Planning and Inference, 143(10):1689–1707, October 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375813001390.

Fischer:2011:ETP

[FK11a] T. Fischer and U. Kamps. On the existence of transformations preserving the structure of order statistics in lower dimensions. Journal of Statistical Planning and Inference, 141(1):536–548, January 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375810003216.

Fromkorth:2011:ALS

[FK11b] Andreas Fromkorth and Michael Kohler. Analysis of least squares regression estimates in case of additional errors in the variables. Journal of Statistical Planning and Inference, 141(1):172–188, January 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375810002788.

Farmakis:2019:TNO

[FK19] Nikos Farmakis and Stratis Kounias. Two new D-optimal designs (83, 56, 12), (83, 55, 12). Journal of Statistical Planning and Inference, 15(??):247–257, ???? 2019. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0378375886901011.

Fort:2013:ESI

[FKLL13] Jean-Claude Fort, Thierry Klein, Agnès Lagnoux, and Béatrice Laurent. Estimation of the Sobol indices in a linear functional multidimensional model. Journal of Statistical Planning and Inference, 143(9):1590–1605, September 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837581300089X.

Filippou:2019:TAR

[FKMR19] Panagiota Filippou, Thomas Kneib, Giampiero Marra, and Rosalba Radice. A trivariate additive regression model with arbitrary link functions and varying correlation matrix. Journal of Statistical Planning and Inference, 199(??):236–248, March 2019. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375818301332.

Franco:2011:MEM

[FKV11] Manuel Franco, Debasis Kundu, and Juana-Maria Vivo. Multivariate extension of modified Sarhan–Balakrishnan bivariate distribution. Journal of Statistical Planning and Inference, 141(11):3400–3412, November 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375811001145.

Fang:2017:PEL

[FLL17a] Jianglin Fang, Wanrong Liu, and Xuewen Lu. Penalized empirical likelihood for semiparametric models with a diverging number of parameters. Journal of Statistical Planning and Inference, 186(??):42–57, July 2017. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375817300290.

Figueroa-Lopez:2017:ENS

[FLL17b] José E. Figueroa-López and Kiseop Lee. Estimation of a noisy subordinated Brownian motion via two-scales power variations. Journal of Statistical Planning and Inference, 189(??):16–37, October 2017. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375817300848.

Ferraty:2010:RUC

[FLTV10] Frédéric Ferraty, Ali Laksaci, Amel Tadj, and Philippe Vieu. Rate of uniform consistency for nonparametric estimates with functional variables. Journal of Statistical Planning and Inference, 140(2):335–352, February 2010. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809002316.

Fan:2013:SIP

[FLW13] Guo-Liang Fan, Han-Ying Liang, and Jiang-Feng Wang. Statistical inference for partially time-varying coefficient errors-in-variables models. Journal of Statistical Planning and Inference, 143(3):505–519, March 2013. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375812002856.

Fischer:2019:AUI

[FM19] Aurélie Fischer and Mathilde Mougeot. Aggregation using input-output trade-off. Journal of Statistical Planning and Inference, 200(??):1–19, May 2019. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375818302349.

Fonseca:2010:LSG

[FMZ10] Miguel Fonseca, João Tiago Mexia, and Roman Zmyślony. Least squares and generalized least squares in models with orthogonal block structure. Journal of Statistical Planning and Inference, 140(5):1346–1352, May 2010. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809003772.

Fujimori:2017:CDS

[FN17] Kou Fujimori and Yoichi Nishiyama. The l_{q} consistency of the Dantzig selector for Cox’s proportional hazards model. Journal of Statistical Planning and Inference, 181(??):62–70, February 2017. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375816301082.

Franke:2013:ACM

[FO13] Angelika Franke and Gerhard Osius. The asymptotic covariance matrix of the odds ratio parameter estimator in semiparametric log-bilinear odds ratio models. Journal of Statistical Planning and Inference, 143(1):63–81, January 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375812002066.

Fontana:2011:FPA

[Fon11] Roberto Fontana. Fractions of permutations. An application to Sudoku. Journal of Statistical Planning and Inference, 141(12):3697–3704, December 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375811002266.

Fournier:2011:CPE

[Fou11] N. Fournier. Compromise Pitman estimators. Journal of Statistical Planning and Inference, 141(8):2656–2669, August 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375811000759.

Feng:2012:JEL

[FP12a] Huijun Feng and Liang Peng. Jackknife empirical likelihood tests for distribution functions. Journal of Statistical Planning and Inference, 142(6):1571–1585, June 2012. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375812000195.

Ferreira:2012:PPE

[FP12b] Helena Ferreira and Luísa Pereira. Point processes of exceedances by random fields. Journal of Statistical Planning and Inference, 142(3):773–779, March 2012. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811003739.

Franco-Pereira:2011:PRL

[FPLR11] Alba M. Franco-Pereira, Rosa E. Lillo, and Juan Romo. The percentile residual life up to time t_{0}: Ordering and aging properties. Journal of Statistical Planning and Inference, 141(11):3554–3563, November 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375811002011.

Filipiak:2013:OCB

[FR13] Katarzyna Filipiak and Rafal Rózański. On the E-optimality of complete block designs under a mixed interference model. Journal of Statistical Planning and Inference, 143(3):583– 592, March 2013. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375812002972.

Fraysse:2014:REC

[Fra14] Philippe Fraysse. Recursive estimation in a class of models of deformation. Journal of Statistical Planning and Inference, 147(??):132–158, April 2014. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375813002620.

Franey:2011:BBA

[FRC11] Mark Franey, Pritam Ranjan, and Hugh Chipman. Branch and bound algorithms for maximizing expected improvement functions. Journal of Statistical Planning and Inference, 141(1):42–55, January 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S037837581000248X.

Frederic:2011:MSS

[Fre11a] Patrizio Frederic. Modeling skew-symmetric distributions using B-spline and penalties. Journal of Statistical Planning and Inference, 141(8):2878–2890, August 2011. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837581100108X.

Frey:2011:NRS

[Fre11b] Jesse Frey. A note on ranked-set sampling using a covariate. Journal of Statistical Planning and Inference, 141(2):809–816, February 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375810003708.

Frey:2011:RCI

[Fre11c] Jesse Frey. Recursive computation of inclusion probabilities in ranked-set sampling. Journal of Statistical Planning and Inference, 141(11):3632–3639, November 2011. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811002229.

Frey:2012:PBE

[Fre12] Jesse Frey. Prediction bands for the EDF of a future sample. Journal of Statistical Planning and Inference, 142(2):506–515, February 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375811003193.

Frey:2013:DDN

[Fre13] Jesse Frey. Data-driven nonparametric prediction intervals. Journal of Statistical Planning and Inference, 143(6):1039– 1048, June 2013. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375813000219.

Fremdt:2014:ADD

[Fre14] Stefan Fremdt. Asymptotic distribution of the delay time in page’s sequential procedure. Journal of Statistical Planning and Inference, 145(??):74–91, February 2014. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375813002139.

Fluss:2012:ARC

[FRF12] Ronen Fluss, Benjamin Reiser, and David Faraggi. Adjusting ROC curves for covariates in the presence of verification bias. Journal of Statistical Planning and Inference, 142(1):1–11, January 2012. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375811001133.

Fruth:2014:TII

[FRK14] J. Fruth, O. Roustant, and S. Kuhnt. Total interaction index: a variance-based sensitivity index for secondorder interaction screening. Journal of Statistical Planning and Inference, 147(??):212–223, April 2014. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375813002930.

Frias:2011:CFE

[FRM11] M. P. Frías and M. D. Ruiz-Medina. Computing functional estimators of spatiotemporal long-range dependence parameters in the spectral-wavelet domain. Journal of Statistical Planning and Inference, 141(7):2417–2427, July 2011. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811000577.

Frolov:2014:ETB

[Fro14] Andrei N. Frolov. Esseen type bounds of the remainder in a combinatorial CLT. Journal of Statistical Planning and Inference, 149(??):90–97, June 2014. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375814000068.

Fontana:2014:CSD

[FRR14] Roberto Fontana, Fabio Rapallo, and Maria Piera Rogantin. A characterization of saturated designs for factorial experiments. Journal of Statistical Planning and Inference, 147(??):204–211, April 2014. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375813002772.

Fontana:2016:AQM

[FRR16] Roberto Fontana, Fabio Rapallo, and Maria Piera Rogantin. Aberration in qualitative multilevel designs. Journal of Statistical Planning and Inference, 174(??):1–10, July 2016. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375815300938.

Forman:2014:TAM

[FS14] Julie Lyng Forman and Michael Sørensen. A transformation approach to modelling multi-modal diffusions. Journal of Statistical Planning and Inference, 146(??):56–69, March 2014. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375813002486.

Fakoor:2018:MFI

[FSS18] Vahid Fakoor, Ali Shariati, and Majid Sarmad. The MRL function inference through empirical likelihood in length-biased sampling. Journal of Statistical Planning and Inference, 196(??):115–131, August 2018. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375817301970.

Fabrizi:2010:RLM

[FT10] Enrico Fabrizi and Carlo Trivisano. Robust linear mixed models for small area estimation. Journal of Statistical Planning and Inference, 140(2):433–443, February 2010. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809002304.

Fushiki:2010:BBP

[Fus10] Tadayoshi Fushiki. Bayesian bootstrap prediction. Journal of Statistical Planning and Inference, 140(1):65–74, January 1, 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375809001876.

Franco:2011:RPG

[FVB11] M. Franco, J. M. Vivo, and N. Balakrishnan. Reliability properties of generalized mixtures of Weibull distributions with a common shape parameter. Journal of Statistical Planning and Inference, 141(8):2600–2613, August 2011. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811000693.

Feng:2018:NIS

[FWS18] Yang Feng, Yichao Wu, and Leonard A. Stefanski. Nonparametric independence screening via favored smoothing bandwidth. Journal of Statistical Planning and Inference, 197(??):1–14, December 2018. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375817302112.

Fan:2014:APE

[FYX14] Jun Fan, Ailing Yan, and Naihua Xiu. Asymptotic properties for M-estimators in linear models with dependent random errors. Journal of Statistical Planning and Inference, 148(??): 49–66, May 2014. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375813003078.

Fan:2013:EGS

[FZ13] Yan Fan and Lixing Zhu. Estimation of general semiparametric quantile regression. Journal of Statistical Planning and Inference, 143(5):896–910, May 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375812003461.

Fang:2016:JAL

[FZS16] Sha Fang, Haixiang Zhang, and Liuquan Sun. Joint analysis of longitudinal data with additive mixed effect model for informative observation times. Journal of Statistical Planning and Inference, 169(??):43–55, February 2016. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375815001421.

Zhang:2016:CAO

[fZyZD16] Tian fang Zhang, Yang yang Zong, and Aloke Dey. On the construction of asymmetric orthogonal arrays. Journal of Statistical Planning and Inference, 170(??):77–82, March 2016. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375815001597.

Gupta:2010:P

[GA10] Sat Gupta and Manohar Aggarwal. Preface. Journal of Statistical Planning and Inference, 140(10):2800, October 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810001138.

Gaffke:2019:OEL

[Gaf19] N. Gaffke. On D-optimality of exact linear regression designs with minimum support. Journal of Statistical Planning and Inference, 15(??):189–204, ???? 2019. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0378375886900972.

Gamrot:2013:ECM

[Gam13] Wojciech Gamrot. On exact computation of minimum sample size for restricted estimation of a binomial parameter. Journal of Statistical Planning and Inference, 143(4):852– 866, April 2013. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375812003072.

Gasmi:2011:PEA

[Gas11a] Soufiane Gasmi. Parameter estimation in an alternating repair model. Journal of Statistical Planning and Inference, 141(11):3605–3616, November 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375811002059.

Gassem:2011:CMT

[Gas11b] Anis Gassem. On Cramér-von Mises type test based on local time of switching diffusion process. Journal of Statistical Planning and Inference, 141(4):1355–1361, April 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810005069.

Gadrich:2012:OAO

[GB12] Tamar Gadrich and Emil Bashkansky. ORDANOVA: Analysis of ordinal variation. Journal of Statistical Planning and Inference, 142(12):3174–3188, December 2012. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375812002054.

Godichon-Baggioni:2019:OEA

[GB19] Antoine Godichon-Baggioni. Online estimation of the asymptotic variance for averaged stochastic gradient algorithms. Journal of Statistical Planning and Inference, 203(??):1–19, December 2019. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375819300011.

Goegebeur:2010:KES

[GBdW10] Yuri Goegebeur, Jan Beirlant, and Tertius de Wet. Kernel estimators for the second order parameter in extreme value statistics. Journal of Statistical Planning and Inference, 140(9):2632–2652, September 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375810001424.

Ghosh:2015:RDB

[GBP15] Abhik Ghosh, Ayanendranath Basu, and Leandro Pardo. On the robustness of a divergence based test of simple statistical hypotheses. Journal of Statistical Planning and Inference, 161(??):91–108, June 2015. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375815000051.

Gilardoni:2011:SOP

[GC11] Gustavo L. Gilardoni and Enrico A. Colosimo. On the superposition of overlapping Poisson processes and nonparametric estimation of their intensity function. Journal of Statistical Planning and Inference, 141(9):3075–3083, September 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375811001261.

Ghosh:2017:CER

[GC17] Subir Ghosh and Shrabanti Chowdhury. CV, ECV, and Robust CV designs for replications under a class of linear models in factorial experiments. Journal of Statistical Planning and Inference, 188(??):1–7, September 2017. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375817300393.

Golzy:2019:GFM

[GC19] Mojgan Golzy and Randy L. Carter. Generalized frailty models for analysis of recurrent events. Journal of Statistical Planning and Inference, 200(??):213–222, May 2019. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375818303355.

Gupta:2012:ARL

[GCDK12] V. K. Gupta, Kashinath Chatterjee, Ashish Das, and Basudev Kole. Addition of runs to an s-level supersaturated design. Journal of Statistical Planning and Inference, 142(8):2402–2408, August 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375812000675.

Gebregziabher:2010:LCB

[GD10] Mulugeta Gebregziabher and Stacia M. DeSantis. Latent class based multiple imputation approach for missing categorical data. Journal of Statistical Planning and Inference, 140(11):3252–3262, November 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375810002004.

Ghosh:2015:MCC

[GD15] Palash Ghosh and Anup Dewanji. Matched case-control study with reporting bias. Journal of Statistical Planning and Inference, 157–158(??):121–136, February/March 2015. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375814001748.

Garg:2018:ELV

[GD18] Mansi Garg and Isha Dewan. On estimation of limiting variance of partial sums of functions of associated random variables. Journal of Statistical Planning and Inference, 192(??):1–17, January 2018. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375817301398.

Grasshoff:2012:ODL

[GDHS12] Ulrike Graßhoff, Anna Doebler, Heinz Holling, and Rainer Schwabe. Optimal design for linear regression models in the presence of heteroscedasticity caused by random coefficients. Journal of Statistical Planning and Inference, 142(5):1108– 1113, May 2012. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375811004307.

Georgiou:2011:ODC

[Geo11] S. D. Georgiou. Orthogonal designs for computer experiments. Journal of Statistical Planning and Inference, 141(4):1519–1525, April 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375810005148.

Georgiou:2014:SDR

[Geo14] Stelios D. Georgiou. Supersaturated designs: a review of their construction and analysis. Journal of Statistical Planning and Inference, 144(??):92–109, January 2014. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375812003084.

Gao:2011:NGD

[GF11] Xiaoli Gao and Yixin Fang. A note on the generalized degrees of freedom under the L _{1} loss function. Journal of Statistical Planning and Inference, 141(2):677–686, February 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375810003307.

Ghosh:2013:CVF

[GF13] Subir Ghosh and Analisa Flores. Common variance fractional factorial designs and their optimality to identify a class of models. Journal of Statistical Planning and Inference, 143(10):1807–1815, October 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375813001456.

Gontscharuk:2018:ANB

[GF18] Veronika Gontscharuk and Helmut Finner. On the asymptotics of a normal beta-transformed empirical process. Journal of Statistical Planning and Inference, 193(??):103–108, February 2018. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S037837581730143X.

Goegebeur:2010:GFT

[GG10] Yuri Goegebeur and Armelle Guillou. Goodness-of-fit testing for Weibull-type behavior. Journal of Statistical Planning and Inference, 140(6):1417–1436, June 2010. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809003784.

Godolphin:2015:RRB

[GG15] J. D. Godolphin and E. J. Godolphin. The robustness of resolvable block designs against the loss of whole blocks or replicates. Journal of Statistical Planning and Inference, 163(??):34–42, August 2015. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375815000257.

Gardes:2011:WTD

[GGG11] Laurent Gardes, Stéphane Girard, and Armelle Guillou. Weibull tail-distributions revisited: a new look at some tail estimators. Journal of Statistical Planning and Inference, 141(1):429–444, January 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375810003113.

Grasshoff:2013:ODD

[GGHS13] Ulrike Graßhoff, Heiko Großmann, Heinz Holling, and Rainer Schwabe. Optimal design for discrete choice experiments. Journal of Statistical Planning and Inference, 143(1):167–175, January 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375812002327.

Gardes:2012:ECT

[GGS12] L. Gardes, A. Guillou, and A. Schorgen. Estimating the conditional tail index by integrating a kernel conditional quantile estimator. Journal of Statistical Planning and Inference, 142(6):1586–1598, June 2012. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375812000201.

Grunewald:2010:EAS

[GH10] Maria Grünewald and Ola Hössjer. Efficient ascertainment schemes for maximum likelihood estimation. Journal of Statistical Planning and Inference, 140(7):2078–2088, July 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810000765.

Gupta:2011:OCS

[GHM11] Sudhir Gupta, Katsuo Hisano, and Luis B. Morales. Optimal k-circulant supersaturated designs. Journal of Statistical Planning and Inference, 141(2):782–786, February 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810003666.

Ghosh:2011:JSJ

[Gho11] Subir Ghosh. Jagdish N. Srivastava: June 20, 1933– November 18, 2010. Journal of Statistical Planning and Inference, 141(5):1633–1634, May 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S037837581000546X.

Ghosh:2013:NTL

[Gho13] Sucharita Ghosh. Normality testing for a long-memory sequence using the empirical moment generating function. Journal of Statistical Planning and Inference, 143(5):944– 954, May 2013. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375812003412.

Ghosh:2014:LSE

[Gho14] Sucharita Ghosh. On local slope estimation in partial linear models under Gaussian subordination. Journal of Statistical Planning and Inference, 155(??):42–53, December 2014. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375814001049.

Ghiglietti:2017:SIS

[GIP17] Andrea Ghiglietti, Francesca Ieva, and Anna Maria Paganoni. Statistical inference for stochastic processes: Two-sample hypothesis tests. Journal of Statistical Planning and Inference, 180(??):49–68, January 2017. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375816300908.

Gaffke:2019:LOD

[GIS19] Norbert Gaffke, Osama Idais, and Rainer Schwabe. Locally optimal designs for gamma models. Journal of Statistical Planning and Inference, 203(??):199–214, December 2019. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375819300345.

Giuliani:2016:OOS

[Giu16] Marc-Antoine Giuliani. Orthogonal One Step Greedy Procedure for heteroscedastic linear models. Journal of Statistical Planning and Inference, 171(??):147–174, April 2016. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375815001925.

Gessel:2010:MGF

[GJ10] Ira M. Gessel and Pallavi Jayawant. Multilinear generating functions for Charlier polynomials. Journal of Statistical Planning and Inference, 140(8):2230–2236, August 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810000303.

Groeneboom:2012:IPT

[GJ12] Piet Groeneboom and Geurt Jongbloed. Isotonic L _{2}- projection test for local monotonicity of a hazard. Journal of Statistical Planning and Inference, 142(7):1644–1658, July 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375812000481.

Glavas:2018:MSC

[GJM18] Lenka Glavas, Jelena Jocković, and Pavle Mladenović. Maximum of the sum of consecutive terms in random permutations. Journal of Statistical Planning and Inference, 197(??):15–20, December 2018. CODEN JSPIDN.

ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375817302124.

Goroncy:2012:RGO

[GK12] Agnieszka Goroncy and Udo Kamps. Relations for mgeneralized order statistics via an Opial-type inequality. Journal of Statistical Planning and Inference, 142(6):1457–1463, June 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375812000043.

Geng:2017:MCT

[GK17] Pei Geng and Hira L. Koul. Model checking in tobit regression with measurement errors using validation data. Journal of Statistical Planning and Inference, 190(??):15–31, November 2017. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375817300824.

Godbole:2010:CCA

[GKM10] A. P. Godbole, M. V. Koutras, and F. S. Milienos. Consecutive covering arrays and a new randomness test. Journal of Statistical Planning and Inference, 140(5):1292–1305, May 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S037837580900370X.

Ghoshal:2015:SIB

[GKvdVvZ15] Subhashis Ghoshal, Bas Kleijn, Aad van der Vaart, and Harry van Zanten. Special issue on Bayesian nonparametrics. Journal of Statistical Planning and Inference, 166(??):1, November 2015. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375815000920.

Gheriballah:2010:RNE

[GLR10] Abdelkader Gheriballah, Ali Laksaci, and Rachida Rouane. Robust nonparametric estimation for spatial regression. Journal of Statistical Planning and Inference, 140(7):1656–1670, July 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375810000662.

Gamboa:2011:MEE

[GLR11] Fabrice Gamboa, Jean-Michel Loubes, and Paul Rochet. Maximum entropy estimation for survey sampling. Journal of Statistical Planning and Inference, 141(1):305–317, January 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375810002806.

Guilloux:2016:AEB

[GLT16] Agathe Guilloux, Sarah Lemler, and Marie-Luce Taupin. Adaptive estimation of the baseline hazard function in the Cox model by model selection, with high-dimensional covariates. Journal of Statistical Planning and Inference, 171(??): 38–62, April 2016. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375815002116.

Gai:2011:CIB

[GLW11] Yujie Gai, Lu Lin, and Xiuli Wang. Consistent inference for biased sub-model of high-dimensional partially linear model. Journal of Statistical Planning and Inference, 141(5):1888– 1898, May 2011. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375810005410.

Guo:2014:FDA

[GLZ14] Shaojun Guo, Shiqing Ling, and Ke Zhu. Factor double autoregressive models with application to simultaneous causality testing. Journal of Statistical Planning and Inference, 148(??):82–94, May 2014. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375813003091.

Gupta:2012:COM

[GM12] Sudhir Gupta and Luis B. Morales. Constructing E(s^{2})- optimal and minimax-optimal k-circulant supersaturated designs via multi-objective tabu search. Journal of Statistical Planning and Inference, 142(6):1415–1420, June 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811004666.

Gupta:2019:OTS

[GM19] Shanti S. Gupta and Klaus J. Miescke. Optimum two-stage selection procedures for Weibull populations. Journal of Statistical Planning and Inference, 15(??):147–156, ???? 2019. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0378375886900935.

Gonzalez-Manteiga:2011:BFL

[GMMC11] Wenceslao González-Manteiga and Adela Martínez-Calvo. Bootstrap in functional linear regression. Journal of Statistical Planning and Inference, 141(1):453–461, January 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810003204.

Giuliano:2015:ARR

[GMP15] Rita Giuliano, Claudio Macci, and Barbara Pacchiarotti. Asymptotic results for runs and empirical cumulative entropies. Journal of Statistical Planning and Inference, 157– 158(??):77–89, February/March 2015. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375814001736.

Galvao:2010:PQR

[GMR10] Antonio F. Galvao and Gabriel V. Montes-Rojas. Penalized quantile regression for dynamic panel data. Journal of Statistical Planning and Inference, 140(11):3476–3497, November 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375810002454.

Ghosh:2010:BLB

[GMSK10] M. Ghosh, B. Mukherjee, U. Santra, and D. Kim. Bayesian and likelihood-based inference for the bivariate normal correlation coefficient. Journal of Statistical Planning and Inference, 140(6):1410–1416, June 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S037837580900367X.

Ganguly:2012:EIT

[GMSK12] A. Ganguly, S. Mitra, D. Samanta, and D. Kundu. Exact inference for the two-parameter exponential distribution under Type-II hybrid censoring. Journal of Statistical Planning and Inference, 142(3):613–625, March 2012. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811003120.

Ghosh:2016:MFO

[GN16] Subir Ghosh and Hans Nyquist. Model fitting and optimal design for a class of binary response models. Journal of Statistical Planning and Inference, 179(??):22–35, December 2016. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375816300647.

Golosnoy:2011:NME

[GO11] Vasyl Golosnoy and Yarema Okhrin. Nonparametric monitoring of equal predictive ability. Journal of Statistical Planning and Inference, 141(9):3170–3180, September 2011. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811001364.

Gob:2012:VBD

[Göb12] Rainer Göb. Variance bounds for the design of audit sampling. Journal of Statistical Planning and Inference, 142(9):2629–2645, September 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375812000511.

Godet:2010:PLM

[God10] Fanny Godet. Prediction of long memory processes on same-realisation. Journal of Statistical Planning and Inference, 140(4):907–926, April 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375809003061.

Good:2019:SDF

[Goo19] I. J. Good. Statistical discussion forum. Journal of Statistical Planning and Inference, 15(??):129, ???? 2019. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0378375886900911.

Gong:2010:CAM

[GP10] Yun Gong and Liang Peng. Coverage accuracy for a mean without third moment. Journal of Statistical Planning and Inference, 140(4):1082–1088, April 2010. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809003437.

Greenshtein:2012:RTD

[GP12] Eitan Greenshtein and Junyong Park. Robust test for detecting a signal in a high dimensional sparse normal vector. Journal of Statistical Planning and Inference, 142(6):1445– 1456, June 2012. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375812000031.

Guillotte:2015:IAC

[GP15a] Simon Guillotte and Julien Perron. Inference on Archimedean copulas using mixtures of Pólya trees. Journal of Statistical Planning and Inference, 166(??):2–13, November 2015. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375814000056.

Guttorp:2015:PEE

[GP15b] Peter Guttorp and Michael D. Perlman. Predicting extinction or explosion in a Galton–Watson branching process with power series offspring distribution. Journal of Statistical Planning and Inference, 167(??):193–215, December 2015. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375815000932.

Goncalves:2016:DBP

[GP16] Sílvia Gonçalves and Benoît Perron. Discussion of “Bootstrap prediction intervals for linear, nonlinear, and nonparametric autoregressions”, by Li Pan and Dimitris Politis. Journal of Statistical Planning and Inference, 177(??):31–34, October 2016. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375815001974. See [PP16a].

Giurcanu:2019:BLT

[GP19] Mihai Giurcanu and Brett Presnell. Bootstrapping LASSOtype estimators in regression models. Journal of Statistical Planning and Inference, 199(??):114–125, March 2019. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375818300648.

Ghattas:2011:DDS

[GPRY11] B. Ghattas, D. Pommeret, L. Reboul, and A. F. Yao. Data driven smooth test for paired populations. Journal of Statistical Planning and Inference, 141(1):262–275, January 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810002867.

Goroncy:2011:LBE

[GR11] Agnieszka Goroncy and Tomasz Rychlik. Lower bounds on the expectations of upper record values. Journal of Statistical Planning and Inference, 141(8):2726–2737, August 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811000814.

Gamboa:2014:OVS

[GR14] Fabrice Gamboa and Alain Rouault. Operator-valued spectral measures and large deviations. Journal of Statistical Planning and Inference, 154(??):72–86, November 2014. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375814000251.

Guo:2015:SCD

[GR15] Wenge Guo and Joseph P. Romano. On stepwise control of directional errors under independence and some dependence. Journal of Statistical Planning and Inference, 163(??):21–33, August 2015. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375815000397.

Guedj:2018:PBH

[GR18] Benjamin Guedj and Sylvain Robbiano. PAC-Bayesian high dimensional bipartite ranking. Journal of Statistical Planning and Inference, 196(??):70–86, August 2018. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375817301945.

Grafstrom:2010:GPS

[Gra10] Anton Grafström. On a generalization of Poisson sampling. Journal of Statistical Planning and Inference, 140(4):982– 991, April 2010. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375809003140.

Graczyk:2012:NAO

[Gra12a] Małgorzata Graczyk. Notes about A-optimal spring balance weighing design. Journal of Statistical Planning and Inference, 142(4):781–784, April 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375811004137.

Grafstrom:2012:SCP

[Gra12b] Anton Grafström. Spatially correlated Poisson sampling. Journal of Statistical Planning and Inference, 142(1):139– 147, January 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375811002734.

Gonzalez-Rodriguez:2012:NWQ

[GRCGL12] Gil González-Rodríguez, Ana Colubi, M. Ángeles Gil, and M. Asunción Lubiano. A new way of quantifying the symmetry of a random variable: Estimation and hypothesis testing. Journal of Statistical Planning and Inference, 142(12):3061–3072, December 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375812002078.

Ginsbourger:2016:DIR

[GRD16] David Ginsbourger, Olivier Roustant, and Nicolas Durrande. On degeneracy and invariances of random fields paths with applications in Gaussian process modelling. Journal of Statistical Planning and Inference, 170(??):117–128, March 2016. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375815001640.

Griffith:2010:MCN

[Gri10] Daniel A. Griffith. The Moran coefficient for nonnormal data. Journal of Statistical Planning and Inference, 140(11):2980–2990, November 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375810001667.

Griffin:2011:OUD

[Gri11] J. E. Griffin. The Ornstein–Uhlenbeck Dirichlet process and other time-varying processes for Bayesian nonparametric inference. Journal of Statistical Planning and Inference, 141(11):3648–3664, November 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375811002242.

Goldberg:2014:PCF

[GRM14] Y. Goldberg, Y. Ritov, and A. Mandelbaum. Predicting the continuation of a function with applications to call center data. Journal of Statistical Planning and Inference, 147(??): 53–65, April 2014. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375813002929.

Groeneboom:2013:NSL

[Gro13a] Piet Groeneboom. Nonparametric (smoothed) likelihood and integral equations. Journal of Statistical Planning and Inference, 143(12):2039–2065, December 2013. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375813001535. See discussion [Jon13a,

vdGM13, Mam13].

Groeneboom:2013:R

[Gro13b] Piet Groeneboom. Rejoinder. Journal of Statistical Planning and Inference, 143(12):2074–2078, December 2013. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375813001572.

Gromping:2018:CIF

[Grö18] Ulrike Grömping. Coding invariance in factorial linear models and a new tool for assessing combinatorial equivalence of factorial designs. Journal of Statistical Planning and Inference, 193(??):1–14, February 2018. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375817301350.

Gunzel:2012:GAS

[GRS^{+}12] Thomas Günzel, Wolf-Dieter Richter, Stefan Scheutzow, Kay Schicker, and John Venz. Geometric approach to the skewed normal distribution. Journal of Statistical Planning and Inference, 142(12):3209–3224, December 2012. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375812002108.

Grunwald:2018:SP

[Grü18] Peter Grünwald. Safe probability. Journal of Statistical Planning and Inference, 195(??):47–63, May 2018. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375817301714.

Gegler:2010:ECL

[GS10] Achim Gegler and Ulrich Stadtmüller. Estimation of the characteristics of a Lévy process. Journal of Statistical Planning and Inference, 140(6):1481–1496, June 2010. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809003711.

Georgiou:2011:BCM

[GS11a] S. D. Georgiou and S. Stylianou. Block-circulant matrices for constructing optimal Latin hypercube designs. Journal of Statistical Planning and Inference, 141(5):1933–1943, May 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375810005641. See corrigen-

dum [GS12a].

Gill:2011:SAU

[GS11b] Paramjit S. Gill and Tim B. Swartz. Stylometric analyses using Dirichlet process mixture models. Journal of Statistical Planning and Inference, 141(11):3665–3674, November 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375811002254.

Georgiou:2012:CBC

[GS12a] S. D. Georgiou and S. Stylianou. Corrigendum to: “Blockcirculant matrices for constructing optimal Latin hypercube designs” [J. Statist. Plann. Inference 141 (2011) 1933–1943]. Journal of Statistical Planning and Inference, 142(6):1623, June 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375812000122. See [GS11a].

Gudendorf:2012:NEM

[GS12b] Gordon Gudendorf and Johan Segers. Nonparametric estimation of multivariate extreme-value copulas. Journal of Statistical Planning and Inference, 142(12):3073–3085, December 2012. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375812001978.

Gaffke:2019:QNA

[GS19] N. Gaffke and R. Schwabe. Quasi-Newton algorithm for optimal approximate linear regression design: Optimization in matrix space. Journal of Statistical Planning and Inference, 198(??):62–78, January 2019. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375818300314.

Grossmann:2012:DFO

[GSG12] Heiko Großmann, Rainer Schwabe, and Steven G. Gilmour. Designs for first-order interactions in paired comparison experiments with two-level factors. Journal of Statistical Planning and Inference, 142(8):2395–2401, August 2012. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375812000808.

Gupta:2010:ART

[GSKP10] V. K. Gupta, Poonam Singh, Basudev Kole, and Rajender Parsad. Addition of runs to a two-level supersaturated design. Journal of Statistical Planning and Inference, 140(9):2531–2535, September 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375810001394.

Gupta:2010:MSE

[GSS10] Sat Gupta, Javid Shabbir, and Supriti Sehra. Mean and sensitivity estimation in optional randomized response models. Journal of Statistical Planning and Inference, 140(10):2870–2874, October 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375810001229.

Gibson:2011:GDA

[GSZ11] Gavin J. Gibson, George Streftaris, and Stan Zachary. Generalised data augmentation and posterior inferences. Journal of Statistical Planning and Inference, 141(1):156–171, January 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375810002727.

Guggemos:2010:PCS

[GT10] Fabien Guggemos and Yves Tillé. Penalized calibration in survey sampling: Design-based estimation assisted by mixed models. Journal of Statistical Planning and Inference, 140(11):3199–3212, November 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375810001771.

Ghahramani:2012:NRE

[GT12a] M. Ghahramani and A. Thavaneswaran. Nonlinear recursive estimation of volatility via estimating functions. Journal of Statistical Planning and Inference, 142(1):171–180, January 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S037837581100276X.

Guan:2012:OSS

[GT12b] Qiang Guan and Yincai Tang. Optimal step-stress test under Type-I censoring for multivariate exponential distribution. Journal of Statistical Planning and Inference, 142(7):1908– 1923, July 2012. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375812000766.

Guan:2012:GSC

[Gua12] Yu Guan. A generalized score confidence interval for a binomial proportion. Journal of Statistical Planning and Inference, 142(4):785–793, April 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375811003569.

Guiasu:2019:GDU

[Gui19] Silviu Guiasu. Grouping data by using the weighted entropy. Journal of Statistical Planning and Inference, 15(??):63–69, ???? 2019. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375886900856.

Gupta:2012:RMR

[Gup12] Ramesh C. Gupta. Regression mean residual life of a parallel system of dependent components. Journal of Statistical Planning and Inference, 142(5):1063–1072, May 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811004150.

Goos:2010:OCC

[GVV10] Peter Goos, Bart Vermeulen, and Martina Vandebroek. D-optimal conjoint choice designs with no-choice options for a nested logit model. Journal of Statistical Planning and Inference, 140(4):851–861, April 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375809002870.

Godolphin:2011:ICR

[GW11] J. D. Godolphin and H. R. Warren. Improved conditions for the robustness of binary block designs against the loss of whole blocks. Journal of Statistical Planning and Inference, 141(11):3498–3505, November 2011. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811001959.

Guo:2012:GFT

[GX12] Xu Guo and Wangli Xu. Goodness-of-fit tests for general linear models with covariates missed at random. Journal of Statistical Planning and Inference, 142(7):2047–2058, July 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375812000869.

Gao:2010:RWE

[GZ10] Shesheng Gao and Yongmin Zhong. Random weighting estimation of kernel density. Journal of Statistical Planning and Inference, 140(9):2403–2407, September 2010. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810000820.

Gao:2014:NOD

[GZ14a] Lucy L. Gao and Julie Zhou. New optimal design criteria for regression models with asymmetric errors. Journal of Statistical Planning and Inference, 149(??):140–151, June 2014. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837581400007X.

Guo:2014:OSM

[GZ14b] Shaojun Guo and Donglin Zeng. An overview of semiparametric models in survival analysis. Journal of Statistical Planning and Inference, 151–152(??):1–16, August/ September 2014. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375813002619.

Gwise:2011:ORA

[GZH11] Thomas E. Gwise, Jianhui Zhou, and Feifang Hu. An optimal response adaptive biased coin design with k heteroscedastic treatments. Journal of Statistical Planning and Inference, 141(1):235–242, January 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375810002922.

Hable:2010:MDE

[Hab10] Robert Hable. Minimum distance estimation in imprecise probability models. Journal of Statistical Planning and Inference, 140(2):461–479, February 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375809002420.

Habiger:2012:MMM

[Hab12] Joshua D. Habiger. A method for modifying multiple testing procedures. Journal of Statistical Planning and Inference, 142(7):2227–2231, July 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375812000572.

Habibi:2013:FSD

[Hab13] Reza Habibi. Finite sample distributions of statistics in change point analysis. Journal of Statistical Planning and Inference, 143(2):434–440, February 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375812002297.

Habiger:2015:MTF

[Hab15] Joshua D. Habiger. Multiple test functions and adjusted pvalues for test statistics with discrete distributions. Journal of Statistical Planning and Inference, 167(??):1–13, December 2015. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375815001172.

Han:2011:VLS

[Han11] Jun Han. Variance least squares estimators for multivariate linear mixed model. Journal of Statistical Planning and Inference, 141(10):3345–3355, October 2011. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811001613.

Han:2014:FSM

[Han14] Peisong Han. A further study of the multiply robust estimator in missing data analysis. Journal of Statistical Planning and Inference, 148(??):101–110, May 2014. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837581300308X.

Harman:2014:MMC

[Har14] Radoslav Harman. Multiplicative methods for computing Doptimal stratified designs of experiments. Journal of Statistical Planning and Inference, 146(??):82–94, March 2014. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375813002541.

Hashimoto:2019:MMP

[Has19] Shintaro Hashimoto. Moment matching priors for nonregular models. Journal of Statistical Planning and Inference, 203(??):169–177, December 2019. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S037837581930031X.

Hayakawa:2016:EWM

[Hay16] Kazuhiko Hayakawa. On the effect of weighting matrix in GMM specification test. Journal of Statistical Planning and Inference, 178(??):84–98, November 2016. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837581630060X.

Hyun:2012:RAE

[HBP12] Jung Won Hyun, Prabir Burman, and Debashis Paul. A regression approach for estimating the parameters of the covariance function of a stationary spatial random process. Journal of Statistical Planning and Inference, 142(8):2330–2344, August 2012. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375812001218.

Houlding:2011:AUT

[HC11a] B. Houlding and F. P. A. Coolen. Adaptive utility and trial aversion. Journal of Statistical Planning and Inference, 141(2):734–747, February 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375810003733.

Hung:2011:NMT

[HC11b] Hung Hung and Chin-Tsang Chiang. Nonparametric methodology for the time-dependent partial area under the ROC curve. Journal of Statistical Planning and Inference, 141(12):3829–3838, December 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375811002680.

Hu:2013:ENV

[HC13] Chih-Chi Hu and Ying Kuen Cheung. On the efficiency of nonparametric variance estimation in sequential dose-finding. Journal of Statistical Planning and Inference, 143(3):593– 602, March 2013. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375812002820.

Huang:2016:ODQ

[HC16] Shih-Hao Huang and Ching-Shui Cheng. Optimal designs for quadratic regression with random block effects: the case of block size two. Journal of Statistical Planning and Inference, 175(??):67–77, August 2016. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375816000409.

Huang:2012:CNG

[HCMP12] Jianhua Z. Huang, Min Chen, Mehdi Maadooliat, and Mohsen Pourahmadi. A cautionary note on generalized linear models for covariance of unbalanced longitudinal data.

Journal of Statistical Planning and Inference, 142(3):743– 751, March 2012. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375811003727.

He:2016:EHC

[HCT16] Fengyang He, Yebin Cheng, and Tiejun Tong. Estimation of high conditional quantiles using the Hill estimator of the tail index. Journal of Statistical Planning and Inference, 176(??):64–77, September 2016. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375816300155.

Hyun:2019:GRM

[HCZ19] Noorie Hyun, David J. Couper, and Donglin Zeng. Gumbel regression models for a monotone increasing continuous biomarker subject to measurement error. Journal of Statistical Planning and Inference, 203(??):160–168, December 2019. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375819300308.

Hore:2014:DIR

[HDC14] Samrat Hore, Anup Dewanji, and Aditya Chatterjee. Design issues related to allocation of experimental units with known covariates into two treatment groups. Journal of Statistical Planning and Inference, 155(??):117–126, December 2014. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375814001025.

Hall:2012:MPI

[HdGT12] Andreia Hall and Maria da Graça Temido. On the maximum of periodic integer-valued sequences with exponential type tails via max-semistable laws. Journal of Statistical Planning and Inference, 142(7):1824–1836, July 2012. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811004654.

He:2011:SEP

[He11] Heping He. Selection of error probability laws by generalized modified profile likelihood. Journal of Statistical Planning and Inference, 141(3):1208–1213, March 2011. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810004325.

Hannig:2018:FLI

[HFI^{+}18] Jan Hannig, Qing Feng, Hari Iyer, C. M. Wang, and Xuhua Liu. Fusion learning for inter-laboratory comparisons. Journal of Statistical Planning and Inference, 195(??):64–79, May 2018. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375817301684.

Hu:2011:DRK

[HFQ11] Zonghui Hu, Dean A. Follmann, and Jing Qin. Dimension reduced kernel estimation for distribution function with incomplete data. Journal of Statistical Planning and Inference, 141(9):3084–3093, September 2011. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811001273.

Hinich:2010:SUP

[HFW10] Melvin J. Hinich, John Foster, and Phillip Wild. A statistical uncertainty principle for estimating the time of a discrete shift in the mean of a continuous time random process. Journal of Statistical Planning and Inference, 140(12):3688–3692, December 2010. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375810002144.

Hossain:2011:ALS

[HG11] Shahadut Hossain and Paul Gustafson. Assessing large sample bias in misspecified model scenarios with reference to exposure model misspecification in errors-in-variable regression: a new computational approach. Journal of Statistical Planning and Inference, 141(3):1161–1169, March 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810004283.

Haslett:2012:DCD

[HG12] S. J. Haslett and K. Govindaraju. Data cloning: Data visualisation, smoothing, confidentiality, and encryption. Journal of Statistical Planning and Inference, 142(2):410–422, February 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375811003053.

Hornik:2013:CFJ

[HG13] Kurt Hornik and Bettina Grün. On conjugate families and Jeffreys priors for von Mises–Fisher distributions. Journal of Statistical Planning and Inference, 143(5):992–999, May 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375812003448.

Heller:2012:CMD

[HGH12] R. Heller, M. Gorfine, and Y. Heller. A class of multivariate distribution-free tests of independence based on graphs. Journal of Statistical Planning and Inference, 142(12):3097–3106, December 2012. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375812002042.

Honda:2014:VSC

[HH14a] Toshio Honda and Wolfgang Karl Härdle. Variable selection in Cox regression models with varying coefficients. Journal of Statistical Planning and Inference, 148(??):67–81, May 2014. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375813003042.

Huang:2014:RDP

[HH14b] Shih-Hao Huang and Mong-Na Lo Huang. Robust designs for probability estimation in binary response experiments. Journal of Statistical Planning and Inference, 154(??):116–132, November 2014. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375813003030.

Huang:2014:IAR

[HH14c] Yufen Huang and Chao-Yen Hsieh. Influence analysis in response surface methodology. Journal of Statistical Planning and Inference, 147(??):188–203, April 2014. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375813002942.

Hermansen:2015:BMT

[HH15] Gudmund Horn Hermansen and Nils Lid Hjort. Bernshtein– von Mises theorems for nonparametric function analysis via locally constant modelling: a unified approach. Journal of Statistical Planning and Inference, 166(??):138–157, November 2015. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375815000373.

Huang:2013:EOD

[HHJ13] Miao-Kuan Huang, Mong-Na Lo Huang, and Baisuo Jin. Exact D-optimal designs for a linear log contrast model with mixture experiment for three and four ingredients. Journal of Statistical Planning and Inference, 143(7):1221–1232, July 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375813000268.

Habtesllassie:2015:ABS

[HHMO15] Y. G. Habtesllassie, Linda M. Haines, H. G. Mwambi, and J. W. Odhiambo. Array-based schemes for group screening with test errors which incorporate a concentration effect. Journal of Statistical Planning and Inference, 167(??):41–57, December 2015. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375815001147.

Hill:2011:EMS

[Hil11] Jonathan B. Hill. Extremal memory of stochastic volatility with an application to tail shape inference. Journal of Statistical Planning and Inference, 141(2):663–676, February 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375810003319.

Hirose:2019:CGA

[Hir19] Masayo Y. Hirose. A class of general adjusted maximum likelihood methods for desirable mean squared error estimation of EBLUP under the Fay–Herriot small area model. Journal of Statistical Planning and Inference, 199(??):302–310, March 2019. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375818301526.

Hung:2014:DTP

[HJ14] Ying Hung and V. Roshan Joseph. Discussion of “Threephase optimal design of sensitivity experiments” by Wu and Tian. Journal of Statistical Planning and Inference, 149(??): 16–19, June 2014. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375813003133. See [WT14b].

Heesen:2016:DAM

[HJ16] Philipp Heesen and Arnold Janssen. Dynamic adaptive multiple tests with finite sample FDR control. Journal of Statistical Planning and Inference, 168(??):38–51, January 2016. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375815001330.

Hartung:2010:ACI

[HK10] Joachim Hartung and Guido Knapp. Adaptive confidence intervals of desired length and power for normal means. Journal of Statistical Planning and Inference, 140(11):3317–3325, November 2010. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S037837581000234X.

Hansmann:2019:ECQ

[HK19] Matthias Hansmann and Michael Kohler. Estimation of conditional quantiles from data with additional measurement errors. Journal of Statistical Planning and Inference, 200(??):176–195, May 2019. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375818303070.

Haines:2018:ODT

[HKN18] Linda M. Haines, Gaëtan M. Kabera, and Principal Ndlovu. D-optimal designs for the two-variable binary logistic regression model with interaction. Journal of Statistical Planning and Inference, 193(??):136–150, February 2018. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375817301441.

Hemmati:2011:NTP

[HKR11] Fariba Hemmati, Esmaile Khorram, and Saeid Rezakhah. A new three-parameter ageing distribution. Journal of Statistical Planning and Inference, 141(7):2266–2275, July 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811000243.

Hallin:2013:DLQ

[HL13] Marc Hallin and Zudi Lu. Discussion of “Local quantile regression” by Spokoiny, Wang, and Härdle. Journal of Statistical Planning and Inference, 143(7):1130–1133, July 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375813000542. See [SWH13a].

Han:2016:SCE

[HLB^{+}16] Yang Han, Wei Liu, Frank Bretz, Fang Wan, and Ping Yang. Statistical calibration and exact one-sided simultaneous tolerance intervals for polynomial regression. Journal of Statistical Planning and Inference, 168(??):90–96, January 2016. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375815001391.

Ho:2012:SRT

[HLCW12] Hsiu J. Ho, Tsung-I. Lin, Hsuan-Yu Chen, and Wan-Lun Wang. Some results on the truncated multivariate t distribution. Journal of Statistical Planning and Inference, 142(1):25–40, January 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S037837581100231X.

Huang:2013:LCA

[HLH13] Tao Huang, Zhongqiang Liu, and Feifang Hu. Longitudinal covariate-adjusted response-adaptive randomized designs. Journal of Statistical Planning and Inference, 143(10):1816–1827, October 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375813000761.

Huang:2012:OCS

[HLL12] Chao Huang, Dennis K. J. Lin, and Min-Qian Liu. An optimality criterion for supersaturated designs with quantitative factors. Journal of Statistical Planning and Inference, 142(7):1780–1788, July 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S037837581200047X.

Hsieh:2010:LRA

[HLS10] S. H. Hsieh, S. M. Lee, and P. S. Shen. Logistic regression analysis of randomized response data with missing covariates. Journal of Statistical Planning and Inference, 140(4):927– 940, April 2010. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375809003103.

Hyun:2012:PHM

[HLS12] Seunggeun Hyun, Jimin Lee, and Yanqing Sun. Proportional hazards model for competing risks data with missing cause of failure. Journal of Statistical Planning and Inference, 142(7):1767–1779, July 2012. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375812000845.

He:2017:MCD

[HLSL17] Yuanzhen He, C. Devon Lin, Fasheng Sun, and Benjian Lv. Marginally coupled designs for two-level qualitative factors. Journal of Statistical Planning and Inference, 187(??):103– 108, August 2017. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375817300332.

Hu:2012:EPG

[HLY12] Jianhua Hu, Fuxiang Liu, and Jinhong You. Estimation of parameters in a generalized GMANOVA model based on an outer product analogy and least squares. Journal of Statistical Planning and Inference, 142(7):2017–2031, July 2012.

CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375812000493.

Holan:2010:TEE

[HM10] Scott H. Holan and Tucker S. McElroy. Tail exponent estimation via broadband log density-quantile regression. Journal of Statistical Planning and Inference, 140(12):3693–3708, December 2010. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375810002156.

Hillier:2011:SCA

[HM11a] Grant Hillier and Federico Martellosio. Spatial circulants, with applications. Journal of Statistical Planning and Inference, 141(7):2368–2385, July 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375811000401.

Hosseinian:2011:RBR

[HM11b] Sahar Hosseinian and Stephan Morgenthaler. Robust binary regression. Journal of Statistical Planning and Inference, 141(4):1497–1509, April 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S037837581000515X.

Huda:2013:TLM

[HM13] S. Huda and Rahul Mukerjee. Two-level minimum aberration designs in N =2 (mod 4) runs. Journal of Statistical Planning and Inference, 143(11):1872–1879, November 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375813001778.

Huang:2012:IAR

[HMW12] Kai Huang, Jie Mi, and Zeyi Wang. Inference about reliability parameter with gamma strength and stress. Journal of Statistical Planning and Inference, 142(4):848–854, April 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S037837581100382X.

He:2012:FIC

[HN12] Qinying He and H. N. Nagaraja. Fisher information in censored samples from Downton’s bivariate exponential distribution. Journal of Statistical Planning and Inference, 142(7):1888–1898, July 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375812000626.

Ho:2010:MPE

[Ho10] Kwok-Wah Ho. A matching prior for extreme quantile estimation of the generalized Pareto distribution. Journal of Statistical Planning and Inference, 140(6):1513–1518, June 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375809003917.

Hochberg:2019:MCR

[Hoc19] Yosef Hochberg. Multiple classification rules for signs of parameters. Journal of Statistical Planning and Inference, 15(??):177–188, ???? 2019. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
0378375886900960.

Hyodo:2019:AMA

[HOK19] Masashi Hyodo, Akihiro Onobuchi, and Hiroyuki Kurakami. An approximate multivariate asymptotic expansion-based test for population bioequivalence. Journal of Statistical Planning and Inference, 200(??):74–86, May 2019. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375818302611.

Honda:2010:NRD

[Hon10] Toshio Honda. Nonparametric regression for dependent data in the errors-in-variables problem. Journal of Statistical Planning and Inference, 140(11):3409–3424, November 2010. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810002478.

Hafner:2010:DBG

[HP10a] Christian M. Hafner and Arie Preminger. Deciding between GARCH and stochastic volatility via strong decision rules. Journal of Statistical Planning and Inference, 140(3):791– 805, March 2010. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375809002894.

Haslett:2010:EAR

[HP10b] Stephen J. Haslett and Simo Puntanen. Effect of adding regressors on the equality of the BLUEs under two linear models. Journal of Statistical Planning and Inference, 140(1):104–110, January 1, 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375809001943.

Hu:2010:ALE

[HP10c] Guikai Hu and Ping Peng. Admissibility for linear estimators of regression coefficient in a general Gauss-Markoff model under balanced loss function. Journal of Statistical Planning and Inference, 140(11):3365–3375, November 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810002417.

Hedayat:2019:CTB

[HP19] A. S. Hedayat and H. Pesotan. Characterization of triply balanced matrices with applications to survey sampling. Journal of Statistical Planning and Inference, 15(??):11–17, ???? 2019. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375886900819.

Hable:2010:ORI

[HRR10] R. Hable, P. Ruckdeschel, and H. Rieder. Optimal robust influence functions in semiparametric regression. Journal of Statistical Planning and Inference, 140(1):226–245, January 1, 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375809002213.

Harman:2011:OES

[HS11a] Radoslav Harman and Frantisek Stulajter. Optimality of equidistant sampling designs for the Brownian motion with a quadratic drift. Journal of Statistical Planning and Inference, 141(8):2750–2758, August 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S037837581100098X.

Harris:2011:MDE

[HS11b] Ian R. Harris and Shuyi Shen. The minimum L _{2} distance estimator for Poisson mixture models. Journal of Statistical Planning and Inference, 141(3):1088–1101, March 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810004131.

Harari:2014:ODG

[HS14] Ofir Harari and David M. Steinberg. Optimal designs for Gaussian process models via spectral decomposition. Journal of Statistical Planning and Inference, 154(??):87–101, November 2014. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375813002991.

Hjort:2018:CDR

[HS18] Nils Lid Hjort and Tore Schweder. Confidence distributions and related themes. Journal of Statistical Planning and Inference, 195(??):1–13, May 2018. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S037837581730174X.

He:2014:SIP

[HTZ14] Hua He, Wan Tang, and Guoxin Zuo. Statistical inference in the partial linear models with the double smoothing local linear regression method. Journal of Statistical Planning and Inference, 146(??):102–112, March 2014. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375813002577.

Hu:2013:ANH

[Hu13] Hongchang Hu. Asymptotic normality of Huber–Dutter estimators in a linear model with AR(1) processes. Journal of Statistical Planning and Inference, 143(3):548–562, March 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375812002807.

Huang:2011:SIP

[Hua11] Zhensheng Huang. Statistical inferences for partially linear single-index models with error-prone linear covariates. Journal of Statistical Planning and Inference, 141(2):899–909, February 2011. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375810004052.

Huckemann:2014:CSM

[Huc14] Stephan Huckemann. A comment to “Statistics on manifolds and landmark based image analysis: a nonparametric theory with applications”. Journal of Statistical Planning and Inference, 145(??):33–36, February 2014. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375813001833.

Hudecova:2013:SCA

[Hud13] Sárka Hudecová. Structural changes in autoregressive models for binary time series. Journal of Statistical Planning and Inference, 143(10):1744–1752, October 2013. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375813001122.

Humphreys:2010:HSL

[Hum10] Katherine Humphreys. A history and a survey of lattice path enumeration. Journal of Statistical Planning and Inference, 140(8):2237–2254, August 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375810000315.

Hung:2012:OBS

[Hun12] Ying-Chao Hung. Optimal Bayesian strategies for the infinite-armed Bernoulli bandit. Journal of Statistical Planning and Inference, 142(1):86–94, January 2012. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811002692.

Harvey:2012:BCI

[HvdM12] J. Harvey and A. J. van der Merwe. Bayesian confidence intervals for means and variances of lognormal and bivariate lognormal distributions. Journal of Statistical Planning and Inference, 142(6):1294–1309, June 2012. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811004502.

Hoff:2013:BSP

[HW13] Peter Hoff and Jon Wakefield. Bayesian sandwich posteriors for pseudo-true parameters: a discussion of “Bayesian inference with misspecified models” by Stephen Walker. Journal of Statistical Planning and Inference, 143(10):1638–1642, October 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375813001171. See [Wal13a].

Hoga:2017:SMT

[HW17] Yannick Hoga and Dominik Wied. Sequential monitoring of the tail behavior of dependent data. Journal of Statistical Planning and Inference, 182(??):29–49, March 2017. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375816301227.

Huang:2011:BEB

[HWZ11] Chu Huang, Hanchao Wang, and Lixin Zhang. Berry–Esseen bounds for kernel estimates of stationaryprocesses. Journal of Statistical Planning and Inference, 141(3):1290–1296, March 2011. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375810004404.

Harrar:2016:CRL

[HX16] Solomon W. Harrar and Jin Xu. Confidence regions for level differences in growth curve models. Journal of Statistical Planning and Inference, 175(??):11–24, August 2016. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375816000264.

Hu:2012:OLQ

[HXZ12] Ting Hu, Dao-Hong Xiang, and Ding-Xuan Zhou. Online learning for quantile regression and support vector regression. Journal of Statistical Planning and Inference, 142(12):3107–3122, December 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S037837581200211X.

Hattori:2010:MHE

[HY10a] Satoshi Hattori and Takashi Yanagawa. Mantel–Haenszel estimators for irregular sparse K2 × J tables. Journal of Statistical Planning and Inference, 140(9):2425–2432, September 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375810000881.

Hyodo:2010:APE

[HY10b] Masashi Hyodo and Takayuki Yamada. Asymptotic properties of the EPMC for modified linear discriminant analysis when sample size and dimension are both large. Journal of Statistical Planning and Inference, 140(9):2739–2748, September 2010. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S037837581000159X.

He:2019:OCR

[HY19] Lei He and Rong-Xian Yue. R-optimality criterion for regression models with asymmetric errors. Journal of Statistical Planning and Inference, 199(??):318–326, March 2019. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375818301654.

Hyun:2011:ODR

[HYF11] Seung Won Hyun, Min Yang, and Nancy Flournoy. Optimal designs for response functions with a downturn. Journal of Statistical Planning and Inference, 141(1):559–575, January 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375810003265.

Hamad:2010:NBN

[HYH10] Naqvi Hamad, Mohammad Zafar Yab, and Muhammad Hanif. Non-binary neighbor balance circular designs for ν = 2n and λ = 2. Journal of Statistical Planning and Inference, 140(11):3013–3016, November 2010. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810001734.

Hyodo:2012:MSC

[HYS12] Masashi Hyodo, Takayuki Yamada, and Muni S. Srivastava. A model selection criterion for discriminant analysis of highdimensional data with fewer observations. Journal of Statistical Planning and Inference, 142(12):3134–3145, December 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375812002030.

Hamadeh:2011:APL

[HZ11a] Tawfik Hamadeh and Jean-Michel Zakoïan. Asymptotic properties of LS and QML estimators for a class of nonlinear GARCH processes. Journal of Statistical Planning and Inference, 141(1):488–507, January 2011. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810003198.

Hu:2011:SRT

[HZ11b] Jianwei Hu and Runchu Zhang. Some results on two-level regular designs with general minimum lower-order confounding. Journal of Statistical Planning and Inference, 141(5):1774– 1782, May 2011. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375810005276.

Huang:2013:VSL

[HZ13] Xianzheng Huang and Hongmei Zhang. Variable selection in linear measurement error models via penalized score functions. Journal of Statistical Planning and Inference, 143(12):2101–2111, December 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375813001869.

Huang:2018:ELB

[HZ18] Haitao Huang and Yichuan Zhao. Empirical likelihood for the bivariate survival function under univariate censoring. Journal of Statistical Planning and Inference, 194(??):32–46, March 2018. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375817301775.

Hedayat:2014:ODS

[HZY14] A. S. Hedayat, Ying Zhou, and Min Yang. Optimal designs for some selected nonlinear models. Journal of Statistical Planning and Inference, 154(??):102–115, November 2014. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375814000792.

Inoue:2011:DSP

[IAH11] Kiyoshi Inoue, Sigeo Aki, and Katuomi Hirano. Distributions of simple patterns in some kinds of exchangeable sequences. Journal of Statistical Planning and Inference, 141(8):2532–2544, August 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375811000590.

Iliopoulos:2011:ELI

[IB11] G. Iliopoulos and N. Balakrishnan. Exact likelihood inference for Laplace distribution based on Type-II censored samples. Journal of Statistical Planning and Inference, 141(3):1224– 1239, March 2011. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375810004350.

Isogai:2010:SEL

[IF10] Eiichi Isogai and Andreas Futschik. Sequential estimation of a linear function of location parameters of two negative exponential distributions. Journal of Statistical Planning and Inference, 140(9):2416–2424, September 2010. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810000819.

Izadi:2012:TMI

[IK12] Muhyiddin Izadi and Baha-Eldin Khaledi. On testing more IFRA ordering. Journal of Statistical Planning and Inference, 142(4):840–847, April 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375811003806.

Igarashi:2015:BCS

[IK15] Gaku Igarashi and Yoshihide Kakizawa. Bias corrections for some asymmetric kernel estimators. Journal of Statistical Planning and Inference, 159(??):37–63, April 2015. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837581400192X.

Isogai:2011:HOA

[IKU11] Eiichi Isogai, Kana Kobayashi, and Chikara Uno. Higher order approximations by a two-stage procedure for a negative exponential distribution. Journal of Statistical Planning and Inference, 141(10):3304–3312, October 2011. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811001558.

Imhof:2014:OED

[Imh14] Lorens A. Imhof. G-optimal exact designs for quadratic regression. Journal of Statistical Planning and Inference, 154(??):133–140, November 2014. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375814000445.

Inglot:2010:IES

[Ing10] Tadeusz Inglot. Intermediate efficiency by shifting alternatives and evaluation of power. Journal of Statistical Planning and Inference, 140(11):3263–3281, November 2010. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810001990.

Iverson:2019:LSP

[IR19] Harald K. Iverson and Ronald H. Randles. Large sample properties of cross-validation assessment statistics. Journal of Statistical Planning and Inference, 15(??):43–62, ???? 2019. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0378375886900844.

Iqbal:2010:EGP

[IT10] Ijaz Iqbal and M. H. Tahir. Erratum “Generalized polygonal designs with block size 3 and λ = 1” by I. Iqbal et al. [J. Statist. Plann. Inference 139 (2009) 3200–3219]. Journal of Statistical Planning and Inference, 140(1):334, January 1, 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375809002195. See [ITA^{+}09].

Iqbal:2009:GPD

[ITA^{+}09] Ijaz Iqbal, M. H. Tahir, Munir Akhtar, S. S. A. Ghazali, Javid Shabbir, and N. S. Bukhari. Generalized polygonal designs with block size 3 and λ = 1. Journal of Statistical Planning and Inference, 139(9):3200–3219, September 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809000706. See erratum [IT10].

Iqbal:2012:GND

[ITA^{+}12] Ijaz Iqbal, M. H. Tahir, M. L. Aggarwal, Asghar Ali, and Iftikhar Ahmed. Generalized neighbor designs with block size 3. Journal of Statistical Planning and Inference, 142(3):626– 632, March 2012. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S037837581100348X.

Ishii:2016:APF

[IYA16] Aki Ishii, Kazuyoshi Yata, and Makoto Aoshima. Asymptotic properties of the first principal component and equality tests of covariance matrices in high-dimension, low-sample-size context. Journal of Statistical Planning and Inference, 170(??):186–199, March 2016. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375815001913.

Ishii:2019:ETH

[IYA19] Aki Ishii, Kazuyoshi Yata, and Makoto Aoshima. Equality tests of high-dimensional covariance matrices under the strongly spiked eigenvalue model. Journal of Statistical Planning and Inference, 202(??):99–111, September 2019. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375819300163.

Jiang:2017:CUD

[JA17] Bochuan Jiang and Mingyao Ai. Construction of uniform U-designs. Journal of Statistical Planning and Inference, 181(??):1–10, February 2017. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375816300891.

Jacroux:2011:UPC

[Jac11] Mike Jacroux. On the use of partial confounding for the construction of alternative regular two-level blocked fractional factorial designs. Journal of Statistical Planning and Inference, 141(7):2201–2208, July 2011. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811000310.

Jacroux:2019:DCM

[Jac19] Mike Jacroux. On the determination and construction of MV-optimal block designs for comparing test treatments with a standard treatment. Journal of Statistical Planning and Inference, 15(??):205–225, ???? 2019. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0378375886900984.

Jones:2011:POS

[JAI11] M. C. Jones and Karim Anaya-Izquierdo. On parameter orthogonality in symmetric and skew models. Journal of Statistical Planning and Inference, 141(2):758–770, February 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837581000368X.

Jansen:2015:GCV

[Jan15] Maarten Jansen. Generalized Cross Validation in variable selection with and without shrinkage. Journal of Statistical Planning and Inference, 159(??):90–104, April 2015. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375814001797.

Jaruskova:2010:ABT

[Jar10] Daniela Jarusková. Asymptotic behaviour of a test statistic for detection of change in mean of vectors. Journal of Statistical Planning and Inference, 140(3):616–625, March 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809002535.

Jaruskova:2013:TCC

[Jar13] Daniela Jarusková. Testing for a change in covariance operator. Journal of Statistical Planning and Inference, 143(9):1500–1511, September 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375813000931.

Jamalizadeh:2012:COS

[JB12] Ahad Jamalizadeh and N. Balakrishnan. Concomitants of order statistics from multivariate elliptical distributions. Journal of Statistical Planning and Inference, 142(2):397–409, February 2012. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375811002801.

Jacques:2014:MBC

[JB14] Julien Jacques and Christophe Biernacki. Model-based clustering for multivariate partial ranking data. Journal of Statistical Planning and Inference, 149(??):201–217, June 2014. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375814000330.

Jackson:2010:HDD

[JBB10] Dan Jackson, Jack Bowden, and Rose Baker. How does the DerSimonian and Laird procedure for random effects meta-analysis compare with its more efficient but harder to compute counterparts? Journal of Statistical Planning and Inference, 140(4):961–970, April 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375809003073.

Jiang:2016:RMU

[JD16] Zhichao Jiang and Peng Ding. Robust modeling using nonelliptically contoured multivariate t distributions. Journal of Statistical Planning and Inference, 177(??):50–63, October 2016. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375815301026.

Jwaid:2013:LSC

[JDB13] T. Jwaid, H. De Meyer, and B. De Baets. Lower semiquadratic copulas with a given diagonal section. Journal of Statistical Planning and Inference, 143(8):1355–1370, August 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375813000451.

Jones:2016:BLA

[JGJR16] Matthew Jones, Michael Goldstein, Philip Jonathan, and David Randell. Bayes linear analysis for Bayesian optimal experimental design. Journal of Statistical Planning and Inference, 171(??):115–129, April 2016. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375815001950.

Johnson:2011:GPE

[JGK11] J. S. Johnson, J. P. Gosling, and M. C. Kennedy. Gaussian process emulation for second-order Monte Carlo simulations. Journal of Statistical Planning and Inference, 141(5):1838– 1848, May 2011. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375810005343.

Jozani:2012:RNS

[JJ12] Mohammad Jafari Jozani and Brad C. Johnson. Randomized nomination sampling for finite populations. Journal of Statistical Planning and Inference, 142(7):2103–2115, July 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375812000456.

Jiang:2012:LRT

[JJY12] Dandan Jiang, Tiefeng Jiang, and Fan Yang. Likelihood ratio tests for covariance matrices of high-dimensional normal distributions. Journal of Statistical Planning and Inference, 142(8):2241–2256, August 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375812001048.

Jhong:2017:PBS

[JKL17] Jae-Hwan Jhong, Ja-Yong Koo, and Seong-Whan Lee. Penalized B-spline estimator for regression functions using total variation penalty. Journal of Statistical Planning and Inference, 184(??):77–93, May 2017. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375816301690.

Johnstone:2011:FPF

[JL11] David Johnstone and Yan-Xia Lin. Fitting probability forecasting models by scoring rules and maximum likelihood. Journal of Statistical Planning and Inference, 141(5):1832– 1837, May 2011. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375810005331.

Lin:2011:GSP

[jLP11] Chen ju Lin and Wen Lea Pearn. Group selection for production yield among k manufacturing lines. Journal of Statistical Planning and Inference, 141(4):1510–1518, April 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810005136.

Joo:2010:SPH

[JM10a] Sui Joo and Jie Mi. Some properties of hazard rate functions of systems with two components. Journal of Statistical Planning and Inference, 140(2):444–453, February 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809002341.

Jozani:2010:IAA

[JM10b] Mohammad Jafari Jozani and Sayed Jamal Mirkamali. Improved attribute acceptance sampling plans based on maxima nomination sampling. Journal of Statistical Planning and Inference, 140(9):2448–2460, September 2010. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810000959.

Jockovic:2011:CCP

[JM11a] Jelena Jocković and Pavle Mladenović. Coupon collector’s problem and generalized Pareto distributions. Journal of Statistical Planning and Inference, 141(7):2348–2352, July 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811000371.

Jozani:2011:CCA

[JM11b] Mohammad Jafari Jozani and Sayed Jamal Mirkamali. Control charts for attributes with maxima nominated samples. Journal of Statistical Planning and Inference, 141(7):2386– 2398, July 2011. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375811000413.

Jozwiak:2013:OTA

[JM13] Katarzyna Jóźwiak and Mirjam Moerbeek. Optimal treatment allocation and study duration for trials with discretetime survival endpoints. Journal of Statistical Planning and Inference, 143(5):971–982, May 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375812003473.

Jin:2013:NLM

[JMW13] Yinghua Jin, Ruixing Ming, and Yaohua Wu. Nonadditivity in loglinear models using φ-divergences and MφE s under product-multinomial sampling. Journal of Statistical Planning and Inference, 143(2):356–367, February 2013. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837581200273X.

Jongbloed:2013:DNS

[Jon13a] Geurt Jongbloed. Discussion of nonparametric (smoothed) likelihood and integral equations by Piet Groeneboom. Journal of Statistical Planning and Inference, 143(12):2066–2067, December 2013. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375813001559. See [Gro13a].

Jonsson:2013:COA

[Jon13b] Fredrik Jonsson. Characterizing optimality among threedecision procedures for directional conclusions. Journal of Statistical Planning and Inference, 143(2):392–399, February 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375812002418.

Jones:2015:UCD

[Jon15] M. C. Jones. Univariate continuous distributions: Symmetries and transformations. Journal of Statistical Planning and Inference, 161(??):119–124, June 2015. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375815000026.

Jorgensen:2012:EEI

[JP12] Bent Jørgensen and Hans Chr. Petersen. Efficient estimation for incomplete multivariate data. Journal of Statistical Planning and Inference, 142(5):1215–1224, May 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811004290.

Jensen:2010:SMI

[JR10] D. R. Jensen and D. E. Ramirez. Surrogate models in ill-conditioned systems. Journal of Statistical Planning and Inference, 140(7):2069–2077, July 2010. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810000728.

Jasinski:2012:BDO

[JR12] Krzysztof Jasiński and Tomasz Rychlik. Bounds on dispersion of order statistics based on dependent symmetrically distributed random variables. Journal of Statistical Planning and Inference, 142(8):2421–2429, August 2012. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375812000663.

Jokiel-Rokita:2010:EPD

[JRM10] Alicja Jokiel-Rokita and Ryszard Magiera. Estimation procedures with delayed observations. Journal of Statistical Planning and Inference, 140(4):992–1002, April 2010. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809003139.

Janssen:2012:LSB

[JSV12] Paul Janssen, Jan Swanepoel, and Noël Veraverbeke. Large sample behavior of the Bernstein copula estimator. Journal of Statistical Planning and Inference, 142(5):1189–1197, May 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375811004253.

Johannes:2013:IRN

[JVV13] Jan Johannes, Sébastien Van Bellegem, and Anne Vanhems. Iterative regularisation in nonparametric instrumental regression. Journal of Statistical Planning and Inference, 143(1):24–39, January 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375812002479.

Jiang:2016:ENC

[JWY16] Hui Jiang, Weigang Wang, and Lei Yu. An exponential nonuniform convergence rate for a class of normalized L-statistics. Journal of Statistical Planning and Inference, 171(??):135–146, April 2016. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375815001937.

Jiang:2011:LMD

[JZ11] Hui Jiang and Shoujiang Zhao. Large and moderate deviations in testing time inhomogeneous diffusions. Journal of Statistical Planning and Inference, 141(9):3160–3169, September 2011. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375811001352.

Jayasinghe:2013:NSE

[JZ13] Chathuri L. Jayasinghe and P. Zeephongsekul. Nonparametric smooth estimation of the expected inactivity time function. Journal of Statistical Planning and Inference, 143(5):911–928, May 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375812003515.

Khaledi:2011:MRL

[KA11] Baha-Eldin Khaledi and Leila Amiri. On the mean residual life order of convolutions of independent uniform random variables. Journal of Statistical Planning and Inference, 141(12):3716–3724, December 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375811002321.

Kakizawa:2018:NDE

[Kak18] Yoshihide Kakizawa. Nonparametric density estimation for nonnegative data, using symmetrical-based inverse and reciprocal inverse Gaussian kernels through dual transformation. Journal of Statistical Planning and Inference, 193(??):117–135, February 2018. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375817301453.

Kashikar:2017:EGR

[Kas17] Akanksha S. Kashikar. Estimation of growth rate in second order branching process. Journal of Statistical Planning and Inference, 191(??):1–12, December 2017. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837581730109X.

Katona:2011:FLO

[Kat11] Gyula O. H. Katona. Finding at least one excellent element in two rounds. Journal of Statistical Planning and Inference, 141(8):2946–2952, August 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375811001169.

Katsaounis:2012:EFD

[Kat12] Tena I. Katsaounis. Equivalence of factorial designs with qualitative and quantitative factors. Journal of Statistical Planning and Inference, 142(1):79–85, January 2012. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811002667.

Katayama:2019:CSA

[Kat19] Shota Katayama. Computational and statistical analyses for robust non-convex sparse regularized regression problem. Journal of Statistical Planning and Inference, 201(??):20–31, July 2019. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375818303434.

Kaul:2014:LLM

[Kau14] Abhishek Kaul. Lasso with long memory regression errors. Journal of Statistical Planning and Inference, 153(??):11–26, October 2014. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375814000779.

Konstantinou:2015:ODF

[KBK15] Maria Konstantinou, Stefanie Biedermann, and Alan C. Kimber. Optimal designs for full and partial likelihood information — with application to survival models. Journal of Statistical Planning and Inference, 165(??):27–37, October 2015. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375815000580.

Kundu:2011:SEF

[KBNB11] Debasis Kundu, Zhidong Bai, Swagata Nandi, and Li Bai. Super efficient frequency estimation. Journal of Statistical Planning and Inference, 141(8):2576–2588, August 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811000632.

Kieffer:2012:PSC

[KBPW12] David Kieffer, Laurent Bianchetti, Olivier Poch, and Nicolas Wicker. Perfect sampling on 2× ⋯ ×2×k contingency tables with an application to SAGE data. Journal of Statistical Planning and Inference, 142(4):896–901, April 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811003995.

Kaiser:2012:CPD

[KCF12] Mark S. Kaiser, Petruta C. Caragea, and Kyoji Furukawa. Centered parameterizations and dependence limitations in Markov random field models. Journal of Statistical Planning and Inference, 142(7):1855–1863, July 2012. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375812000778.

Kim:2013:ATM

[KCZ13] Dong-Yun Kim, Yuehua Cui, and Ou Zhao. Asymptotic test of mixture model and its applications to QTL interval mapping. Journal of Statistical Planning and Inference, 143(8):1320–1329, August 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375813000657.

Kateri:2012:CSM

[KD12] Maria Kateri and Petros Dellaportas. Conditional symmetry models for three-way contingency tables. Journal of Statistical Planning and Inference, 142(8):2430–2439, August 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375812000730.

Katsaounis:2013:EFF

[KDJ13] T. I. Katsaounis, A. M. Dean, and Bradley Jones. On equivalence of fractional factorial designs based on singular value decomposition. Journal of Statistical Planning and Inference, 143(11):1950–1953, November 2013. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837581300150X.

Kemp:2010:FPS

[Kem10] Adrienne W. Kemp. Families of power series distributions, with particular reference to the Lerch family. Journal of Statistical Planning and Inference, 140(8):2255–2259, August 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375810000327.

Karacsony:2010:ANK

[KF10] Zsolt Karácsony and Peter Filzmoser. Asymptotic normality of kernel type regression estimators for random fields. Journal of Statistical Planning and Inference, 140(4):872– 886, April 2010. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S037837580900305X.

Kabaila:2012:MCP

[KF12] Paul Kabaila and Davide Farchione. The minimum coverage probability of confidence intervals in regression after a preliminary F test. Journal of Statistical Planning and Inference, 142(4):956–964, April 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375811004071.

Khaledi:2011:SCO

[KFK11] Baha-Eldin Khaledi, Sepideh Farsinezhad, and Subhash C. Kochar. Stochastic comparisons of order statistics in the scale model. Journal of Statistical Planning and Inference, 141(1):276–286, January 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375810002855.

Kedem:2010:SDF

[KG10a] Benjamin Kedem and Richard E. Gagnon. Semiparametric distribution forecasting. Journal of Statistical Planning and Inference, 140(12):3734–3741, December 2010. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810002181.

Kundu:2010:MSB

[KG10b] Debasis Kundu and Rameshwar D. Gupta. Modified Sarhan– Balakrishnan singular bivariate distribution. Journal of Statistical Planning and Inference, 140(2):526–538, February 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375809002468.

Kyung:2010:MLE

[KG10c] M. Kyung and S. K. Ghosh. Maximum likelihood estimation for directional conditionally autoregressive models. Journal of Statistical Planning and Inference, 140(11):3160–3179, November 2010. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375810001795.

Kessels:2010:OTL

[KGV10] Roselinde Kessels, Peter Goos, and Martina Vandebroek. Optimal two-level conjoint designs with constant attributes in the profile sets. Journal of Statistical Planning and Inference, 140(11):3035–3046, November 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375810001746.

Khattree:2010:APB

[Kha10] Ravindra Khattree. Antieigenvalues provide a bound on realized signal to noise ratio. Journal of Statistical Planning and Inference, 140(10):2846–2848, October 2010. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810001199.

Khademnoe:2016:PPB

[KHN16] Omid Khademnoe and S. Mohammad E. Hosseini-Nasab. On properties of percentile bootstrap confidence intervals for prediction in functional linear regression. Journal of Statistical Planning and Inference, 170(??):129–143, March 2016. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375815001639.

Kuiper:2012:GOR

[KHS12] R. M. Kuiper, H. Hoijtink, and M. J. Silvapulle. Generalization of the order-restricted information criterion for multivariate normal linear models. Journal of Statistical Planning and Inference, 142(8):2454–2463, August 2012. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375812001231.

Kim:2010:BRR

[Kim10] Hyune-Ju Kim. Bounding the resampling risk for sequential Monte Carlo implementation of hypothesis tests. Journal of Statistical Planning and Inference, 140(7):1834–1843, July 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375810000145.

Kim:2016:BCE

[Kim16] Jae H. Kim. Bias-correction and endogenous lag order algorithm for bootstrap prediction intervals. Journal of Statistical Planning and Inference, 177(??):41–44, October 2016. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375815001998.

Kim:2011:PCR

[KK11] Yongdai Kim and Dohyun Kim. Posterior consistency of random effects models for binary data. Journal of Statistical Planning and Inference, 141(11):3391–3399, November 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810005793.

Kohler:2012:PAO

[KK12] Michael Kohler and Adam Krzyzak. Pricing of American options in discrete time using least squares estimates with complexity penalties. Journal of Statistical Planning and Inference, 142(8):2289–2307, August 2012. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837581200078X.

Kirch:2015:UEF

[KK15] Claudia Kirch and Joseph Tadjuidje Kamgaing. On the use of estimating functions in monitoring time series for change points. Journal of Statistical Planning and Inference, 161(??): 25–49, June 2015. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375814002067.

Kim:2016:CMS

[KK16] Jeankyung Kim and Hyune-Ju Kim. Consistent model selection in segmented line regression. Journal of Statistical Planning and Inference, 170(??):106–116, March 2016. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375815001627.

Kulkarni:2019:OBP

[KK19] R. V. Kulkarni and V. G. Kulkarni. Optimal Bayes procedures for selecting the better of two Bernoulli populations. Journal of Statistical Planning and Inference, 15(??):311–330, ???? 2019. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375886901060.

Kalyagin:2017:RIR

[KKK17] Valery A. Kalyagin, Alexander P. Koldanov, and Petr A. Koldanov. Robust identification in random variable networks. Journal of Statistical Planning and Inference, 181(??):30–40, February 2017. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375816301033.

Kalyagin:2019:LFU

[KKKP19] Valery A. Kalyagin, Alexander P. Koldanov, Petr A. Koldanov, and Panos M. Pardalos. Loss function, unbiasedness, and optimality of Gaussian graphical model selection. Journal of Statistical Planning and Inference, 201(??):32– 39, July 2019. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375818303446.

Konomi:2015:BTM

[KKL15] Bledar Konomi, Georgios Karagiannis, and Guang Lin. On the Bayesian treed multivariate Gaussian process with linear model of coregionalization. Journal of Statistical Planning and Inference, 157–158(??):1–15, February/March 2015. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375814001608.

Kuruwita:2010:DEU

[KKP10] C. N. Kuruwita, K. B. Kulasekera, and W. J. Padgett. Density estimation using asymmetric kernels and Bayes bandwidths with censored data. Journal of Statistical Planning and Inference, 140(7):1765–1774, July 2010. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810000030.

Krishnamoorthy:2010:IFP

[KL10a] K. Krishnamoorthy and Meesook Lee. Inference for functions of parameters in discrete distributions based on fiducial approach: Binomial and Poisson cases. Journal of Statistical Planning and Inference, 140(5):1182–1192, May 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809003589.

Krishnamoorthy:2010:CLS

[KL10b] K. Krishnamoorthy and Yin Lin. Confidence limits for stressstrength reliability involving Weibull models. Journal of Statistical Planning and Inference, 140(7):1754–1764, July 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810000042.

Kulik:2011:SRR

[KL11] Rafał Kulik and Paweł Lorek. Some results on random design regression with long memory errors and predictors. Journal of Statistical Planning and Inference, 141(1):508–523, January 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S037837581000323X.

Kang:2012:BIP

[KL12] Moonsu Kang and Jaewon Lee. Bayesian inference for the proportion of true null hypotheses using minimum Hellinger distance. Journal of Statistical Planning and Inference, 142(4):820–825, April 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375811003788.

Kwon:2013:DLQ

[KL13] Koohyun Kwon and Sokbae Lee. Discussion of “Local quantile regression” by Spokoiny, Wang, and Härdle. Journal of Statistical Planning and Inference, 143(7):1136–1138, July 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375813000608. See [SWH13a].

Kline:2011:BFA

[Kli11] Brendan Kline. The Bayesian and frequentist approaches to testing a one-sided hypothesis about a multivariate mean. Journal of Statistical Planning and Inference, 141(9):3131–3141, September 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375811001315.

Kao:2011:RMS

[KLK11] Fei-Fei Kao, Ching-Ho Leu, and Chien-Hao Ko. Remainder Markov systematic sampling. Journal of Statistical Planning and Inference, 141(11):3595–3604, November 2011. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811002035.

Kim:2015:SBR

[KLKO15] Gwangsu Kim, Jeongran Lee, Yongdai Kim, and Hee-Seok Oh. Sparse Bayesian representation in time-frequency domain. Journal of Statistical Planning and Inference, 166(??):126–137, November 2015. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375815000385.

Khardani:2011:URS

[KLS11] Salah Khardani, Mohamed Lemdani, and Elias Ould Saïd. Uniform rate of strong consistency for a smooth kernel estimator of the conditional mode for censored time series. Journal of Statistical Planning and Inference, 141(11):3426–3436, November 2011. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375811002060.

Karunamuni:2010:REB

[KLW10] R. J. Karunamuni, J. Li, and J. Wu. Robust empirical Bayes tests for continuous distributions. Journal of Statistical Planning and Inference, 140(1):268–282, January 1, 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809002237.

Krinik:2010:BQS

[KM10] Alan Krinik and Sri Gopal Mohanty. On batch queueing systems: a combinatorial approach. Journal of Statistical Planning and Inference, 140(8):2271–2284, August 2010. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810000340.

Kunert:2011:ODI

[KM11] Joachim Kunert and Sabine Mersmann. Optimal designs for an interference model. Journal of Statistical Planning and Inference, 141(4):1623–1632, April 2011. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810005203.

Kortbi:2012:TLE

[KM12a] Othmane Kortbi and Éric Marchand. Truncated linear estimation of a bounded multivariate normal mean. Journal of Statistical Planning and Inference, 142(9):2607–2618, September 2012. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375812001462.

Koutras:2012:EAR

[KM12b] M. V. Koutras and F. S. Milienos. Exact and asymptotic results for pattern waiting times. Journal of Statistical Planning and Inference, 142(6):1464–1479, June 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375812000067.

Kunert:2010:OBD

[KME10] J. Kunert, R. J. Martin, and J. Eccleston. Optimal block designs comparing treatments with a control when the errors are correlated. Journal of Statistical Planning and Inference, 140(9):2719–2738, September 2010. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810001588.

Koukouvinos:2011:ASD

[KMMP11] C. Koukouvinos, E. Massou, K. Mylona, and C. Parpoula. Analyzing supersaturated designs with entropic measures. Journal of Statistical Planning and Inference, 141(3):1307– 1312, March 2011. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375810004702.

Karlsen:2010:NRE

[KMT10] Hans Arnfinn Karlsen, Terje Myklebust, and Dag Tjøstheim. Nonparametric regression estimation in a null recurrent time series. Journal of Statistical Planning and Inference, 140(12):3619–3626, December 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375810002090.

Kohler:2015:ERF

[KMW15] Michael Kohler, Florian Müller, and Harro Walk. Estimation of a regression function corresponding to latent variables. Journal of Statistical Planning and Inference, 162(??):88–109, July 2015. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375814002031.

Kustosz:2016:SSD

[KMW16] Christoph P. Kustosz, Christine H. Müller, and Martin Wendler. Simplified simplicial depth for regression and autoregressive growth processes. Journal of Statistical Planning and Inference, 173(??):125–146, June 2016. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375816000069.

Kundu:2010:SDR

[KNM10] Chanchal Kundu, Asok K. Nanda, and Sudhansu S. Maiti. Some distributional results through past entropy. Journal of Statistical Planning and Inference, 140(5):1280–1291, May 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375809003656.

Koenker:2013:CLQ

[Koe13] Roger Koenker. Comment on “Local quantile regression”. Journal of Statistical Planning and Inference, 143(7):1134– 1135, July 2013. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375813000591.

Koltchinskii:2013:DGC

[Kol13] Vladimir Koltchinskii. Discussion: “Generic chaining and the l_{1}-penalty” by Sara van de Geer. Journal of Statistical Planning and Inference, 143(6):1019–1021, June 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375812003667. See [vdG13b].

Komaki:2011:BPD

[Kom11] Fumiyasu Komaki. Bayesian predictive densities based on latent information priors. Journal of Statistical Planning and Inference, 141(12):3705–3715, December 2011. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811002345.

Kong:2012:CIJ

[Kon12] Xin-Bing Kong. Confidence interval of the jump activity index based on empirical likelihood using high frequency data. Journal of Statistical Planning and Inference, 142(6):1378– 1387, June 2012. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375811004617.

Kong:2013:PJP

[Kon13] Xin-Bing Kong. Is a pure jump process fitting the high frequency data better than a jump-diffusion process? Journal of Statistical Planning and Inference, 143(2):315–320, February 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375812002753.

Koul:2011:MDL

[Kou11] Hira L. Koul. Minimum distance lack-of-fit tests for fixed design. Journal of Statistical Planning and Inference, 141(1):65–79, January 2011. CODEN JSPIDN.

ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375810002661.

Krishnamoorthy:2011:ICF

[KP11] K. Krishnamoorthy and Jie Peng. Improved closed-form prediction intervals for binomial and Poisson distributions. Journal of Statistical Planning and Inference, 141(5):1709–1718, May 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375810005215.

Kozubowski:2011:NMM

[KPQ11] Tomasz J. Kozubowski, Anna K. Panorska, and Fares Qeadan. A new multivariate model involving geometric sums and maxima of exponentials. Journal of Statistical Planning and Inference, 141(7):2353–2367, July 2011. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811000437.

Kim:2014:FAM

[KPR14] Peter T. Kim, Shaun Pinder, and Stephen T. A. Rush. Fréchet analysis and the microbiome. Journal of Statistical Planning and Inference, 145(??):37–41, February 2014. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375813001845.

Kim:2013:SPS

[KPZ13] Inyoung Kim, Herbert Pang, and Hongyu Zhao. Statistical properties on semiparametric regression for evaluating pathway effects. Journal of Statistical Planning and Inference, 143(4):745–763, April 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375812003035.

Kundu:2012:BIP

[KR12] Debasis Kundu and Mohammad Z. Raqab. Bayesian inference and prediction of order statistics for a Type-II censored Weibull distribution. Journal of Statistical Planning and Inference, 142(1):41–47, January 2012. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811002448.

Krattenthaler:2010:DGC

[Kra10] C. Krattenthaler. Determinants of (generalised) Catalan numbers. Journal of Statistical Planning and Inference, 140(8):2260–2270, August 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375810000339.

Krajina:2012:MME

[Kra12] Andrea Krajina. A method of moments estimator of tail dependence in meta-elliptical models. Journal of Statistical Planning and Inference, 142(7):1811–1823, July 2012. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375812000298.

Kreiss:2016:DBP

[Kre16] Jens-Peter Kreiss. Discussion: “Bootstrap prediction intervals for linear, nonlinear and nonparametric autoregressions”. Journal of Statistical Planning and Inference, 177(??):28–30, October 2016. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375815001962. See [PP16a].

Krebs:2018:OSE

[Kre18] Johannes T. N. Krebs. Orthogonal series estimates on strong spatial mixing data. Journal of Statistical Planning and Inference, 193(??):15–41, February 2018. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375817301362.

Kristensen:2010:USR

[Kri10] Kristian Kristensen. The use of spatial and randomisationbased methods for analysis of trials with treatments randomised into rows and columns. Journal of Statistical Planning and Inference, 140(6):1542–1549, June 2010. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809003942.

Kroll:2019:NPP

[Kro19] Martin Kroll. Non-parametric Poisson regression from independent and weakly dependent observations by model selection. Journal of Statistical Planning and Inference, 199(??):249–270, March 2019. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375818301356.

Kvam:1993:IEA

[KS93] Paul H. Kvam and Francisco J. Samaniego. On the inadmissibility of empirical averages as estimators in ranked set sampling. Journal of Statistical Planning and Inference, 36(1):39–55, July 1993. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
037837589390100K. See [KS11b].

Koul:2010:CVM

[KS10a] Hira L. Koul and Weixing Song. Conditional variance model checking. Journal of Statistical Planning and Inference, 140(4):1056–1072, April 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375809003450.

Koul:2010:GFT

[KS10b] Hira L. Koul and Donatas Surgailis. Goodness-of-fit testing under long memory. Journal of Statistical Planning and Inference, 140(12):3742–3753, December 2010. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810002193.

Kubokawa:2011:NML

[KS11a] Tatsuya Kubokawa and William E. Strawderman. Nonminimaxity of linear combinations of restricted location estimators and related problems. Journal of Statistical Planning and Inference, 141(6):2141–2155, June 2011. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811000188.

Kvam:2011:EIS

[KS11b] P. H. Kvam and F. J. Samaniego. Erratum to “On the inadmissibility of standard estimators based on ranked set sampling” [Journal of Statistical Planning and Inference 36 (1993) 39–55]. Journal of Statistical Planning and Inference, 141(5):2001, May 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375810005021. See [KS93].

Kibria:2012:IEP

[KS12] B. M. Golam Kibria and A. K. Md. E. Saleh. Improving the estimators of the parameters of a probit regression model: a ridge regression approach. Journal of Statistical Planning and Inference, 142(6):1421–1435, June 2012. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837581100468X.

Koul:2016:MDP

[KS16] Hira L. Koul and Weixing Song. Minimum distance partial linear regression model checking with Berkson measurement errors. Journal of Statistical Planning and Inference, 174(??):129–152, July 2016. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375816000082.

Kariya:2019:ORT

[KS19a] Takeaki Kariya and Bimal K. Sinha. Optimality robustness of tests in two population problems. Journal of Statistical Planning and Inference, 15(??):167–176, ???? 2019. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0378375886900959.

Klaassen:2019:SEE

[KS19b] Chris A. J. Klaassen and Nanang Susyanto. Semiparametrically efficient estimation of Euclidean parameters under equality constraints. Journal of Statistical Planning and Inference, 201(??):120–132, July 2019. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375818303598.

Kupresanin:2010:RFF

[KSKE10] Ana Kupresanin, Hyejin Shin, David King, and R. L. Eubank. An RKHS framework for functional data analysis. Journal of Statistical Planning and Inference, 140(12):3627–3637, December 2010. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375810002107.

Kim:2011:RHF

[KSL11] Kion Kim, Damla Sentürk, and Runze Li. Recent history functional linear models for sparse longitudinal data. Journal of Statistical Planning and Inference, 141(4):1554–1566, April 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375810005033.

Keshavarz:2018:OCP

[KSN18] Hossein Keshavarz, Clayton Scott, and XuanLong Nguyen. Optimal change point detection in Gaussian processes. Journal of Statistical Planning and Inference, 193(??):151–178, February 2018. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S037837581730160X.

Karunanayaka:2017:CDU

[KT17] Ruwan Chamara Karunanayaka and Boxin Tang. Compromise designs under baseline parameterization. Journal of Statistical Planning and Inference, 190(??):32–38, November 2017. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375817300794.

Kitagawa:2014:ATS

[KU14] Hayato Kitagawa and Masayuki Uchida. Adaptive test statistics for ergodic diffusion processes sampled at discrete times. Journal of Statistical Planning and Inference, 150(??):84–110, July 2014. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S037837581400038X.

Kubokawa:2014:GDP

[Kub14] Tatsuya Kubokawa. General dominance properties of double shrinkage estimators for ratio of positive parameters. Journal of Statistical Planning and Inference, 147(??):224–234, April 2014. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375813002954.

Kutoyants:2011:GFT

[Kut11] Yury A. Kutoyants. Goodness-of-fit tests for perturbed dynamical systems. Journal of Statistical Planning and Inference, 141(5):1655–1666, May 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375810005057.

Kyriakoussis:2010:DDB

[KV10] Andreas Kyriakoussis and Malvina G. Vamvakari. q-discrete distributions based on q-Meixner and q-Charlier orthogonal polynomials — asymptotic behaviour. Journal of Statistical Planning and Inference, 140(8):2285–2294, August 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810000352.

Kharin:2011:REA

[KV11] Yuriy S. Kharin and Valeriy A. Voloshko. Robust estimation of ar coefficients under simultaneously influencing outliers and missing values. Journal of Statistical Planning and Inference, 141(9):3276–3288, September 2011. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811001583.

Karabatsos:2012:BNC

[KW12] George Karabatsos and Stephen G. Walker. A Bayesian nonparametric causal model. Journal of Statistical Planning and Inference, 142(4):925–934, April 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375811004022.

Karami:2014:RSD

[KW14] Jamil Hasan Karami and Douglas P. Wiens. Robust static designs for approximately specified nonlinear regression models. Journal of Statistical Planning and Inference, 144(??):55–62, January 2014. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375812002959.

Kremer:2014:MLE

[KWL14] Alexander Kremer, Rafael Weißbach, and Friedrich Liese. Maximum likelihood estimation for left-censored survival times in an additive hazard model. Journal of Statistical Planning and Inference, 149(??):33–45, June 2014. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375814000354.

Kim:2019:QSE

[KWS19] Hyungwoo Kim, Yichao Wu, and Seung Jun Shin. Quantileslicing estimation for dimension reduction in regression. Journal of Statistical Planning and Inference, 198(??):1–12, January 2019. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375818300272.

Kochar:2011:TBC

[KX11a] Subhash Kochar and Maochao Xu. The tail behavior of the convolutions of gamma random variables. Journal of Statistical Planning and Inference, 141(1):418–428, January 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810003125.

Krishnamoorthy:2011:TIS

[KX11b] K. Krishnamoorthy and Fang Xie. Tolerance intervals for symmetric location-scale families based on uncensored or censored samples. Journal of Statistical Planning and Inference, 141(3):1170–1182, March 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375810004295.

Kutoyants:2014:ABS

[KZ14] Yury A. Kutoyants and Li Zhou. On approximation of the backward stochastic differential equation. Journal of Statistical Planning and Inference, 150(??):111–123, July 2014. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375814000378. See corrigendum [KZ15].

Kutoyants:2015:CAB

[KZ15] Yury A. Kutoyants and Li Zhou. Corrigendum to “On approximation of the backward stochastic differential equation” [J. Statist. Plann. Inference 150 (2014) 111–123]. Journal of Statistical Planning and Inference, 163(??):48, August 2015. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375814001037. See [KZ14].

Korolev:2016:NVM

[KZ16] V. Yu. Korolev and A. I. Zeifman. On normal variance-mean mixtures as limit laws for statistics with random sample sizes. Journal of Statistical Planning and Inference, 169(??):34–42, February 2016. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S037837581500141X.

Kim:2010:NSS

[KZW10] Kun Ho Kim, Zhou Zhou, and Wei Biao Wu. Non-stationary structural model with time-varying demand elasticities. Journal of Statistical Planning and Inference, 140(12):3809–3819, December 2010. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375810002259.

Lwin:2014:SCF

[LA14] Thaung Lwin and David Albrecht. Sampling of a continuous flow in the plane based on two-dimensional variogram models. Journal of Statistical Planning and Inference, 146(??):70–81, March 2014. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375813002097.

Lacko:2014:UEE

[Lac14] Vladimír Lacko. Ultimate efficiency of experimental designs for Ornstein–Uhlenbeck type processes. Journal of Statistical Planning and Inference, 149(??):77–89, June 2014. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375814000111.

Liu:2010:OSC

[LAk10] Wei Liu and Pascal Ah-kine. Optimal simultaneous confidence bands in simple linear regression. Journal of Statistical Planning and Inference, 140(5):1225–1235, May 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809003590.

Lagos-Alvarez:2019:KFM

[LÁPMF19] Bernardo Lagos-Álvarez, Leonardo Padilla, Jorge Mateu, and Guillermo Ferreira. A Kalman filter method for estimation and prediction of space-time data with an autoregressive structure. Journal of Statistical Planning and Inference, 203(??):117–130, December 2019. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375819300278.

Laurent:2012:IBI

[Lau12] Stéphane Laurent. Intrinsic Bayesian inference on a Poisson rate and on the ratio of two Poisson rates. Journal of Statistical Planning and Inference, 142(9):2656–2671, September 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375812000870.

Lavancier:2011:AIA

[Lav11] Frédéric Lavancier. Aggregation of isotropic autoregressive fields. Journal of Statistical Planning and Inference, 141(12):3862–3866, December 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S037837581100228X.

Li:2011:ODT

[LB11] Gang Li and N. Balakrishnan. Optimal designs for tumor regrowth models. Journal of Statistical Planning and Inference, 141(2):644–654, February 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375810003332.

Li:2015:EEL

[LB15] Huiqin Li and Zhidong Bai. Extreme eigenvalues of large dimensional quaternion sample covariance matrices. Journal of Statistical Planning and Inference, 159(??):1–14, April 2015. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375814001773.

Liu:2017:GCA

[LBH17] Zhongqiang Liu, Tao Ban, and Tao Huang. General covariateadaptive randomization targeting unequal allocation ratio. Journal of Statistical Planning and Inference, 191(??):68–80, December 2017. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375817300885.

Lee:2012:SML

[LBL12] Chihoon Lee, Jaya P. N. Bishwal, and Myung Hee Lee. Sequential maximum likelihood estimation for reflected Ornstein–Uhlenbeck processes. Journal of Statistical Planning and Inference, 142(5):1234–1242, May 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375811004320.

Lee:2012:JTA

[LBR12] Hyunsook Lee, G. Jogesh Babu, and C. R. Rao. A jackknife type approach to statistical model selection. Journal of Statistical Planning and Inference, 142(1):301–311, January 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375811003028.

Lin:2011:GFT

[LC11] Kuo-Chin Lin and Yi-Ju Chen. A goodness-of-fit test for logistic-normal models using nonparametric smoothing method. Journal of Statistical Planning and Inference, 141(2):1069–1076, February 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375810004271.

Luo:2013:EBL

[LC13] Shan Luo and Zehua Chen. Extended BIC for linear regression models with diverging number of relevant features and high or ultra-high feature spaces. Journal of Statistical Planning and Inference, 143(3):494–504, March 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375812002832.

Li:2015:CWU

[LC15] Chang Li and Daniel C. Coster. Construction of weak universal optimal block designs with various correlation structures and block sizes. Journal of Statistical Planning and Inference, 160(??):1–10, May 2015. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375814001931.

Li:2016:ASS

[LC16] Gengxin Li and Yuehua Cui. Assessing statistical significance in variance components linkage analysis: a theoretical justification. Journal of Statistical Planning and Inference, 178(??):70–83, November 2016. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375816300593.

Lai:2017:SIB

[LCC^{+}17] Yi-Hsuan Lai, Hung-Chia Chen, Lin-An Chen, Dung-Tsa Chen, and Hui-Nien Hung. Statistical inferences based on outliers for gene expression analysis. Journal of Statistical Planning and Inference, 187(??):130–142, August 2017. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375817300356.

Li:2011:SIP

[LCL11] Degui Li, Jia Chen, and Zhengyan Lin. Statistical inference in partially time-varying coefficient models. Journal of Statistical Planning and Inference, 141(2):995–1013, February 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375810004155.

Li:2013:EPS

[LCQ^{+}13] Weiming Li, Jiaqi Chen, Yingli Qin, Zhidong Bai, and Jianfeng Yao. Estimation of the population spectral distribution from a large dimensional sample covariance matrix. Journal of Statistical Planning and Inference, 143(11):1887–1897, November 2013. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375813001626.

Lai:2010:LTM

[LCT10] Yi-Hsuan Lai, Lin-An Chen, and Chau-Shyun Tang. Linear trimmed means for the linear regression with AR(1) errors model. Journal of Statistical Planning and Inference, 140(11):3457–3467, November 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375810002521.

Liang:2012:EBS

[LCY12] TaChen Liang, Lee-Shen Chen, and Ming-Chung Yang. Efficient Bayesian sampling plans for exponential distributions with random censoring. Journal of Statistical Planning and Inference, 142(2):537–551, February 2012. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811003235.

Lorenz:2012:CWT

[LD12] Douglas J. Lorenz and Somnath Datta. Comparing waiting times in a multi-stage model: a log-rank approach. Journal of Statistical Planning and Inference, 142(10):2832–2843, October 2012. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S037837581200153X.

Leu:2014:BSA

[LD14] Richard Leu and Paul Damien. Bayesian shape analysis of the complex Bingham distribution. Journal of Statistical Planning and Inference, 149(??):183–200, June 2014. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375813003029.

Liang:2011:CQE

[LdUÁ11] Han-Ying Liang and Jacobo de Uña-Álvarez. Conditional quantile estimation with auxiliary information for lefttruncated and dependent data. Journal of Statistical Planning and Inference, 141(11):3475–3488, November 2011. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811002047.

Le:2013:IDS

[Le13] Trung Kien Le. Information dependency: Strong consistency of Darbellay–Vajda partition estimators. Journal of Statistical Planning and Inference, 143(12):2089–2100, December 2013. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375813001870.

Leblanc:2012:BPB

[Leb12] Alexandre Leblanc. On the boundary properties of Bernstein polynomial estimators of density and distribution functions. Journal of Statistical Planning and Inference, 142(10):2762–2778, October 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375812001322.

Lecue:2013:CGC

[Lec13] Guillaume Lecué. Comment to “Generic chaining and the l_{1}-penalty” by Sara van de Geer. Journal of Statistical Planning and Inference, 143(6):1022–1025, June 2013. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375812003655.

Lee:2018:LTE

[Lee18] Ji Hyung Lee. Limit theory for explosive autoregression under conditional heteroskedasticity. Journal of Statistical Planning and Inference, 196(??):30–55, August 2018. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375817301921.

Lehner:2011:CII

[Leh11] Franz Lehner. Cumulants as iterated integrals. Journal of Statistical Planning and Inference, 141(4):1448–1454, April 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375810004805.

Lemonte:2011:LPS

[Lem11] Artur J. Lemonte. Local power of some tests in exponential family nonlinear models. Journal of Statistical Planning and Inference, 141(5):1981–1989, May 2011. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810005689.

Lemonte:2012:LPP

[Lem12] Artur J. Lemonte. Local power properties of some asymptotic tests in symmetric linear regression models. Journal of Statistical Planning and Inference, 142(5):1178–1188, May 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375811004241.

Lewis:2018:CIB

[Lew18] Nicholas Lewis. Combining independent Bayesian posteriors into a confidence distribution, with application to estimating climate sensitivity. Journal of Statistical Planning and Inference, 195(??):80–92, May 2018. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375817301702.

Lemonte:2011:SLR

[LF11] Artur J. Lemonte and Silvia L. P. Ferrari. Signed likelihood ratio tests in the Birnbaum-Saunders regression model. Journal of Statistical Planning and Inference, 141(2):1031–1040, February 2011. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375810004180.

Liu:2012:RFP

[LG12] Juxin Liu and Paul Gustafson. Regression with fractional polynomials when interactions are erroneously omitted. Journal of Statistical Planning and Inference, 142(6):1348–1355, June 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375811004575.

Lyles:2012:RBM

[LGG12] Robert H. Lyles, Ying Guo, and Sander Greenland. Reducing bias and mean squared error associated with regressionbased odds ratio estimators. Journal of Statistical Planning and Inference, 142(12):3235–3241, December 2012. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375812001954.

Labra:2012:EDH

[LGLO12] Filidor V. Labra, Aldo M. Garay, Victor H. Lachos, and Edwin M. M. Ortega. Estimation and diagnostics for heteroscedastic nonlinear regression models based on scale mixtures of skew-normal distributions. Journal of Statistical Planning and Inference, 142(7):2149–2165, July 2012. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375812000651.

Lin:2011:MCG

[LH11] G. D. Lin and J. S. Huang. Maximum correlation for the generalized Sarmanov bivariate distributions. Journal of Statistical Planning and Inference, 141(8):2738–2749, August 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811000826.

Li:2014:DEL

[LH14] Yiou Li and Fred J. Hickernell. Design of experiments for linear regression models when gradient information is available. Journal of Statistical Planning and Inference, 144(??):141–151, January 2014. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S037837581300027X.

Lin:2016:GPL

[LHH16] Huazhen Lin, Ye He, and Jian Huang. A global partial likelihood estimation in the additive Cox proportional hazards model. Journal of Statistical Planning and Inference, 169(??):71–87, February 2016. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375815001433.

Li:2011:ELM

[Li11] Zaixing Li. Estimation in linear mixed models for longitudinal data under linear restricted conditions. Journal of Statistical Planning and Inference, 141(2):869–876, February 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375810003770.

Li:2012:LFT

[Li12] Xiaoyu Li. Lack-of-fit testing of a regression model with response missing at random. Journal of Statistical Planning and Inference, 142(1):155–170, January 2012. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811002758.

Lian:2011:PDB

[Lia11a] Heng Lian. On posterior distribution of Bayesian wavelet thresholding. Journal of Statistical Planning and Inference, 141(1):318–324, January 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375810002958.

Lian:2011:STP

[Lia11b] Heng Lian. Shrinkage tuning parameter selection in precision matrices estimation. Journal of Statistical Planning and Inference, 141(8):2839–2848, August 2011. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811001054.

Liang:2011:CSE

[Lia11c] TaChen Liang. Comments on “Simultaneous estimation of Poisson means of the selected subset”. Journal of Statistical Planning and Inference, 141(9):3289–3291, September 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375811001339.

Lian:2012:ELC

[Lia12] Heng Lian. Empirical likelihood confidence intervals for nonparametric functional data analysis. Journal of Statistical Planning and Inference, 142(7):1669–1677, July 2012. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375812000559.

Lian:2014:ARC

[Lia14a] Heng Lian. Adaptive rates of contraction of posterior distributions in Bayesian wavelet regression. Journal of Statistical Planning and Inference, 145(??):92–101, February 2014. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375813002255.

Lian:2014:SAP

[Lia14b] Heng Lian. Some asymptotic properties for functional canonical correlation analysis. Journal of Statistical Planning and Inference, 153(??):1–10, October 2014. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375814000949.

Lian:2015:FSD

[Lia15] Heng Lian. Functional sufficient dimension reduction: Convergence rates and multiple functional case. Journal of Statistical Planning and Inference, 167(??):58–68, December 2015. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375815001135.

Lin:2010:WEE

[Lin10] Pei-Sheng Lin. A working estimating equation for spatial count data. Journal of Statistical Planning and Inference, 140(9):2470–2477, September 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375810000923.

Lin:2011:CDD

[Lin11] Wei Lin. On completely data-driven bandwidth selection for single-index models. Journal of Statistical Planning and Inference, 141(5):1673–1683, May 2011. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810005185.

Lin:2012:ASF

[Lin12] Pei-Sheng Lin. Analysis of spatial frailty models by a weighted estimating equation. Journal of Statistical Planning and Inference, 142(6):1436–1444, June 2012. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837581200002X.

Lin:2014:OBO

[Lin14] Chang-Yun Lin. Optimal blocked orthogonal arrays. Journal of Statistical Planning and Inference, 145(??):139–147, February 2014. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375813002115.

Liu:2011:ELE

[Liu11a] Qiang Liu. Estimation of the linear EV model with censored data. Journal of Statistical Planning and Inference, 141(7):2463–2471, July 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375811000656.

Liu:2011:ILE

[Liu11b] Xu-Qing Liu. Improved Liu estimator in a linear regression model. Journal of Statistical Planning and Inference, 141(1):189–196, January 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375810002776.

Lorchirachoonkul:2012:MSS

[LJ12] Vichit Lorchirachoonkul and Jirawan Jitthavech. A modified C_{p} statistic in a system-of-equations model. Journal of Statistical Planning and Inference, 142(8):2386–2394, August 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375812001012.

Li:2015:SSF

[LJA15] Kang Li, Bochuan Jiang, and Mingyao Ai. Sliced spacefilling designs with different levels of two-dimensional uniformity. Journal of Statistical Planning and Inference, 157–158(??):90–99, February/March 2015. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375814001682.

Li:2012:SAM

[LJK12] Judy X. Li, Daniel R. Jeske, and Jeffrey A. Klein. Sequential analysis methodology for a Poisson GLMM with applications to multicenter randomized clinical trials. Journal of Statistical Planning and Inference, 142(12):3225–3234, December 2012. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375812001760.

Laib:2011:RSC

[LL11a] Naâmane Laib and Djamal Louani. Rates of strong consistencies of the regression function estimator for functional stationary ergodic data. Journal of Statistical Planning and Inference, 141(1):359–372, January 2011. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810002880.

Li:2011:TRC

[LL11b] Shuang Li and James Lynch. On a threshold representation for complex load-sharing systems. Journal of Statistical Planning and Inference, 141(8):2811–2823, August 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811001030.

Liao:2011:BCN

[LL11c] Wenhui Liao and Yan Liu. Bound and convergence of the non-constant dynamic linear model. Journal of Statistical Planning and Inference, 141(1):602–609, January 2011. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810003599.

Liu:2011:COS

[LL11d] Yan Liu and Min-Qian Liu. Construction of optimal supersaturated design with large number of levels. Journal of Statistical Planning and Inference, 141(6):2035–2043, June 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810005732.

Lee:2013:REM

[LL13] Woojoo Lee and Youngjo Lee. Random-effect models with singular precision. Journal of Statistical Planning and Inference, 143(12):2128–2141, December 2013. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375813001602.

Liu:2015:VCM

[LL15] Shu Liu and Guodong Li. Varying-coefficient meancovariance regression analysis for longitudinal data. Journal of Statistical Planning and Inference, 160(??):89–106, May 2015. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S037837581400202X.

Li:2019:TPC

[LL19] Wei Li and Johannes Lederer. Tuning parameter calibration for ℓ_{1}-regularized logistic regression. Journal of Statistical Planning and Inference, 202(??):80–98, September 2019. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375819300060.

Li:2013:AIS

[LLJK13] Cui-Xia Li, Xiao-Lin Liang, Bing-Yi Jing, and Xin-Bing Kong. The asymptotics of the integrated self-weighted cross volatility estimator. Journal of Statistical Planning and Inference, 143(10):1708–1718, October 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375813001067.

Lee:2012:OTH

[LLL^{+}12] Sang H. Lee, Johan Lim, Erning Li, Marina Vannucci, and Eva Petkova. Order test for high-dimensional twosample means. Journal of Statistical Planning and Inference, 142(9):2719–2725, September 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S037837581200105X.

Lu:2018:EEP

[LLL18] Minggen Lu, Tao Lu, and Chin-Shang Li. Efficient estimation of partially linear additive Cox model under monotonicity constraint. Journal of Statistical Planning and Inference, 192(??):18–34, January 2018. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375817301349.

Liu:2019:CMU

[LLLP19] Guanfu Liu, Pengfei Li, Yukun Liu, and Xiaolong Pu. On consistency of the MLE under finite mixtures of location-scale distributions with a structural parameter. Journal of Statistical Planning and Inference, 199(??):29–44, March 2019. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837581830051X.

Laurent:2011:TIP

[LLM11] B. Laurent, J.-M. Loubes, and C. Marteau. Testing inverse problems: a direct or an indirect problem? Journal of Statistical Planning and Inference, 141(5):1849–1861, May 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810005355.

Lloyd:2010:PVB

[Llo10] Chris J. Lloyd. P-values based on approximate conditioning and p*. Journal of Statistical Planning and Inference, 140(4):1073–1081, April 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375809003449.

Liu:2014:SCB

[LLP14] Lin Liu, Chu-In Charles Lee, and Jianan Peng. Simultaneous confidence bounds in monotone dose-response studies. Journal of Statistical Planning and Inference, 145(??):113–124, February 2014. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375813002085.

Liu:2019:CMC

[LLS19] Zhifan Liu, Chunling Liu, and Zhihua Sun. Consistent model check of errors-in-variables varying-coefficient model with auxiliary variable. Journal of Statistical Planning and Inference, 198(??):13–28, January 2019. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375818300284.

Ling:2015:NRE

[LLV15] Nengxiang Ling, Longlong Liang, and Philippe Vieu. Nonparametric regression estimation for functional stationary ergodic data with missing at random. Journal of Statistical Planning and Inference, 162(??):75–87, July 2015. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375815000075.

Liu:2016:GVS

[LLW16] Xianhui Liu, Yuanyuan Lin, and Zhanfeng Wang. Group variable selection for relative error regression. Journal of Statistical Planning and Inference, 175(??):40–50, August 2016. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375816000276.

Liu:2012:SEW

[LLX12] Wanrong Liu, Xuewen Lu, and Changchun Xie. Semiparametric estimation for weighted average derivatives with responses missing at random. Journal of Statistical Planning and Inference, 142(1):347–357, January 2012. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811003090.

Lin:2011:SBC

[LLZ11a] Lu Lin, Feng Li, and Lixing Zhu. Simulation-based consistent inference for biased working model of non-sparse highdimensional linear regression. Journal of Statistical Planning and Inference, 141(12):3780–3792, December 2011. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811002394.

Liu:2011:DEA

[LLZ11b] Xu Liu, Peixin Liu, and Yong Zhou. Distribution estimation with auxiliary information for missing data. Journal of Statistical Planning and Inference, 141(2):711–724, February 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375810003629.

Liu:2017:STM

[LLZS17] Zhongying Liu, Baisen Liu, Shurong Zheng, and Ning-Zhong Shi. Simultaneous testing of mean vector and covariance matrix for high-dimensional data. Journal of Statistical Planning and Inference, 188(??):82–93, September 2017. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375817300551.

Laha:2011:SRD

[LM11a] Arnab Kumar Laha and K. C. Mahesh. SB-robustness of directional mean for circular distributions. Journal of Statistical Planning and Inference, 141(3):1269–1276, March 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810004386.

Lee:2011:CBE

[LM11b] Juhee Lee and Steven N. MacEachern. Consistency of Bayes estimators without the assumption that the model is correct. Journal of Statistical Planning and Inference, 141(2):748–757, February 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375810003721.

Li:2013:EIV

[LM13] Tizheng Li and Changlin Mei. Estimation and inference for varying coefficient partially nonlinear models. Journal of Statistical Planning and Inference, 143(11):2023–2037, November 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375813001146.

Lin:2014:DVR

[LM14] C. Devon Lin and Susanne Morrill. Design of variable resolution for model selection. Journal of Statistical Planning and Inference, 155(??):127–134, December 2014. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375814001013.

Li:2016:EBD

[LM16] Yan Li and Yajun Mei. Effect of bivariate data’s correlation on sequential tests of circular error probability. Journal of Statistical Planning and Inference, 171(??):99–114, April 2016. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375815002001.

Liao:2017:CPE

[LM17a] Xiyue Liao and Mary C. Meyer. Change-point estimation using shape-restricted regression splines. Journal of Statistical Planning and Inference, 188(??):8–21, September 2017. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375817300423.

Lopuhaa:2017:ISE

[LM17b] Hendrik P. Lopuhaä and Eni Musta. Isotonized smooth estimators of a monotone baseline hazard in the Cox model. Journal of Statistical Planning and Inference, 191(??):43–67, December 2017. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375817301064.

Li:2019:ODM

[LM19a] Qiyu Li and Ilya Molchanov. Optimal design for multivariate multiple linear regression with set-identified response. Journal of Statistical Planning and Inference, 203(??):215–223, December 2019. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375819300357.

Loonis:2019:DSD

[LM19b] V. Loonis and X. Mary. Determinantal sampling designs. Journal of Statistical Planning and Inference, 199(??):60–88, March 2019. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375818300533.

LMoudden:2017:PDE

[LMKS17] Aziz L’Moudden, Éric Marchand, Othmane Kortbi, and William E. Strawderman. On predictive density estimation for gamma models with parametric constraints. Journal of Statistical Planning and Inference, 185(??):56–68, June 2017. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375817300113.

Li:2014:ASB

[LMT14] Pei Li, Arden Miller, and Boxin Tang. Algorithmic search for baseline minimum aberration designs. Journal of Statistical Planning and Inference, 149(??):172–182, June 2014. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375814000299.

Loeppky:2010:BSD

[LMW10] Jason L. Loeppky, Leslie M. Moore, and Brian J. Williams. Batch sequential designs for computer experiments. Journal of Statistical Planning and Inference, 140(6):1452–1464, June 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375809003747.

Loeppky:2012:PAB

[LMW12] Jason L. Loeppky, Leslie M. Moore, and Brian J. Williams. Projection array based designs for computer experiments. Journal of Statistical Planning and Inference, 142(6):1493– 1505, June 2012. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375812000080.

Li:2018:CQC

[LMZ18] Xiangjie Li, Xuejun Ma, and Jingxiao Zhang. Conditional quantile correlation screening procedure for ultrahighdimensional varying coefficient models. Journal of Statistical Planning and Inference, 197(??):69–92, December 2018. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375818300016.

Ley:2015:EIA

[LN15] Christophe Ley and Anouk Neven. Efficient inference about the tail weight in multivariate student t distributions. Journal of Statistical Planning and Inference, 167(??):123–134, December 2015. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375815000981.

Luts:2011:AHT

[LNdDLdCH11] Jan Luts, José Antonio Roldán Nofuentes, Juan de Dios Luna del Castillo, and Sabine Van Huffel. Asymptotic hypothesis test to compare likelihood ratios of multiple diagnostic tests in unpaired designs. Journal of Statistical Planning and Inference, 141(11):3578–3594, November 2011. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811002023.

Laitila:2011:TPS

[LO11a] Thomas Laitila and Jens Olofsson. A two-phase sampling scheme and πps designs. Journal of Statistical Planning and Inference, 141(5):1646–1654, May 2011. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810005045.

Lampio:2011:CDM

[LÖ11b] Pekka H. J. Lampio and Patric R. J. Östergård. Classification of difference matrices over cyclic groups. Journal of Statistical Planning and Inference, 141(3):1194–1207, March 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375810004349.

Loh:2010:BIP

[Loh10] Ji Meng Loh. Bootstrapping an inhomogeneous point process. Journal of Statistical Planning and Inference, 140(3):734– 749, March 2010. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375809002845.

Lombardo:2011:TWA

[Lom11] Rosaria Lombardo. Three-way association measure decompositions: the Delta index. Journal of Statistical Planning and Inference, 141(5):1789–1799, May 2011. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837581000529X.

Long:2012:NGF

[Lon12] Qi Long. A note on generalized functional linear model and its application. Journal of Statistical Planning and Inference, 142(9):2599–2606, September 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375812000742.

Longla:2017:RLT

[Lon17] Martial Longla. Remarks on limit theorems for reversible Markov processes and their applications. Journal of Statistical Planning and Inference, 187(??):28–43, August 2017. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375817300320.

Lefevre:2011:PSR

[LP11a] Claude Lefèvre and Philippe Picard. Polynomial structures in rank statistics distributions. Journal of Statistical Planning and Inference, 141(4):1380–1393, April 2011. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810004763.

Li:2011:ELT

[LP11b] Minqiang Li and Liang Peng. Empirical likelihood test via estimating equations. Journal of Statistical Planning and Inference, 141(7):2428–2439, July 2011. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811000589.

Lee:2012:GAS

[LP12] Jinsuk Lee and Rong Pan. A GLM approach to step-stress accelerated life testing with interval censoring. Journal of Statistical Planning and Inference, 142(4):810–819, April 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811003764.

Li:2010:BRH

[LPY10] Deyuan Li, Liang Peng, and Jingping Yang. Bias reduction for high quantiles. Journal of Statistical Planning and Inference, 140(9):2433–2441, September 2010. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810000984.

Lei:2011:ELQ

[LQ11] Qingzhu Lei and Yongsong Qin. Empirical likelihood for quantiles under negatively associated samples. Journal of Statistical Planning and Inference, 141(3):1325–1332, March 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S037837581000474X.

Lu:2016:SAH

[LQ16] Qiyi Lu and Xingye Qiao. Significance analysis of highdimensional, low-sample size partially labeled data. Journal of Statistical Planning and Inference, 176(??):78–94, September 2016. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375815300537.

Li:2017:CTL

[LQ17] Hongyi Li and Hong Qin. Construction on two-level factorials with flexible partially replicated runs via quarter foldover. Journal of Statistical Planning and Inference, 188(??):55–66, September 2017. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S037837581730054X.

Li:2012:CIP

[LQL12] Yinghua Li, Yongsong Qin, and Qingzhu Lei. Confidence intervals for probability density functions under associated samples. Journal of Statistical Planning and Inference, 142(6):1516–1524, June 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375812000109.

Leiva:2011:LDM

[LR11] Ricardo Leiva and Anuradha Roy. Linear discrimination for multi-level multivariate data with separable means and jointly equicorrelated covariance structure. Journal of Statistical Planning and Inference, 141(5):1910–1924, May 2011.

CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810005434.

Le-Rademacher:2011:LFS

[LRB11] J. Le-Rademacher and L. Billard. Likelihood functions and some maximum likelihood estimators for symbolic data. Journal of Statistical Planning and Inference, 141(4):1593–1602, April 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375810005161.

Leonenko:2010:SIR

[LS10a] N. N. Leonenko and N. Suvak. Statistical inference for reciprocal gamma diffusion process. Journal of Statistical Planning and Inference, 140(1):30–51, January 1, 2010. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837580900189X.

Liu:2010:NEF

[LS10b] Fang Liu and Sanat K. Sarkar. A note on estimating the false discovery rate under mixture model. Journal of Statistical Planning and Inference, 140(6):1601–1609, June 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809003796.

Litvinova:2012:TWF

[LS12] Svetlana Litvinova and Mervyn J. Silvapulle. A test when the Fisher information may be infinite, exemplified by a test for marginal independence in extreme value distributions. Journal of Statistical Planning and Inference, 142(6):1506–1515, June 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375812000092.

Lim:2012:AUH

[LSP12a] Changwon Lim, Pranab K. Sen, and Shyamal D. Peddada. Accounting for uncertainty in heteroscedasticity in nonlinear regression. Journal of Statistical Planning and Inference, 142(5):1047–1062, May 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375811004083.

Lopez:2012:BCR

[LSP12b] Olivier Lopez and Philippe Saint-Pierre. Bivariate censored regression relying on a new estimator of the joint distribution function. Journal of Statistical Planning and Inference, 142(8):2440–2453, August 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375812000936.

Lin:2012:CCT

[LST12] C. Devon Lin, Randy R. Sitter, and Boxin Tang. Creating catalogues of two-level nonregular fractional factorial designs based on the criteria of generalized aberration. Journal of Statistical Planning and Inference, 142(2):445–456, February 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375811003144.

Lu:2013:EMS

[LSW13] Xiaomin Lu, Anqi Sun, and Samuel S. Wu. On estimating the mean of the selected normal population in two-stage adaptive designs. Journal of Statistical Planning and Inference, 143(7):1215–1220, July 2013. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375813000281.

Lin:2016:SUE

[LSWY16] Lu Lin, Yufeng Shi, Xin Wang, and Shuzhen Yang. ksample upper expectation linear regression — modeling, identifiability, estimation and prediction. Journal of Statistical Planning and Inference, 170(??):15–26, March 2016. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375815001573.

Lazariv:2013:CCM

[LSZ13] Taras Lazariv, Wolfgang Schmid, and Svitlana Zabolotska. On control charts for monitoring the variance of a time series. Journal of Statistical Planning and Inference, 143(9):1512–1526, September 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375813000736.

Liu:2016:NAM

[LSZ16] Yanping Liu, Sanat K. Sarkar, and Zhigen Zhao. A new approach to multiple testing of grouped hypotheses. Journal of Statistical Planning and Inference, 179(??):1–14, December 2016. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375816300854.

Liu:2010:PSC

[LT10] Xiao Liu and Loon-Ching Tang. Planning sequential constant-stress accelerated life tests with stepwise loaded auxiliary acceleration factor. Journal of Statistical Planning and Inference, 140(7):1968–1985, July 2010. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810000698.

Langel:2011:SIQ

[LT11] Matti Langel and Yves Tillé. Statistical inference for the quintile share ratio. Journal of Statistical Planning and Inference, 141(8):2976–2985, August 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375811001200.

Liu:2013:VMS

[LT13] Y. Liu and J. A. Tawn. Volatility model selection for extremes of financial time series. Journal of Statistical Planning and Inference, 143(3):520–530, March 2013. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375812002777.

Lekivetz:2014:MLO

[LT14a] Ryan Lekivetz and Boxin Tang. Multi-level orthogonal arrays for estimating main effects and specified interactions. Journal of Statistical Planning and Inference, 144(??):123–132, January 2014. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375812003308.

Liao:2014:MBT

[LT14b] Chen-Tuo Liao and Shin-Fu Tsai. Minimum breakdown twolevel orthogonal main-effect plans. Journal of Statistical Planning and Inference, 155(??):135–140, December 2014. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375814001219.

Lindqvist:2018:PTI

[LT18] Bo H. Lindqvist and Gunnar Taraldsen. On the proper treatment of improper distributions. Journal of Statistical Planning and Inference, 195(??):93–104, May 2018. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375817301659.

Lu:2013:ISE

[Lu13] Lu Lu. On the integrated squared error of the linear wavelet density estimator. Journal of Statistical Planning and Inference, 143(9):1548–1565, September 2013. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375813000682.

Lukas:2014:PCD

[Luk14] Mark A. Lukas. Performance criteria and discrimination of extreme undersmoothing in nonparametric regression. Journal of Statistical Planning and Inference, 153(??):56–74, October 2014. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375814000809.

Li:2010:FMC

[LVK10] Feng Li, Mattias Villani, and Robert Kohn. Flexible modeling of conditional distributions using smooth mixtures of asymmetric student t densities. Journal of Statistical Planning and Inference, 140(12):3638–3654, December 2010. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810002119.

Landsman:2013:NSL

[LVY13] Zinoviy Landsman, Steven Vanduffel, and Jing Yao. A note on Stein’s lemma for multivariate elliptical distributions. Journal of Statistical Planning and Inference, 143(11):2016–2022, November 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375813001407.

Liu:2010:SAR

[LW10] Wei Liu and Lang Wu. Some asymptotic results for semiparametric nonlinear mixed-effects models with incomplete data. Journal of Statistical Planning and Inference, 140(1):52–64, January 1, 2010. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375809001888.

Lai:2011:PLS

[LW11a] Peng Lai and Qihua Wang. Partially linear single-index model with missing responses at random. Journal of Statistical Planning and Inference, 141(2):1047–1058, February 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375810004234.

Lin:2011:BIJ

[LW11b] Tsung-I Lin and Wan-Lun Wang. Bayesian inference in joint modelling of location and scale parameters of the t distribution for longitudinal data. Journal of Statistical Planning and Inference, 141(4):1543–1553, April 2011. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837581000501X.

Li:2012:WLS

[LW12a] Xiayan Li and Qihua Wang. The weighted least square based estimators with censoring indicators missing at random. Journal of Statistical Planning and Inference, 142(11):2913–2925, November 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375812001784.

Li:2012:VPR

[LW12b] Xuan Li and Xikui Wang. Variance-penalized responseadaptive randomization with mismeasurement. Journal of Statistical Planning and Inference, 142(7):2128–2135, July 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375812000638.

Luo:2015:ELB

[LW15] Ruimiao Luo and Qihua Wang. Empirical likelihood based weighted GMM estimation with missing response at random. Journal of Statistical Planning and Inference, 156(??):64–79, January 2015. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375814001384.

Lee:2019:DEC

[LW19] Kim May Lee and James Wason. Design of experiments for a confirmatory trial of precision medicine. Journal of Statistical Planning and Inference, 199(??):179–187, March 2019. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375818300880.

Lwin:2011:PEF

[Lwi11] Thaung Lwin. Parameter estimation in first-order autoregressive model for statistical process monitoring in the presence of data autocorrelation. Journal of Statistical Planning and Inference, 141(8):2556–2575, August 2011. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811000620.

Liu:2013:APM

[LWZ13] Guangying Liu, Zhengyuan Wei, and Xinsheng Zhang. Asymptotic properties for multipower variation of semimartingales and Gaussian integral processes with jumps. Journal of Statistical Planning and Inference, 143(8):1307– 1319, August 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375813000323.

Liu:2018:EEL

[LX18] Huilan Liu and Xiaochao Xia. Estimation and empirical likelihood for single-index multiplicative models. Journal of Statistical Planning and Inference, 193(??):70–88, February 2018. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375817301404.

Lai:2012:VTI

[LYC12] Yi-Hsuan Lai, Ya-Fen Yen, and Lin-An Chen. Validation of tolerance interval. Journal of Statistical Planning and Inference, 142(4):902–907, April 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375811004010.

Liu:2014:NOD

[LYC14] Xin Liu, Rong-Xian Yue, and Kashinath Chatterjee. A note on R-optimal designs for multi-factor models. Journal of Statistical Planning and Inference, 146(??):139–144, March 2014. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375813002590.

Liu:2015:MRO

[LYC15] Xin Liu, Rong-Xian Yue, and Kashinath Chatterjee. Modelrobust R-optimal designs in linear regression models. Journal of Statistical Planning and Inference, 167(??):135–143, December 2015. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S037837581500097X.

Lane:2014:ITS

[LYF14] Adam Lane, Ping Yao, and Nancy Flournoy. Information in a two-stage adaptive optimal design. Journal of Statistical Planning and Inference, 144(??):173–187, January 2014. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837581300164X.

Li:2012:PIW

[LYH12] Ying Li, Li Yu, and Taizhong Hu. Probability inequalities for weighted distributions. Journal of Statistical Planning and Inference, 142(5):1272–1278, May 2012. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811004289.

Liu:2011:ICF

[LYL11] Yang Liu, Jian-Feng Yang, and Min-Qian Liu. Isomorphism check in fractional factorial designs via letter interaction pattern matrix. Journal of Statistical Planning and Inference, 141(9):3055–3062, September 2011. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811001248.

Liu:2013:ODP

[LYL13] Xin Liu, Rong-Xian Yue, and Dennis K. J. Lin. Optimal design for prediction in multiresponse linear models based on rectangular confidence region. Journal of Statistical Planning and Inference, 143(11):1954–1967, November 2013. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375813001493.

Liu:2012:RBE

[LYLZ12] Tianqing Liu, Xiaohui Yuan, Nan Lin, and Baoxue Zhang. Rank-based empirical likelihood inference on medians of k populations. Journal of Statistical Planning and Inference, 142(4):1009–1026, April 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375811004149.

Liu:2016:ACO

[LYXC16] Xin Liu, Rong-Xian Yue, Jing Xu, and Kashinath Chatterjee. Algorithmic construction of R-optimal designs for second-order response surface models. Journal of Statistical Planning and Inference, 178(??):61–69, November 2016. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375816300428.

Li:2012:GPL

[LZ12a] Jianbo Li and Riquan Zhang. General partially linear varying-coefficient transformation model with right censored data. Journal of Statistical Planning and Inference, 142(6):1285–1293, June 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375811004484.

Liu:2012:SEL

[LZ12b] Xiaoxia Liu and Yichuan Zhao. Semi-empirical likelihood inference for the ROC curve with missing data. Journal of Statistical Planning and Inference, 142(12):3123–3133, December 2012. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375812002121.

Li:2013:TCA

[LZ13a] Xiaoye Li and Zhibiao Zhao. Testing for changes in autocovariances of nonparametric time series models. Journal of Statistical Planning and Inference, 143(2):237–250, February 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375812002492.

Li:2013:NTR

[LZ13b] Zaixing Li and Lixing Zhu. A new test for random effects in linear mixed models with longitudinal data. Journal of Statistical Planning and Inference, 143(1):82–95, January 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375812002364.

Lv:2011:NEV

[LZH11] Yazhao Lv, Riquan Zhang, and Zhensheng Huang. Nonparametric estimation of varying coefficient error-in-variable models with validation sampling. Journal of Statistical Planning and Inference, 141(10):3323–3344, October 2011. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811001601.

Li:2010:LBE

[LZL^{+}10] Qizhai Li, Gang Zheng, Aiyi Liu, Shifeng Xiong, Zhaohai Li, and Kai Yu. The limiting bound of Efron’s W - formula for hypothesis testing when a nuisance parameter is present only under the alternative. Journal of Statistical Planning and Inference, 140(6):1610–1617, June 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809004017.

Ling:2012:SNC

[LZL12] Xiaoliang Ling, Peng Zhao, and Xiaohu Li. On a system of nonhomogeneous components sharing a common frailty. Journal of Statistical Planning and Inference, 142(6):1330– 1338, June 2012. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375811004551.

Liu:2016:CDA

[LZL16] Xuhua Liu, Junlong Zhao, and Na Li. A confidence distribution approach to inferring the among-group variance component in one-way random effects model with unequal error variances. Journal of Statistical Planning and Inference, 171(??): 79–91, April 2016. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375815002025.

Li:2015:EDB

[LZZ15] William Li, Qi Zhou, and Runchu Zhang. Effective designs based on individual word length patterns. Journal of Statistical Planning and Inference, 163(??):43–47, August 2015. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375815000130.

Mitra:2008:SAT

[MA08] Murari Mitra and M. Z. Anis. An L-statistic approach to a test of exponentiality against IFR alternatives. Journal of Statistical Planning and Inference, 138(10):3144–3148, October 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375807004272.

Mudholkar:2010:DTA

[MA10] Govind S. Mudholkar and Kobby O. Asubonteng. Datatransformation approach to lifetimes data analysis: an overview. Journal of Statistical Planning and Inference, 140(10):2904–2917, October 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375810001266.

Mai:2017:PBQ

[MA17] The Tien Mai and Pierre Alquier. Pseudo-Bayesian quantum tomography with rank-adaptation. Journal of Statistical Planning and Inference, 184(??):62–76, May 2017. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375816301501.

Mabon:2016:ADL

[Mab16] Gwennaëlle Mabon. Adaptive deconvolution of linear functionals on the nonnegative real line. Journal of Statistical Planning and Inference, 178(??):1–23, November 2016. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375816300283.

Macci:2010:LDB

[Mac10] Claudio Macci. Large deviations for Bayesian estimators in first-order autoregressive processes. Journal of Statistical Planning and Inference, 140(9):2790–2797, September 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810000960.

Machete:2013:CPS

[Mac13] Reason L. Machete. Contrasting probabilistic scoring rules. Journal of Statistical Planning and Inference, 143(10):1781–1790, October 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375813001158.

Madi:2010:NEE

[Mad10] Mohamed T. Madi. A note on the equivariant estimation of an exponential scale using progressively censored data. Journal of Statistical Planning and Inference, 140(6):1437–1440, June 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375809003760.

Mainassara:2011:MPT

[Mai11] Y. Boubacar Mainassara. Multivariate portmanteau test for structural VARMA models with uncorrelated but nonindependent error terms. Journal of Statistical Planning and Inference, 141(8):2961–2975, August 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375811001194.

Majumdar:2014:GPS

[Maj14] Anandamayee Majumdar. Gaussian processes on the support of cylindrical surfaces, with application to periodic spatio-temporal data. Journal of Statistical Planning and Inference, 153(??):27–41, October 2014. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375814000755.

Mammen:2013:DNS

[Mam13] Enno Mammen. Discussion of “Nonparametric (smoothed) likelihood and integral equations”, by Piet Groeneboom. Journal of Statistical Planning and Inference, 143(12):2072– 2073, December 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375813001560.

See [Gro13a].

Martin:2018:IMC

[Mar18] Ryan Martin. On an inferential model construction using generalized associations. Journal of Statistical Planning and Inference, 195(??):105–115, May 2018. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375816301537.

Mukherjee:2011:SPS

[MB11] Amitava Mukherjee and Uttam Bandyopadhyay. Some partially sequential nonparametric tests for detecting linear trend. Journal of Statistical Planning and Inference, 141(8):2645–2655, August 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375811000747.

Mandal:2011:MDE

[MBB11] A. Mandal, S. K. Bhandari, and A. Basu. Minimum disparity estimation based on combined disparities: Asymptotic results. Journal of Statistical Planning and Inference, 141(2):701–710, February 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375810003642.

Mojirsheibani:2011:CWC

[MC11] Majid Mojirsheibani and Shojaeddin Chenouri. Classification when the covariate vectors have unequal dimensions. Journal of Statistical Planning and Inference, 141(5):1944–1957, May 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375810005653.

Martinez-Camblor:2012:GBA

[MCC12] Pablo Martínez-Camblor and Norberto Corral. A general bootstrap algorithm for hypothesis testing. Journal of Statistical Planning and Inference, 142(2):589–600, February 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811003478.

McNicholas:2010:MBC

[McN10] Paul D. McNicholas. Model-based classification using latent Gaussian mixture models. Journal of Statistical Planning and Inference, 140(5):1175–1181, May 2010. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809003607.

Maturi:2010:NPC

[MCSC10] Tahani A. Maturi, Pauline Coolen-Schrijner, and Frank P. A. Coolen. Nonparametric predictive comparison of lifetime data under progressive censoring. Journal of Statistical Planning and Inference, 140(2):515–525, February 2010. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809002456.

McGree:2012:ABC

[MDT^{+}12] J. M. McGree, C. C. Drovandi, M. H. Thompson, J. A. Eccleston, S. B. Duffull, K. Mengersen, A. N. Pettitt, and T. Goggin. Adaptive Bayesian compound designs for dose finding studies. Journal of Statistical Planning and Inference, 142(6):1480–1492, June 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375812000079.

Meister:2011:GCD

[Mei11] Alexander Meister. On general consistency in deconvolution mode estimation. Journal of Statistical Planning and Inference, 141(2):771–781, February 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375810003678.

Meinshausen:2013:DGS

[Mei13] Nicolai Meinshausen. Discussion of “Grouping Strategies and Thresholding for High Dimension Linear Models”. Journal of Statistical Planning and Inference, 143(9):1439–1440, September 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375813000505. See [MPT13a].

Mendo:2012:AOE

[Men12] L. Mendo. Asymptotically optimum estimation of a probability in inverse binomial sampling under general loss functions. Journal of Statistical Planning and Inference, 142(10):2862–2870, October 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375812001504.

Mendy:2013:PES

[Men13] Ibrahima Mendy. Parametric estimation for sub-fractional Ornstein–Uhlenbeck process. Journal of Statistical Planning and Inference, 143(4):663–674, April 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S037837581200328X.

McElroy:2010:SBM

[MF10] Tucker S. McElroy and David F. Findley. Selection between models through multi-step-ahead forecasting. Journal of Statistical Planning and Inference, 140(12):3655–3675, December 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375810002120.

Martinez-Florez:2013:MAP

[MFABG13] Guillermo Martínez-Flórez, Barry C. Arnold, Heleno Bolfarine, and Héctor W. Gómez. The multivariate alphapower model. Journal of Statistical Planning and Inference, 143(7):1244–1255, July 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375813000256.

Farahmand:2012:RLS

[mFS12] Amir massoud Farahmand and Csaba Szepesvári. Regularized least-squares regression: Learning from a βmixing sequence. Journal of Statistical Planning and Inference, 142(2):493–505, February 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375811003181.

Mee:2011:STS

[MFW11] Robert W. Mee, Joseph J. Ford, and Samuel S. Wu. Step-up test with a single cutoff for unreplicated orthogonal designs. Journal of Statistical Planning and Inference, 141(1):325–334, January 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375810002946.

Menendez:2010:RCP

[MGB10] Patricia Menéndez, Sucharita Ghosh, and Jan Beran. On rapid change points under long memory. Journal of Statistical Planning and Inference, 140(11):3343–3354, November 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375810002314.

Morgan:2014:JSE

[MGD14] J. P. Morgan, Subir Ghosh, and A. M. Dean. J. N. Srivastava and experimental design. Journal of Statistical Planning and Inference, 144(??):3–18, January 2014. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375812003011.

Maghsoodloo:2010:COT

[MH10] Saeed Maghsoodloo and Ching-Ying Huang. Comparing the overlapping of two independent confidence intervals with a single confidence interval for two normal population parameters. Journal of Statistical Planning and Inference, 140(11):3295–3305, November 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375810002405.

McGinniss:2016:MIT

[MH16] J. McGinniss and O. Harel. Multiple imputation in three or more stages. Journal of Statistical Planning and Inference, 176(??):33–51, September 2016. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375815300720.

Mulder:2010:EIC

[MHK10] Joris Mulder, Herbert Hoijtink, and Irene Klugkist. Equality and inequality constrained multivariate linear models: Objective model selection using constrained posterior priors. Journal of Statistical Planning and Inference, 140(4):887– 906, April 2010. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375809003127.

Mirakhmedov:2010:GGF

[Mir10] Sherzod M. Mirakhmedov. On Greenwood goodness-of-fit test. Journal of Statistical Planning and Inference, 140(11):3017–3025, November 2010. CODEN JSPIDN.

ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375810001692.

Miyata:2010:LAM

[Miy10] Yoichi Miyata. Laplace approximations to means and variances with asymptotic modes. Journal of Statistical Planning and Inference, 140(2):382–392, February 2010. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809002705.

Miyata:2011:MLE

[Miy11] Yoichi Miyata. Maximum likelihood estimators in finite mixture models with censored data. Journal of Statistical Planning and Inference, 141(1):56–64, January 2011. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810002430.

Miyata:2018:LAU

[Miy18] Yoichi Miyata. Laplace approximations using n α-consistent estimators. Journal of Statistical Planning and Inference, 194(??):25–31, March 2018. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375817301787.

Ma:2012:NEC

[MJS12] Tiefeng Ma, Lijie Jia, and Yingsheng Su. A new estimator of covariance matrix. Journal of Statistical Planning and Inference, 142(2):529–536, February 2012. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811003223.

Maaziz:2018:PEM

[MK18] M. Maaziz and S. Kharfouchi. Parameter estimation of Markov switching bilinear model using the (em) algorithm. Journal of Statistical Planning and Inference, 192(??):35–44, January 2018. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375817301222.

Matsuura:2015:OEP

[MKT15] Shun Matsuura, Hiroshi Kurata, and Thaddeus Tarpey. Optimal estimators of principal points for minimizing expected mean squared distance. Journal of Statistical Planning and Inference, 167(??):102–122, December 2015. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837581500110X.

Meyer:2018:CRC

[MKW18] Mary C. Meyer, Soo-Young Kim, and Haonan Wang. Convergence rates for constrained regression splines. Journal of Statistical Planning and Inference, 193(??):179–188, February 2018. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375817301593.

Meemark:2011:CBM

[MLa11] Yotsanan Meemark and Worrawate Leela-apiradee. A change of basis matrix and integrals of power of cosine. Journal of Statistical Planning and Inference, 141(3):1319–1324, March 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375810004738.

Maistre:2017:PNC

[MLP17] Samuel Maistre, Pascal Lavergne, and Valentin Patilea. Powerful nonparametric checks for quantile regression. Journal of Statistical Planning and Inference, 180(??):13–29, January 2017. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375816300921.

Misra:2012:SPC

[MM12] Amit Kumar Misra and Neeraj Misra. Stochastic properties of conditionally independent mixture models. Journal of Statistical Planning and Inference, 142(6):1599–1607, June 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375812000213.

McElroy:2016:TIE

[MN16] Tucker McElroy and Chaitra H. Nagaraja. Tail index estimation with a fixed tuning parameter fraction. Journal of Statistical Planning and Inference, 170(??):27–45, March 2016. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837581500155X.

Mexia:2010:ALL

[MO10a] João T. Mexia and Manuela M. Oliveira. Asymptotic linearity and limit distributions, approximations. Journal of Statistical Planning and Inference, 140(2):353–357, February 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S037837580900295X.

Moala:2010:EMP

[MO10b] Fernando A. Moala and Anthony O’Hagan. Elicitation of multivariate prior distributions: a nonparametric Bayesian approach. Journal of Statistical Planning and Inference, 140(7):1635–1655, July 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375810000157.

Mohanty:2010:P

[Moh10] Sri Gopal Mohanty. Preface. Journal of Statistical Planning and Inference, 140(8):2156, August 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375810000212.

Mojirsheibani:2012:CRF

[Moj12] Majid Mojirsheibani. On the correct regression function (in L _{2} ) and its applications when the dimension of the covariate vector is random. Journal of Statistical Planning and Inference, 142(9):2586–2598, September 2012. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375812001334.

Montanari:2017:MMR

[Mon17] Giorgio E. Montanari. Mixed model regression estimation of a spatial total in the continuous plane paradigm. Journal of Statistical Planning and Inference, 185(??):1–14, June 2017. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375816301513.

Morris:2014:MDO

[Mor14] Max D. Morris. Maximin distance optimal designs for computer experiments with time-varying inputs and outputs. Journal of Statistical Planning and Inference, 144(??):63–68, January 2014. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375812002996.

Meintanis:2011:STM

[MP11] Simos G. Meintanis and Stephen Portnoy. Specification tests in mixed effects models. Journal of Statistical Planning and Inference, 141(8):2545–2555, August 2011. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811000619.

Makri:2012:CCB

[MP12a] Frosso S. Makri and Zaharias M. Psillakis. Counting certain binary strings. Journal of Statistical Planning and Inference, 142(4):908–924, April 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375811004046.

Matthews:2012:ODT

[MP12b] J. N. S. Matthews and P. M. Philipson. Optimal designs for threshold-determining limiting dilution assays. Journal of Statistical Planning and Inference, 142(6):1388–1395, June 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375811004629.

Marella:2019:MIT

[MP19] Daniela Marella and Danny Pfeffermann. Matching information from two independent informative samples. Journal of Statistical Planning and Inference, 203(??):70–81, December 2019. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375818303082.

Marcon:2017:MNE

[MPN^{+}17] G. Marcon, S. A. Padoan, P. Naveau, P. Muliere, and J. Segers. Multivariate nonparametric estimation of the Pickands dependence function using Bernstein polynomials. Journal of Statistical Planning and Inference, 183(??):1–17, April 2017. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375816301276.

Mulero:2010:NAS

[MPRG10] Julio Mulero, Franco Pellerey, and Rosario Rodríguez-Griñolo. Negative aging and stochastic comparisons of residual lifetimes in multivariate frailty models. Journal of Statistical Planning and Inference, 140(6):1594–1600, June 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810000029.

Mougeot:2013:GSTa

[MPT13a] Mathilde Mougeot, Dominique Picard, and Karine Tribouley. Grouping strategies and thresholding for high dimensional linear models. Journal of Statistical Planning and Inference, 143(9):1417–1438, September 2013. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375813000463. See discussion [Mei13,

Obo13, vdG13a, MY13].

Mougeot:2013:GSTb

[MPT13b] Mathilde Mougeot, Dominique Picard, and Karine Tribouley. Grouping strategies and thresholding for high dimensional linear models rejoinder. Journal of Statistical Planning and Inference, 143(9):1457–1465, September 2013. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375813000529.

Muller:2010:RPM

[MQ10] Peter Müller and Fernando Quintana. Random partition models with regression on covariates. Journal of Statistical Planning and Inference, 140(10):2801–2808, October 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837581000114X.

Montes-Rojas:2010:TRE

[MR10] Gabriel V. Montes-Rojas. Testing for random effects and serial correlation in spatial autoregressive models. Journal of Statistical Planning and Inference, 140(4):1013–1020, April 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375809003164.

Ma:2011:LAC

[MR11a] Chunsheng Ma and John Robinson. Lot acceptance and compliance testing based on the sample mean and minimum/maximum. Journal of Statistical Planning and Inference, 141(7):2440–2448, July 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375811000607.

Malinovsky:2011:BIM

[MR11b] Yaakov Malinovsky and Yosef Rinott. Best invariant and minimax estimation of quantiles in finite populations. Journal of Statistical Planning and Inference, 141(8):2633–2644, August 2011. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375811000735.

Mena:2011:GSB

[MRW11] Ramsés H. Mena, Matteo Ruggiero, and Stephen G. Walker. Geometric stick-breaking processes for continuoustime Bayesian nonparametric modeling. Journal of Statistical Planning and Inference, 141(9):3217–3230, September 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811001510.

Mokhtari:2011:IWD

[MRY11] Elham Bayat Mokhtari, A. Habibi Rad, and F. Yousefzadeh. Inference for Weibull distribution based on progressively Type-II hybrid censored data. Journal of Statistical Planning and Inference, 141(8):2824–2838, August 2011. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811001042.

Moradi:2010:AAS

[MS10] Mohammad Moradi and Mohammad Salehi. An adaptive allocation sampling design for which the conventional stratified estimator is an appropriate estimator. Journal of Statistical Planning and Inference, 140(4):1030–1037, April 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809003188.

Majerski:2011:NAE

[MS11] P. Majerski and Z. Szkutnik. A note on asymptotic expansions for the power of perturbed tests. Journal of Statistical Planning and Inference, 141(12):3736–3743, December 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375811002461.

Maity:2012:TSI

[MS12a] Arnab Maity and Michael Sherman. Testing for spatial isotropy under general designs. Journal of Statistical Planning and Inference, 142(5):1081–1091, May 2012. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811004186.

McNicholas:2012:CGE

[MS12b] Paul D. McNicholas and Sanjeena Subedi. Clustering gene expression time course data using mixtures of multivariate t-distributions. Journal of Statistical Planning and Inference, 142(5):1114–1127, May 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375811004332.

Murugiah:2012:SPP

[MS12c] Siva Murugiah and Trevor Sweeting. Selecting the precision parameter prior in Dirichlet process mixture models. Journal of Statistical Planning and Inference, 142(7):1947–1959, July 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375812000602.

Maruyama:2013:IRB

[MS13a] Yuzo Maruyama and William E. Strawderman. Improved robust Bayes estimators of the error variance in linear models. Journal of Statistical Planning and Inference, 143(6):1091– 1097, June 2013. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375813000244.

Mazumder:2013:RSS

[MS13b] Satyaki Mazumder and Robert Serfling. A robust sample spatial outlyingness function. Journal of Statistical Planning and Inference, 143(1):144–159, January 2013. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375812002303.

Muller:2018:EHR

[MS18] Ursula U. Müller and Anton Schick. Efficiency for heteroscedastic regression with responses missing at random. Journal of Statistical Planning and Inference, 196(??):132– 143, August 2018. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375817302070.

Meintanis:2014:PWC

[MSA14] Simos G. Meintanis, Jan Swanepoel, and James Allison. The probability weighted characteristic function and goodness-of-fit testing. Journal of Statistical Planning and Inference, 146(??):122–132, March 2014. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375813002358.

Mallik:2016:AVB

[MSBM16] Atul Mallik, Bodhisattva Sen, Moulinath Banerjee, and George Michailidis. Asymptotics for p-value based threshold estimation under repeated measurements. Journal of Statistical Planning and Inference, 174(??):85–103, July 2016. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375816000197.

Monteiro:2010:IVA

[MSP10] Magda Monteiro, Manuel G. Scotto, and Isabel Pereira. Integer-valued autoregressive processes with periodic structure. Journal of Statistical Planning and Inference, 140 (6):1529–1541, June 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375809003954.

Manes:2011:CSH

[MSTT11] K. Manes, A. Sapounakis, I. Tasoulas, and P. Tsikouras. Counting strings at height j in Dyck paths. Journal of Statistical Planning and Inference, 141(6):2100–2107, June 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837581000580X.

Muller:2012:EED

[MSW12] Ursula U. Müller, Anton Schick, and Wolfgang Wefelmeyer. Estimating the error distribution function in semiparametric additive regression models. Journal of Statistical Planning and Inference, 142(2):552–566, February 2012. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811003430.

Muhlenstadt:2010:SBS

[Müh10] Thomas Mühlenstädt. Simplex based space filling designs. Journal of Statistical Planning and Inference, 140(3):585– 596, March 2010. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375809002481.

Muller:2012:EDP

[Mül12] Ursula U. Müller. Estimating the density of a possibly missing response variable in nonlinear regression. Journal of Statistical Planning and Inference, 142(5):1198–1214, May 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811004265.

Molenberghs:2011:WGF

[MV11a] Geert Molenberghs and Geert Verbeke. On the Weibullgamma frailty model, its infinite moments, and its connection to generalized log-logistic, logistic, Cauchy, and extreme-value distributions. Journal of Statistical Planning and Inference, 141(2):861–868, February 2011. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810003782.

Montanari:2011:DRM

[MV11b] Angela Montanari and Cinzia Viroli. Dimensionally reduced mixtures of regression models. Journal of Statistical Planning and Inference, 141(5):1744–1752, May 2011. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810005240.

Mameli:2015:HOA

[MV15] Valentina Mameli and Laura Ventura. Higher-order asymptotics for scoring rules. Journal of Statistical Planning and Inference, 165(??):13–26, October 2015. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375815000567.

Ma:2015:SNR

[MX15] Shujie Ma and Shizhong Xu. Semiparametric nonlinear regression for detecting gene and environment interactions. Journal of Statistical Planning and Inference, 156(??):31–47, January 2015. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375814001554.

Ma:2010:EIE

[MY10] Tiefeng Ma and Rendao Ye. Efficient improved estimation of the parameters in two seemingly unrelated regression models. Journal of Statistical Planning and Inference, 140(9):2749–2754, September 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375810001576.

Ma:2011:SBK

[MY11] Shujie Ma and Lijian Yang. Spline-backfitted kernel smoothing of partially linear additive model. Journal of Statistical Planning and Inference, 141(1):204–219, January 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810002752.

Mairal:2013:DAG

[MY13] J. Mairal and B. Yu. Discussion about “Grouping strategies and thresholding for high dimensional linear models”. Journal of Statistical Planning and Inference, 143(9):1451–1453, September 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375813000475. See [MPT13a].

Ma:2011:FSP

[MYJ11] Tiefeng Ma, Rendao Ye, and Lijie Jia. Finite-sample properties of the Graybill–Deal estimator. Journal of Statistical Planning and Inference, 141(11):3675–3680, November 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811002291.

Matsumoto:2011:IQC

[MYW11] C. Matsumoto, H. Yanagihara, and H. Wakaki. Improvement of the quality of the chi-square approximation for the ADF test on a covariance matrix with a linear structure. Journal of Statistical Planning and Inference, 141(4):1535–1542, April 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375810005124.

Miranda:2010:CMC

[MZ10] Enrique Miranda and Marco Zaffalon. Conditional models: Coherence and inference through sequences of joint mass functions. Journal of Statistical Planning and Inference, 140(7):1805–1833, July 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375810000169.

Mytalas:2013:CLT

[MZ13] George C. Mytalas and Michael A. Zazanis. Central Limit Theorem approximations for the number of runs in Markovdependent binary sequences. Journal of Statistical Planning and Inference, 143(2):321–333, February 2013. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375812002625.

Moghaddass:2012:ODM

[MZP12] Ramin Moghaddass, Ming J. Zuo, and Mayank Pandey. Optimal design and maintenance of a repairable multi-state system with standby components. Journal of Statistical Planning and Inference, 142(8):2409–2420, August 2012. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375812000924.

Maronge:2017:ODS

[MZWF17] Jacob M. Maronge, Yi Zhai, Douglas P. Wiens, and Zhide Fang. Optimal designs for spline wavelet regression models. Journal of Statistical Planning and Inference, 184(??):94–104, May 2017. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375816301525.

Nagy:2015:CMD

[Nag15] Stanislav Nagy. Consistency of h-mode depth. Journal of Statistical Planning and Inference, 165(??):91–103, October 2015. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375815000786.

Nandi:2012:EPT

[Nan12] Swagata Nandi. Estimation of parameters of two-dimensional sinusoidal signal in heavy-tailed errors. Journal of Statistical Planning and Inference, 142(10):2799–2808, October 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375812001528.

Ng:2010:PTT

[NB10] H. K. T. Ng and N. Balakrishnan. Precedence-type test based on Kaplan–Meier estimator of cumulative distribution function. Journal of Statistical Planning and Inference, 140(8):2295–2311, August 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375810000364.

Nagatsuka:2012:CPE

[NB12a] Hideki Nagatsuka and N. Balakrishnan. A consistent parameter estimation in the three-parameter lognormal distribution. Journal of Statistical Planning and Inference, 142(7):2071– 2086, July 2012. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375812000286.

Nagatsuka:2012:PQE

[NB12b] Hideki Nagatsuka and N. Balakrishnan. Parameter and quantile estimation for the three-parameter gamma distribution based on statistics invariant to unknown location. Journal of Statistical Planning and Inference, 142(7):2087–2102, July 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375812000274.

Nandram:2013:BTI

[NBSB13] Balgobin Nandram, Dilli Bhatta, Joe Sedransk, and Dhiman Bhadra. A Bayesian test of independence in a two-way contingency table using surrogate sampling. Journal of Statistical Planning and Inference, 143(8):1392–1408, August 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375813000566.

Nie:2011:AVM

[NCC11] Lei Nie, Yong Chen, and Haitao Chu. Asymptotic variances of maximum likelihood estimator for the correlation coefficient from a BVN distribution with one variable subject to censoring. Journal of Statistical Planning and Inference, 141(1):392–401, January 2011. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810003149.

Nanda:2011:DPH

[ND11] Asok K. Nanda and Suchismita Das. Dynamic proportional hazard rate and reversed hazard rate models. Journal of Statistical Planning and Inference, 141(6):2108–2119, June 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811000024.

Navarro:2010:SNR

[NdAA10] Jorge Navarro, Yolanda del Aguila, and Majid Asadi. Some new results on the cumulative residual entropy. Journal of Statistical Planning and Inference, 140(1):310–322, January 1, 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375809002250.

Nofuentes:2010:CAM

[NdCJ10] José Antonio Roldán Nofuentes, Juan Dios Luna del Castillo, and Ana Eugenia Marín Jimenez. Comparison of the accuracy of multiple binary tests in the presence of partial disease verification. Journal of Statistical Planning and Inference, 140(9):2504–2519, September 2010. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810001291.

Ndao:2016:NEC

[NDD16] Pathé Ndao, Aliou Diop, and Jean-François Dupuy. Nonparametric estimation of the conditional extreme-value index with random covariates and censoring. Journal of Statistical Planning and Inference, 168(??):20–37, January 2016.

CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375815001305.

Niu:2016:CNC

[NGXZ16] Cuizhen Niu, Xu Guo, Wangli Xu, and Lixing Zhu. Checking nonparametric component for partial linear regression model with missing response. Journal of Statistical Planning and Inference, 168(??):1–19, January 2016. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375815001317.

Nishiyama:2013:TLH

[NHSP13] Takahiro Nishiyama, Masashi Hyodo, Takashi Seo, and Tatjana Pavlenko. Testing linear hypotheses of mean vectors for high-dimension data with unequal covariance matrices. Journal of Statistical Planning and Inference, 143(11):1898–1911, November 2013. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375813001614.

Ni:2012:PLS

[Ni12]  Huey-Fan Ni.  Penalized least-squares estimation for

regression coefficients in high-dimensional partially linear models. Journal of Statistical Planning and Inference, 142(2):379–389, February 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375811001029.

Niki:2011:OSO

[NIHY11] N. Niki, M. Iwata, H. Hashiguchi, and S. Yamada. Optimal selection and ordering of columns in supersaturated designs. Journal of Statistical Planning and Inference, 141(7):2449– 2462, July 2011. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375811000644.

Nikolov:2012:SAW

[Nik12] Nikolai Nikolov. Self-avoiding walks on Z×0,1. Journal of Statistical Planning and Inference, 142(1):376–377, January 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S037837581100303X.

Nikoloulopoulos:2013:ENC

[Nik13] Aristidis K. Nikoloulopoulos. On the estimation of normal copula discrete regression models using the continuous extension and simulated likelihood. Journal of Statistical Planning and Inference, 143(11):1923–1937, November 2013. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375813001523.

Nilson:2011:PYD

[Nil11] Tomas Nilson. Pseudo-Youden designs balanced for intersection. Journal of Statistical Planning and Inference, 141(6):2030–2034, June 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375810005720.

Nishiyama:2011:ERD

[Nis11] Yoichi Nishiyama. On Z-estimation by rounded data. Journal of Statistical Planning and Inference, 141(1):287–292, January 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375810002843.

Ng:2016:CNN

[NJ16] Chi Tim Ng and Harry Joe. Comparison of non-nested models under a general measure of distance. Journal of Statistical Planning and Inference, 170(??):166–185, March 2016. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375815001883.

Nkurunziza:2010:TCH

[Nku10] Sévérien Nkurunziza. Testing concerning the homogeneity of some predator-prey populations. Journal of Statistical Planning and Inference, 140(1):323–333, January 1, 2010. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809002274.

Nkwanta:2010:RMH

[Nkw10] Asamoah Nkwanta. Riordan matrices and higher-dimensional lattice walks. Journal of Statistical Planning and Inference, 140(8):2321–2334, August 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375810000388.

Nordman:2014:REL

[NL14] Daniel J. Nordman and Soumendra N. Lahiri. A review of empirical likelihood methods for time series. Journal of Statistical Planning and Inference, 155(??):1–18, December 2014. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375813002528.

Nordman:2016:DBP

[NL16] Daniel J. Nordman and Soumendra N. Lahiri. A discussion of “Bootstrap prediction intervals for linear, nonlinear, and nonparametric autoregressions” by L. Pan and D. N. Politis. Journal of Statistical Planning and Inference, 177(??):35–40, October 2016. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375815001986. See [PP16a].

Nematollahi:2012:EPS

[NMS12] N. Nematollahi and F. Motamed-Shariati. Estimation of the parameter of the selected uniform population under the entropy loss function. Journal of Statistical Planning and Inference, 142(7):2190–2202, July 2012. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375812000250.

Novikov:2010:LMP

[NN10] Andrey Novikov and Petr Novikov. Locally most powerful sequential tests of a simple hypothesis vs. one-sided alternatives. Journal of Statistical Planning and Inference, 140(3):750–765, March 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375809002857.

Nogales:2013:IMK

[Nog13] A. G. Nogales. On independence of Markov kernels and a generalization of two theorems of Basu. Journal of Statistical Planning and Inference, 143(3):603–610, March 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375812002765.

Nowak:2012:OPP

[Now12] Piotr Nowak. Order preserving property of moment estimators. Journal of Statistical Planning and Inference, 142(7):1849–1854, July 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375812000584.

Nandi:2010:EFA

[NPK10] Swagata Nandi, Anurag Prasad, and Debasis Kundu. An efficient and fast algorithm for estimating the parameters of two-dimensional sinusoidal signals. Journal of Statistical Planning and Inference, 140(1):153–168, January 1, 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809002171.

Nayak:2010:CRU

[NQ10] Tapan K. Nayak and Min Qin. The concept of risk unbiasedness in statistical prediction. Journal of Statistical Planning and Inference, 140(7):1923–1938, July 2010. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810000467.

Nguimkeu:2011:TOI

[NR11] P. E. Nguimkeu and M. Rekkas. Third-order inference for autocorrelation in nonlinear regression models. Journal of Statistical Planning and Inference, 141(11):3413–3425, November 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375811001625.

Navarro:2012:CCS

[NR12] Jorge Navarro and Rafael Rubio. Comparisons of coherent systems with non-identically distributed components. Journal of Statistical Planning and Inference, 142(6):1310–1319, June 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375811004526.

Nair:2010:NMA

[NS10a] N. Unnikrishnan Nair and P. G. Sankaran. A new measure of association for bivariate survival data. Journal of Statistical Planning and Inference, 140(9):2569–2581, September 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375810001321.

Navarro:2010:ADB

[NS10b] Jorge Navarro and José María Sarabia. Alternative definitions of bivariate equilibrium distributions. Journal of Statistical Planning and Inference, 140(7):2046–2056, July 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810000686.

Niederhausen:2010:PAB

[NS10c] Heinrich Niederhausen and Shaun Sullivan. Pattern avoiding ballot paths and finite operator calculus. Journal of Statistical Planning and Inference, 140(8):2312–2320, August 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810000376.

Niaparast:2013:ODQ

[NS13] Mehrdad Niaparast and Rainer Schwabe. Optimal design for quasi-likelihood estimation in Poisson regression with random coefficients. Journal of Statistical Planning and Inference, 143(2):296–306, February 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375812002467.

Nakakita:2019:ATE

[NU19] Shogo H. Nakakita and Masayuki Uchida. Adaptive test for ergodic diffusions plus noise. Journal of Statistical Planning and Inference, 203(??):131–150, December 2019. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837581930028X.

Nair:2010:MRL

[NV10] N. Unnikrishnan Nair and B. Vineshkumar. L-moments of residual life. Journal of Statistical Planning and Inference, 140(9):2618–2631, September 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375810001382.

Nakayama:2017:SVM

[NYA17] Yugo Nakayama, Kazuyoshi Yata, and Makoto Aoshima. Support vector machine and its bias correction in highdimension, low-sample-size settings. Journal of Statistical Planning and Inference, 191(??):88–100, December 2017. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837581730085X.

Ning:2011:DUD

[NZF11] Jian-Hui Ning, Yong-Dao Zhou, and Kai-Tai Fang. Discrepancy for uniform design of experiments with mixtures. Journal of Statistical Planning and Inference, 141(4):1487–1496, April 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375810004842.

Obozinski:2013:GST

[Obo13] Guillaume Obozinski. Grouping strategies and thresholding for high dimensional linear models: Discussion. Journal of Statistical Planning and Inference, 143(9):1441–1446, September 2013. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375813000487. See [MPT13a].

Ogden:2010:WMF

[OG10] R. Todd Ogden and Ernest Greene. Wavelet modeling of functional random effects with application to human vision data. Journal of Statistical Planning and Inference, 140(12):3797–3808, December 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375810002247.

Ogasawara:2013:ACE

[Oga13] Haruhiko Ogasawara. Asymptotic cumulants of the estimator of the canonical parameter in the exponential family. Journal of Statistical Planning and Inference, 143(12):2142–2150, December 2013. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375813001092.

Ogasawara:2017:ADM

[Oga17] Haruhiko Ogasawara. Accurate distributions of Mallows’ C_{p} and its unbiased modifications with applications to shrinkage estimation. Journal of Statistical Planning and Inference, 184(??):105–116, May 2017. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375816301380.

Ogihara:2019:APB

[Ogi19] Teppei Ogihara. On the asymptotic properties of Bayes-type estimators with general loss functions. Journal of Statistical Planning and Inference, 199(??):136–150, March 2019. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375818300661.

OHagan:2013:BIM

[O’H13] Anthony O’Hagan. Bayesian inference with misspecified models: Inference about what? Journal of Statistical Planning and Inference, 143(10):1643–1648, October 2013. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375813001195. See [Wal13a].

Olmo-Jimenez:2011:GBB

[OJMRCSRA11] M. J. Olmo-Jiménez, A. M. Martínez-Rodríguez, A. Conde-Sánchez, and J. Rodríguez-Avi. A generalization of the bivariate Beta–Binomial distribution. Journal of Statistical Planning and Inference, 141(7):2303–2311, July 2011. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811000280.

Oganian:2011:MMP

[OK11] Anna Oganian and Alan F. Karr. Masking methods that preserve positivity constraints in microdata. Journal of Statistical Planning and Inference, 141(1):31–41, January 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810002326.

Offen:2019:DPC

[OL19] Walter W. Offen and Ramon C. Littell. Design of paired comparison experiments when treatments are levels of a single quantitative variable. Journal of Statistical Planning and Inference, 15(??):331–346, ???? 2019. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0378375886901072.

Omelka:2012:TEC

[OP12] M. Omelka and M. Pauly. Testing equality of correlation coefficients in two populations via permutation methods. Journal of Statistical Planning and Inference, 142(6):1396–1406, June 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375811004630.

Overgaard:2019:POU

[OPP19] Morten Overgaard, Erik Thorlund Parner, and Jan Pedersen. Pseudo-observations under covariate-dependent censoring. Journal of Statistical Planning and Inference, 202(??):112–122, September 2019. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375819300175.

Ogunyemi:2010:MAS

[OQ10] Theophilus Ogunyemi and Xianggui Qu. Minimal augmented square designs and their extensions. Journal of Statistical Planning and Inference, 140(3):697–704, March 2010. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809002791.

Ou:2013:PRT

[OQC13] Zujun Ou, Hong Qin, and Xu Cai. Partially replicated twolevel fractional factorial designs via semifoldover. Journal of Statistical Planning and Inference, 143(4):809–817, April 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375812003254.

Ou:2011:OBF

[OQL11] Zujun Ou, Hong Qin, and Hongyi Li. Optimal blocking and foldover plans for nonregular two-level designs. Journal of Statistical Planning and Inference, 141(5):1635–1645, May 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375810005665.

Okamoto:2010:DMS

[OS10] Naoya Okamoto and Takashi Seo. On the distributions of multivariate sample skewness. Journal of Statistical Planning and Inference, 140(10):2809–2816, October 2010. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810001151.

Ouadah:2013:UBK

[Oua13]  Sarah Ouadah.  Uniform-in-bandwidth kernel estima-

tion for censored data. Journal of Statistical Planning and Inference, 143(8):1273–1284, August 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375813000621.

Omelka:2013:BCC

[OVG13] Marek Omelka, Noël Veraverbeke, and Irène Gijbels. Bootstrapping the conditional copula. Journal of Statistical Planning and Inference, 143(1):1–23, January 2013. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375812002029.

Ou:2016:QRM

[OZC16] Fang-Shu Ou, Donglin Zeng, and Jianwen Cai. Quantile regression models for current status data. Journal of Statistical Planning and Inference, 178(??):112–127, November 2016. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375816300581.

Ou:2019:TFF

[OZQ19] Zujun Ou, Minghui Zhang, and Hong Qin. Tripling of fractional factorial designs. Journal of Statistical Planning and Inference, 199(??):151–159, March 2019. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375818300740.

Ozturk:2011:PEL

[Ozt11] Omer Ozturk. Parametric estimation of location and scale parameters in ranked set sampling. Journal of Statistical Planning and Inference, 141(4):1616–1622, April 2011. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810005197.

Ozturk:2012:QIB

[Ozt12] Omer Ozturk. Quantile inference based on partially rank-ordered set samples. Journal of Statistical Planning and Inference, 142(7):2116–2127, July 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375812000468.

Paul:2012:THC

[PA12] Sudhir Paul and Kazi Azad. Testing homogeneity in clustered (longitudinal) count data regression model with over-dispersion. Journal of Statistical Planning and Inference, 142(6):1608–1618, June 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375812000225.

Park:2013:TMV

[PA13] Junyong Park and Deepak Nag Ayyala. A test for the mean vector in large dimension and small samples. Journal of Statistical Planning and Inference, 143(5):929–943, May 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375812003424.

Paul:2014:RMT

[PA14] Debashis Paul and Alexander Aue. Random matrix theory in statistics: a review. Journal of Statistical Planning and Inference, 150(??):1–29, July 2014. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375813002280.

Palma:2010:SML

[Pal10] Wilfredo Palma. On the sample mean of locally stationary long-memory processes. Journal of Statistical Planning and Inference, 140(12):3764–3774, December 2010. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810002211.

Panny:2010:DRB

[Pan10] Wolfgang Panny. Deletions in random binary search trees: a story of errors. Journal of Statistical Planning and Inference, 140(8):2335–2345, August 2010. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837581000039X.

Park:2010:BTW

[Par10a] Cheolwoo Park. Block thresholding wavelet regression using SCAD penalty. Journal of Statistical Planning and Inference, 140(9):2755–2770, September 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375810001564.

Park:2010:GVO

[Par10b] Junyong Park. The generalized P -value in one-sided testing in two sample multivariate normal populations. Journal of Statistical Planning and Inference, 140(4):1044–1055, April 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375809003462.

Pascual:2010:ALT

[Pas10] Francis Pascual. Accelerated life test planning with independent lognormal competing risks. Journal of Statistical Planning and Inference, 140(4):1089–1100, April 2010. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809003577.

Patriota:2012:SMB

[Pat12] Alexandre G. Patriota. On scale-mixture Birnbaum-Saunders distributions. Journal of Statistical Planning and Inference, 142(7):2221–2226, July 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375812000596.

Pavlides:2012:LAM

[Pav12] Marios G. Pavlides. Local asymptotic minimax theory for block-decreasing densities. Journal of Statistical Planning and Inference, 142(8):2322–2329, August 2012. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837581200119X.

Park:2011:ECD

[PBND11] D. K. Park, Mausumi Bose, W. I. Notz, and A. M. Dean. Efficient crossover designs in the presence of interactions between direct and carry-over treatment effects. Journal of Statistical Planning and Inference, 141(2):846–860, February 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375810003757.

Price:2016:EDO

[PBRT16] David J. Price, Nigel G. Bean, Joshua V. Ross, and Jonathan Tuke. On the efficient determination of optimal Bayesian experimental designs using ABC: a case study in optimal observation of epidemics. Journal of Statistical Planning and Inference, 172(??):1–15, May 2016. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375815002177.

Pan:2013:SWQ

[PC13] Baoguo Pan and Min Chen. Self-weighted quasi-maximum exponential likelihood estimator for ARFIMA-GARCH models. Journal of Statistical Planning and Inference, 143(4):716– 729, April 2013. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375812003175.

Praskova:2014:MPD

[PC14] Zuzana Prásková and Ondrej Chochola. M-procedures for detection of a change under weak dependence. Journal of Statistical Planning and Inference, 149(??):60–76, June 2014. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837581400010X.

Pourahmadi:2010:P

[PD10] Mohsen Pourahmadi and R. A. Davis. Preface. Journal of Statistical Planning and Inference, 140(12):3578–3579, December 2010. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375810002041.

Park:2011:ETC

[PD11] Junyong Park and J. Wade Davis. Estimating and testing conditional sums of means in high dimensional multivariate binary data. Journal of Statistical Planning and Inference, 141(2):1021–1030, February 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375810004192.

Pereira:2010:EBN

[Per10] Luísa Pereira. On the extremal behavior of a nonstationary normal random field. Journal of Statistical Planning and Inference, 140(11):3567–3576, November 2010. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810002296.

Papageorgiou:2010:MLT

[PG10] Georgios Papageorgiou and Malay Ghosh. Multivariate limited translation empirical Bayes estimators. Journal of Statistical Planning and Inference, 140(5):1320–1329, May 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809003681.

Papageorgiou:2012:ESA

[PG12] Georgios Papageorgiou and Malay Ghosh. Estimation of small area event rates and of the associated standard errors. Journal of Statistical Planning and Inference, 142(7):2009– 2016, July 2012. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S037837581200095X.

Peng:2011:PLS

[PH11] Heng Peng and Tao Huang. Penalized least squares for single index models. Journal of Statistical Planning and Inference, 141(4):1362–1379, April 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375810004726.

Park:2013:NEL

[PH13] Cheolwoo Park and Jib Huh. Nonparametric estimation of a log-variance function in scale-space. Journal of Statistical Planning and Inference, 143(10):1766–1780, October 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375813001419.

Park:2017:HTR

[PH17] Seyoung Park and Xuming He. Hypothesis testing for regional quantiles. Journal of Statistical Planning and Inference, 191(??):13–24, December 2017. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375817301088.

Porter:2015:BSH

[PHW15] Aaron T. Porter, Scott H. Holan, and Christopher K. Wikle. Bayesian semiparametric hierarchical empirical likelihood spatial models. Journal of Statistical Planning and Inference, 165(??):78–90, October 2015. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375815000749.

Pradhan:2013:IOC

[PK13] Biswabrata Pradhan and Debasis Kundu. Inference and optimal censoring schemes for progressively censored Birnbaum-Saunders distribution. Journal of Statistical Planning and Inference, 143(6):1098–1108, June 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375812003485.

Park:2012:EVE

[PKL12] Chun Gun Park, Inyoung Kim, and Yung-Seop Lee. Error variance estimation via least squares for small sample nonparametric regression. Journal of Statistical Planning and Inference, 142(8):2369–2385, August 2012. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375812000973.

Park:2012:OPS

[PKMK12] Changyi Park, Kwang-Rae Kim, Rangmi Myung, and Ja-Yong Koo. Oracle properties of SCAD-penalized support vector machine. Journal of Statistical Planning and Inference, 142(8):2257–2270, August 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375812001061.

Pang:2010:IFB

[PL10] Fang Pang and Min-Qian Liu. Indicator function based on complex contrasts and its application in general factorial designs. Journal of Statistical Planning and Inference, 140(1):189–197, January 1, 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375809002146.

Proietti:2011:LPF

[PL11] Tommaso Proietti and Alessandra Luati. Low-pass filter design using locally weighted polynomial regression and discrete prolate spheroidal sequences. Journal of Statistical Planning and Inference, 141(2):831–845, February 2011. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810003769.

Paik:2014:HNN

[PL14] Minhui Paik and Michael D. Larsen. Handling nonignorable nonresponse with respondent modeling and the SIR algorithm. Journal of Statistical Planning and Inference, 145(??):179–189, February 2014. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375813002279.

Park:2016:ITH

[PL16] Sang-Hyeon Park and Kiseop Lee. Informed traders’ hedging with news arrivals. Journal of Statistical Planning and Inference, 175(??):1–10, August 2016. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375815300252.

Plancade:2013:AEC

[Pla13] Sandra Plancade. Adaptive estimation of the conditional cumulative distribution function from current status data. Journal of Statistical Planning and Inference, 143(9):1466–1485, September 2013. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375813000876.

Puri:2019:PCP

[PMK19] P. D. Puri, B. D. Mehta, and Sanpei Kageyama. Patterned constructions of partially efficiency-balanced designs. Journal of Statistical Planning and Inference, 15(??):365–378, ???? 2019. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375886901096.

Peng:2016:GAC

[PMW^{+}16] Limin Peng, Amita Manatunga, Ming Wang, Ying Guo, and AKM Fazlur Rahman. A general approach to categorizing a continuous scale according to an ordinal outcome. Journal of Statistical Planning and Inference, 172(??):23–35, May 2016. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375815002153.

Phoa:2012:OEO

[PMX12] Frederick K. H. Phoa, Rahul Mukerjee, and Hongquan Xu. One-eighth- and one-sixteenth-fraction quaternary code designs with high resolution. Journal of Statistical Planning and Inference, 142(5):1073–1080, May 2012. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811004174.

Pagendam:2010:LOD

[PP10] D. E. Pagendam and P. K. Pollett. Locally optimal designs for the simple death process. Journal of Statistical Planning and Inference, 140(11):3096–3105, November 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810001849.

Pagendam:2013:ODE

[PP13] D. E. Pagendam and P. K. Pollett. Optimal design of experimental epidemics. Journal of Statistical Planning and Inference, 143(3):563–572, March 2013. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375812003059.

Pan:2016:BPI

[PP16a] Li Pan and Dimitris N. Politis. Bootstrap prediction intervals for linear, nonlinear and nonparametric autoregressions. Journal of Statistical Planning and Inference, 177(??):1–27, October 2016. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S037837581400175X. See dis-

cussion [Kre16, GP16, NL16].

Pan:2016:RBP

[PP16b] Li Pan and Dimitris N. Politis. Rejoinder — bootstrap prediction intervals for linear, nonlinear and nonparametric autoregressions. Journal of Statistical Planning and Inference, 177(??):45–49, October 2016. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375814001979.

Puetz:2016:EAM

[PPB16] Stephen J. Puetz, Andreas Prokoph, and Glenn Borchardt. Evaluating alternatives to the Milankovitch theory. Journal of Statistical Planning and Inference, 170(??):158–165, March 2016. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375815001901.

Penev:2011:BTB

[PR11] Spiridon Penev and Avraham Ruderman. On the behaviour of tests based on sample spacings for moderate samples. Journal of Statistical Planning and Inference, 141(3):1240–1249, March 2011. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375810004362.

Patilea:2012:AEV

[PR12] Valentin Patilea and Hamdi Raïssi. Adaptive estimation of vector autoregressive models with time-varying variance: Application to testing linear causality in mean. Journal of Statistical Planning and Inference, 142(11):2891–2912, November 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375812001553.

Papathomas:2016:EDB

[PR16] Michail Papathomas and Sylvia Richardson. Exploring dependence between categorical variables: Benefits and limitations of using variable selection within Bayesian clustering in relation to log-linear modelling with interaction terms. Journal of Statistical Planning and Inference, 173(??):47–63, June 2016. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375816000033.

Pronzato:2010:POD

[Pro10] Luc Pronzato. Penalized optimal designs for dose-finding. Journal of Statistical Planning and Inference, 140(1):283– 296, January 1, 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375809002225.

Prendergast:2014:BTY

[PS14] Luke A. Prendergast and Robert G. Staudte. Better than you think: Interval estimators of the difference of binomial proportions. Journal of Statistical Planning and Inference, 148(??):38–48, May 2014. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S037837581300298X.

Piffl:2015:DDE

[PS15] Martin Piffl and Ernst Stadlober. The depth-design: an efficient generation of high dimensional computer experiments. Journal of Statistical Planning and Inference, 164(??):10–26, September 2015. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375815000427.

Petry:2012:SVS

[PT12] Sebastian Petry and Gerhard Tutz. Shrinkage and variable selection by polytopes. Journal of Statistical Planning and Inference, 142(1):48–64, January 2012. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837581100245X.

Papastathopoulos:2013:EGP

[PT13] Ioannis Papastathopoulos and Jonathan A. Tawn. Extended generalised Pareto models for tail estimation. Journal of Statistical Planning and Inference, 143(1):131–143, January 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375812002388.

Philippe:2010:RSL

[PV10] Anne Philippe and Marie-Claude Viano. Random sampling of long-memory stationary processes. Journal of Statistical Planning and Inference, 140(5):1110–1124, May 2010. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809003486.

Pumi:2019:BAF

[PVB^{+}19] Guilherme Pumi, Marcio Valk, Cleber Bisognin, Fábio Mariano Bayer, and Taiane Schaedler Prass. Beta autoregressive fractionally integrated moving average models. Journal of Statistical Planning and Inference, 200(??):196–212, May 2019. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375818303288.

Park:2011:BRA

[PY11] Cheolwoo Park and Young Joo Yoon. Bridge regression: Adaptivity and group selection. Journal of Statistical Planning and Inference, 141(11):3506–3519, November 2011. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811001960.

Podobnik:2012:PPI

[PZ12] Tomaz Podobnik and Tomi Zivko. On probabilistic parametric inference. Journal of Statistical Planning and Inference, 142(12):3152–3166, December 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375812001991.

Pronzato:2014:ACO

[PZ14] Luc Pronzato and Anatoly A. Zhigljavsky. Algorithmic construction of optimal designs on compact sets for concave and differentiable criteria. Journal of Statistical Planning and Inference, 154(??):141–155, November 2014. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375814000731.

Qin:2015:EML

[QCG15] Hong Qin, Kashinath Chatterjee, and Sutapa Ghosh. Extended mixed-level supersaturated designs. Journal of Statistical Planning and Inference, 157–158(??):100–107, February/March 2015. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375814001542.

Qi:2011:SPS

[Qi11] Houduo Qi. A semidefinite programming study of the Elfving theorem. Journal of Statistical Planning and Inference, 141(9):3117–3130, September 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375811001303.

Qingguo:2010:ESM

[Qin10] Tang Qingguo. L_{1}-estimation in a semiparametric model with longitudinal data. Journal of Statistical Planning and Inference, 140(2):393–405, February 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375809002742.

Qiu:2010:CSM

[Qiu10] Dingxi Qiu. A comparative study of the K-means algorithm and the normal mixture model for clustering: Bivariate homoscedastic case. Journal of Statistical Planning and Inference, 140(7):1701–1711, July 2010. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809004054.

Qin:2010:ELL

[QL10] Yongsong Qin and Qingzhu Lei. On empirical likelihood for linear models with missing responses. Journal of Statistical Planning and Inference, 140(11):3399–3408, November 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375810002375.

Quinlan:2015:ROC

[QL15] Kevin R. Quinlan and Dennis K. J. Lin. Run order considerations for Plackett and Burman designs. Journal of Statistical Planning and Inference, 165(??):56–62, October 2015. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375815000725.

Qin:2011:ELP

[QLL11] Yongsong Qin, Yinghua Li, and Qingzhu Lei. Empirical likelihood for probability density functions under negatively associated samples. Journal of Statistical Planning and Inference, 141(1):373–381, January 2011. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810002879.

Quessy:2011:APT

[QM11] Jean-François Quessy and Mélina Mailhot. Asymptotic power of tests of normality under local alternatives. Journal of Statistical Planning and Inference, 141(8):2787–2802, August 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375811001005.

Qu:2010:CSR

[QOK10] Xianggui Qu, Theophilus Ogunyemi, and Robert Kushler. A class of saturated row-column designs. Journal of Statistical Planning and Inference, 140(3):781–790, March 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809002882.

Quesada-Rubio:2012:SAP

[QRGLDMA^{+}12] J. M. Quesada-Rubio, J. Garcia-Leal, M. J. Del-Moral-Avila, E. Navarrete-Alvarez, and M. J. Rosales-Moreno. Study of asymptotic properties in a semiparametric additivemultiplicative hazard model. Journal of Statistical Planning and Inference, 142(4):974–985, April 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375811004101.

Qu:2011:RCD

[Qu11] Xianggui Qu. Row-column designs with minimal units. Journal of Statistical Planning and Inference, 141(9):3193–3200, September 2011. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375811001492.

Quessy:2010:AAP

[Que10] Jean-François Quessy. Applications and asymptotic power of marginal-free tests of stochastic vectorial independence. Journal of Statistical Planning and Inference, 140(11):3058–3075, November 2010. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375810001710.

Quinn:2010:VVD

[Qui10] Jennifer J. Quinn. Visualizing Vandermonde’s determinant through nonintersecting lattice paths. Journal of Statistical Planning and Inference, 140(8):2346–2350, August 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810000406.

Qin:2012:UPR

[QWC12] Hong Qin, Zhenghong Wang, and Kashinath Chatterjee. Uniformity pattern and related criteria for q-level factorials. Journal of Statistical Planning and Inference, 142(5):1170– 1177, May 2012. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S037837581100423X.

Qin:2011:SPT

[QZZ11] Huaizhen Qin, Guohua Zou, and Sherry Z. F. Zhou. On the sensitivity of pre-test estimators to covariance misspecification. Journal of Statistical Planning and Inference, 141(1):293–304, January 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375810002818.

Rodriguez:2010:TSN

[RA10] Paulino Pérez Rodríguez and José A. Villaseñor Alva. On testing the skew normal hypothesis. Journal of Statistical Planning and Inference, 140(11):3148–3159, November 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810001801.

Raqab:2012:PCR

[RA12] Mohammad Z. Raqab and Jafar Ahmadi. Pitman closeness of record values from two sequences to population quantiles. Journal of Statistical Planning and Inference, 142(4):855– 862, April 2012. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375811003818.

Rabier:2014:EPQ

[Rab14a] C. E. Rabier. On empirical processes for quantitative trait locus mapping under the presence of a selective genotyping and an interference phenomenon. Journal of Statistical Planning and Inference, 153(??):42–55, October 2014. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375814000974.

Rabier:2014:SIS

[Rab14b] C. E. Rabier. On statistical inference for selective genotyping. Journal of Statistical Planning and Inference, 147(??):24–52, April 2014. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375813002966.

Rocchetti:2014:REM

[RAB14c] Irene Rocchetti, Marco Alfó, and Dankmar Böhning. A regression estimator for mixed binomial capture-recapture data. Journal of Statistical Planning and Inference, 145(??): 165–178, February 2014. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375813001900.

Reynaud-Bouret:2011:ADE

[RBRTM11] Patricia Reynaud-Bouret, Vincent Rivoirard, and Christine Tuleau-Malot. Adaptive density estimation: a curse of support? Journal of Statistical Planning and Inference, 141(1):115–139, January 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375810002648.

Robinson:2010:SPA

[RBS10] Timothy J. Robinson, Jeffrey B. Birch, and B. Alden Starnes. A semi-parametric approach to dual modeling when no replication exists. Journal of Statistical Planning and Inference, 140(10):2860–2869, October 2010. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810001217.

Ramprasath:2012:SOT

[RD12] L. Ramprasath and T. M. Durairajan. Some observations in testing for nonstandard mixtures. Journal of Statistical Planning and Inference, 142(11):2985–2992, November 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375812001826.

Regnault:2011:EUP

[Reg11] Philippe Regnault. Estimation using plug-in of the stationary distribution and Shannon entropy of continuous time Markov processes. Journal of Statistical Planning and Inference, 141(8):2711–2725, August 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375811000802.

Rabie:2013:ODC

[RF13] Huwaida Rabie and Nancy Flournoy. Optimal designs for contingent response models with application to toxicityefficacy studies. Journal of Statistical Planning and Inference, 143(8):1371–1379, August 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375813000700.

Robertson:2010:GLC

[RGW10] Stephen D. Robertson, Henry L. Gray, and Wayne A. Woodward. The generalized linear chirp process. Journal of Statistical Planning and Inference, 140(12):3676–3687, December 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375810002132.

Ren:2011:EGF

[RH11] Jian-Jian Ren and Bin He. Estimation and goodness-of-fit for the Cox model with various types of censored data. Journal of Statistical Planning and Inference, 141(2):961–971, February 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375810004179.

Ruwet:2013:CPR

[RH13] C. Ruwet and G. Haesbroeck. Classification performance resulting from a 2-means. Journal of Statistical Planning and Inference, 143(2):408–418, February 2013. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375812002728.

Rosalsky:2017:CLT

[RL17] Andrew Rosalsky and Deli Li. A central limit theorem for bootstrap sample sums from non-i.i.d. models. Journal of Statistical Planning and Inference, 180(??):69–80, January 2017. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S037837581630088X.

Reisen:2017:ELM

[RLLT17] V. A. Reisen, C. Lévy-Leduc, and M. S. Taqqu. An Mestimator for the long-memory parameter. Journal of Statistical Planning and Inference, 187(??):44–55, August 2017. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375817300319.

Ren:2012:OST

[RLZ12] Junbai Ren, Peng Li, and Runchu Zhang. An optimal selection of two-level regular single arrays for robust parameter experiments. Journal of Statistical Planning and Inference, 142(11):3037–3046, November 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375812001619.

Raqab:2011:IGR

[RM11] Mohammad Z. Raqab and Mohamed T. Madi. Inference for the generalized Rayleigh distribution based on progressively censored data. Journal of Statistical Planning and Inference, 141(10):3313–3322, October 2011. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811001595.

Rigat:2012:NSR

[RM12] Fabio Rigat and Pietro Muliere. Nonparametric survival regression using the beta-Stacy process. Journal of Statistical Planning and Inference, 142(9):2688–2700, September 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375812001036.

Rosenblatt:2010:SAP

[Ros10] Murray Rosenblatt. Spectral analysis for processes with almost periodic covariances. Journal of Statistical Planning and Inference, 140(12):3608–3612, December 2010. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810002077.

Rosenblum:2014:UMP

[Ros14] Michael Rosenblum. Uniformly most powerful tests for simultaneously detecting a treatment effect in the overall population and at least one subpopulation. Journal of Statistical Planning and Inference, 155(??):107–116, December 2014. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375814001153.

Ravi:2011:MTC

[RP11a] S. Ravi and A. S. Praveena. On von Mises type conditions for p-max stable laws, rates of convergence and generalized log Pareto distributions. Journal of Statistical Planning and Inference, 141(9):3021–3034, September 2011. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811001212.

Rukhin:2011:CPA

[RP11b] Andrey Rukhin and Carey E. Priebe. A comparative power analysis of the maximum degree and size invariants for random graph inference. Journal of Statistical Planning and Inference, 141(2):1041–1046, February 2011. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810004246.

Rodriguez:2015:SSS

[RQ15] Abel Rodríguez and Fernando A. Quintana. On species sampling sequences induced by residual allocation models. Journal of Statistical Planning and Inference, 157– 158(??):108–120, February/March 2015. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S037837581400158X.

Rodger:2010:GCG

[RR10] C. A. Rodger and Julie Rogers. Generalizing Clatworthy group divisible designs. Journal of Statistical Planning and Inference, 140(9):2442–2447, September 2010. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810000972.

Rao:2012:SNT

[RR12] K. S. Madhava Rao and M. Raghunath. A simple nonparametric test for bivariate symmetry about a line. Journal of Statistical Planning and Inference, 142(2):430–444, February 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375811003107.

Ray:2014:CTP

[RRN14] Douglas M. Ray, Paul A. Roediger, and Barry T. Neyer. Commentary: Three-phase optimal design for sensitivity experiments. Journal of Statistical Planning and Inference, 149(??): 20–25, June 2014. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375813003121.

Rebafka:2011:IBM

[RRS11] Tabea Rebafka, François Roueff, and Antoine Souloumiac. Information bounds and MCMC parameter estimation for the pile-up model. Journal of Statistical Planning and Inference, 141(1):1–16, January 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375810002740.

Rao:2010:CLU

[RS10] T. J. Rao and S. Sengupta. Calibrated linear unbiased estimators in finite population sampling. Journal of Statistical Planning and Inference, 140(3):652–658, March 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809002560.

Rho:2013:IBF

[RS13a] Yeonwoo Rho and Xiaofeng Shao. Improving the bandwidthfree inference methods by prewhitening. Journal of Statistical Planning and Inference, 143(11):1912–1922, November 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375813001596.

Ritov:2013:RCV

[RS13b] Ya’acov Ritov and Gershon Saar. Remarks on ‘Correlated variables in regression: Clustering and sparse estimation’ by Bühlmann, Rütimann, van de Geer and Zhang. Journal of Statistical Planning and Inference, 143(11):1863–1865, November 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375813001249. See [BRvdGZ13a].

Racz:2015:IIO

[RS15] Michael J. Racz and J. Sedransk. Inference for identifying outlying health care providers. Journal of Statistical Planning and Inference, 160(??):51–59, May 2015. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375814001992.

Rikimaru:2016:GAG

[RS16] Yuuki Rikimaru and Ritei Shibata. A good approximation of the Gaussian likelihood of simultaneous autoregressive model which yields us an asymptotically efficient estimate of parameters. Journal of Statistical Planning and Inference, 173(??): 31–46, June 2016. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375816000045.

Radloff:2019:LOD

[RS19] Martin Radloff and Rainer Schwabe. Locally D-optimal designs for non-linear models on the k-dimensional ball. Journal of Statistical Planning and Inference, 203(??):106–116, December 2019. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375818301502.

Ren:2012:OBA

[RSH12] Cuirong Ren, Dongchu Sun, and Chong He. Objective Bayesian analysis for a spatial model with nugget effects. Journal of Statistical Planning and Inference, 142(7):1933– 1946, July 2012. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S037837581200081X.

Rodger:2019:CDN

[RSS19] C. A. Rodger, Dinesh G. Sarvate, and Jennifer Seberry. Coloured designs, new group divisible designs and pairwise balanced designs. Journal of Statistical Planning and Inference, 15(??):379–389, ???? 2019. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0378375886901102.

Rezaei:2010:EGP

[RTM10] Sadegh Rezaei, Rasool Tahmasbi, and Manijeh Mahmoodi. Estimation of P [Y < X] for generalized Pareto distribution. Journal of Statistical Planning and Inference, 140(2):480–494, February 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375809002432.

Rukhin:2011:ECP

[Ruk11] Andrew L. Rukhin. Estimating common parameters in heterogeneous random effects models. Journal of Statistical Planning and Inference, 141(9):3181–3192, September 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811001376.

Rukhin:2012:CIT

[Ruk12] Andrew L. Rukhin. Confidence intervals for treatment effect from restricted maximum likelihood. Journal of Statistical Planning and Inference, 142(7):1999–2008, July 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375812000791.

Rukhin:2014:BEH

[Ruk14] Andrew L. Rukhin. Bayes estimators of heterogeneity variance and T -systems. Journal of Statistical Planning and Inference, 154(??):156–165, November 2014. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375814000020.

Rukhin:2017:OBM

[Ruk17] A. L. Rukhin. Objective Bayes models for compatibility assessment and bias estimation. Journal of Statistical Planning and Inference, 182(??):74–87, March 2017. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375816301100.

Rukhin:2019:EHV

[Ruk19a] Andrew L. Rukhin. Estimating heterogeneity variances to select a random effects model. Journal of Statistical Planning and Inference, 202(??):1–13, September 2019. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375818303574.

Rukhin:2019:QEQ

[Ruk19b] Andrew L. Rukhin. Quadratic estimators of quadratic functions of normal parameters. Journal of Statistical Planning and Inference, 15(??):301–310, ???? 2019. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0378375886901059.

Ruschendorf:2019:ADC

[Rüs19] Ludger Rüschendorf. On attainable distribution and classification vectors. Journal of Statistical Planning and Inference, 15(??):259–264, ???? 2019. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
0378375886901023.

Ramirez:2010:WDE

[RV10] Pepa Ramírez and Brani Vidakovic. Wavelet density estimation for stratified size-biased sample. Journal of Statistical Planning and Inference, 140(2):419–432, February 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809002298.

Rajaratnam:2016:NCE

[RV16] Bala Rajaratnam and Dario Vincenzi. A note on covariance estimation in the unbiased estimator of risk framework. Journal of Statistical Planning and Inference, 175(??):25–39, August 2016. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375816000252.

Renaud:2010:RCD

[RVF10] Olivier Renaud and Maria-Pia Victoria-Feser. A robust coefficient of determination for regression. Journal of Statistical Planning and Inference, 140(7):1852–1862, July 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810000194.

Radulovic:2018:WCS

[RW18] Dragan Radulović and Marten Wegkamp. Weak convergence of stationary empirical processes. Journal of Statistical Planning and Inference, 194(??):75–84, March 2018. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375817301635.

RoyChoudhury:2015:CPT

[RWB15] Arindam RoyChoudhury, Amy Willis, and John Bunge. Consistency of a phylogenetic tree maximum likelihood estimator. Journal of Statistical Planning and Inference, 161(??): 73–80, June 2015. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375815000038.

Singh:2010:BAC

[SA10] Sarjinder Singh and Raghunath Arnab. Bias-adjustment and calibration of jackknife variance estimator in the presence of non-response. Journal of Statistical Planning and Inference, 140(4):862–871, April 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375809002948.

Sadegh:2011:NMR

[Sad11] Mohammad Khanjari Sadegh. A note on the mean residual life function of a coherent system with exchangeable or nonidentical components. Journal of Statistical Planning and Inference, 141(9):3267–3275, September 2011. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837581100156X.

Salehi:2011:RAC

[SAE11] E. T. Salehi, M. Asadi, and S. Eryilmaz. Reliability analysis of consecutive k-out-of-n systems with nonidentical components lifetimes. Journal of Statistical Planning and Inference, 141(8):2920–2932, August 2011. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837581100111X.

Sagnol:2011:COD

[Sag11] Guillaume Sagnol. Computing optimal designs of multiresponse experiments reduces to second-order cone programming. Journal of Statistical Planning and Inference, 141(5):1684–1708, May 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375810005318.

Saha:2011:APC

[Sah11] Krishna K. Saha. Analysis of proportions for clustered binary data in the presence of unequal dispersions. Journal of Statistical Planning and Inference, 141(8):2933–2945, August 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375811001157.

Sailer:2010:EUL

[Sai10] Oliver Sailer. An exact upper limit for the variance bias in a carry-over model with correlated errors. Journal of Statistical Planning and Inference, 140(7):2057–2068, July 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810000741.

Sarkar:2014:SCB

[Sar14] Palash Sarkar. On some connections between statistics and cryptology. Journal of Statistical Planning and Inference, 148(??):20–37, May 2014. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375813001110.

Sayyareh:2012:TIS

[Say12] Abdolreza Sayyareh. Tracking interval for selecting between non-nested models: an investigation for Type II right censored data. Journal of Statistical Planning and Inference, 142(12):3201–3208, December 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375812001577.

Stoimenova:2011:CET

[SB11] Eugenia Stoimenova and N. Balakrishnan. A class of exceedance-type statistics for the two-sample problem. Journal of Statistical Planning and Inference, 141(9):3244–3255, September 2011. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375811001534.

Schnell:2019:SPL

[SB19] Patrick M. Schnell and Maitreyee Bose. Spectral parameterization for linear mixed models applied to confounding of fixed effects by random effects. Journal of Statistical Planning and Inference, 200(??):47–62, May 2019. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837581830257X.

Sun:2019:GLR

[SBAC19] Wei Sun, Huybrechts F. Bindele, Ash Abebe, and Hannah Correia. General local rank estimation for singleindex varying coefficient models. Journal of Statistical Planning and Inference, 202(??):57–79, September 2019. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375819300059.

Schenk:2011:BEP

[SBCK11] N. Schenk, M. Burkschat, E. Cramer, and U. Kamps. Bayesian estimation and prediction with multiply Type-II censored samples of sequential order statistics from one- and two-parameter exponential distributions. Journal of Statistical Planning and Inference, 141(4):1575–1587, April 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810005094.

Sarkar:2010:MPF

[SC10] Angshuman Sarkar and Kashinath Chatterjee. An MEP.2 plan in 3^{n} factorial experiment and its probability of correct identification. Journal of Statistical Planning and Inference, 140(11):3531–3539, November 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S037837581000265X.

Shang:2011:CBL

[SC11] Zuofeng Shang and Murray K. Clayton. Consistency of Bayesian linear model selection with a growing number of parameters. Journal of Statistical Planning and Inference, 141(11):3463–3474, November 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375811001947.

Strydom:2013:MLE

[SC13] H. F. Strydom and N. A. S. Crowther. Maximum likelihood estimation of parameter structures for the Wishart distribution using constraints. Journal of Statistical Planning and Inference, 143(4):783–794, April 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375812003187.

Song:2016:EVE

[SC16] Guangjing Song and Haixia Chang. Equalities of various estimators in the general growth curve model and the restricted growth curve model. Journal of Statistical Planning and Inference, 169(??):88–100, February 2016. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375815001585.

Schochet:2010:RAS

[Sch10a] Peter Z. Schochet. Is regression adjustment supported by the Neyman model for causal inference? Journal of Statistical Planning and Inference, 140(1):246–259, January 1, 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375809002201.

Schott:2010:RRE

[Sch10b] James R. Schott. Reduced-rank estimation of the difference between two covariance matrices. Journal of Statistical Planning and Inference, 140(4):1038–1043, April 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809003206.

Schweder:2018:CEP

[Sch18] Tore Schweder. Confidence is epistemic probability for empirical science. Journal of Statistical Planning and Inference, 195(??):116–125, May 2018. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375817301738.

Sarkar:2019:GSB

[SCHG19] Sanat K. Sarkar, Aiying Chen, Li He, and Wenge Guo. Group sequential BH and its adaptive versions controlling the FDR. Journal of Statistical Planning and Inference, 199(??):219– 235, March 2019. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375818301320.

Choi:2011:GLL

[sCK11] Hwan sik Choi and Nicholas M. Kiefer. Geometry of the log-likelihood ratio statistic in misspecified models. Journal of Statistical Planning and Inference, 141(6):2091–2099, June 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S037837581000577X.

Simas:2010:SML

[SCR10] Alexandre B. Simas, Gauss M. Cordeiro, and Andréa V. Rocha. Skewness of maximum likelihood estimators in dispersion models. Journal of Statistical Planning and Inference, 140(7):2111–2121, July 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375810000807.

Scricciolo:2015:BA

[Scr15] Catia Scricciolo. Bayesian adaptation. Journal of Statistical Planning and Inference, 166(??):87–101, November 2015. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375814001980.

Shrivastava:2010:GBI

[SD10] Abhishek K. Shrivastava and Yu Ding. Graph based isomorph-free generation of two-level regular fractional factorial designs. Journal of Statistical Planning and Inference, 140(1):169–179, January 1, 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S037837580900216X.

Sun:2016:OED

[SD16] Fangfang Sun and Angela Dean. A-optimal and A-efficient designs for discrete choice experiments. Journal of Statistical Planning and Inference, 170(??):144–157, March 2016. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375815001615.

Stylianou:2015:CON

[SDGK15] S. Stylianou, K. Drosou, S. D. Georgiou, and C. Koukouvinos. Column-orthogonal and nearly column-orthogonal designs for models with second-order terms. Journal of Statistical Planning and Inference, 161(??):81–90, June 2015. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837581500004X.

Svensson:2010:APS

[SdL10] Ingrid Svensson and Sara Sjöstedt de Luna. Asymptotic properties of a stochastic EM algorithm for mixtures with censored data. Journal of Statistical Planning and Inference, 140(1):111–127, January 1, 2010. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809002092.

Shirazi:2013:NRE

[SDNH13] E. Shirazi, H. Doosti, H. A. Niroumand, and N. Hosseinioun. Nonparametric regression estimates with censored data based on block thresholding method. Journal of Statistical Planning and Inference, 143(7):1150–1165, July 2013. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375813000190.

Steinberg:2014:DTP

[SDV14] David M. Steinberg, Hovav Dror, and Liel Villa. Discussion of “Three-phase optimal design of sensitivity experiments” by Wu and Tian. Journal of Statistical Planning and Inference, 149(??):26–29, June 2014. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375813003145. See [WT14b].

Sang:2019:JEL

[SDZ19] Yongli Sang, Xin Dang, and Yichuan Zhao. Jackknife empirical likelihood methods for Gini correlations and their equality testing. Journal of Statistical Planning and Inference, 199(??):45–59, March 2019. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375818300521.

Shah:2019:SAR

[SE19] K. R. Shah and J. A. Eccleston. On some aspects of row-column designs. Journal of Statistical Planning and Inference, 15(??):87–95, ???? 2019. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
037837588690087X.

Seberry:2019:GHM

[Seb19] Jennifer Seberry. Generalized Hadamard matrices and colourable designs in the construction of regular GDDs with two and three association classes. Journal of Statistical Planning and Inference, 15(??):237–246, ???? 2019. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/037837588690100X.

Segers:2015:HCE

[Seg15] Johan Segers. Hybrid copula estimators. Journal of Statistical Planning and Inference, 160(??):23–34, May 2015. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375814001955.

Sei:2014:IIB

[Sei14] Tomonari Sei. Infinitely imbalanced binomial regression and deformed exponential families. Journal of Statistical Planning and Inference, 149(??):116–124, June 2014. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375814000044.

Sekido:2012:ACO

[Sek12] Hiroto Sekido. An algorithm for calculating D-optimal designs for polynomial regression through a fixed point. Journal of Statistical Planning and Inference, 142(4):935–943, April 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375811004034.

Staicu:2010:SOA

[SF10] Ana-Maria Staicu and Donald A. S. Fraser. The second order ancillary is rotation based. Journal of Statistical Planning and Inference, 140(3):831–836, March 2010. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809002924.

Shen:2011:DNB

[SG11] Gang Shen and Jayanta K. Ghosh. Developing a new BIC for detecting change-points. Journal of Statistical Planning and Inference, 141(4):1436–1447, April 2011. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810004866.

Shan:2012:CSI

[SG12a] Na Shan and Jianhua Guo. Covariate selection for identifying the causal effects of stochastic interventions using causal networks. Journal of Statistical Planning and Inference, 142(1):212–220, January 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375811002795.

Shaw:2012:FPC

[SG12b] Simon C. Shaw and Michael Goldstein. Finite population corrections for multivariate Bayes sampling. Journal of Statistical Planning and Inference, 142(10):2844–2861, October 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375812001516.

Sarkar:2012:APC

[SGF12] Sanat K. Sarkar, Wenge Guo, and Helmut Finner. On adaptive procedures controlling the familywise error rate. Journal of Statistical Planning and Inference, 142(1):65–78, January 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375811002655.

Sun:2013:WLL

[SGL13] Jing Sun, Yujie Gai, and Lu Lin. Weighted local linear composite quantile estimation for the case of general error distributions. Journal of Statistical Planning and Inference, 143(6):1049–1063, June 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375813000050.

Sartono:2012:CTL

[SGS12] Bagus Sartono, Peter Goos, and Eric D. Schoen. Classification of three-level strength-3 arrays. Journal of Statistical Planning and Inference, 142(4):794–809, April 2012. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811003752.

Shen:2011:NML

[She11] Pao-Sheng Shen. The nonparametric maximum likelihood estimator for middle-censored data. Journal of Statistical Planning and Inference, 141(7):2494–2499, July 2011. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811000711.

Sheena:2014:IEC

[She14] Yo Sheena. Inference on the eigenvalues of the covariance matrix of a multivariate normal distribution — geometrical view. Journal of Statistical Planning and Inference, 150(??): 66–83, July 2014. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375814000408.

Stoffer:2010:SSA

[SHQG10] David S. Stoffer, Sangdae Han, Li Qin, and Wensheng Guo. Smoothing spline ANOPOW. Journal of Statistical Planning and Inference, 140(12):3789–3796, December 2010. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810002235.

Shutoh:2011:AER

[Shu11] Nobumichi Shutoh. Asymptotic expansions relating to discrimination based on two-step monotone missing samples. Journal of Statistical Planning and Inference, 141(3):1297– 1306, March 2011. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375810004416.

Shutoh:2012:AAE

[Shu12] Nobumichi Shutoh. An asymptotic approximation for EPMC in linear discriminant analysis based on monotone missing data. Journal of Statistical Planning and Inference, 142(1):110–125, January 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375811002710.

Shen:2016:ODB

[SHW16] Gang Shen, Seung Won Hyun, and Weng Kee Wong. Optimal designs based on the maximum quasi-likelihood estimator. Journal of Statistical Planning and Inference, 178(??):128–139, November 2016. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375816300659.

Stoklosa:2016:ACO

[SHYH16] Jakub Stoklosa, Wen-Han Hwang, Paul S. F. Yip, and Richard M. Huggins. Accounting for contamination and outliers in covariates for open population capturerecapture models. Journal of Statistical Planning and Inference, 176(??):52–63, September 2016. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375816300167.

Singh:2010:IEP

[SK10] Housila P. Singh and Sunil Kumar. Improved estimation of population mean under two-phase sampling with subsampling the non-respondents. Journal of Statistical Planning and Inference, 140(9):2536–2550, September 2010. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810001369.

Shoaee:2012:NAC

[SK12] Shirin Shoaee and Esmaile Khorram. A new absolute continuous bivariate generalized exponential distribution. Journal of Statistical Planning and Inference, 142(7):2203–2220, July 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375812000249.

Schiopu-Kratina:2014:MCO

[SKFMR14] Ioana Schiopu-Kratina, Jean-Marc Fillion, Lenka Mach, and Philip T. Reiss. Maximizing the conditional overlap in business surveys. Journal of Statistical Planning and Inference, 149(??):98–115, June 2014. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375814000093.

Strzalkowska-Kominiak:2010:PHT

[SKS10] Ewa Strzalkowska-Kominiak and Winfried Stute. On the probability of holes in truncated samples. Journal of Statistical Planning and Inference, 140(6):1519–1528, June 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809003978.

Shang:2014:BHD

[SL14a] Zuofeng Shang and Ping Li. Bayesian high-dimensional screening via MCMC. Journal of Statistical Planning and Inference, 155(??):54–78, December 2014. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375814001165.

Soleymani:2014:BPL

[SL14b] Mehdi Soleymani and Stephen M. S. Lee. A bootstrap procedure for local semiparametric density estimation amid model uncertainties. Journal of Statistical Planning and Inference, 153(??):75–86, October 2014. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375814000780.

Son:2019:MPA

[SL19a] Won Son and Johan Lim. Modified path algorithm of fused lasso signal approximator for consistent recovery of change points. Journal of Statistical Planning and Inference, 200(??):223–238, May 2019. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375818303380.

Sun:2019:BOB

[SL19b] Shaun Zheng Sun and Richard A. Lockhart. Bayesian optimality for Beran’s class of tests of uniformity around the circle. Journal of Statistical Planning and Inference, 198(??):79– 90, January 2019. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375818300326.

Staicu:2015:STF

[SLC15] Ana-Maria Staicu, Soumen N. Lahiri, and Raymond J. Carroll. Significance tests for functional data with complex dependence structure. Journal of Statistical Planning and Inference, 156(??):1–13, January 2015. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375814001566.

Shin:2018:SPT

[SLC18] Yongyun Shin, Jennifer Elston Lafata, and Yu Cao. Statistical power in two-level hierarchical linear models with arbitrary number of factor levels. Journal of Statistical Planning and Inference, 194(??):106–121, March 2018. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375817301611.

Stoyanov:2014:HMP

[SLD14] Jordan Stoyanov, Gwo Dong Lin, and Anirban Das-Gupta. Hamburger moment problem for powers and products of random variables. Journal of Statistical Planning and Inference, 154(??):166–177, November 2014. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375813002796.

Szkutnik:2012:NCE

[SlG12] Zbigniew Szkutnik and lukasz Grzyb. A note on consistent estimation of Kullback–Leibler discrepancy in Poisson regression. Journal of Statistical Planning and Inference, 142(6):1619–1622, June 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375811004678.

Shen:2018:PCG

[SLgX18] Jieli Shen, Regina Y. Liu, and Min ge Xie. Prediction with confidence — a general framework for predictive inference. Journal of Statistical Planning and Inference, 195(??):126– 140, May 2018. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375817301696.

Sun:2010:COL

[SLL10] Fasheng Sun, Min-Qian Liu, and Dennis K. J. Lin. Construction of orthogonal Latin hypercube designs with flexible run sizes. Journal of Statistical Planning and Inference, 140(11):3236–3242, November 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S037837581000203X.

Song:2013:RWN

[SLW13] Yuping Song, Zhengyan Lin, and Hanchao Wang. Reweighted Nadaraya–Watson estimation of second-order jumpdiffusion model. Journal of Statistical Planning and Inference, 143(4):730–744, April 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375812003047.

Shows:2010:SEI

[SLZ10] Justin Hall Shows, Wenbin Lu, and Hao Helen Zhang. Sparse estimation and inference for censored median regression. Journal of Statistical Planning and Inference, 140(7):1903– 1917, July 2010. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375810000674.

Slud:2011:SAE

[SM11] Eric V. Slud and Tapabrata Maiti. Small-area estimation based on survey data from a left-censored Fay– Herriot model. Journal of Statistical Planning and Inference, 141(11):3520–3535, November 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375811001972.

Stefanov:2013:DAE

[SM13] Valeri T. Stefanov and Raimondo Manca. Distributions associated with (k_{1}, k_{2}) events on semi-Markov binary trials. Journal of Statistical Planning and Inference, 143(7):1233– 1243, July 2013. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S037837581200362X.

Shen:2016:APB

[SM16] Lu Shen and Max D. Morris. Augmented Plackett-burman designs with replication and improved bias properties. Journal of Statistical Planning and Inference, 179(??):15–21, December 2016. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375816300660.

Syring:2017:GPI

[SM17] Nicholas Syring and Ryan Martin. Gibbs posterior inference on the minimum clinically important difference. Journal of Statistical Planning and Inference, 187(??):67–77, August 2017. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375817300368.

Sabzikar:2019:PEA

[SMM19] Farzad Sabzikar, A. Ian McLeod, and Mark M. Meerschaert. Parameter estimation for ARTFIMA time series. Journal of Statistical Planning and Inference, 200(??):129–145, May 2019. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375818302829.

Sebastiao:2010:CUH

[SMPF10] J. R. Sebastião, A. P. Martins, L. Pereira, and H. Ferreira. Clustering of upcrossings of high values. Journal of Statistical Planning and Inference, 140(4):1003–1012, April 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809003176.

Silva:2010:IDE

[SN10] Jorge Silva and Shrikanth S. Narayanan. Information divergence estimation based on data-dependent partitions. Journal of Statistical Planning and Inference, 140(11):3180–3198, November 2010. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375810001783.

Sun:2012:BTR

[SN12] Dongchu Sun and Shawn Ni. Bayesian testing of restrictions on vector autoregressive models. Journal of Statistical Planning and Inference, 142(11):3008–3022, November 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375812001814.

Soicher:2013:OES

[Soi13] Leonard H. Soicher. Optimal and efficient semi-Latin squares. Journal of Statistical Planning and Inference, 143(3):573– 582, March 2013. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375812002789.

Song:2011:BLU

[Son11a] Guang-Jing Song. On the best linear unbiased estimator and the linear sufficiency of a general growth curve model. Journal of Statistical Planning and Inference, 141(8):2700–2710, August 2011. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375811000784.

Song:2011:DFT

[Son11b] Weixing Song. Distribution-free test in Tobit mean regression model. Journal of Statistical Planning and Inference, 141(8):2891–2901, August 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375811001091.

Sorensen:2019:ISC

[Sør19] Helle Sørensen. Independence, successive and conditional likelihood for time series of counts. Journal of Statistical Planning and Inference, 200(??):20–31, May 2019. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375818302556.

Silva:2012:CSD

[SP12a] Jorge F. Silva and Patricio Parada. On the convergence of Shannon differential entropy, and its connections with density and entropy estimation. Journal of Statistical Planning and Inference, 142(7):1716–1732, July 2012. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375812000705.

Sonksen:2012:RPC

[SP12b] Michael Sonksen and Mario Peruggia. Reference priors for constrained rate models of count data. Journal of Statistical Planning and Inference, 142(11):3023–3036, November 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375812001772.

Shu:2010:BPB

[SR10] Xiaohua Shu and Damaraju Raghavarao. Balanced and partially balanced incomplete block designs with autocorrelation errors. Journal of Statistical Planning and Inference, 140(11):3230–3235, November 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375810002016.

SanMartin:2013:IPR

[SR13a] Ernesto San Martín and Jean-Marie Rolin. Identification of parametric Rasch-type models. Journal of Statistical Planning and Inference, 143(1):116–130, January 2013. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375812002273.

Szekely:2013:ESC

[SR13b] Gábor J. Székely and Maria L. Rizzo. Energy statistics: a class of statistics based on distances. Journal of Statistical Planning and Inference, 143(8):1249–1272, August 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375813000633.

Steland:2014:DCP

[SR14] Ansgar Steland and Ewaryst Rafajlowicz. Decoupling changepoint detection based on characteristic functions: Methodology, asymptotics, subsampling and application. Journal of Statistical Planning and Inference, 145(??):49–73, February 2014. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375813001894.

Simas:2011:BCE

[SRBS11] Alexandre B. Simas, Andréa V. Rocha, and Wagner Barreto-Souza. Bias-corrected estimators for dispersion models with dispersion covariates. Journal of Statistical Planning and Inference, 141(9):3063–3074, September 2011. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837581100125X.

Srivastava:2010:BV

[Sri10] Jagdish N. Srivastava. Bon voyage. Journal of Statistical Planning and Inference, 140(7):1619, July 2010. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810000893.

Sarkar:2000:API

[SS00] Jyotirmoy Sarkar and Sahadeb Sarkar. Availability of a periodically inspected system under perfect repair. Journal of Statistical Planning and Inference, 91(1):77–90, November 1, 2000. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375800001282. See erra-

tum [TLBJ11].

Shen:2010:IPW

[sS10a] Pao sheng Shen. An inverse-probability-weighted approach to the estimation of distribution function with middle-censored data. Journal of Statistical Planning and Inference, 140(7):1844–1851, July 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375810000200.

Shen:2010:JML

[sS10b] Pao sheng Shen. Jackknife methods for left-truncated data. Journal of Statistical Planning and Inference, 140(11):3468–3475, November 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375810002508.

Shen:2011:EBC

[sS11] Pao sheng Shen. Estimation of the bivariate causespecific distribution function with doubly censored competing risks data. Journal of Statistical Planning and Inference, 141(8):2614–2621, August 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S037837581100070X.

Samarakoon:2012:ESL

[SS12a] Nithansha Samarakoon and Weixing Song. Empirical smoothing lack-of-fit tests for variance function. Journal of Statistical Planning and Inference, 142(5):1128–1140, May 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375811004277.

Shen:2012:MRM

[sS12b] Pao sheng Shen. Median regression model with left truncated and right censored data. Journal of Statistical Planning and Inference, 142(7):1757–1766, July 2012. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375812000614.

Shen:2012:MSC

[sS12c] Pao sheng Shen. Modified self-consistent estimators of the survival function with twice censored data. Journal of Statistical Planning and Inference, 142(6):1549–1556, June 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837581200016X.

Shah:2013:DCV

[SS13a] Rajen D. Shah and Richard J. Samworth. Discussion of ‘Correlated variables in regression: Clustering and sparse estimation’ by Peter Bühlmann, Philipp Rütimann, Sara van de Geer and Cun-Hui Zhang. Journal of Statistical Planning and Inference, 143(11):1866–1868, November 2013. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375813001250. See [BRvdGZ13a].

Shen:2013:EAB

[sS13b] Pao sheng Shen. Estimation of the association for bivariate doubly censored data. Journal of Statistical Planning and Inference, 143(3):630–636, March 2013. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375812002698.

Simecek:2013:MTA

[SS13c] Petr Simecek and Marie Simecková. Modification of Tukey’s additivity test. Journal of Statistical Planning and Inference, 143(1):197–201, January 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S037837581200239X.

Srivastava:2013:ACI

[SS13d] Radhendushka Srivastava and Debasis Sengupta. Asymptotic confidence interval of power spectrum of a continuous time process through progressively faster sampling. Journal of Statistical Planning and Inference, 143(11):1938–1949, November 2013. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375813001511.

Shi:2016:SPC

[SS16] Jianhong Shi and Weixing Song. Sparse principal component analysis with measurement errors. Journal of Statistical Planning and Inference, 175(??):87–99, August 2016. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375816000422.

Sommer:2019:MAS

[SS19] Andreas Sommer and Ansgar Steland. Multistage acceptance sampling under nonparametric dependent sampling designs. Journal of Statistical Planning and Inference, 199(??):89–113, March 2019. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375818300545.

Stingo:2011:EBR

[SSC11] Francesco Claudio Stingo, Elena Stanghellini, and Rosa Capobianco. On the estimation of a binary response model in a selected population. Journal of Statistical Planning and Inference, 141(10):3293–3303, October 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375811001571.

Shen:2010:SEC

[sScY10] Pao sheng Shen and Yi chen Yang. Semiparametric estimation of the cumulative incidence function under competing risks and left-truncated sampling. Journal of Statistical Planning and Inference, 140(7):1863–1873, July 2010. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810000431.

Svenson:2014:ESI

[SSDM14] Joshua Svenson, Thomas Santner, Angela Dean, and Hyejung Moon. Estimating sensitivity indices based on Gaussian process metamodels with compactly supported correlation functions. Journal of Statistical Planning and Inference, 144(??):160–172, January 2014. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S037837581300075X.

Sun:2017:FMA

[SSL^{+}17] Zhimeng Sun, Liuquan Sun, Xiaoling Lu, Ji Zhu, and Yongzhuang Li. Frequentist model averaging estimation for the censored partial linear quantile regression model. Journal of Statistical Planning and Inference, 189(??):1–15, October 2017. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375817300678.

Shlomo:2012:EIR

[SSS12] Natalie Shlomo, Chris Skinner, and Barry Schouten. Estimation of an indicator of the representativeness of survey response. Journal of Statistical Planning and Inference, 142(1):201–211, January 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375811002783.

Sun:2011:RAL

[SSZ11] Liuquan Sun, Xinyuan Song, and Jie Zhou. Regression analysis of longitudinal data with time-dependent covariates in the presence of informative observation and censoring times. Journal of Statistical Planning and Inference, 141(8):2902–2919, August 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375811001108.

Schuhmacher:2016:QRL

[SSZ16] Dominic Schuhmacher, Anja Sturm, and Henryk Zähle. On qualitative robustness of the Lotka–Nagaev estimator for the offspring mean of a supercritical Galton–Watson process. Journal of Statistical Planning and Inference, 169(??):56–70, February 2016. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375815001445.

Steinebach:2011:STG

[ST11] Josef G. Steinebach and Hella Timmermann. Sequential testing of gradual changes in the drift of a stochastic process. Journal of Statistical Planning and Inference, 141(8):2682–2699, August 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375811000772.

Shi:2019:SDR

[ST19] Chenlu Shi and Boxin Tang. Supersaturated designs robust to two-factor interactions. Journal of Statistical Planning and Inference, 200(??):119–128, May 2019. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375818302775.

Stefanov:2011:CPG

[Ste11a] Valeri T. Stefanov. Compound patterns and generating functions: From basic waiting times to counts of occurrence. Journal of Statistical Planning and Inference, 141(7):2298–2302, July 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375811000279.

Stepniak:2011:PSC

[Ste11b] Czesław Stepniak. On a problem with singularity in comparison of linear experiments. Journal of Statistical Planning and Inference, 141(7):2489–2493, July 2011. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811000395.

Song:2010:MSO

[STS10] Hai-Yan Song, Jian Tao, and Ning-Zhong Shi. S − −S method for stochastically ordered multinomial populations with missing data. Journal of Statistical Planning and Inference, 140(3):684–696, March 2010. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809002584.

Sumetkijakan:2017:NAB

[Sum17] Songkiat Sumetkijakan. Non-atomic bivariate copulas and implicitly dependent random variables. Journal of Statistical Planning and Inference, 186(??):1–12, July 2017. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375817300137.

Sniekers:2015:CSF

[SvdV15] Suzanne Sniekers and Aad van der Vaart. Credible sets in the fixed design model with Brownian motion prior. Journal of Statistical Planning and Inference, 166(??):78–86, November 2015. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375814001220.

Szabo:2015:HBC

[SvdVvZ15] Botond Szabó, Aad van der Vaart, and Harry van Zanten. Honest Bayesian confidence sets for the L^{2}-norm. Journal of Statistical Planning and Inference, 166(??):36–51, November 2015. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375814001244.

Spokoiny:2013:LQR

[SWH13a] Vladimir Spokoiny, Weining Wang, and Wolfgang Karl Härdle. Local quantile regression. Journal of Statistical Planning and Inference, 143(7):1109–1129, July 2013. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375813000530. See discussion [HL13,

KL13].

Spokoiny:2013:RLQ

[SWH13b] Vladimir Spokoiny, Weining Wang, and Wolfgang Karl Härdle. Rejoinder: Local quantile regression. Journal of Statistical Planning and Inference, 143(7):1145–1149, July 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837581300061X.

Sinha:2011:SOD

[SX11] Sanjoy K. Sinha and Xiaojian Xu. Sequential D-optimal designs for generalized linear mixed models. Journal of Statistical Planning and Inference, 141(4):1394–1402, April 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810004775.

Shao:2011:EUP

[SXX11] Jun Shao, Zhiguo Xiao, and Ruifeng Xu. Estimation with unbalanced panel data having covariate measurement error. Journal of Statistical Planning and Inference, 141(2):800–808, February 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S037837581000371X.

Shao:2012:PSC

[SY12] Q. Shao and L. Yang. Polynomial spline confidence bands for time series trend. Journal of Statistical Planning and Inference, 142(7):1678–1689, July 2012. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375812000754.

Sun:2018:CLT

[SY18] Xichao Sun and Litan Yan. Central limit theorems and parameter estimation associated with a weightedfractional Brownian motion. Journal of Statistical Planning and Inference, 192(??):45–64, January 2018. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375817301210.

Shehzad:2011:MND

[SYA11a] Farrukh Shehzad, Mohammad Zafar Yab, and Rashid Ahmed. Minimal neighbor designs in circular blocks of unequal sizes. Journal of Statistical Planning and Inference, 141(12):3681–3685, December 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375811002278.

Shehzad:2011:SSP

[SYA11b] Farrukh Shehzad, Mohammad Zafar Yab, and Rashid Ahmed. Some series of proper generalized neighbor designs. Journal of Statistical Planning and Inference, 141(12):3808–3813, December 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375811002412.

Sun:2013:CNS

[SYL13] Fasheng Sun, Yuhui Yin, and Min-Qian Liu. Construction of nested space-filling designs using difference matrices. Journal of Statistical Planning and Inference, 143(1):160–166, January 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375812002340.

Sun:2010:CMR

[SZ10] Liuquan Sun and Qiang Zhao. A class of mean residual life regression models with censored survival data. Journal of Statistical Planning and Inference, 140(11):3425–3441, November 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375810002491.

Shinki:2012:ATF

[SZ12] Kazuhiko Shinki and Zhengjun Zhang. Asymptotic theory for fractionally integrated asymmetric power ARCH models. Journal of Statistical Planning and Inference, 142(11):2871–2890, November 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375812001541.

Subramanian:2013:MBC

[SZ13] Sundarraman Subramanian and Peixin Zhang. Model-based confidence bands for survival functions. Journal of Statistical Planning and Inference, 143(7):1166–1185, July 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375813000293.

Sohrabi:2019:ATE

[SZ19] Maryam Sohrabi and Mahmoud Zarepour. Asymptotic theory for M-estimates in unstable AR(p) processes with infinite variance innovations. Journal of Statistical Planning and Inference, 198(??):105–118, January 2019. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837581830034X.

Shao:2016:SMN

[SZZ16] Qi-Man Shao, Kan Zhang, and Wen-Xin Zhou. Stein’s method for nonlinear statistics: a brief survey and recent progress. Journal of Statistical Planning and Inference, 168(??):68–89, January 2016. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375815001342.

Takada:2010:SBN

[Tak10] Yoshikazu Takada. Selecting the best normal population better than a standard under unequal variances. Journal of Statistical Planning and Inference, 140(9):2693–2705, September 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375810001473.

Tan:2010:ECD

[Tan10] Zhiqiang Tan. On estimation of conditional density models with two-phase sampling. Journal of Statistical Planning and Inference, 140(7):1986–2002, July 2010. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810000650.

Tan:2013:CPS

[Tan13] Zhiqiang Tan. Calibrated path sampling and stepwise bridge sampling. Journal of Statistical Planning and Inference, 143(4):675–690, April 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375812003163.

Taraldsen:2010:MLE

[Tar10] Gunnar Taraldsen. A maximum-likelihood estimator with infinite error. Journal of Statistical Planning and Inference, 140(1):1–8, January 1, 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375809002730.

Tang:2011:RSW

[TB11] Liansheng Larry Tang and N. Balakrishnan. A randomsum Wilcoxon statistic and its application to analysis of ROC and LROC data. Journal of Statistical Planning and Inference, 141(1):335–344, January 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375810002909.

Talke:2012:GSF

[TB12] Ismael S. Talke and John J. Borkowski. Generation of spacefilling uniform designs in unit hypercubes. Journal of Statistical Planning and Inference, 142(12):3189–3200, December 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375812002261.

Tu:2012:NCB

[TCC12] Yi-Hsuan Tu, Bin Cheng, and Ying Kuen Cheung. A note on confidence bounds after fixed-sequence multiple tests. Journal of Statistical Planning and Inference, 142(11):2993–2998, November 2012. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S037837581200184X.

Tsunemi:2012:BNM

[TCE^{+}12] Miriam Harumi Tsunemi, Thiago Feitosa Campos, Luís Gustavo Esteves, José Galvão Leite, and Sergio Wechsler. A Bayesian nonparametric model for Taguchi’s on-line quality monitoring procedure for attributes. Journal of Statistical Planning and Inference, 142(9):2701–2709, September 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375812001371.

Tang:2010:CDT

[TDW10] Liansheng Tang, Pang Du, and Chengqing Wu. Compare diagnostic tests using transformation-invariant smoothed ROC curves. Journal of Statistical Planning and Inference, 140(11):3540–3551, November 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375810002739.

Terres:2016:SPG

[TG16] Maria A. Terres and Alan E. Gelfand. Spatial process gradients and their use in sensitivity analysis for environmental processes. Journal of Statistical Planning and Inference, 168(??):106–119, January 2016. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375815001378.

Takai:2014:EUD

[TH14] Keiji Takai and Kenichi Hayashi. Effects of unlabeled data on classification error in normal discriminant analysis. Journal of Statistical Planning and Inference, 147(??):66–83, April 2014. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375813002905.

Tang:2012:KSD

[THG12] Wan Tang, Hua He, and Douglas Gunzler. Kernel smoothing density estimation when group membership is subject to missing. Journal of Statistical Planning and Inference, 142(3):685–694, March 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375811003557.

Thulin:2014:DTJ

[Thu14] Måns Thulin. Decision-theoretic justifications for Bayesian hypothesis testing using credible sets. Journal of Statistical Planning and Inference, 146(??):133–138, March 2014. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375813002498.

Tian:2013:PBU

[Tia13] Yongge Tian. On properties of BLUEs under general linear regression models. Journal of Statistical Planning and Inference, 143(4):771–782, April 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375812003205.

Tung:2012:SBS

[TJ12] David D. Tung and S. Rao Jammalamadaka. U-statistics based on spacings. Journal of Statistical Planning and Inference, 142(3):673–684, March 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375811003533.

Tyssedal:2011:SPD

[TKB11] John Tyssedal, Murat Kulahci, and Søren Bisgaard. Splitplot designs with mirror image pairs as sub-plots. Journal of Statistical Planning and Inference, 141(12):3686–3696, December 2011. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375811002357.

Tsai:2011:OPR

[TL11] Shin-Fu Tsai and Chen-Tuo Liao. Optimal partially replicated parallel-flats designs for 2^{n_{1}} × 3^{n_{2}} mixed factorial experiments. Journal of Statistical Planning and Inference, 141(8):2803–2810, August 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375811001017.

Tsai:2013:MBD

[TL13] Shin-Fu Tsai and Chen-Tuo Liao. Minimum breakdown designs in blocks of size two. Journal of Statistical Planning and Inference, 143(1):202–208, January 2013. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837581200242X.

Taraldsen:2018:CFM

[TL18a] G. Taraldsen and B. H. Lindqvist. Conditional fiducial models. Journal of Statistical Planning and Inference, 195(??):141–152, May 2018. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375817301647.

Tsai:2018:OTL

[TL18b] Shin-Fu Tsai and Chen-Tuo Liao. D-optimal two-level parallel-flats block designs with partial replication. Journal of Statistical Planning and Inference, 193(??):89–102, February 2018. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375817301416.

Tang:2011:EAP

[TLBJ11] Tianqiao Tang, Daming Lin, Dragan Banjevic, and Andrew K. S. Jardine. Erratum to “Availability of a periodically inspected system under perfect repair” [J. Statist. Plann. Inference 91 (2000) 77–90]. Journal of Statistical Planning and Inference, 141(1):610, January 2011. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810003186. See [SS00].

Tang:2013:ASS

[TLBJ13] Tianqiao Tang, Daming Lin, Dragan Banjevic, and Andrew K. S. Jardine. Availability of a system subject to hidden failure inspected at constant intervals with nonnegligible downtime due to inspection and downtime due to repair/replacement. Journal of Statistical Planning and Inference, 143(1):176–185, January 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375812002017.

Tichon:2012:GMA

[TLM12] Jenna G. Tichon, William Li, and Robert G. Mcleod. Generalized minimum aberration two-level split-plot designs. Journal of Statistical Planning and Inference, 142(6):1407–1414, June 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375811004642.

Tateishi:2010:NRM

[TMK10] Shohei Tateishi, Hidetoshi Matsui, and Sadanori Konishi. Nonlinear regression modeling via the lasso-type regularization. Journal of Statistical Planning and Inference, 140(5):1125–1134, May 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375809003528.

Toto:2010:BPI

[TN10] Ma. Criselda S. Toto and Balgobin Nandram. A Bayesian predictive inference for small area means incorporating covariates and sampling weights. Journal of Statistical Planning and Inference, 140(11):2963–2979, November 2010. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810001643.

Tsukuda:2014:SAC

[TN14] Koji Tsukuda and Yoichi Nishiyama. On L ^{2} space approach to change point problems. Journal of Statistical Planning and Inference, 149(??):46–59, June 2014. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375814000275.

Thall:2012:HGP

[TNWW12] Peter F. Thall, Hoang Q. Nguyen, Xuemei Wang, and Johannes E. Wolff. A hybrid geometric phase II/III clinical trial design based on treatment failure time and toxicity. Journal of Statistical Planning and Inference, 142(4):944– 955, April 2012. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375811004058.

Tone:2013:KDE

[Ton13] Cristina Tone. Kernel density estimators for random fields satisfying an interlaced mixing condition. Journal of Statistical Planning and Inference, 143(8):1285–1294, August 2013.

CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375813000438.

Torabi:2012:HBE

[Tor12] Mahmoud Torabi. Hierarchical Bayes estimation of spatial statistics for rates. Journal of Statistical Planning and Inference, 142(1):358–365, January 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375811003119.

Torgovitski:2015:PDS

[Tor15] L. Torgovitski. Panel data segmentation under finite time horizon. Journal of Statistical Planning and Inference, 167(??):69–89, December 2015. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375815001123.

Tarpey:2010:PPC

[TP10] Thaddeus Tarpey and Eva Petkova. Principal point classification: Applications to differentiating drug and placebo responses in longitudinal studies. Journal of Statistical Planning and Inference, 140(2):539–550, February 2010. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809002523.

Tang:2012:HTS

[TQ12] Nian-Sheng Tang and Shi-Fang Qiu. Homogeneity test, sample size determination and interval construction of difference of two proportions in stratified bilateral-sample designs. Journal of Statistical Planning and Inference, 142(5):1243–1251, May 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375811004496.

Thavaneswaran:2013:ILN

[TRL13] A. Thavaneswaran, N. Ravishanker, and Y. Liang. Inference for linear and nonlinear stable error processes via estimating functions. Journal of Statistical Planning and Inference, 143(4):827–841, April 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375812003291.

Tan:2010:FLP

[TRWP10] Fei Tan, Gibson Johnston Rayner, Xiaodong Wang, and Hanxiang Peng. A full likelihood procedure of exchangeable negative binomials for modelling correlated and overdispersed count data. Journal of Statistical Planning and Inference, 140(10):2849–2859, October 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375810001205.

Tyssedal:2010:SDP

[TS10] John Tyssedal and Oddgeir Samset. Supersaturated designs of projectivity P =3 or near P =3. Journal of Statistical Planning and Inference, 140(4):1021–1029, April 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837580900319X.

Trutschnig:2012:IMC

[TS12] Wolfgang Trutschnig and Juan Fernández Sánchez. Idempotent and multivariate copulas with fractal support. Journal of Statistical Planning and Inference, 142(12):3086–3096, December 2012. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375812002133.

Trutschnig:2013:SRS

[TS13] Wolfgang Trutschnig and Juan Fernández Sánchez. Some results on shuffles of two-dimensional copulas. Journal of Statistical Planning and Inference, 143(2):251–260, February 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375812002650.

Tsai:2011:NLT

[Tsa11] Min-Hsiao Tsai. A note on Läuter-type robust design for polynomial regression models. Journal of Statistical Planning and Inference, 141(5):1958–1967, May 2011. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810005719.

Tsai:2017:CNO

[Tsa17] Shin-Fu Tsai. Construction of nested orthogonal arrays of parallel-flats type. Journal of Statistical Planning and Inference, 181(??):41–51, February 2017. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375816301057.

Tsao:2019:EGE

[Tsa19] Min Tsao. Estimable group effects for strongly correlated variables in linear models. Journal of Statistical Planning and Inference, 198(??):29–42, January 2019. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375818300296.

Taskinen:2010:SSE

[TSO10] S. Taskinen, S. Sirkiä, and H. Oja. k-step shape estimators based on spatial signs and ranks. Journal of Statistical Planning and Inference, 140(11):3376–3388, November 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810002399.

Tsou:2011:DMV

[Tso11] Tsung-Shan Tsou. Determining the mean-variance relationship in generalized linear models — a parametric robust way. Journal of Statistical Planning and Inference, 141(1):197–203, January 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375810002764.

Tsukuma:2010:PBM

[Tsu10] Hisayuki Tsukuma. Proper Bayes minimax estimators of the normal mean matrix with common unknown variances. Journal of Statistical Planning and Inference, 140(9):2596–2606, September 2010. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375810001448.

Tang:2013:SMM

[TSZ13] Rui Tang, Jun Shao, and Zhengjun Zhang. Sparse moving maxima models for tail dependence in multivariate financial time series. Journal of Statistical Planning and Inference, 143(5):882–895, May 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375812003497.

Tsai:2011:MSA

[TTB11] Chih-Chun Tsai, Sheng-Tsaing Tseng, and N. Balakrishnan. Mis-specification analyses of gamma and Wiener degradation processes. Journal of Statistical Planning and Inference, 141(12):3725–3735, December 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375811002333.

Torrado:2012:ART

[TV12] Nuria Torrado and J. J. P. Veerman. Asymptotic reliability theory of k-out-of-n systems. Journal of Statistical Planning and Inference, 142(9):2646–2655, September 2012. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375812001310.

Timmermans:2015:NSD

[TvS15] Catherine Timmermans and Rainer von Sachs. A novel semi-distance for measuring dissimilarities of curves with sharp local patterns. Journal of Statistical Planning and Inference, 160(??):35–50, May 2015. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375814001967.

Troendle:2011:PMT

[TW11] James F. Troendle and Peter H. Westfall. Permutational multiple testing adjustments with multivariate multiple group data. Journal of Statistical Planning and Inference, 141(6):2021–2029, June 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375810005707.

Tang:2015:IEE

[TWLQ15] Yanlin Tang, Yinfeng Wang, Jingru Li, and Weimin Qian. Improving estimation efficiency in quantile regression with longitudinal data. Journal of Statistical Planning and Inference, 165(??):38–55, October 2015. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375815000683.

Tang:2013:ECM

[TX13] Yu Tang and Hongquan Xu. An effective construction method for multi-level uniform designs. Journal of Statistical Planning and Inference, 143(9):1583–1589, September 2013. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375813000918.

Tian:2011:ECI

[TXLV11] Lili Tian, Chengjie Xiong, Chin-Ying Lai, and Albert Vexler. Exact confidence interval estimation for the difference in diagnostic accuracy with three ordinal diagnostic groups. Journal of Statistical Planning and Inference, 141(1):549–558, January 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375810003277.

Tianyao:2010:OEB

[TY10] Sun Tianyao and Tang Yu. Optimal efficiency balanced designs and their constructions. Journal of Statistical Planning and Inference, 140(9):2771–2777, September 2010. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810001606.

Unkel:2017:CPA

[Unk17] Steffen Unkel. On the conditional probability for assessing time dependence of association in shared frailty models with bivariate current status data. Journal of Statistical Planning and Inference, 182(??):1–11, March 2017. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375816301252.

Ursu:2014:RMP

[UP14] Eugen Ursu and Jean-Christophe Pereau. Robust modelling of periodic vector autoregressive time series. Journal of Statistical Planning and Inference, 155(??):93–106, December 2014. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375814001190.

Vaggelatou:2010:PAM

[Vag10] Eutichia Vaggelatou. Poisson approximation of the mixed Poisson distribution with infinitely divisible mixing law. Journal of Statistical Planning and Inference, 140(1):128– 138, January 1, 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375809002080.

Vellaisamy:2010:SEP

[VAM10] P. Vellaisamy and Riyadh Al-Mosawi. Simultaneous estimation of Poisson means of the selected subset. Journal of Statistical Planning and Inference, 140(11):3355–3364, November 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375810002302.

Vareschi:2015:NLD

[Var15] T. Vareschi. Noisy Laplace deconvolution with error in the operator. Journal of Statistical Planning and Inference, 157–158(??):16–35, February/March 2015. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375814001591.

Vasudeva:2013:ASB

[Vas13] R. Vasudeva. Almost sure behaviour of near moving maxima. Journal of Statistical Planning and Inference, 143(1):96–106, January 2013. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375812002406.

Vazquez:2010:CPE

[VB10a] Emmanuel Vazquez and Julien Bect. Convergence properties of the expected improvement algorithm with fixed mean and covariance functions. Journal of Statistical Planning and Inference, 140(11):3088–3095, November 2010. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810001850.

Volterman:2010:ENC

[VB10b] William Volterman and N. Balakrishnan. Exact nonparametric confidence, prediction and tolerance intervals based on multi-sample Type-II right censored data. Journal of Statistical Planning and Inference, 140(11):3306–3316, November 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375810002351.

Vock:2011:JTT

[VB11] Michael Vock and N. Balakrishnan. A Jonckheere– Terpstra-type test for perfect ranking in balanced ranked set sampling. Journal of Statistical Planning and Inference, 141(2):624–630, February 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375810003290.

vanBerkum:2019:OPC

[vB19] E. E. M. van Berkum. Optimal paired comparison designs for factorial and quadratic models. Journal of Statistical Planning and Inference, 15(??):265–278, ???? 2019. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0378375886901035.

VanHala:2017:NSD

[VBLN17] Matthew Van Hala, Soutir Bandyopadhyay, Soumendra N. Lahiri, and Daniel J. Nordman. On the non-standard distribution of empirical likelihood estimators with spatial data. Journal of Statistical Planning and Inference, 187(??):109– 114, August 2017. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375817300307.

Variyath:2010:ELB

[VCA10] Asokan Mulayath Variyath, Jiahua Chen, and Bovas Abraham. Empirical likelihood based variable selection. Journal of Statistical Planning and Inference, 140(4):971–981, April 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375809003152.

Ventura:2010:DPD

[VCR10] L. Ventura, S. Cabras, and W. Racugno. Default prior distributions from quasi- and quasi-profile likelihoods. Journal of Statistical Planning and Inference, 140(11):2937–2942, November 2010. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375810001709.

vandeGeer:2013:DPM

[vdG13a] Sara van de Geer. Discussion of the paper by M. Mougeot, D. Picard and K. Tribouley: Grouping strategies and thresholding for high dimension linear models. Journal of Statistical Planning and Inference, 143(9):1447–1450, September 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375813000517. See [MPT13a].

vandeGeer:2013:GCP

[vdG13b] Sara van de Geer. Generic chaining and the ℓ_{1}-penalty. Journal of Statistical Planning and Inference, 143(6):1001–1012, June 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375812003631. See discussion

[BC13, Kol13].

vandeGeer:2013:GCL

[vdG13c] Sara van de Geer. Generic chaining and the l1-penalty rejoinder. Journal of Statistical Planning and Inference, 143(6):1026–1028, June 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375812003643.

vandeGeer:2013:DPP

[vdGM13] Sara van de Geer and Marloes H. Maathuis. Discussion of the paper by Piet Groeneboom: Nonparametric (smoothed) likelihood and integral equations. Journal of Statistical Planning and Inference, 143(12):2068–2071, December 2013. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375813001547. See [Gro13a].

vandeGeer:2015:PLS

[vdGM15] Sara van de Geer and Alan Muro. Penalized least squares estimation in the additive model with different smoothness for the components. Journal of Statistical Planning and Inference, 162(??):43–61, July 2015. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375815000117.

vanderMerwe:2010:PCI

[vdMC10] Abrie J. van der Merwe and Delson Chikobvu. A process capability index for averages of observations from new batches in the case of the balanced random effects model. Journal of Statistical Planning and Inference, 140(1):20–29, January 1, 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375809001852.

VanKeilegom:2011:NLS

[VdUÁMM11] Ingrid Van Keilegom, Jacobo de Uña-Álvarez, and Luis Meira-Machado. Nonparametric location-scale models for censored successive survival times. Journal of Statistical Planning and Inference, 141(3):1118–1131, March 2011. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810004210.

vandeVen:2014:OBM

[vdVW14] P. M. van de Ven and D. C. Woods. Optimal blocked minimum-support designs for non-linear models. Journal of Statistical Planning and Inference, 144(??):152–159, January 2014. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375813000311.

Vexler:2013:NBF

[VDW13] Albert Vexler, Wei Deng, and Gregory E. Wilding. Nonparametric Bayes factors based on empirical likelihood ratios.

Journal of Statistical Planning and Inference, 143(3):611– 620, March 2013. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375812002790.

Verdier:2013:ACM

[Ver13] Ghislain Verdier. Application of copulas to multivariate control charts. Journal of Statistical Planning and Inference, 143(12):2151–2159, December 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375813001080.

Verdebout:2017:ESR

[Ver17] Thomas Verdebout. On the efficiency of some rank-based test for the homogeneity of concentrations. Journal of Statistical Planning and Inference, 191(??):101–109, December 2017. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375817300897.

Vexler:2013:EDB

[VGH13] Albert Vexler, Gregory Gurevich, and Alan D. Hutson. An exact density-based empirical likelihood ratio test for paired data. Journal of Statistical Planning and Inference, 143(2):334–345, February 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375812002662.

Vermeulen:2011:ROC

[VGV11] Bart Vermeulen, Peter Goos, and Martina Vandebroek. Rank-order choice-based conjoint experiments: Efficiency and design. Journal of Statistical Planning and Inference, 141(8):2519–2531, August 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S037837581100036X.

Veronese:2018:FCO

[VM18] Piero Veronese and Eugenio Melilli. Fiducial, confidence and objective Bayesian posterior distributions for a multidimensional parameter. Journal of Statistical Planning and Inference, 195(??):153–173, May 2018. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375817301726.

Miakonkana:2014:IRE

[vMMA14] Guy vanie M. Miakonkana and Asheber Abebe. Iterative rank estimation for generalized linear models. Journal of Statistical Planning and Inference, 151–152(??):60–72, August/ September 2014. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375814000573.

Viljoen:2010:CSS

[VN10] H. Viljoen and D. G. Nel. Common singular spectrum analysis of several time series. Journal of Statistical Planning and Inference, 140(1):260–267, January 1, 2010. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809002183.

Vishwakarma:2012:GPE

[VS12] Gajendra K. Vishwakarma and Housila P. Singh. A general procedure for estimating the mean using double sampling for stratification and multi-auxiliary information. Journal of Statistical Planning and Inference, 142(5):1252–1261, May 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375811004514.

Vexler:2011:ELR

[VSK^{+}11] Albert Vexler, Guogen Shan, Seongeun Kim, Wan-Min Tsai, Lili Tian, and Alan D. Hutson. An empirical likelihood ratio based goodness-of-fit test for inverse Gaussian distributions. Journal of Statistical Planning and Inference, 141(6):2128– 2140, June 2011. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375810005823.

Velilla:2018:GFT

[VT18] Santiago Velilla and Huong Nguyen Thu. A goodness-of-fit test for VARMA(p, q) models. Journal of Statistical Planning and Inference, 197(??):126–140, December 2018. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375818300041.

Vellaisamy:2010:CCT

[VV10] P. Vellaisamy and V. Vijay. Collapsibility of contingency tables based on conditional models. Journal of Statistical Planning and Inference, 140(5):1243–1255, May 2010. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809003632.

Valdora:2014:REG

[VY14] Marina Valdora and Víctor J. Yohai. Robust estimators for generalized linear models. Journal of Statistical Planning and Inference, 146(??):31–48, March 2014. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375813002516.

Vexler:2017:EEL

[VZH17] Albert Vexler, Li Zou, and Alan D. Hutson. An extension to empirical likelihood for evaluating probability weighted moments. Journal of Statistical Planning and Inference, 182(??):50–60, March 2017. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375816301215.

Wang:2016:ODM

[WA16] Yaping Wang and Mingyao Ai. Optimal designs for multiple treatments with unequal variances. Journal of Statistical Planning and Inference, 171(??):175–183, April 2016. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375815001895.

Waernbaum:2010:PSM

[Wae10] Ingeborg Waernbaum. Propensity score model specification for estimation of average treatment effects. Journal of Statistical Planning and Inference, 140(7):1948–1956, July 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810000443.

Wagener:2011:MDB

[Wag11] Jens Wagener. The maximization of determinants of block Toeplitz matrices with applications in optimal design theory. Journal of Statistical Planning and Inference, 141(9):3105–3116, September 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375811001297.

Wahba:2010:EDD

[Wah10] Grace Wahba. Encoding dissimilarity data for statistical model building. Journal of Statistical Planning and Inference, 140(12):3580–3596, December 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375810002053.

Walker:2013:BIM

[Wal13a] Stephen G. Walker. Bayesian inference with misspecified models. Journal of Statistical Planning and Inference, 143(10):1621–1633, October 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/

S037837581300116X. See discussion [De 13, HW13, O’H13] and reply [Wal13b].

Walker:2013:RDB

[Wal13b] Stephen G. Walker. Reply to the discussion: Bayesian inference with misspecified models. Journal of Statistical Planning and Inference, 143(10):1649–1652, October 2013. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375813001201. See [Wal13a].

Wang:2017:URS

[WAL17] Xinlei Wang, Soohyun Ahn, and Johan Lim. Unbalanced ranked set sampling in cluster randomized studies. Journal of Statistical Planning and Inference, 187(??):1–16, August 2017. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375817300186.

Wang:2010:IEE

[Wan10a] Bing Xing Wang. Interval estimation for exponential progressive Type-II censored step-stress accelerated lifetesting. Journal of Statistical Planning and Inference, 140(9):2706–2718, September 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375810001461.

Wang:2010:MBP

[Wan10b] Hsiuying Wang. The monotone boundary property and the full coverage property of confidence intervals for a binomial proportion. Journal of Statistical Planning and Inference, 140(2):495–501, February 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375809002717.

Wang:2011:OMN

[Wan11] Zhaoliang Wang. One mixed negative binomial distribution with application. Journal of Statistical Planning and Inference, 141(3):1153–1160, March 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375810004313.

Wang:2010:APN

[WC10] Lihong Wang and Haiyan Cai. Asymptotic properties of nonparametric regression for long memory random fields. Journal of Statistical Planning and Inference, 140(3):837–850, March 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375809002936.

Wang:2017:IOA

[WCZ17] Lu Wang, Yong Chen, and Hongjian Zhu. Implementing optimal allocation in clinical trials with multiple endpoints. Journal of Statistical Planning and Inference, 182(??):88–99, March 2017. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375816301069.

Wang:2018:LTM

[WDRB18] Huan Wang, Hongsheng Dai, Marialuisa Restaino, and Yanchun Bao. Linear transformation models for censored data under truncation. Journal of Statistical Planning and Inference, 193(??):42–54, February 2018. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375817301386.

Weiss:2011:GCM

[Wei11] Christian H. Weiß. Generalized choice models for categorical time series. Journal of Statistical Planning and Inference, 141(8):2849–2862, August 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375811001066.

Wei:2012:SIR

[Wei12] Chuanhua Wei. Statistical inference for restricted partially linear varying coefficient errors-in-variables models. Journal of Statistical Planning and Inference, 142(8):2464–2472, August 2012. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375812000882.

Weng:2015:EMD

[Wen15] Ruby C. Weng. Expansions for multivariate densities. Journal of Statistical Planning and Inference, 167(??):174–181, December 2015. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375815000956.

Westerlund:2016:STN

[Wes16] Joakim Westerlund. A simple test for nonstationarity in mixed panels: a further investigation. Journal of Statistical Planning and Inference, 173(??):1–30, June 2016. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375816000057.

Wang:2011:BAT

[WF11] Yi-Fu Wang and Tsai-Hung Fan. A Bayesian analysis on time series structural equation models. Journal of Statistical Planning and Inference, 141(6):2071–2078, June 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810005756.

Waters:2015:BNA

[WFG^{+}15] Andrew Waters, Kassandra Fronczyk, Michele Guindani, Richard G. Baraniuk, and Marina Vannucci. A Bayesian nonparametric approach for the analysis of multiple categorical item responses. Journal of Statistical Planning and Inference, 166(??):52–66, November 2015. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375814001189.

Wu:2012:BCE

[WG12] Yueqin Wu and Minggao Gu. Bias corrected estimation for generalized probit regression with covariate measurement error and censored responses. Journal of Statistical Planning and Inference, 142(1):221–231, January 2012. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811002813.

Wied:2013:MCC

[WG13] Dominik Wied and Pedro Galeano. Monitoring correlation change in a sequence of random variables. Journal of Statistical Planning and Inference, 143(1):186–196, January 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837581200208X.

Wang:2011:FOS

[WH11] Dongliang Wang and Alan D. Hutson. A fractional order statistic towards defining a smooth quantile function for discrete data. Journal of Statistical Planning and Inference, 141(9):3142–3150, September 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375811001327.

Wandler:2012:FAM

[WH12] Damian V. Wandler and Jan Hannig. A fiducial approach to multiple comparisons. Journal of Statistical Planning and Inference, 142(4):878–895, April 2012. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811004009.

Wang:2012:FPI

[WHI12] C. M. Wang, Jan Hannig, and Hari K. Iyer. Fiducial prediction intervals. Journal of Statistical Planning and Inference, 142(7):1980–1990, July 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375812000687.

Wang:2011:BRS

[WHY11] Xuejun Wang, Shuhe Hu, and Wenzhi Yang. The Bahadur representation for sample quantiles under strongly mixing sequence. Journal of Statistical Planning and Inference, 141(2):655–662, February 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375810003320.

Wiens:2019:MPD

[Wie19a] Douglas P. Wiens. Maximin power designs in testing lack of fit. Journal of Statistical Planning and Inference, 199(??):311–317, March 2019. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375818301642.

Wierich:2019:ODC

[Wie19b] W. Wierich. On optimal designs and complete class theorems for experiments with continuous and discrete factors of influence. Journal of Statistical Planning and Inference, 15(??): 19–27, ???? 2019. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0378375886900820.

Wishart:2013:WDP

[Wis13] Justin Rory Wishart. Wavelet deconvolution in a periodic setting with long-range dependent errors. Journal of Statistical Planning and Inference, 143(5):867–881, May 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375812003503.

Wang:2010:PDU

[WK10] Xinlei Wang and Alan F. Karr. Preserving data utility via BART. Journal of Statistical Planning and Inference, 140(9):2551–2561, September 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375810001357.

Williams:2013:TAC

[WK13] Matthew R. Williams and Dong-Yun Kim. A test for abrupt change in hazard regression models with Weibull baselines. Journal of Statistical Planning and Inference, 143(9):1566–1574, September 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375813000888.

Wang:2011:SEP

[WL11] Qihua Wang and Ruimiao Luo. Semi-empirical pseudolikelihood for estimating equations in the presence of missing responses. Journal of Statistical Planning and Inference, 141(8):2589–2599, August 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375811000668.

Wiens:2014:VOD

[WL14] Douglas P. Wiens and Pengfei Li. V-optimal designs for heteroscedastic regression. Journal of Statistical Planning and Inference, 145(??):125–138, February 2014. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375813002310.

Wang:2016:RSI

[WL16] Kangning Wang and Lu Lin. Robust structure identification and variable selection in partial linear varying coefficient models. Journal of Statistical Planning and Inference, 174(??):153–168, July 2016. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375816000070.

Wang:2011:AEL

[WLCK11] Ni Wang, Jye-Chi Lu, Di Chen, and Paul H. Kvam. Adjusted empirical likelihood models with estimating equations for accelerated life tests. Journal of Statistical Planning and Inference, 141(1):140–155, January 2011. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810002636.

Wang:2017:TSW

[WLL^{+}17] Pin-Wen Wang, Shou-En Lu, Yong Lin, Weichung J. Shih, and K. K. Gordon Lan. Two-stage winner designs for noninferiority trials with pre-specified non-inferiority margin. Journal of Statistical Planning and Inference, 183(??):44–61, April 2017. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375816301239.

Wang:2013:CEG

[WLS13] Molin Wang, Xiaomei Liao, and Donna Spiegelman. Can efficiency be gained by correcting for misclassification? Journal of Statistical Planning and Inference, 143(11):1980–1987, November 2013. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S037837581300147X.

Wang:2012:RC

[WM12] Xiaowei Wang and J. P. Morgan. RCBDs with a control. Journal of Statistical Planning and Inference, 142(5):1225– 1233, May 2012. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375811004319.

Wang:2018:PCO

[WM18a] Fangfang Wang and Chunsheng Ma. Peakedness and convex ordering for elliptically contoured random fields. Journal of Statistical Planning and Inference, 197(??):21–34, December 2018. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375817302136.

Wang:2018:PCP

[WM18b] Min Wang and Yuzo Maruyama. Posterior consistency of g-prior for variable selection with a growing number of parameters. Journal of Statistical Planning and Inference, 196(??):19–29, August 2018. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S037837581730191X.

Wang:2010:NSO

[WMB10] Bo Wang, Robert G. McLeod, and John F. Brewster. A note on the selection of optimal foldover plans for 16- and 32-run fractional factorial designs. Journal of Statistical Planning and Inference, 140(6):1497–1500, June 2010. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809003905.

Wang:2012:DDE

[WMH12] Dongliang Wang, Jeffrey C. Miecznikowski, and Alan D. Hutson. Direct density estimation of L-estimates via characteristic functions with applications. Journal of Statistical Planning and Inference, 142(2):567–578, February 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811003442.

Wang:2010:CBG

[WMM^{+}10] Bin Wang, Satya N. Mishra, Madhuri S. Mulekar, Nutan Mishra, and Kun Huang. Comparison of bootstrap and generalized bootstrap methods for estimating high quantiles. Journal of Statistical Planning and Inference, 140(10):2926–2935, October 2010. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S037837581000128X.

Wu:2015:PIR

[WMO15] Jiwen Wu, Mary C. Meyer, and Jean D. Opsomer. Penalized isotonic regression. Journal of Statistical Planning and Inference, 161(??):12–24, June 2015. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375814002055.

Wang:2010:DOS

[WN10] Ke Wang and H. N. Nagaraja. Distribution of order statistics from selected subsets of concomitants. Journal of Statistical Planning and Inference, 140(11):3076–3087, November 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375810001758.

Withers:2011:BRE

[WN11a] Christopher S. Withers and Saralees Nadarajah. Biasreduced estimates for skewness, kurtosis, L-skewness and L-kurtosis. Journal of Statistical Planning and Inference, 141(12):3839–3861, December 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375811002679.

Withers:2011:UEM

[WN11b] Christopher S. Withers and Saralees Nadarajah. Unbiased estimates for moments and cumulants in linear regression. Journal of Statistical Planning and Inference, 141(12):3867–3875, December 2011. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375811002308.

Withers:2013:EIP

[WN13] Christopher S. Withers and Saralees Nadarajah. Estimators for the inverse powers of a normal mean. Journal of Statistical Planning and Inference, 143(2):441–455, February 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375812002315.

Wong:2012:IEG

[Won12] Augustine Wong. Interval estimation of P (Y < X) for generalized Pareto distribution. Journal of Statistical Planning and Inference, 142(2):601–607, February 2012. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811003211.

Woodroofe:2019:CIF

[Woo19] Michael Woodroofe. Confidence intervals with fixed proportional accuracy. Journal of Statistical Planning and Inference, 15(??):131–146, ???? 2019. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
0378375886900923.

Wang:2010:ELC

[WQ10] Qihua Wang and Yongsong Qin. Empirical likelihood confidence bands for distribution functions with missing responses. Journal of Statistical Planning and Inference, 140(9):2778–2789, September 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375810001655.

Wu:2011:ABI

[WQ11] Lang Wu and Jin Qiu. Approximate bounded influence estimation for longitudinal data with outliers and measurement errors. Journal of Statistical Planning and Inference, 141(7):2321–2330, July 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375811000383.

Wang:2014:BFC

[WS14] Min Wang and Xiaoqian Sun. Bayes factor consistency for nested linear models with a growing number of parameters. Journal of Statistical Planning and Inference, 147(??):95–105, April 2014. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375813002784.

Wang:2015:MSR

[WS15] Shanshan Wang and Robert Serfling. On masking and swamping robustness of leading nonparametric outlier identifiers for univariate data. Journal of Statistical Planning and Inference, 162(??):62–74, July 2015. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375815000087.

Wang:2010:ESI

[WSHW10] Yanhua Wang, Junshan Shen, Shuyuan He, and Qihua Wang. Estimation of single index model with missing response at random. Journal of Statistical Planning and Inference, 140(7):1671–1690, July 2010. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810000753.

Wang:2017:EPR

[WSL17] Zhanfeng Wang, Jian Qing Shi, and Youngjo Lee. Extended t-process regression models. Journal of Statistical Planning and Inference, 189(??):38–60, October 2017. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375817300861.

Wang:2019:NER

[WSSG19] Xiaoguang Wang, Lixin Song, Leyuan Sun, and Hang Gao. Nonparametric estimation of the ROC curve based on the Bernstein polynomial. Journal of Statistical Planning and Inference, 203(??):39–56, December 2019. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375819300187.

Wang:2010:SCA

[WT10] Shufang Wang and Alex Tsodikov. A self-consistency approach to multinomial logit model with random effects. Journal of Statistical Planning and Inference, 140(7):1939–1947, July 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375810000455.

Wu:2014:R

[WT14a] C. F. Jeff Wu and Yubin Tian. Rejoinder. Journal of Statistical Planning and Inference, 149(??):30–32, June 2014. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375814000032.

Wu:2014:TPO

[WT14b] C. F. Jeff Wu and Yubin Tian. Three-phase optimal design of sensitivity experiments. Journal of Statistical Planning and Inference, 149(??):1–15, June 2014. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/

pii/S0378375813002607. See discussion [HJ14, SDV14].

Wang:2011:ETM

[WTL11] Chi Wang, Zhiqiang Tan, and Thomas A. Louis. Exponential tilt models for two-group comparison with censored data. Journal of Statistical Planning and Inference, 141(3):1102– 1117, March 2011. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375810004209.

Wang:2012:OAE

[WTZ12] Xiaodi Wang, Yincai Tang, and Yingshan Zhang. Orthogonal arrays for estimating global sensitivity indices of non-parametric models based on ANOVA high-dimensional model representation. Journal of Statistical Planning and Inference, 142(7):1801–1810, July 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375812000900.

Wu:2016:IPC

[Wu16] Yanhong Wu. Inference for post-change parameters after sequential CUSUM test under AR(1) model. Journal of Statistical Planning and Inference, 168(??):52–67, January 2016. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375815001329.

Wu:2017:ADS

[Wu17] Tung-Lung Wu. Approximations of distributions of scan statistics of inhomogeneous Poisson processes. Journal of Statistical Planning and Inference, 190(??):52–59, November 2017. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375817300770.

Wu:2019:RCS

[Wu19] Tiee-Jian Wu. On the L_{p} rates of convergence of simple linear rank statistics under contiguous alternatives. Journal of Statistical Planning and Inference, 15(??):29–41, ???? 2019. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0378375886900832.

Wu:2012:DKL

[WV12] Qiang Wu and Paul Vos. Decomposition of Kullback–Leibler risk and unbiasedness for parameter-free estimators. Journal of Statistical Planning and Inference, 142(6):1525–1536, June 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375812000110.

Wang:2011:PDF

[WW11a] Xin Wang and Jinhua Wang. Partitioned difference families and almost difference sets. Journal of Statistical Planning and Inference, 141(5):1899–1909, May 2011. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810005446.

Worms:2011:ELB

[WW11b] Julien Worms and Rym Worms. Empirical likelihood based confidence regions for first order parameters of heavy-tailed distributions. Journal of Statistical Planning and Inference, 141(8):2769–2786, August 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375811000991.

Wu:2012:PSE

[WW12a] Chien-Hua Wu and Shu-Mei Wan. Power series estimator for binary data. Journal of Statistical Planning and Inference, 142(2):579–588, February 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375811003466.

Wu:2012:SEL

[WW12b] Rongning Wu and Qin Wang. Shrinkage estimation for linear regression with ARMA errors. Journal of Statistical Planning and Inference, 142(7):2136–2148, July 2012. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375812000948.

Wang:2013:FBW

[WW13] Xue Wang and Stephen G. Walker. Full Bayesian wavelet inference with a nonparametric prior. Journal of Statistical Planning and Inference, 143(1):55–62, January 2013. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375812002005.

Wang:2015:SEV

[WW15a] Guannan Wang and Li Wang. Spline estimation and variable selection for single-index prediction models with diverging number of index parameters. Journal of Statistical Planning and Inference, 162(??):1–19, July 2015. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375815000245.

Wang:2015:EST

[WW15b] Xuan Wang and Qihua Wang. Estimation for semiparametric transformation models with length-biased sampling. Journal of Statistical Planning and Inference, 156(??):80–89, January 2015. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375814001402.

Wang:2010:EA

[WWC10] X. Wang, D. Wu, and M. Cheng. The existence of (q, k, λ, t)- ADFs for k = 4, 5, 6. Journal of Statistical Planning and Inference, 140(11):3243–3251, November 2010. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810002028.

Wang:2019:PEL

[WWZ19] Dongliang Wang, Tong Tong Wu, and Yichuan Zhao. Penalized empirical likelihood for the sparse Cox regression model. Journal of Statistical Planning and Inference, 201(??):71–85, July 2019. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375818303550.

Wang:2015:CPS

[WX15] Lu Wang and Lan Xue. Constrained polynomial spline estimation of monotone additive models. Journal of Statistical Planning and Inference, 167(??):27–40, December 2015. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375815001159.

Wang:2019:FHD

[WX19] Rui Wang and Xingzhong Xu. A feasible high dimensional randomization test for the mean vector. Journal of Statistical Planning and Inference, 199(??):160–178, March 2019. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375818300879.

Wang:2010:NCI

[WY10] Xiao-Feng Wang and Deping Ye. On nonparametric comparison of images and regression surfaces. Journal of Statistical Planning and Inference, 140(10):2875–2884, October 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375810001230.

Wang:2014:AOD

[WY14] Tianhua Wang and Min Yang. Adaptive optimal designs for dose-finding studies based on sigmoid E_{max} models. Journal of Statistical Planning and Inference, 144(??):188–197, January 2014. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375813002267.

Wu:2010:MCC

[WYH10] Yong Wu, Qifan Yang, and Yikun Huang. Machine covering with combined partial information. Journal of Statistical Planning and Inference, 140(8):2351–2354, August 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810000418.

Wei:2010:BRL

[WYY^{+}10] Xianglan Wei, Shanchao Yang, Keming Yu, Xin Yang, and Guodong Xing. Bahadur representation of linear kernel quantile estimator of VaR under α-mixing assumptions. Journal of Statistical Planning and Inference, 140(7):1620–1634, July 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375810000054.

Wieczorek:2010:OEN

[WZ10] B. Wieczorek and K. Ziegler. On optimal estimation of a non-smooth mode in a nonparametric regression model with α-mixing errors. Journal of Statistical Planning and Inference, 140(2):406–418, February 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375809002286.

Wang:2011:ELQ

[WZ11a] Huixia Judy Wang and Zhongyi Zhu. Empirical likelihood for quantile regression models with longitudinal data. Journal of Statistical Planning and Inference, 141(4):1603–1615, April 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375810005173.

Wilmut:2011:OMD

[WZ11b] Michael Wilmut and Julie Zhou. D-optimal minimax design criterion for two-level fractional factorial designs. Journal of Statistical Planning and Inference, 141(1):576–587, January 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375810003253.

Wang:2018:SML

[WZ18] Xuan Wang and Xiao-Hua Zhou. Semiparametric maximum likelihood estimation for the Cox model with lengthbiased survival data. Journal of Statistical Planning and Inference, 196(??):163–173, August 2018. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375817302094.

Wang:2013:ALV

[WZW13] HaiYing Wang, Guohua Zou, and Alan T. K. Wan. Adaptive LASSO for varying-coefficient partially linear measurement error models. Journal of Statistical Planning and Inference, 143(1):40–54, January 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375812002455.

Wang:2019:OFF

[WZZ19] Congcong Wang, Qianqian Zhao, and Shengli Zhao. Optimal fractional factorial split-plot designs when the whole plot factors are important. Journal of Statistical Planning and Inference, 199(??):1–13, March 2019. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375818300491.

Xiang:2016:CBF

[XGK16] Ruoxuan Xiang, Malay Ghosh, and Kshitij Khare. Consistency of Bayes factors under hyper g-priors with growing model size. Journal of Statistical Planning and Inference, 173(??):64–86, June 2016. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375816000021.

Xiao:2012:AIE

[XI12] Changfu Xiao and Anastasia Ivanova. Adaptive isotonic estimation of the minimum effective and peak doses in the presence of covariates. Journal of Statistical Planning and Inference, 142(7):1899–1907, July 2012. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375812000547.

Xia:2013:CLQ

[Xia13] Yingcun Xia. Comment on: “Local quantile regression” by Spokoiny, Wang, and Härdle. Journal of Statistical Planning and Inference, 143(7):1139, July 2013. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375813000669.

Xing:2011:CRN

[Xin11] Yang Xing. Convergence rates of nonparametric posterior distributions. Journal of Statistical Planning and Inference, 141(11):3382–3390, November 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375810004787.

Xiong:2013:ROS

[XJ13] Shifeng Xiong and V. Roshan Joseph. Regression with outlier shrinkage. Journal of Statistical Planning and Inference, 143(11):1988–2001, November 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375813001444.

Xu:2011:PRE

[XL11] Jing Xu and Michael LaValley. Pearson residual and efficiency of parameter estimates in generalized linear model. Journal of Statistical Planning and Inference, 141(2):1014–1020, February 2011. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375810004143.

Xiong:2012:ELI

[XL12] Xianzhu Xiong and Zhengyan Lin. Empirical likelihood inference for probability density functions under association. Journal of Statistical Planning and Inference, 142(4):986– 992, April 2012. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375811004113.

Xu:2011:OPF

[XLJ11] Li-Wen Xu, Min Lu, and Chun-Fu Jiang. Optimal prediction in finite populations under matrix loss. Journal of Statistical Planning and Inference, 141(8):2503–2512, August 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810005379.

Xu:2010:MAB

[XM10] Hongquan Xu and Robert W. Mee. Minimum aberration blocking schemes for 128-run designs. Journal of Statistical Planning and Inference, 140(11):3213–3229, November 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837581000176X.

Xu:2011:RMC

[XR11] Ximing Xu and N. Reid. On the robustness of maximum composite likelihood estimate. Journal of Statistical Planning and Inference, 141(9):3047–3054, September 2011. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811001236.

Xu:2014:AMR

[XSJ14] Zhibin Xu, Liuquan Sun, and Hong Ji. Additive mean residual life model with covariate measurement errors. Journal of Statistical Planning and Inference, 153(??):87–99, October 2014. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375814000810.

Xu:2015:ROC

[XT15] Ancha Xu and Yincai Tang. Reference optimality criterion for planning accelerated life testing. Journal of Statistical Planning and Inference, 167(??):14–26, December 2015. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375815001160.

Xu:2017:MFF

[Xu17] Kai Xu. Model-free feature screening via a modified composite quantile correlation. Journal of Statistical Planning and Inference, 188(??):22–35, September 2017. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375817300411.

Xiang:2013:BCR

[XW13a] Fei Xiang and Stephen G. Walker. Bayesian consistency for regression models under a supremum distance. Journal of Statistical Planning and Inference, 143(3):468–478, March 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375812002960.

Xu:2013:EMF

[XW13b] Tu Xu and Junhui Wang. An efficient model-free estimation of multiclass conditional probability. Journal of Statistical Planning and Inference, 143(12):2079–2088, December 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375813001882.

Xue:2018:UAS

[XWY18] Yuan Xue, Qin Wang, and Xiangrong Yin. A unified approach to sufficient dimension reduction. Journal of Statistical Planning and Inference, 197(??):168–179, December 2018. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375818300168.

Xue:2011:ELG

[XXC11] Dong Xue, Liugen Xue, and Weihu Cheng. Empirical likelihood for generalized linear models with missing responses. Journal of Statistical Planning and Inference, 141(6):2007– 2020, June 2011. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S037837581100022X.

Xu:2012:RVS

[XXWL12] Ganggang Xu, Yanbiao Xiang, Suojin Wang, and Zhengyan Lin. Regularization and variable selection for infinite variance autoregressive models. Journal of Statistical Planning and Inference, 142(9):2545–2553, September 2012. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375812001309.

Xiong:2019:RQL

[XZ19] Lanyu Xiong and Fukang Zhu. Robust quasi-likelihood estimation for the negative binomial integer-valued GARCH(1,1) model with an application to transaction counts. Journal of Statistical Planning and Inference, 203(??):178–198, December 2019. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375819300321.

Xiao:2018:LSE

[XZZ18] Weilin Xiao, Xili Zhang, and Ying Zuo. Least squares estimation for the drift parameters in the sub-fractional Vasicek processes. Journal of Statistical Planning and Inference, 197(??):141–155, December 2018. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375818300053.

Yang:2015:CWT

[Yan15] Yuping Yang. A CLT for weighted time-dependent uniform empirical processes. Journal of Statistical Planning and Inference, 157–158(??):36–53, February/March 2015. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375814001700.

Yang:2017:TMA

[Yan17] Xiangdong Yang. A theorem of Malliavin applied to the uniqueness of probabilistic moments. Journal of Statistical Planning and Inference, 184(??):18–24, May 2017. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375816301392.

Yao:2010:PLM

[Yao10] Weixin Yao. A profile likelihood method for normal mixture with unequal variance. Journal of Statistical Planning and Inference, 140(7):2089–2098, July 2010. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810000777.

Yu:2018:ODL

[YAW18] Jun Yu, Mingyao Ai, and Yaping Wang. Optimal designs for linear models with Fredholm-type errors. Journal of Statistical Planning and Inference, 194(??):65–74, March 2018. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375817301751.

Yuan:2010:CAG

[YBB10] Ao Yuan, Paul Bezandry, and George Bonney. Compound adaptive GPU design for clinical trials. Journal of Statistical Planning and Inference, 140(11):3505–3519, November 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375810002697.

Yang:2011:MDD

[YC11a] Su-Fen Yang and Wan-Yun Chen. Monitoring and diagnosing dependent process steps using VSI control charts. Journal of Statistical Planning and Inference, 141(5):1808–1816, May 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375810005306.

Yu:2011:MLD

[YC11b] Chi Wai Yu and Bertrand Clarke. Median loss decision theory. Journal of Statistical Planning and Inference, 141(2):611–623, February 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375810004088.

Yue:2012:SSS

[YC12] Jack C. Yue and Murray K. Clayton. Sequential sampling in the search for new shared species. Journal of Statistical Planning and Inference, 142(5):1031–1039, May 2012. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811003958.

Yu:2016:SDB

[YCB16] Xiaoli Yu, Jiahua Chen, and Rollin Brant. Sequential design for binary dose-response experiments. Journal of Statistical Planning and Inference, 177(??):64–73, October 2016. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837581530152X.

Yu:2015:ODR

[YCY15] Chun Yu, Kun Chen, and Weixin Yao. Outlier detection and robust mixture modeling using nonconvex penalized likelihood. Journal of Statistical Planning and Inference, 164(??):27–38, September 2015. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375815000543.

Yu:2013:CAS

[YDZA13] Keming Yu, Wei Dang, Huiming Zhu, and Rahim Al Hamzawi. Comment on article by Spokoiny, Wang and Härdle. Journal of Statistical Planning and Inference, 143(7):1140–1144, July 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S037837581300058X.

Yung:2012:CPB

[YH12] Wesley Yung and David Haziza. Comments on the paper “Bias-adjustment and calibration of jackknife variance estimator in the presence of non-response”. Journal of Statistical Planning and Inference, 142(7):2232–2240, July 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811004472.

Yu:2019:ASE

[YH19] Chaoyu Yu and Peter D. Hoff. Adaptive sign error control. Journal of Statistical Planning and Inference, 201(??):133– 145, July 2019. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375819300023.

Yamamoto:2019:CFO

[YHS19] Hirotaka Yamamoto, Masatake Hirao, and Masanori Sawa. A construction of the fourth order rotatable designs invariant under the hyperoctahedral group. Journal of Statistical Planning and Inference, 200(??):63–73, May 2019. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375818302581.

Yang:2012:BET

[YHWL12] Wenzhi Yang, Shuhe Hu, Xuejun Wang, and Nengxiang Ling. The Berry-Esséen type bound of sample quantiles for strong mixing sequence. Journal of Statistical Planning and Inference, 142(3):660–672, March 2012. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811003491.

Yamazaki:2015:IFS

[YK15] Daisuke Yamazaki and Eiji Kurozumi. Improving the finite sample performance of tests for a shift in mean. Journal of Statistical Planning and Inference, 167(??):144–173, December 2015. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375815000968.

Yousef:2013:SNB

[YKH13] A. S. Yousef, A. C. Kimber, and H. I. Hamdy. Sensitivity of normal-based triple sampling sequential point estimation to the normality assumption. Journal of Statistical Planning and Inference, 143(9):1606–1618, September 2013. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375813000724.

Yan:2010:CDC

[YL10a] Zhiwu Yan and Charles Locke. Crossover designs for comparing test treatments to a control treatment when subject effects are random. Journal of Statistical Planning and Inference, 140(5):1214–1224, May 2010. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809003620.

Yi:2010:SPP

[YL10b] Wende Yi and Stephen Shaoyi Liao. Statistical properties of parametric estimators for Markov chain vectors based on copula models. Journal of Statistical Planning and Inference, 140(6):1465–1480, June 2010. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809003723.

Yin:2013:OLH

[YL13] Yuhui Yin and Min-Qian Liu. Orthogonal Latin hypercube designs for Fourier-polynomial models. Journal of Statistical Planning and Inference, 143(2):307–314, February 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375812002716.

Yang:2014:BTL

[YL14]  Po Yang and William Li.  Blocked two-level semi-

foldover designs. Journal of Statistical Planning and Inference, 144(??):133–140, January 2014. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375813000037.

Yu:2016:ODS

[YLC^{+}16] Jichang Yu, Yanyan Liu, Jianwen Cai, Dale P. Sandler, and Haibo Zhou. Outcome-dependent sampling design and inference for Cox’s proportional hazards model. Journal of Statistical Planning and Inference, 178(??):24–36, November 2016. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375816300404.

Yin:2014:SLH

[YLL14] Yuhui Yin, Dennis K. J. Lin, and Min-Qian Liu. Sliced Latin hypercube designs via orthogonal arrays. Journal of Statistical Planning and Inference, 149(??):162–171, June 2014. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375814000287.

Yuan:2017:NCO

[YLL17] Ru Yuan, Dennis K. J. Lin, and Min-Qian Liu. Nearly column-orthogonal designs based on leave-one-out good lattice point sets. Journal of Statistical Planning and Inference, 185(??):29–40, June 2017. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375817300101.

Yanagimoto:2011:SCP

[YO11] Takemi Yanagimoto and Toshio Ohnishi. Saddlepoint condition on a predictor to reconfirm the need for the assumption of a prior distribution. Journal of Statistical Planning and Inference, 141(5):1990–2000, May 2011. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810005690.

Yue:2011:OTD

[YQC11] Rong-Xian Yue, Hong Qin, and Kashinath Chatterjee. Optimal U-type design for Bayesian nonparametric multiresponse prediction. Journal of Statistical Planning and Inference, 141(7):2472–2479, July 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S037837581100067X.

Yan:2011:IIL

[YS11a] Guofen Yan and J. Sedransk. Improved inference for a linear mixed-effects model when the subpopulation effects are clustered. Journal of Statistical Planning and Inference, 141(11):3489–3497, November 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375811001984.

Yothers:2011:ESC

[YS11b] Greg Yothers and Allan R. Sampson. Erratum to “Simultaneous confidence band for the difference of segmented linear models” [Journal of Statistical Planning and Inference 141(2) (February) (2011) 1059–1068]. Journal of Statistical Planning and Inference, 141(5):2004–2005, May 2011. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810005422. See [YS11c].

Yothers:2011:SCB

[YS11c] Greg Yothers and Allan R. Sampson. Simultaneous confidence band for the difference of segmented linear models. Journal of Statistical Planning and Inference, 141(2):1059–1068, February 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/

S0378375810004222. See erratum [YS11b].

Suen:2010:SOS

[ySD10] Chung yi Suen and Ashish Das. E(s2)-optimal supersaturated designs with odd number of runs. Journal of Statistical Planning and Inference, 140(6):1398–1409, June 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809003115.

Suen:2013:OFF

[ySDM13] Chung yi Suen, Ashish Das, and C. K. Midha. Optimal fractional factorial designs and their construction. Journal of Statistical Planning and Inference, 143(10):1828–1834, October 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375813001079.

You:2010:IEQ

[YSGN10] Alexandre You, Ulrike Schneider, Armelle Guillou, and Philippe Naveau. Improving extreme quantile estimation via a folding procedure. Journal of Statistical Planning and Inference, 140(7):1775–1787, July 2010. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810000182.

Yang:2011:TMH

[YSH11] Zhao Yang, Xuezheng Sun, and James W. Hardin. Testing marginal homogeneity in clustered matched-pair data. Journal of Statistical Planning and Inference, 141(3):1313–1318, March 2011. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375810004714.

Yao:2010:EAT

[YSW10] Lili Yao, Zhihua Sun, and Qihua Wang. Estimation of average treatment effects based on parametric propensity score model. Journal of Statistical Planning and Inference, 140(3):806–816, March 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375809002900.

Yu:2013:MAP

[YTHL13] Yao Yu, Sally W. Thurston, Russ Hauser, and Hua Liang. Model averaging procedure for partially linear single-index models. Journal of Statistical Planning and Inference, 143(12):2160–2170, December 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375813001055.

Yao:2013:NTP

[YU13] Feng Yao and Aman Ullah. A nonparametric R ^{2} test for the presence of relevant variables. Journal of Statistical Planning and Inference, 143(9):1527–1547, September 2013. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375813000712.

Yuan:2013:CGS

[Yua13] Ming Yuan. Comments on “Grouping strategies and thresholding for high dimension linear models”. Journal of Statistical Planning and Inference, 143(9):1454–1456, September 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375813000499.

Yuan:2010:LBM

[YW10] De Mei Yuan and Xiu Shan Wu. Limiting behavior of the maximum of the partial sum for asymptotically negatively associated random variables under residual Cesàro alpha-integrability assumption. Journal of Statistical Planning and Inference, 140(9):2395–2402, September 2010. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810000844.

Yang:2016:NAL

[YW16] Yuehan Yang and Lan Wu. Nonnegative adaptive lasso for ultra-high dimensional regression models and a two-stage method applied in financial modeling. Journal of Statistical Planning and Inference, 174(??):52–67, July 2016. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375816000215.

Ye:2010:SIM

[YY10] Zhishen Ye and Jie Yang. Sliced inverse moment regression using weighted chi-squared tests for dimension reduction. Journal of Statistical Planning and Inference, 140(11):3121–3131, November 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375810001825.

Yang:2014:ALP

[YY14] Hu Yang and Jing Yang. The adaptive L1-penalized LAD regression for partially linear single-index models. Journal of Statistical Planning and Inference, 151–152 (??):73–89, August/September 2014. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375814000585.

Yang:2019:SIS

[YYX19] Guangren Yang, Weixin Yao, and Sijia Xiang. Sure independence screening in ultrahigh dimensional generalized additive models. Journal of Statistical Planning and Inference, 199(??):126–135, March 2019. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S037837581830065X.

Yang:2014:SEL

[YYZ14] Hanfang Yang, Crystal Yau, and Yichuan Zhao. Smoothed empirical likelihood inference for the difference of two quantiles with right censoring. Journal of Statistical Planning and Inference, 146(??):95–101, March 2014. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375813002346.

Yang:2012:NEL

[YZ12] Hanfang Yang and Yichuan Zhao. New empirical likelihood inference for linear transformation models. Journal of Statistical Planning and Inference, 142(7):1659–1668, July 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375812000523.

Yang:2017:AUD

[YZZ17] Feng Yang, Yong-Dao Zhou, and Xue-Ru Zhang. Augmented uniform designs. Journal of Statistical Planning and Inference, 182(??):61–73, March 2017. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375816301112.

Zhao:2011:MOP

[ZB11] Peng Zhao and N. Balakrishnan. MRL ordering of parallel systems with two heterogeneous components. Journal of Statistical Planning and Inference, 141(2):631–638, February 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375810003605.

Zhang:2010:GML

[ZC10] Runchu Zhang and Yi Cheng. General minimum lower order confounding designs: an overview and a construction theory. Journal of Statistical Planning and Inference, 140(7):1719– 1730, July 2010. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375809004042.

Zhang:2011:SSG

[ZC11] Yong Zhang and Kejun Chen. Super-simple group divisible designs with block size 4 and index 9. Journal of Statistical Planning and Inference, 141(9):3231–3243, September 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375811001522.

Zhao:2012:MTF

[ZC12] Shengli Zhao and Xiangfei Chen. Mixed two- and four-level fractional factorial split-plot designs with clear effects. Journal of Statistical Planning and Inference, 142(7):1789–1793, July 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375812000894.

Zhang:2017:SIM

[ZCLZ17] Jun Zhang, Qian Chen, Bingqing Lin, and Yan Zhou. On the single-index model estimate of the conditional density function: Consistency and implementation. Journal of Statistical Planning and Inference, 187(??):56–66, August 2017. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375817300149.

Zhao:2011:PWS

[ZCN11] Peng Zhao, Ping Shing Chan, and Hon Keung Tony Ng. Peakedness for weighted sums of symmetric random variables. Journal of Statistical Planning and Inference, 141(5):1737– 1743, May 2011. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375810005239.

Zhang:2013:CVT

[ZCP13] Shunpu Zhang, Huann-Sheng Chen, and Ruth M. Pfeiffer. A combined p-value test for multiple hypothesis testing. Journal of Statistical Planning and Inference, 143(4):764–770, April 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375812003199.

Zou:2011:LPT

[ZCX11] Bin Zou, Rong Chen, and Zongben Xu. Learning performance of Tikhonov regularization algorithm with geometrically beta-mixing observations. Journal of Statistical Planning and Inference, 141(3):1077–1087, March 2011. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810003587.

Zhang:2019:MAS

[ZDLW19] Qiao-Zhen Zhang, Hong-Sheng Dai, Min-Qian Liu, and Ya Wang. A method for augmenting supersaturated designs. Journal of Statistical Planning and Inference, 199(??):207– 218, March 2019. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375818301137.

Zheng:2012:NAS

[ZG12] Shimin Zheng and A. K. Gupta. A new approach to statistical efficiency of weighted least squares fitting algorithms for reparameterization of nonlinear regression models. Journal of Statistical Planning and Inference, 142(4):1001–1008, April 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375811004162.

Zheng:2013:APQ

[ZGK13] Qi Zheng, Colin Gallagher, and K. B. Kulasekera. Adaptive penalized quantile regression for high dimensional data. Journal of Statistical Planning and Inference, 143(6):1029– 1038, June 2013. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375812003680.

Zheng:2012:LBA

[ZGST12] Shurong Zheng, Jianhua Guo, Ning-Zhong Shi, and Guo-Liang Tian. Likelihood-based approaches for multivariate linear models under inequality constraints for incomplete data. Journal of Statistical Planning and Inference, 142(11):2926–2942, November 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375812001802.

Zhu:2017:SPS

[ZGW^{+}17] Dekang Zhu, Dan P. Guralnik, Xuezhi Wang, Xiang Li, and Bill Moran. Statistical properties of the single linkage hierarchical clustering estimator. Journal of Statistical Planning and Inference, 185(??):15–28, June 2017. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375816301689.

Zhou:2017:HDG

[ZGZ17] Bu Zhou, Jia Guo, and Jin-Ting Zhang. High-dimensional general linear hypothesis testing under heteroscedasticity. Journal of Statistical Planning and Inference, 188(??):36–54, September 2017. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S037837581730040X.

Zahid:2012:RSE

[ZH12] Faisal Maqbool Zahid and Christian Heumann. Response shrinkage estimation in multinomial logit models. Journal of Statistical Planning and Inference, 142(1):95–109, January 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375811002709.

Zhang:2010:OCG

[Zha10] Zhengcheng Zhang. Ordering conditional general coherent systems with exchangeable components. Journal of Statistical Planning and Inference, 140(2):454–460, February 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809002511.

Zhang:2012:ADF

[Zha12] Jin-Ting Zhang. An approximate degrees of freedom test for heteroscedastic two-way ANOVA. Journal of Statistical Planning and Inference, 142(1):336–346, January 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811003089.

Zhao:2013:ACD

[Zha13] Junlong Zhao. Asymptotic convergence of dimension reduction based boosting in classification. Journal of Statistical Planning and Inference, 143(4):651–662, April 2013. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375812003229.

Zhang:2015:NFF

[Zha15] Jian Zhang. On nonparametric feature filters in electromagnetic imaging. Journal of Statistical Planning and Inference, 164(??):39–53, September 2015. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375815000555.

Zhigljavsky:2010:NGT

[Zhi10] Anatoly Zhigljavsky. Nonadaptive group testing with lies: Probabilistic existence theorems. Journal of Statistical Planning and Inference, 140(10):2885–2893, October 2010. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810001242.

Zhou:2010:MEA

[Zho10] Weihua Zhou. M-estimation of the accelerated failure time model under a convex discrepancy function. Journal of Statistical Planning and Inference, 140(9):2669–2681, September 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375810001412.

Zhou:2019:AMC

[Zho19] Quan Zhou. Asymptotics of multivariate contingency tables with fixed marginals. Journal of Statistical Planning and Inference, 198(??):165–170, January 2019. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375818300363.

Zhu:2012:ZIP

[Zhu12] Fukang Zhu. Zero-inflated Poisson and negative binomial integer-valued GARCH models. Journal of Statistical Planning and Inference, 142(4):826–839, April 2012. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837581100379X.

Zhu:2010:NBT

[ZJ10] Rong Zhu and Harry Joe. Negative binomial time series models based on expectation thinning operators. Journal of Statistical Planning and Inference, 140(7):1874–1888, July 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837581000042X.

Zhou:2015:THD

[ZK15] Cheng Zhou and Xin-Bing Kong. Testing of high dimensional mean vectors via approximate factor model. Journal of Statistical Planning and Inference, 167(??):216–227, December 2015. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375815000762.

Zambom:2017:LSM

[ZK17] Adriano Z. Zambom and Seonjin Kim. Lag selection and model specification testing in nonparametric autoregressive conditional heteroscedastic models. Journal of Statistical Planning and Inference, 186(??):13–27, July 2017. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375817300174.

Zhang:2011:BET

[ZL11a] Jing-Jing Zhang and Han-Ying Liang. Berry–Esseen type bounds in heteroscedastic semi-parametric model. Journal of Statistical Planning and Inference, 141(11):3447–3462, November 2011. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375811001923.

Zhang:2011:FBM

[ZL11b] Qiang Zhang and Dejian Lai. Fractional Brownian motion and long term clinical trial recruitment. Journal of Statistical Planning and Inference, 141(5):1783–1788, May 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810005288.

Zhao:2012:LDP

[ZL12] Shoujiang Zhao and Qiaojing Liu. Large deviations for parameter estimators of α-Brownian bridge. Journal of Statistical Planning and Inference, 142(3):695–707, March 2012.

CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837581100351X.

Zhang:2017:RBL

[ZL17] Feipeng Zhang and Qunhua Li. Robust bent line regression. Journal of Statistical Planning and Inference, 185(??):41–55, June 2017. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375817300095.

Zhou:2018:VCP

[ZL18] Zhiyong Zhou and Zhengyan Lin. Varying coefficient partially nonlinear models with nonstationary regressors. Journal of Statistical Planning and Inference, 194(??):47–64, March 2018. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375817301763.

Zhao:2016:NCB

[ZLL16a] Shengli Zhao, Dennis K. J. Lin, and Pengfei Li. A note on the construction of blocked two-level designs with general minimum lower order confounding. Journal of Statistical Planning and Inference, 172(??):16–22, May 2016. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375815002165.

Zheng:2016:FAC

[ZLL16b] Yao Zheng, Yang Li, and Guodong Li. On Fréchet autoregressive conditional duration models. Journal of Statistical Planning and Inference, 175(??):51–66, August 2016. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375816000410.

Zhao:2017:GAL

[ZLL17] Weihua Zhao, Heng Lian, and Hua Liang. GEE analysis for longitudinal single-index quantile regression. Journal of Statistical Planning and Inference, 187(??):78–102, August 2017. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375817300344.

Zhao:2012:DRE

[ZLM12] Yang Zhao, Jerald F. Lawless, and Donald L. McLeish. Design and relative efficiency in two-phase studies. Journal of Statistical Planning and Inference, 142(11):2953–2964, November 2012. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375812001759.

Zhao:2017:RRR

[ZLM17] Weihua Zhao, Heng Lian, and Shujie Ma. Robust reducedrank modeling via rank regression. Journal of Statistical Planning and Inference, 180(??):1–12, January 2017. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375816301045.

Zhu:2010:MIV

[ZLW10] Fukang Zhu, Qi Li, and Dehui Wang. A mixture integervalued ARCH model. Journal of Statistical Planning and Inference, 140(7):2025–2036, July 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375810000613.

Zhang:2016:NMJ

[ZLZ^{+}16] Bo Zhang, Wei Liu, Hui Zhang, Qihui Chen, and Zhiwei Zhang. A note on misspecification in joint modeling of correlated data with informative cluster sizes. Journal of Statistical Planning and Inference, 170(??):46–63, March 2016. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375815300185.

Zhao:2013:CBT

[ZLZK13] Shengli Zhao, Pengfei Li, Runchu Zhang, and Rohana Karunamuni. Construction of blocked two-level regular designs with general minimum lower order confounding. Journal of Statistical Planning and Inference, 143(6):1082–1090, June 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375813000025.

Zhang:2016:CQR

[ZLZL16] Riquan Zhang, Yazhao Lv, Weihua Zhao, and Jicai Liu. Composite quantile regression and variable selection in single-index coefficient model. Journal of Statistical Planning and Inference, 176(??):1–21, September 2016. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375816300180.

Zhou:2012:PCE

[ZN12] Haojin Zhou and Tapan K. Nayak. Pitman closest equivariant estimators and predictors under location-scale models. Journal of Statistical Planning and Inference, 142(6):1367– 1377, June 2012. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375811004599.

Zhao:2013:TCC

[ZN13] Yichuan Zhao and Duytrac Nguyen. Tests for comparison of competing risks under the additive risk model. Journal of Statistical Planning and Inference, 143(4):842–851, April 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375812003266.

Zou:2014:GKD

[Zou14] Hui Zou. Generalizing Koenker’s distribution. Journal of Statistical Planning and Inference, 148(??):123–127, May 2014. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375813003066.

Zhang:2015:BCS

[ZPS15] Zhengwu Zhang, Debdeep Pati, and Anuj Srivastava. Bayesian clustering of shapes of curves. Journal of Statistical Planning and Inference, 166(??):171–186, November 2015. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375815000798.

Zhang:2010:EAM

[ZS10] Zhengjun Zhang and Richard L. Smith. On the estimation and application of max-stable processes. Journal of Statistical Planning and Inference, 140(5):1135–1153, May 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809003516.

Zhang:2015:ESP

[ZS15] Yuanqing Zhang and Dianmei Shen. Estimation of semiparametric varying-coefficient spatial panel data models with random-effects. Journal of Statistical Planning and Inference, 159(??):64–80, April 2015. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375814001803.

ZafarYab:2010:PGN

[ZSA10] Muhammad ZafarYab, Farrukh Shehzad, and Rashid Ahmed. Proper generalized neighbor designs in circular blocks. Journal of Statistical Planning and Inference, 140(11):3498–3504, November 2010. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375810002685.

Zevallos:2012:NID

[ZSH12] Mauricio Zevallos, Bruno Santos, and Luiz K. Hotta. A note on influence diagnostics in AR(1) time series models. Journal of Statistical Planning and Inference, 142(11):2999–3007, November 2012. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375812001863.

Zhou:2014:OBE

[ZSL14] Rong Zhou, Siva Sivaganesan, and Martial Longla. An objective Bayesian estimation of parameters in a logbinomial model. Journal of Statistical Planning and Inference, 146(??):113–121, March 2014. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375813002309.

Zheng:2010:SID

[ZSM10] Shurong Zheng, Ning-Zhong Shi, and Wenqing Ma. Statistical inference on difference or ratio of means from heteroscedastic normal populations. Journal of Statistical Planning and Inference, 140(5):1236–1242, May 2010. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809003644.

Zhao:2011:SFQ

[ZW11] Wenle Zhao and Yanqiu Weng. A simplified formula for quantification of the probability of deterministic assignments in permuted block randomization. Journal of Statistical Planning and Inference, 141(1):474–478, January 2011. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810003162.

Zhang:2012:SPL

[ZW12a] Tao Zhang and Qihua Wang. Semiparametric partially linear regression models for functional data. Journal of Statistical Planning and Inference, 142(9):2518–2529, September 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375812001206.

Zhao:2012:RPL

[ZW12b] Xin Zhao and Martin T. Wells. Reference priors for linear models with general covariance structures. Journal of Statistical Planning and Inference, 142(8):2473–2484, August 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375812000146.

Zuo:2012:APM

[ZWM12] Jianying Zuo, Huaiqing Wu, and William Q. Meeker. Asymptotic properties of mean cumulative function estimators from window-observation recurrence data. Journal of Statistical Planning and Inference, 142(11):2943–2952, November 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375812001607.

Zhou:2011:CLS

[ZWY11] Yong Zhou, Alan T. K. Wan, and Yuan Yuan. Combining least-squares and quantile regressions. Journal of Statistical Planning and Inference, 141(12):3814–3828, December 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375811002436.

Zhang:2010:IIP

[ZWZ10] Haixiang Zhang, Dehui Wang, and Fukang Zhu. Inference for INAR(p) processes with signed generalized power series thinning operator. Journal of Statistical Planning and Inference, 140(3):667–683, March 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375809002572.

Zhao:2012:NTB

[ZX12] Jianxin Zhao and Xingzhong Xu. New tests based on Rubin’s empirical distribution function. Journal of Statistical Planning and Inference, 142(6):1356–1366, June 2012. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811004587.

Zhang:2019:WAE

[ZX19] Dongxue Zhang and Zhiming Xia. Weighted-averaging estimator for possible threshold in segmented linear regression model. Journal of Statistical Planning and Inference, 200(??):102–118, May 2019. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375818302763.

Zhao:2012:FIU

[ZXD12] Shuran Zhao, Xingzhong Xu, and Xiaobo Ding. Fiducial inference under nonparametric situations. Journal of Statistical Planning and Inference, 142(10):2779–2798, October 2012. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375812001474.

Zheng:2011:ELM

[ZY11] Ming Zheng and Wen Yu. Empirical likelihood method for the multivariate accelerated failure time models. Journal of Statistical Planning and Inference, 141(2):972–983, February 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S037837581000412X.

Zhang:2014:NME

[ZYZL14] Jun Zhang, Yao Yu, Bu Zhou, and Hua Liang. Nonlinear measurement error models subject to additive distortion. Journal of Statistical Planning and Inference, 150(??):49–65, July 2014. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S037837581400041X.

Zhang:2010:RPM

[ZZ10] Zhiyi Zhang and Jun Zhou. Re-parameterization of multinomial distributions and diversity indices. Journal of Statistical Planning and Inference, 140(7):1731–1738, July 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809004030.

Zhao:2014:WVP

[ZZ14a] Haibing Zhao and Jiajia Zhang. Weighted p-value procedures for controlling FDR of grouped hypotheses. Journal of Statistical Planning and Inference, 151–152(??):90–106, August/ September 2014. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375814000597.

Zhou:2014:GGM

[ZZ14b] Zhiyang Zhou and Runchu Zhang. A generalized general minimum lower order confounding criterion for nonregular designs. Journal of Statistical Planning and Inference, 148(??):95–100, May 2014. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0378375813003054.

Zhang:2013:SIP

[ZZOY13] Hong Zhang, Donglin Zeng, Sylviane Olschwang, and Kai Yu. Semiparametric inference on the penetrances of rare genetic mutations based on a case-family design. Journal of Statistical Planning and Inference, 143(2):368–377, February 2013. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375812002741.

Zhao:2012:SMR

[ZZW12] XiaoBing Zhao, Xian Zhou, and JingLong Wang. Semiparametric model for recurrent event data with excess zeros and informative censoring. Journal of Statistical Planning and Inference, 142(1):289–300, January 2012. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375811003016.

Zhao:2019:SNS

[ZZW19] Yanyan Zhao, Changliang Zou, and Zhaojun Wang. A scalable nonparametric specification testing for massive data. Journal of Statistical Planning and Inference, 200(??):161– 175, May 2019. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375818303069.